### v1

# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

Abstract
We study continuous-time portfolio choice in diffusion markets with parameter $\theta\in\Theta$ and uncertainty law $q(d\theta)$. Nature draws latent $\theta\sim q$ at time 0; the investor cannot observe it and must deploy a single $\theta$-blind feedback policy maximizing an ex–ante CRRA objective averaged over diffusion noise and $\theta$. Our methods access $q$ only by sampling and assume no parametric form. We extend Pontryagin–Guided Direct Policy Optimization (PG–DPO) by sampling $\theta$ inside the simulator and computing discrete-time gradients via backpropagation through time (BPTT), and we propose projected PG–DPO (P–PGDPO) that projects costate estimates to satisfy the $q$-aggregated Pontryagin first-order condition, yielding a deployable rule. We prove a BPTT–PMP correspondence uniform on compacts and a residual-based $\theta$-blind policy-gap bound under local stability with explicit discretization/Monte Carlo errors; experiments show projection-driven stability and accurate decision-time benchmark recovery in high dimensions.
1 Introduction
A central problem in quantitative finance is to allocate wealth dynamically across many risky assets in continuous time. In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and the investor solves a Hamilton–Jacobi–Bellman (HJB) equation to obtain closed-form optimal portfolios and value functions; see, for example, merton1969lifetime; merton1971optimum and the subsequent literature. In realistic markets, however, neither expected returns nor volatilities are known: they must be estimated from finite samples, often using many assets and predictors and under nontrivial model selection and regularization. Empirically, expected-return forecasting is fragile and return predictability is unstable across samples and over time, with many proposed predictors delivering limited out-of-sample gains (e.g., goyalwelch2008; campbellthompson2008; rapach2010forecast; danglhalling2012tvp; lettauvanNieuwerburgh2008shifts; pettenuzzo2014constraints). In such settings it is crucial to distinguish diffusion risk (Brownian noise conditional on parameters) from statistical parameter uncertainty (error in estimated coefficients). A long line of portfolio-choice work shows that return predictability, learning, and parameter uncertainty can induce substantial intertemporal hedging effects and more conservative allocations (e.g., kandelstambaugh1996predictability; barberis2000investor; campbell2002strategic; brandt2005portfolio; brennan1998investor; xia2001learning; maenhout2004robust).
We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, where the estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$. We treat $q$ as an input object that encapsulates all statistical information available at time 0: it may be derived from resampling approximations (e.g. bootstrap) (e.g., efron1979bootstrap; efron1994bootstrap), model averaging or Bayesian model uncertainty pipelines (e.g., pastor2000portfolio; avramov2002stock; cremers2002stock), approximate posteriors, or other procedures. Our goal is not to revisit inference, but to optimize decisions given this uncertainty description. Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside the simulator; we do not assume closed-form densities, conjugate updates, or any particular parametric form. Concretely, we seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$. This formulation also supports offline diagnostics that quantify how recommended allocations vary across the statistically plausible models encoded in $q$.
A key modeling choice is that $\theta$ is latent: Nature draws a fixed but unobserved $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed on $[0,T]$. While the investor knows $q$, she does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy. We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a belief/posterior process. This fixed-$q$ commitment is intended as a decision-time benchmark: it targets a single deployable rule given an exogenous uncertainty law, cleanly decoupling how uncertainty is produced (any pipeline yielding $q$) from how decisions are optimized (our solver given $q$). We do not claim that belief-state control is conceptually inappropriate; rather, it defines a different (and typically far more demanding) problem than computing a single $\theta$-blind deployable feedback rule from a fixed uncertainty description (e.g., bensoussan1985optimal; pham2017dynamic). At the same time, fixed-$q$ optimization couples heterogeneous market models: gradient signals can vary substantially across $\theta$ draws and may partially cancel when learning a single global policy end-to-end.
The $\theta$-blind constraint also changes what a first-order optimality condition means. If $\theta$ were observable, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition and an associated $\theta$-conditional full-information feedback map (infeasible under latent $\theta$). Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind. Taking the first variation of the ex–ante objective and using Fubini’s theorem shows that the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the state process, in the standard stochastic maximum principle framework (e.g., yong1999stochastic; fleming2006controlled; pham2009continuous). Because $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$ for our portfolio Hamiltonian, this aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projected portfolio rule. Notably, the condition and resulting projection are agnostic to the internal construction of $q$ and depend only on its role as the ex–ante mixing law.
These features place the problem outside the practical reach of classical dynamic programming in the high-dimensional regime we target. In low-dimensional deterministic-parameter Markov models, DP/HJB is canonical; however, even with several factors it requires solving an HJB equation in the state $(t,X_{t},Y_{t})$, where grid-based PDE methods are quickly defeated by the curse of dimensionality (e.g., bellman1961adaptive; kushner2001numerical). Deep PDE surrogates such as PINNs (e.g., raissi2019physics; sirignano2018dgm) and deep BSDE methods (e.g., han2018solving; beck2019machine) alleviate the need for grids, but fully nonlinear portfolio HJBs with many assets and factors remain numerically delicate, especially when accurate mixed derivatives are required. If one further models parameter uncertainty via belief-state augmentation, the state becomes a posterior measure and the control problem becomes infinite-dimensional (e.g., bensoussan1985optimal; pham2017dynamic).
Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (huh2025breaking; huh2025constraint). PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and employs backpropagation through time (BPTT) to compute exact gradients of terminal utility. Crucially, intermediate pathwise sensitivities computed by BPTT coincide with the stochastic costates (adjoints) in PMP, mirroring the classical duality between backpropagation and adjoint methods (see lecun1988optimal; yong1999stochastic). In the latent-parameter setting, we approximate the ex–ante objective by sampling $\theta\sim q$ inside the simulator and fixing it along each trajectory, while the policy depends only on observable states. To stabilize learning under heterogeneous $\theta$ draws, we extend the projected variant, P–PGDPO, to latent $\theta$: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated Pontryagin first-order condition. This reconstruction yields a robust deployable $\theta$-blind rule obtained from the $q$-aggregated criticality, and can be amortized into a fast-to-evaluate policy.
In high-dimensional scaling experiments under static Gaussian drift uncertainty, the two-stage projected pipeline substantially improves decision-time accuracy relative to end-to-end learning, with clear stabilization effects in aligned regimes. In misaligned regimes, projection gains diminish with dimension; diagnostics indicate that deterioration is driven primarily by growth of aggregated first-order residuals and curvature mismatch rather than by catastrophic numerical inversion. In factor-driven markets with mean-reverting investment opportunities where return–factor correlation induces intertemporal hedging demand, the projected pipeline recovers the analytic decision-time benchmark under the same $\theta$-blind deployability restriction, while a model-free PPO baseline remains far from the reference in the regimes we test.
Our main theoretical guarantee is a residual-based ex–ante $\theta$-blind policy-gap bound for the deployable fixed-$q$ commitment problem: under mild slab-wise local stability conditions for the $q$-aggregated projection map, a small warm-up aggregated first-order residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal interior deployable $\theta$-blind policy, up to discretization and Monte Carlo error.
Our contributions are threefold. (i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated Pontryagin first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$. (ii) We extend PG–DPO to this setting by sampling $\theta$ only inside the simulator and using BPTT to compute exact discrete-time gradients and pathwise sensitivities, and we establish a conditional BPTT–PMP correspondence uniform over $\theta$ on compact subsets of $\Theta$. (iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization and stability gains from projection.
The remainder of the paper is organized as follows. Section 2 formulates the fixed-$q$ ex–ante portfolio problem under a latent parameter and a deployable $\theta$-blind Markov feedback restriction, and derives the $\theta$-conditional versus $q$-aggregated Pontryagin first-order conditions together with Gaussian decision-time reference models. Section 3 develops PG–DPO and uncertainty-aware P–PGDPO for the latent-$\theta$ setting, establishes the conditional BPTT–PMP correspondence, and proves a residual-based ex–ante $\theta$-blind policy-gap bound under local stability of the aggregated projection map. Section 4 reports high-dimensional scaling experiments under static Gaussian drift uncertainty, and Section 5 studies hedging-demand recovery in factor-driven markets with mean-reverting investment opportunities. Technical proofs and implementation details are collected in the appendix.
2 Dynamic Portfolio Choice in Estimated Diffusion Markets with Latent Parameter Uncertainty
In this section we formulate a continuous-time dynamic portfolio choice problem with CRRA preferences in a diffusion market whose coefficients are estimated from data and are therefore statistically uncertain. Rather than committing to a particular estimation architecture, we treat the market as belonging to a (possibly high-dimensional) parameterized family indexed by $\theta\in\Theta$, and we represent the uncertainty in the estimated parameter by an exogenously given probability law $q(d\theta)$ over $\Theta$.
  • Nature draws a fixed but unobserved $\theta\sim q$ at time 0 and keeps it constant on $[0,T]$.
  • The investor knows $q$ but does not observe the realized $\theta$ and must deploy a single $\theta$-blind portfolio policy.
  • Performance is evaluated ex–ante by averaging terminal utility over both diffusion noise and $\theta\sim q$.
  • We restrict to $\theta$-blind Markov feedback rules $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ and do not augment the state by a belief/posterior process.
This fixed-$q$, $\theta$-blind formulation is intentionally algorithm-facing: we view the estimation procedure that produced $q(d\theta)$ as exogenous, and our goal is to compute stable, scalable portfolio rules given this uncertainty description. It is also a commitment model: the investor commits at time 0 to a single feedback map and does not update $q$ during trading. As a result, one must distinguish between (i) $\theta$-conditional (full-information) optimality conditions and objects that would be available if $\theta$ were observable (infeasible under latent $\theta$), and (ii) $q$-aggregated conditions that characterize optimality within the $\theta$-blind admissible class. Throughout, $\theta$-conditional objects are used only for offline diagnostics (e.g., heterogeneity inspection and infeasible upper bounds), whereas our algorithms target a single deployable $\theta$-blind rule; see Section 2.2.
2.1 Model and ex–ante objective in estimated diffusion markets
We work on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ supporting Brownian motions of appropriate dimension. Time is continuous and runs over a fixed finite horizon $[0,T]$.
Deterministic-parameter reference (classical CRRA Merton).
There is one risk-free asset (money market account) with price process $B$ satisfying
$\frac{dB_{t}}{B_{t}}=r\,dt,\hskip 18.49988ptB_{0}=1,$ (1)
where $r\in\mathbb{R}$ is a constant short rate. In the classical Merton model, the $d$ risky assets have prices $S_{t}=(S_{t}^{1},\dots,S_{t}^{d})^{\top}$ solving
$\frac{dS_{t}}{S_{t}}:=\bigg(\frac{dS_{t}^{1}}{S_{t}^{1}},\dots,\frac{dS_{t}^{d}}{S_{t}^{d}}\bigg)^{\top}=r\,\mathbf{1}\,dt+\mu\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (2)
with constant excess returns $\mu\in\mathbb{R}^{d}$, volatility matrix $\Sigma^{1/2}\in\mathbb{R}^{d\times d}$, and a $d$-dimensional Brownian motion $W$. An investor with CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$, chooses a progressively measurable portfolio fraction $\pi_{t}\in\mathbb{R}^{d}$; the wealth process satisfies
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\lambda\big)dt+\pi_{t}^{\top}\Sigma^{1/2}dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (3)
where $\lambda:=\mu$ is the vector of risk premia. In this benchmark setting the optimal policy is constant:
$\pi^{\star}=\frac{1}{\gamma}\,\Sigma^{-1}\lambda,$ (4)
and the corresponding value function is given explicitly by
$V^{\mathrm{Merton}}(t,x;\lambda)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)\Big(r+\tfrac{1}{2\gamma}\lambda^{\top}\Sigma^{-1}\lambda\Big)(T-t)\Big\},$ (5)
see, for example, merton1969lifetime; merton1971optimum. We use this constant-coefficient model only as a deterministic-parameter reference.
Estimated diffusion market family (conditional on a latent parameter).
In our main formulation, drift and volatility are not assumed known. Instead, we consider a general multi-asset, multi-factor diffusion family indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$, where $\theta$ represents the (possibly high-dimensional) parameter produced by an estimation procedure. Conditional on $\theta$, the $d$ risky assets and an $m$-dimensional factor process $Y_{t}$ evolve as
$\frac{dS_{t}}{S_{t}}$ $=r\,\mathbf{1}\,dt+b\big(Y_{t},\theta\big)\,dt+\sigma\big(Y_{t},\theta\big)\,dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (6)
$dY_{t}$ $=a\big(Y_{t},\theta\big)\,dt+\beta\big(Y_{t},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m},$ (7)
where $W$ and $W^{Y}$ are Brownian motions (possibly of different dimension) that may be instantaneously correlated. We write the instantaneous covariance and return–factor cross-covariance as
$\Sigma(y,\theta):=\sigma(y,\theta)\sigma(y,\theta)^{\top},\hskip 18.49988pt\Sigma_{SY}(y,\theta):=\sigma(y,\theta)\,\rho\,\beta(y,\theta)^{\top},$ (8)
where $\rho$ is defined by $d\langle W,W^{Y}\rangle_{t}=\rho\,dt$. Thus $\Sigma(y,\theta)\in\mathbb{R}^{d\times d}$ and $\Sigma_{SY}(y,\theta)\in\mathbb{R}^{d\times m}$.
Uncertainty law $q(d\theta)$ and information structure.
The parameter $\theta$ is estimated from finite samples and is uncertain. We summarize this uncertainty by a probability distribution
$q(d\theta).$ (9)
We deliberately do not tie $q$ to any specific inference paradigm. Concretely, $q$ may represent an empirical/sampling distribution produced by resampling procedures such as the bootstrap (efron1979bootstrap; efron1994bootstrap), a distribution induced by model averaging or sub-sample aggregation procedures such as bagging (breiman1996bagging), an approximate Bayesian posterior (when a prior and likelihood/criterion are specified), or an asymptotic normal (or sandwich) approximation in parametric or semiparametric estimation. For our purposes, $q$ is an input object describing statistically plausible market parameters.
Remark 1 (Latent parameter, observability, and admissible controls).
We interpret $\theta$ as a latent (unobserved) market parameter: Nature draws an $\mathcal{F}_{0}$-measurable random variable $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed over $[0,T]$. The investor knows $q$ but does not observe the realized $\theta$, so deployable portfolio rules cannot take $\theta$ as an input.
We consider the observable market filtration
$\mathcal{F}_{t}^{\mathrm{obs}}:=\sigma\{(S_{s},Y_{s}):0\leq s\leq t\},\hskip 18.49988pt0\leq t\leq T,$ (10)
where $\sigma\{\cdot\}$ denotes the $\sigma$-field generated by the observed asset and factor paths (with the usual augmentation). Admissible portfolio processes are required to be progressively measurable with respect to $(\mathcal{F}_{t}^{\mathrm{obs}})$.
Throughout the paper we restrict attention to the Markov feedback subclass
$\mathcal{A}^{\mathrm{fb}}:=\Big\{\pi\in\mathcal{A}^{\mathrm{obs}}:\ \exists\,\bar{\pi}:[0,T]\times(0,\infty)\times\mathbb{R}^{m}\to\mathbb{R}^{d}\ \text{s.t.}\ \pi_{t}=\bar{\pi}(t,X_{t},Y_{t})\Big\},$ (11)
where $\mathcal{A}^{\mathrm{obs}}$ is defined below. This restriction reflects a fixed-$q$ commitment model: the investor uses historical data to form $q$ prior to trading and does not perform online filtering/belief-state updates during $[0,T]$.
Whenever we display $\theta$-conditional (full-information) controls or sensitivity objects, they are computed under frozen-$\theta$ simulations and are used only for offline diagnostics; the deployed policy class and the learned policy remain $\theta$-blind.
Wealth dynamics and admissibility (given $\theta$).
For any fixed $\theta$, the corresponding wealth dynamics under a portfolio process $\pi_{t}(\omega)\in\mathbb{R}^{d}$ adapted to $\mathcal{F}_{t}^{\mathrm{obs}}$ are
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\Big(r+\pi_{t}^{\top}b(Y_{t},\theta)\Big)dt+\pi_{t}^{\top}\sigma(Y_{t},\theta)\,dW_{t},$ (12)
and we denote by $\mathcal{A}^{\mathrm{obs}}$ the set of progressively measurable portfolio processes adapted to $(\mathcal{F}_{t}^{\mathrm{obs}})$ for which (12) admits a (strictly) positive wealth solution. In the Markovian feedback case $\pi\in\mathcal{A}^{\mathrm{fb}}$ one may think of $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$.
Ex–ante objective under latent $\theta$ (and simulator viewpoint).
The investor evaluates policies under an ex–ante objective that averages over both diffusion noise for fixed $\theta$ and the parametric uncertainty encoded by (9):
$J(\pi):=\mathbb{E}_{\theta\sim q}\bigg[\mathbb{E}\big[U(X_{T}^{\pi})\,\big|\,\theta\big]\bigg]=\int_{\Theta}\mathbb{E}\big[U(X_{T}^{\pi})\,\big|\,\theta\big]\,q(d\theta).$ (13)
The corresponding optimization problem (under our feedback restriction) is
$\sup_{\pi\in\mathcal{A}^{\mathrm{fb}}}J(\pi).$ (14)
Whenever it exists, we denote by
$\pi^{\star,\mathrm{blind}}\in\arg\max_{\pi\in\mathcal{A}^{\mathrm{fb}}}J(\pi)$
an optimal $\theta$-blind feedback for the fixed-$q$ commitment problem (14). For each fixed $\theta$, we also write $\pi^{\star,\theta}$ for the (infeasible) $\theta$-conditional full-information optimal control that would be available if $\theta$ were observed.
The $\theta$-blind constraint makes (14) strictly harder than solving a separate control problem for each fixed $\theta$, since the latter yields a $\theta$-indexed full-information family. Ex–ante averaging in (13) can also create gradient cancellation across heterogeneous parameter draws when one attempts to learn a single global policy end-to-end. While an $\mathcal{F}_{t}^{\mathrm{obs}}$-adapted policy could, in principle, filter $\theta$ online and solve a belief-state control problem (see, e.g., bensoussan1985optimal; pham2017dynamic), we do not pursue that formulation here.
Approximating the outer expectation in (13) amounts to sampling $\theta\sim q$ inside the simulator (once per trajectory or once per update), running (6)–(7) under that frozen draw, and updating a $\theta$-blind feedback policy to perform well on average over such draws. This is the setting targeted by the simulation-based PG–DPO and P–PGDPO methods developed in Section 3.
2.2 Pontryagin optimality under latent parameters: full-information vs. aggregated conditions
This subsection records the Hamiltonian structure underlying our projection step and clarifies what “Pontryagin first-order conditions” mean when the market parameter $\theta$ is latent and admissible controls are $\theta$-blind. In particular, we distinguish between (i) $\theta$-conditional (full-information) criticality conditions that would apply if $\theta$ were observable (and are therefore infeasible under latent $\theta$), and (ii) $q$-aggregated criticality conditions that characterize stationarity within the $\theta$-blind admissible class for the fixed-$q$ ex–ante objective. Our discussion follows standard stochastic control/PMP arguments for diffusion control (e.g. yong1999stochastic; fleming2006controlled; pham2009continuous). We also comment on the relationship to partial-information (belief-state) PMP, but we do not develop that formulation here.
A $\theta$-conditional (full-information) Hamiltonian and first-order condition (infeasible under latent $\theta$).
Fix $\theta\in\Theta$ and suppose, for the moment, that $\theta$ were observable to the controller. In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (8):
$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,V_{x},V_{xx},V_{xy}):=x\,\pi^{\top}b(y,\theta)\,V_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,V_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,V_{xy},$ (15)
where $V_{x},V_{xx}$ are evaluated at $(t,x,y)$ and $V_{xy}(t,x,y)\in\mathbb{R}^{m}$. The last term in (15) is the return–factor hedging term induced by $d\langle W,W^{Y}\rangle\neq 0$.
The pointwise first-order condition for an interior optimizer is
$\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}=x\,V_{x}^{\star,\theta}\,b(y,\theta)+x^{2}\,V_{xx}^{\star,\theta}\,\Sigma(y,\theta)\,\pi+x\,\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}\;=\;0.$ (16)
Assuming $\Sigma(y,\theta)$ is invertible and $V_{xx}^{\star,\theta}<0$, this yields the closed-form $\theta$-conditional full-information portfolio rule
$\pi^{\star,\theta}(t,x,y)=-\,\frac{1}{x\,V_{xx}^{\star,\theta}(t,x,y)}\,\Sigma(y,\theta)^{-1}\Big(V_{x}^{\star,\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}(t,x,y)\Big).$ (17)
This $\theta$-indexed rule is not deployable under latent parameters; we record it only as a full-information benchmark and diagnostic reference. In our setting, deployable policies never take the realized $\theta$ as an input; $\theta$ is accessed only through sampling inside the simulator when approximating $q$-expectations.
$q$-aggregated Pontryagin condition for the $\theta$-blind ex–ante problem (Markov feedback).
We now return to the actual setting: $\theta$ is latent, policies are $\theta$-blind, and we restrict attention to the Markov feedback class $\mathcal{A}^{\mathrm{fb}}$ (Remark 1). Under this restriction we neither perform online filtering of $\theta$ nor replace $q$ by a time-varying posterior distribution. Accordingly, the relevant Pontryagin condition is not the $\theta$-conditional criticality (16) enforced pointwise in $\theta$, but rather a necessary condition for optimality within the $\theta$-blind admissible class for the fixed-$q$ objective (13).
To see why ex–ante aggregation enters the first-order condition, take any $\theta$-blind admissible perturbation $h=\{h_{t}\}_{t\in[0,T]}$ that is progressively measurable with respect to the observation filtration $(\mathcal{F}_{t}^{\mathrm{obs}})$ and square-integrable, and define $\pi^{\varepsilon}:=\pi+\varepsilon h$ for small $\varepsilon$. For each fixed $\theta$, the stochastic maximum principle yields the first-variation identity
$\left.\frac{d}{d\varepsilon}J^{\theta}(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)^{\top}h_{t}\,dt\,\Big|\,\theta\right],$ (18)
where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market. Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
$\left.\frac{d}{d\varepsilon}J(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\mathbb{E}_{\theta\sim q}\!\Big[\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)\Big]^{\top}h_{t}\,dt\right].$ (19)
Hence, for an interior $\theta$-blind optimum $\pi^{\star,\mathrm{blind}}$, the first variation must vanish for all such perturbations $h$, which implies the aggregated first-order condition
$\mathbb{E}_{\theta\sim q}\!\Big[\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)\Big]=0,\hskip 18.49988pt\text{a.s.\ for a.e.\ }t\in[0,T].$ (20)
Equation (20) is the correct necessary condition for the ex–ante problem under the $\theta$-blind constraint. In particular, it is generally distinct from imposing (16) for each $\theta$ separately, because $\theta$-conditional criticality cannot be enforced by a single deployable $\theta$-blind policy.
To operationalize (20) in the Markov feedback class, fix a feedback policy $\pi\in\mathcal{A}^{\mathrm{fb}}$ and, for each frozen $\theta$, consider the corresponding $\theta$-conditional Pontryagin sensitivity objects $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ along the induced state process. In smooth Markov regimes these coincide with spatial derivatives of a decoupling field and, in particular, reduce to $(V_{x},V_{xx},V_{xy})$ in the full-information setting; in our algorithms we estimate them pathwise by automatic differentiation (see Section 3).
For the portfolio Hamiltonian (15), $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ is affine in $\pi$. This motivates defining the $\theta$-conditional “projection inputs”
$A_{t}^{\theta}(t,x,y)$ $:=x\,p_{x,t}^{\theta}(t,x,y)\,\Sigma(y,\theta)\in\mathbb{R}^{d\times d},$ (21)
$G_{t}^{\theta}(t,x,y)$ $:=p_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,p_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d},$ (22)
and their $q$-aggregated counterparts
$A_{t}(t,x,y):=\mathbb{E}_{\theta\sim q}\!\big[A_{t}^{\theta}(t,x,y)\big],\hskip 18.49988ptG_{t}(t,x,y):=\mathbb{E}_{\theta\sim q}\!\big[G_{t}^{\theta}(t,x,y)\big].$ (23)
These objects summarize how the latent parameter affects the first-order stationarity condition through the $\theta$-conditional sensitivities.
Theorem 1 ($q$-aggregated first-order condition under latent $\theta$ (deployable $\theta$-blind stationarity)).
Consider the fixed-$q$ ex–ante objective (13) over the $\theta$-blind Markov feedback class $\mathcal{A}^{\mathrm{fb}}$. Assume standard smoothness/integrability conditions ensuring validity of first variations within $\mathcal{A}^{\mathrm{fb}}$ and existence of the associated $\theta$-conditional Pontryagin objects. If $\pi^{\star,\mathrm{blind}}$ is a locally optimal interior policy in $\mathcal{A}^{\mathrm{fb}}$, then (20) holds. Moreover, in the portfolio setting (15), the aggregated stationarity is equivalent to the statewise linear system
$A_{t}(t,x,y)\,\pi^{\star,\mathrm{blind}}(t,x,y)=-\,G_{t}(t,x,y),\hskip 18.49988pt(t,x,y)\in[0,T]\times(0,\infty)\times\mathbb{R}^{m},$ (24)
(where $A_{t},G_{t}$ are defined by (23) using the $\theta$-conditional Pontryagin objects generated by $\pi^{\star,\mathrm{blind}}$). Whenever $A_{t}(t,x,y)$ is invertible on the working domain, (24) is equivalently expressed as the projected feedback rule
$\pi^{\mathrm{agg}}(t,x,y)=-\,A_{t}(t,x,y)^{-1}\,G_{t}(t,x,y).$ (25)
Proof sketch.
The conditional first-variation identity (18) is standard for diffusion control under a fixed parameter $\theta$ (e.g. yong1999stochastic; fleming2006controlled; pham2009continuous). Taking the outer expectation over $\theta\sim q$ yields (19). Since $h$ is an arbitrary $\theta$-blind admissible perturbation, vanishing of the first variation at an interior optimum implies (20). For the quadratic portfolio Hamiltonian (15), substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ and introducing (21)–(23) yields the linear system (24) and the projected form (25) whenever $A_{t}$ is invertible. ∎
Note that $\pi^{\mathrm{agg}}$ is generally not equal to the naive average $\mathbb{E}_{\theta\sim q}[\pi^{\star,\theta}(t,x,y)]$ of $\theta$-conditional full-information controls, reflecting the noncommutativity between averaging over $\theta$ and solving a first-order condition. In particular, even if one could compute $\pi^{\star,\theta}$ for each $\theta$, averaging these infeasible oracles does not, in general, enforce the deployable $q$-aggregated stationarity (20).
Remark 2 (Relation to belief-state/learning formulations).
If one allows history-dependent policies that explicitly infer $\theta$ from observed returns, a principled partial-information formulation introduces a time-varying posterior/belief state $q_{t}(\cdot)=\mathbb{P}(\theta\in\cdot\mid\mathcal{F}_{t}^{\mathrm{obs}})$. In such belief-state problems, the corresponding PMP/Hamiltonian criticality condition is expressed in terms of conditional expectations under $q_{t}$ (or, equivalently, conditional on $\mathcal{F}_{t}^{\mathrm{obs}}$); see, e.g., haussmann1987maximum; li1995general; baghery2007maximum. We do not pursue that learning/belief-state route here. Our algorithms and theory target the fixed-$q$, $q$-aggregated projection (25) under the $\theta$-blind Markov feedback restriction (52).
2.3 Gaussian references at a fixed decision time
This subsection collects Gaussian benchmarks that isolate decision-time statistical uncertainty and yield closed-form reference allocations. We fix a calendar decision time $t_{0}$ at which an external estimation pipeline outputs an uncertainty law $q_{t_{0}}(d\theta)$ for a risk-premium parameter, and we treat this law as an $\mathcal{F}_{t_{0}}$-measurable input for portfolio choice over the remaining horizon. This interface accommodates both Bayesian posterior/prior-like uncertainty descriptions (e.g., barberis2000investor; pastor2000portfolio) and frequentist sampling/resampling laws conditional on $\mathcal{F}_{t_{0}}$ (e.g., bootstrap or bagging) (e.g., efron1979bootstrap; efron1994bootstrap; breiman1996bagging). Throughout this subsection we work conditionally on $\mathcal{F}_{t_{0}}$, suppress conditioning by writing $q(d\theta)$, and shift the trading clock so that the decision time becomes 0 and the remaining horizon is $T$. These references are used as analytic targets and sanity checks for our numerical sections (Sections 4 and 5), rather than as characterizations of the unrestricted optimum of (14) over the full feedback class.
We present two decision-time references. Section 2.3.1 considers static drift uncertainty: a latent premium is drawn from $q$ once at time 0 and kept fixed on $[0,T]$, providing the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4. Section 2.3.2 considers a mean-reverting (OU) premium with Gaussian initial uncertainty, which induces a horizon-dependent Gaussian law for the time-averaged premium and yields a tractable closed-form reference used in the hedging-demand recovery study of Section 5. For completeness, an online linear–Gaussian illustration that produces a time-varying uncertainty law $q_{t}$ via Kalman–Bucy filtering is deferred to Appendix A; it is included only to motivate a plug-in (receding-horizon) decision-time workflow in which $q_{t}$ is treated as an externally updated input at each decision time, rather than solving the fully optimal belief-state control problem (e.g., bensoussan1985optimal; pham2017dynamic).
2.3.1 Static Gaussian drift uncertainty
We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter drawn at the decision time. The agent commits to a single $\theta$-blind policy, and all ex–ante uncertainty is summarized by the decision-time law $q$.
Market model (static latent drift).
Let $d$ risky assets satisfy
$\frac{dS_{t}}{S_{t}}=r\,\mathbf{1}\,dt+\theta\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (26)
where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and the latent excess-return vector is drawn at time 0 as
$\theta\sim q(d\theta).$ (27)
A $\theta$-blind portfolio fraction process $\pi_{t}\in\mathbb{R}^{d}$ generates wealth
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\theta\big)\,dt+\pi_{t}^{\top}\Sigma^{1/2}\,dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (28)
and we evaluate the ex–ante objective
$J(\pi):=\mathbb{E}_{\theta\sim q}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta\big]\Big].$ (29)
For reference, under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (merton1969lifetime; merton1971optimum), which is infeasible here because $\theta$ is latent.
Analytic $q$-references via constant portfolios. To obtain a transparent closed-form benchmark that depends only on the decision-time law $q$, we temporarily restrict attention to constant portfolio fractions
$\pi_{t}\equiv\pi\in\mathbb{R}^{d}.$ (30)
This restriction is used solely to define an analytic $q$-reference; it is not imposed on the learning problem.
Log utility ($\gamma=1$).
Let $m_{\theta}:=\mathbb{E}_{\theta\sim q}[\theta]$. For constant $\pi$, the objective depends on $q$ only through $m_{\theta}$, and the optimal constant portfolio is
$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\theta},$ (31)
which is independent of $T$ in this static benchmark. In the one-asset case ($d=1$, $\Sigma=\sigma^{2}$),
$\pi_{q,\log}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\sigma^{2}}.$ (32)
CRRA ($\gamma\neq 1$): tilted optimality and Gaussian shrinkage.
For $\gamma\neq 1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal, and
$J(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)rT-\tfrac{1}{2}\gamma(1-\gamma)T\,\pi^{\top}\Sigma\pi\Big\}\,M_{q}\big((1-\gamma)T\pi\big),$ (33)
where $M_{q}(u):=\mathbb{E}_{\theta\sim q}[\exp(u^{\top}\theta)]$ is the moment generating function of $q$. Any interior optimizer $\pi_{q,\gamma}^{\mathrm{const}}(T)$ satisfies the tilted first-order condition
$\gamma\,\Sigma\,\pi_{q,\gamma}^{\mathrm{const}}(T)=\nabla_{u}\log M_{q}(u)\Big|_{u=(1-\gamma)T\,\pi_{q,\gamma}^{\mathrm{const}}(T)}.$ (34)
If $q$ is Gaussian,
$\theta\sim\mathcal{N}(m_{\theta},P),\hskip 18.49988ptP\succeq 0,$ (35)
then $\nabla_{u}\log M_{q}(u)=m_{\theta}+Pu$ and the reference reduces to the linear system
$\big(\gamma\Sigma-(1-\gamma)T\,P\big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=m_{\theta},$ (36)
hence
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma-(1-\gamma)T\,P\big)^{-1}m_{\theta}.$ (37)
For $\gamma>1$, this takes the familiar shrinkage form
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma+(\gamma-1)T\,P\big)^{-1}m_{\theta},\hskip 18.49988pt(\gamma>1),$ (38)
and in one dimension ($d=1$, $\Sigma=\sigma^{2}$, $P=p$),
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\gamma\sigma^{2}+(\gamma-1)T\,p},\hskip 18.49988pt(\gamma>1).$ (39)
2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
We next replace the static premium by a mean-reverting Gaussian premium process, a standard reduced-form device for return predictability and intertemporal hedging (campbell2002strategic; xia2001learning). Our goal here is not to introduce additional information structure, but to obtain a closed-form, decision-time Gaussian reference for the time-averaged premium over the remaining horizon. This induces a horizon-dependent effective premium law that can be used as a controlled analytic input in numerical experiments.
OU premium dynamics and decision-time uncertainty.
Let the uncertain initial state be $\vartheta$ and set $Y_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0})$, so the decision-time law is $q=\mathcal{N}(m_{0},P_{0})$. The premium factor follows
$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0}),$ (40)
and risky excess returns satisfy
$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},$ (41)
allowing instantaneous correlation
$d\langle W,W^{Y}\rangle_{t}=\rho\,dt,\hskip 18.49988pt\rho\in\mathbb{R}^{d\times m}.$ (42)
Integrated premium and induced Gaussian law.
Define the integrated premium
$I_{T}:=\int_{0}^{T}Y_{s}\,ds\in\mathbb{R}^{m}.$ (43)
Since (40) is linear-Gaussian with Gaussian initial condition, $I_{T}$ is Gaussian. Its mean and covariance are
$m_{I}(T)=\mathbb{E}[I_{T}]=T\bar{y}+K^{-1}\big(I-e^{-KT}\big)\,(m_{0}-\bar{y}),$ (44)
$C_{I}(T)=\mathrm{Cov}(I_{T})$ $=K^{-1}\big(I-e^{-KT}\big)\,P_{0}\,\big(I-e^{-KT}\big)^{\top}K^{-\top}$ (45)
$\qquad+\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\Xi^{\top}\,\big(I-e^{-K(T-s)}\big)^{\top}K^{-\top}\,ds.$
Notably, (44)–(45) depend only on the OU dynamics and the decision-time uncertainty $(m_{0},P_{0})$; they do not depend on the return–factor correlation $\rho$.
Horizon-averaged premium and effective Gaussian law.
Define the horizon-averaged effective premium
$\bar{\theta}_{T}:=\frac{1}{T}\,B\,I_{T}\in\mathbb{R}^{d}.$ (46)
Then
$\bar{\theta}_{T}\sim\mathcal{N}(m_{\bar{\theta}}(T),P_{\bar{\theta}}(T)),\hskip 18.49988ptm_{\bar{\theta}}(T)=\frac{1}{T}B\,m_{I}(T),\hskip 18.49988ptP_{\bar{\theta}}(T)=\frac{1}{T^{2}}B\,C_{I}(T)\,B^{\top}.$ (47)
When $\rho=0$, this induced law can be plugged directly into the static Gaussian reference of Section 2.3.1. When $\rho\neq 0$, the marginal law (47) remains valid, but constant-portfolio expected utility involves an additional cross-covariance term capturing the return–state shock linkage that generates hedging demand (campbell2002strategic; xia2001learning).
Closed-form references under constant portfolios.
Restricting to constant fractions $\pi_{t}\equiv\pi$ turns the problem into a transparent decision-time benchmark: only the integrated premium $I_{T}=\int_{0}^{T}Y_{s}\,ds$ enters the drift of $\log X_{T}^{\pi}$, while the risk term remains time-homogeneous. This yields a closed-form target that depends on the decision-time law $q=\mathcal{N}(m_{0},P_{0})$ only through the induced mean $m_{I}(T)=\mathbb{E}[I_{T}]$ (and, for CRRA, through covariances as well).
Log utility ($\gamma=1$).
With $\pi_{t}\equiv\pi$, the log-utility criterion reduces to a strictly concave quadratic in $\pi$ whose drift term depends on the OU factor only through the mean integrated premium $m_{I}(T)=\mathbb{E}\!\left[\int_{0}^{T}Y_{s}\,ds\right]$. Hence the decision-time reference depends on $q=\mathcal{N}(m_{0},P_{0})$ only through $m_{\bar{\theta}}(T)=(1/T)B\,m_{I}(T)$ (and, in particular, does not involve return–factor correlation), giving
$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\bar{\theta}}(T)=\frac{1}{T}\,\Sigma^{-1}B\,m_{I}(T).$ (48)
CRRA ($\gamma>1$).
Define
$C_{IW}(T):=\mathrm{Cov}(I_{T},W_{T})=\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\,\rho^{\top}\,ds\;\in\;\mathbb{R}^{m\times d},$ (49)
and the induced symmetric cross term
$M_{\mathrm{cross}}(T):=B\,C_{IW}(T)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(T)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (50)
Then the Gaussian-$q$ decision-time reference under constant portfolios is characterized by
$\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=B\,m_{I}(T),\hskip 18.49988pt(\gamma>1),$ (51)
equivalently
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)^{-1}B\,m_{I}(T).$
When $\rho=0$, we have $C_{IW}(T)=0$ and $M_{\mathrm{cross}}(T)=0$, recovering the independence-case shrinkage reference.
2.4 Why dynamic programming and deep PDE surrogates break down in high-dimensional uncertain markets
This subsection explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted here. DP is conceptually sound in low-dimensional Markovian settings (fleming2006controlled; pham2009continuous), but two issues dominate in our regime: (i) numerically learning the value-function derivatives required for optimal policies becomes prohibitive as dimension and nonlinearity grow, and (ii) principled parameter uncertainty magnifies these difficulties.
Classical HJB: curse of dimensionality and full nonlinearity. With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (fleming2006controlled). Grid-based solvers scale exponentially in the state dimension (bellman1961adaptive; kushner2001numerical). In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points. Moreover, constraints, transaction costs, and non-affine dynamics typically yield fully nonlinear HJBs, where stable monotone schemes are delicate even in moderate dimension and become impractical in the regime we target (kushner2001numerical).
Deep PDE surrogates: fewer grids, same derivative bottleneck. PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (raissi2019physics; sirignano2018dgm; han2018solving; beck2019machine), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities (e.g. $V_{xy}$) that drive intertemporal hedging. In practice this induces nonconvex, ill-conditioned objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable hedging demands in high dimension.
Latent parameter uncertainty: belief-state blowup and $\theta$-blind aggregation. A principled DP treatment augments the state with a posterior/belief over parameters, leading to a value function $V(t,x,y,\Pi)$ on a space of measures in general (bensoussan1985optimal; pham2017dynamic). Even when finite-dimensional conjugate summaries exist, the enlarged HJB is substantially harder. For deep surrogates, uncertainty either requires solving many $\theta$-conditional problems (expensive) or treating $\theta$ as an extra input (higher effective dimension, worse conditioning). In our deployable $\theta$-blind setting, a single policy must perform well under $\theta\sim q$, coupling heterogeneous models and potentially causing high-variance gradients and cancellation across parameter draws.
We therefore avoid value-function PDE/BSDE baselines in this regime and instead work with a $q$-aggregated Pontryagin stationarity condition and projection map, estimating expectations over $\theta\sim q$ via Monte Carlo inside the simulator. While $\theta$-conditional PMP objects can still be computed under frozen-$\theta$ simulations for inspection, our deployable target and guarantees are expressed in terms of $q$-aggregated stationarity, motivating the simulation-based methods in Section 3.
3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
We study the fixed-$q$ ex–ante portfolio choice problem of Section 2 under latent parameter uncertainty $\theta\sim q$. The investor must deploy a $\theta$-blind policy (Remark 1), so the control can depend on observable states $(t,X_{t},Y_{t})$ but cannot take $\theta$ as an input. We restrict attention to Markov feedback policies parameterized by a neural network $\pi_{\varphi}(t,x,y)$.
Our solution approach follows a two-stage pipeline:
  • Stage 1 (PG–DPO). We perform stochastic gradient ascent on the ex–ante objective $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$, sampling $\theta$ only inside the simulator while keeping $\pi_{\varphi}$ deployable and $\theta$-blind.
  • Stage 2 (Pontryagin projection). Under a warm-up policy, we estimate Pontryagin sensitivity objects by BPTT (conditionally on frozen $\theta$), aggregate them across $\theta\sim q$, and construct a single deployable portfolio by projecting onto the aggregated first-order condition (20).
A practical subtlety is that the $q$-aggregated Pontryagin condition involves mixed moments across $\theta$ (products of $\theta$-dependent costates and $\theta$-dependent coefficients). In moderate to high dimensions, these quantities can be statistically noisy under finite Monte Carlo budgets. In our implementation, the main stabilization mechanisms are (i) estimating stage 2 objects under a warm-up policy (two-time-scale stabilization), (ii) computing the same projection in a residual/control-variate form (Section 3.3.1), and (iii) amortizing projection via interactive distillation (Section 3.3.2).
Section 3.1 reviews baseline PG–DPO and the conditional BPTT–PMP correspondence. Section 3.2 develops the stage 2 $q$-aggregated projection under latent $\theta$, together with a residual-based policy-gap guarantee. Section 3.3 records two practical couplings between stage 1 and stage 2 (residual form and interactive distillation).
3.1 PG–DPO as stochastic gradient ascent and BPTT–PMP correspondence
Setup and objectives (frozen $\theta$, deployable $\theta$-blind feedback).
A latent parameter $\theta\in\Theta$ is sampled from a fixed law $q(d\theta)$ inside the simulator and kept frozen along each simulated trajectory. A deployable portfolio policy is a $\theta$-blind Markov feedback rule represented by a neural network
$\pi_{\varphi}:\ [0,T]\times(0,\infty)\times\mathbb{R}^{m}\to\mathbb{R}^{d},\hskip 18.49988pt(t,x,y)\mapsto\pi_{\varphi}(t,x,y),\hskip 18.49988pt\varphi\in\mathbb{R}^{p},$ (52)
which does not take $\theta$ as an input. For a fixed frozen $\theta$, the $\theta$-conditional state is $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ and evolves as
$\frac{dX_{t}^{\pi,\theta}}{X_{t}^{\pi,\theta}}$ $=\Big(r+\pi_{t}^{\top}b\big(Y_{t}^{\theta},\theta\big)\Big)\,dt+\pi_{t}^{\top}\sigma\big(Y_{t}^{\theta},\theta\big)\,dW_{t},\hskip 18.49988ptX_{0}=x>0,$ (53)
$dY_{t}^{\theta}$ $=a\big(Y_{t}^{\theta},\theta\big)\,dt+\beta\big(Y_{t}^{\theta},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m}.$ (54)
For each fixed $\theta$ we evaluate $\pi_{\varphi}$ by the conditional objective
$J^{\theta}(\varphi):=\mathbb{E}\big[U\big(X_{T}^{\pi_{\varphi},\theta}\big)\,\big|\,\theta\big],$ (55)
where the expectation is over Brownian paths in (53)–(54). The fixed-$q$ ex–ante objective is
$J(\varphi):=\mathbb{E}_{\theta\sim q}\big[J^{\theta}(\varphi)\big]=\mathbb{E}\Big[U\big(X_{T}^{\pi_{\varphi},\theta}\big)\Big],$ (56)
where the last expectation is joint over $\theta\sim q$ and $(W,W^{Y})$. Thus $\sup_{\varphi}J(\varphi)$ is a standard stochastic optimization problem: $\theta$ is sampled inside the simulator while the policy remains $\theta$-blind.
Discretization, sampling over $\theta$, and baseline PG–DPO update.
We discretize $[0,T]$ into $N$ steps of length $\Delta t$ and approximate (53)–(54) by an Euler scheme. For episode $i$ we denote the discrete state by $(X_{k}^{(i)},Y_{k}^{(i)})_{k=0,\dots,N}$ and write $\theta^{(i)}$ for the frozen parameter used to generate that simulated environment. Given $\pi_{\varphi}$ and Brownian increments, the mapping
$\big(x^{(i)},y^{(i)},\theta^{(i)},\{\Delta W_{k}^{(i)},\Delta W_{k}^{Y,(i)}\}_{k=0}^{N-1},\varphi\big)\longmapsto U\big(X_{N}^{(i)}\big)$
is a finite computational graph, so automatic differentiation computes exact discrete gradients $\nabla_{\varphi}U(X_{N}^{(i)})$.
A typical PG–DPO update samples a mini-batch of initial states $\{(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})\}_{i=1}^{M}$ from a user-chosen training distribution $\nu$ on $0,T)\times(0,\infty)\times\mathbb{R}^{m}$, samples $\theta\sim q$ inside the simulator (unseen by the policy) and holds it frozen for the update, and simulates forward:
$Y_{k+1}^{(i)}$ $=Y_{k}^{(i)}+a\big(Y_{k}^{(i)},\theta\big)\Delta t^{(i)}+\beta\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{Y,(i)},$
$X_{k+1}^{(i)}$ $=X_{k}^{(i)}+X_{k}^{(i)}\Big(r+\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}b\big(Y_{k}^{(i)},\theta\big)\Big)\Delta t^{(i)}$
$\hphantom{={}}{}+X_{k}^{(i)}\,\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}\sigma\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{(i)},$
starting from $X_{0}^{(i)}=x_{0}^{(i)}$, $Y_{0}^{(i)}=y_{0}^{(i)}$. The episode reward is
$J^{(i)}(\varphi):=U\big(X_{N}^{(i)}\big),$ (57)
and BPTT computes $\nabla_{\varphi}J^{(i)}(\varphi)$. The policy parameters are then updated (e.g. by Adam) as
$\varphi\leftarrow\varphi+\alpha\,\frac{1}{M}\sum_{i=1}^{M}\nabla_{\varphi}J^{(i)}(\varphi).$ (58)
Sampling $\theta$ independently per episode (i.e. $\theta^{(i)}\sim q$) or sampling one $\theta\sim q$ per update and reusing it across the batch both yield unbiased stochastic gradients for $J(\varphi)$.
Pathwise costates from BPTT and the (conditional) BPTT–PMP correspondence.
BPTT returns not only $\nabla_{\varphi}J^{(i)}(\varphi)$ but also pathwise sensitivities with respect to intermediate state variables, which coincide with discrete-time adjoint variables (costates) in the sense of Pontryagin. For a single episode (suppressing $i$ and $\theta$ in notation), define the pathwise wealth costate
$p_{k}:=\frac{\partial U(X_{N})}{\partial X_{k}},\hskip 18.49988ptk=0,\dots,N,$ (59)
and the additional pathwise sensitivity objects used in projected controls:
$p_{x,k}:=\frac{\partial p_{k}}{\partial X_{k}},\hskip 18.49988ptp_{y,k}:=\frac{\partial p_{k}}{\partial Y_{k}},\hskip 18.49988ptk=0,\dots,N.$ (60)
Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, uniform on compacts)).
Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (53)–(54) under the $\theta$-blind policy $\pi_{\varphi}$ and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle (adjoint) system. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the resulting continuous-time Pontryagin objects under $\pi_{\varphi}$ (and, in smooth Markov regimes, the corresponding spatial derivatives of the decoupling field). Let $(p_{k},p_{x,k},p_{y,k})$ be the discrete pathwise quantities computed by BPTT for the Euler discretization with step $\Delta t$, as defined in (59)–(60).
Then, as $\Delta t\to 0$, the BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories). Moreover, for any compact set $K\subset\Theta$, the constants in the convergence bounds can be chosen uniformly for all $\theta\in K$.
Proof.
See Appendix B. ∎
Across $\theta\sim q$, these Pontryagin objects form a $\theta$-indexed family. Baseline PG–DPO trains against the ex–ante objective (56) by repeatedly sampling $\theta\sim q$ inside the simulator, while the deployable policy remains $\theta$-blind.
3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection and a residual-based policy-gap bound
Stage 2 is a projection step: given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin sensitivity objects by BPTT/Monte Carlo under frozen $\theta\sim q$, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated Pontryagin stationarity condition derived in Section 2.2. The main point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from estimated Pontryagin objects to a portfolio rule.
Working domain and norms. Fix a measurable working state domain $D\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ (e.g. a training/evaluation band) and a reference measure $\mu$ on $D$ (e.g. an empirical state distribution induced by rollouts). For $h:D\to\mathbb{R}^{n}$ we write
$\|h\|_{L^{2}(\mu)}:=\Big(\int_{D}\|h(z)\|^{2}\,\mu(dz)\Big)^{1/2},\hskip 18.49988ptz=(t,x,y),$
and for $\theta$-indexed families (used when tracking frozen-$\theta$ quantities in analysis/inspection),
$\|f\|_{L^{2}(q\otimes\mu)}:=\bigg(\int_{\Theta}\int_{D}\|f^{\theta}(z)\|^{2}\,\mu(dz)\,q(d\theta)\bigg)^{1/2}.$ (61)
Mixed-moment $q$-aggregation under a warm-up policy.
By Theorem 1, any locally optimal interior deployable $\theta$-blind policy $\pi^{\star,\mathrm{blind}}$ for the fixed-$q$ ex–ante problem satisfies the $q$-aggregated stationarity condition (20). In the portfolio Hamiltonian (15), this stationarity is equivalent to a statewise linear system and hence to the projected form (25) on the working domain (under invertibility of the aggregated curvature term). P–PGDPO constructs a practical approximation of this projection by estimating the relevant aggregated Pontryagin objects under a fixed warm-up policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$.
Fix a query state $z=(t,x,y)\in D$ and a frozen parameter $\theta$. We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute pathwise Pontryagin sensitivity objects by autodiff/BPTT; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
$\widehat{p}_{t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{x,t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{y,t}^{\theta}(z).$ (62)
Using these, define the $\theta$-conditional estimated projection inputs
$\widehat{A}_{t}^{\theta}(t,x,y)$ $:=x\,\widehat{p}_{x,t}^{\theta}(t,x,y)\,\Sigma(y,\theta)\in\mathbb{R}^{d\times d},$ (63)
$\widehat{G}_{t}^{\theta}(t,x,y)$ $:=\widehat{p}_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,\widehat{p}_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d}.$ (64)
Aggregating across $\theta\sim q$ (approximated in practice by sampling $M_{\theta}$ frozen parameters) gives
$\widehat{A}_{t}(t,x,y)$ $:=\mathbb{E}_{\theta\sim q}\Big[\widehat{A}_{t}^{\theta}(t,x,y)\Big],$ (65)
$\widehat{G}_{t}^{\mathrm{mix}}(t,x,y)$ $:=\mathbb{E}_{\theta\sim q}\Big[\widehat{G}_{t}^{\theta}(t,x,y)\Big].$ (66)
Whenever $\widehat{A}_{t}(t,x,y)$ is invertible on $D$, we obtain the mixed-moment projected policy
$\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y):=-\,\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{G}_{t}^{\mathrm{mix}}(t,x,y).$ (67)
Residual diagnostic and a slab-wise small-gain policy-gap bound.
To connect the projected policy (67) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity. Let $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ denote the mixed-moment $q$-aggregated projection inputs induced by a policy $\pi$ (i.e. the objects in (23) evaluated using the $\theta$-conditional Pontryagin objects generated by $\pi$). Define the warm-up aggregated stationarity residual on $D$ by
$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}(t,x,y):=A_{\pi^{\mathrm{warm}}}(t,x,y)\,\pi^{\mathrm{warm}}(t,x,y)+G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}(t,x,y),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\big\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\big\|_{L^{2}(\mu)}.$ (68)
In practice we monitor the estimator $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$ computed from the same BPTT/Monte Carlo pipeline.
A technical point is that a global small-gain condition of the form $C_{1}<1$ can be overly restrictive. Following the slab-wise philosophy in our prior PGDPO analysis (e.g. huh2025breaking), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab. Concretely, assume $D$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$. Let
$D_{k}:=D\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
We write $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map. Theorem 3 below shows that, under a mild slab-wise local stability regime (i.e. a short-time contraction of $T$ on each $D_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error. The proof combines a projection-map stability bound (Appendix C.1) with a slab-wise closure (Appendix C.2), in the same spirit as the slab analyses used in huh2025breaking.
Theorem 3 (Residual-based ex–ante $\theta$-blind policy-gap bound for P–PGDPO (mixed-moment, deployable, slab-wise local)).
Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix C.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$.
Let $\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem. Assume there exists a neighborhood $\mathcal{U}$ of $\pi^{\star,\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in\mathcal{U}$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G},$
and assume the slab-wise Lipschitz gain of Appendix C.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G_{\pi_{1}}^{\mathrm{mix}}-G_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k}.$ (69)
Define
$\rho(\tau):=\big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\big)\tau^{1/2},\hskip 18.49988pt\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k}).$ (70)
Assume the slab partition is chosen so that $\rho_{*}<1$.
Let $\widehat{\pi}^{\mathrm{agg,mix}}$ be the mixed-moment projected policy (67) computed from BPTT/Monte Carlo estimates under $\pi^{\mathrm{warm}}$, and let $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ be the population residual (68). Then there exists $C_{2}>0$ such that
$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}).$ (71)
Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
Proof.
See Appendix C.3. ∎
3.3 Coupling stage 1 and stage 2: residual projection and interactive distillation
We keep the ex–ante objective (56) and the $\theta$-blind deployability constraint throughout. Stage 2 is not a separate optimization problem: it reuses the current stage 1 policy as a warm-up control under which the (costate-based) projection ingredients are estimated, and then applies a $q$-aggregated Pontryagin projection as a post-processing map.
This subsection records two couplings between the two stages, each with a distinct role. First, we implement the projected rule in a residual (control-variate) form, which is algebraically equivalent to the direct projection but typically reduces Monte Carlo variance and improves numerical stability in high dimensions. Second, we use the projected output as a teacher signal via interactive distillation. Beyond acting as an optimization aid, distillation serves an amortization purpose: stage 2 projection can be accurate but Monte-Carlo intensive, whereas a distilled student policy can approximate the projected rule with a single forward pass at stage 1 inference cost.
3.3.1 Control-variate (residual) form of the projected rule
Recall the mixed-moment projected rule (67). In high dimensions, Monte Carlo noise in the projection inputs can be non-negligible, and solving a linear system with $\widehat{A}_{t}$ can amplify this noise. A convenient stabilization is to compute the same projected rule in a residual form around the warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$.
Define the $\theta$-conditional residual (under frozen-$\theta$ simulations)
$\widehat{r}_{\mathrm{FOC}}^{\theta}(t,x,y):=\widehat{A}_{t}^{\theta}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\theta}(t,x,y),$ (72)
and the aggregated residual (the quantity we actually solve against)
$\widehat{r}_{\mathrm{FOC}}(t,x,y):=\widehat{A}_{t}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\mathrm{mix}}(t,x,y).$ (73)
Whenever $\widehat{A}_{t}(t,x,y)$ is invertible, the projected rule admits the identity
$\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y)=\pi_{\varphi^{\mathrm{warm}}}(t,x,y)-\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{r}_{\mathrm{FOC}}(t,x,y),$ (74)
which is an algebraic rewriting of (67) (hence it does not change the target). Its practical value is variance reduction: when the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small, and it often concentrates faster because the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and partially cancel.
3.3.2 Interactive distillation: projection-guided training and amortized deployment
Let $\pi_{\varphi}$ be the trainable stage 1 policy network. At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher. This coupling serves two purposes. During training it provides projection-guided targets that can stabilize and accelerate stage 1 optimization; after training it amortizes the expensive projection by distilling it into a fast deployable policy network.
In residual form (74), the teacher is the $\theta$-blind map
$\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y):=\pi_{\varphi^{-}}(t,x,y)-\big(\widehat{A}_{t}^{\varphi^{-}}(t,x,y)\big)^{-1}\,\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}(t,x,y),$ (75)
where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy. We then train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to this teacher on the working domain:
$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(t,x,y)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y)\big)\big\|^{2}\Big],$ (76)
where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance. The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2; once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps: we hold $\varphi^{-}$ fixed for several updates of $\varphi$ under (76), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher. A practical schedule is to start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain (e.g., residual magnitudes and curvature/denominator stability) indicate that the stage 2 map has become reliable. Moreover, to avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we may apply projection guidance only on states where the projection checks certify reliability (an “adaptive teacher selection”); implementation details are deferred to the appendix (Appendix D).
4 Breaking the Dimensional Barrier under Drift Uncertainty
This section instantiates the decision-time static Gaussian drift-uncertainty benchmark in Section 2.3.1 and uses its closed-form constant-portfolio $q$-reference as an analytic target. Nature draws a fixed latent drift $\theta\sim q$ at $t=0$ and keeps it constant over $[0,T]$, while the investor cannot observe $\theta$ and must deploy a single $\theta$-blind policy under an ex–ante CRRA objective. Because the benchmark admits a transparent decision-time reference, we can measure accuracy directly via decision-time RMSE, rather than relying only on realized utility.
Our goal is to test whether Pontryagin-guided learning and projection remain stable as the number of assets grows. We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions. We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection) (and, when applicable, amortized variants via interactive distillation) under matched simulation budgets that scale linearly with $d$.
4.1 Benchmark market and evaluation protocol
This subsection fixes the benchmark and evaluation protocol used in Section 4. Our goal is to provide controlled evidence that the proposed two-stage pipeline remains computationally stable and accurate as the number of assets $d$ grows under decision-time parameter uncertainty. The aligned vs. misaligned uncertainty geometries serve as two representative stress-test regimes; the main message is scalability under uncertainty rather than any specific choice of $P$.
$\theta$-blind deployability (and what uses $\theta$). Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $t=0$ and $X_{0}$), and never takes the realized latent premium $\theta$ as an input. The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (notably in Stage 2 projection). Any $\theta$-indexed objects (when referenced elsewhere) are used only for offline diagnostics and are not part of the deployable decision rule.
Static decision-time uncertainty benchmark. We adopt the static Gaussian drift-uncertainty market of Section 2.3.1, i.e., (26) with (35). Equivalently, we simulate
$\frac{dS_{t}}{S_{t}}=r\,\mathbf{1}\,dt+\theta\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988pt\theta\sim\mathcal{N}(m,P),$
where the latent premium $\theta$ is drawn once at time 0 and kept fixed over $[0,T]$. The deployable policy is $\theta$-blind and interacts with $q$ only through sampling $\theta$ inside the simulator.
APT-style factor construction of $(m,\Sigma)$. We construct the mean premium and covariance via a low-dimensional factor representation. Let $W^{f}$ be a $k_{\Sigma}$-dimensional Brownian motion (factor shocks) and $W^{\varepsilon}$ a $d$-dimensional Brownian motion (idiosyncratic shocks), independent of $W^{f}$. We write excess returns as
$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=\theta\,dt+B\,\Sigma_{f}^{1/2}\,dW_{t}^{f}+\mathrm{diag}(\sqrt{D})\,dW_{t}^{\varepsilon},$ (77)
with $B\in\mathbb{R}^{d\times k_{\Sigma}}$, $\Sigma_{f}\succ 0$, and $D\in(0,\infty)^{d}$. This implies
$\Sigma=B\,\Sigma_{f}\,B^{\top}+\mathrm{diag}(D)\;=\;FF^{\top}+\mathrm{diag}(D),$ (78)
where $F:=B\,\mathrm{chol}(\Sigma_{f})$. We generate the mean premium in an APT-like form by drawing a factor price vector $\lambda_{m}\in\mathbb{R}^{k_{\Sigma}}$ and setting
$m:=B\,\lambda_{m}.$ (79)
One-shot generation and fairness across methods. For each dimension $d$, we generate a single market instance $(B,\Sigma_{f},D,\lambda_{m})$ once (using a fixed random seed) and hold it fixed across all algorithmic comparisons and MC-budget variants. Within a fixed $d$, we change only the uncertainty covariance $P$ (aligned vs. misaligned and the scale $s$ below). This isolates algorithmic effects from instance-to-instance randomness and makes the scaling comparisons controlled.
Uncertainty regimes (aligned vs. misaligned). We consider two geometries for the drift-uncertainty covariance $P$, controlled by a scalar magnitude $s>0$.
Aligned: uncertainty shares market risk directions,
$P=s\,\Sigma,\hskip 18.49988pts>0.$ (80)
Misaligned: uncertainty factors are rotated away from the market factor space,
$P=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}+s\,\mathrm{diag}(D),$ (81)
where $\widetilde{B}$ is generated independently of $B$ (or explicitly orthogonalized against the span of $B$ to enforce large principal angles). The factor term is rescaled so that its overall magnitude matches the aligned case under the same $s$ (e.g., by matching $\mathrm{tr}(P)$ or $\|P\|_{F}$ up to the shared diagonal component). This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
Experiment grid and simulation budgets. We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$, for both aligned and misaligned geometries. To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ sampling budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ paths and a high regime with $N_{\mathrm{MC}}=400\cdot d$ (where $N_{\mathrm{MC}}$ denotes the per-update or per-estimator path budget, depending on the stage). All methods share the same discretization scheme (Euler) and action constraints; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs. Stage 2) are reported in the implementation appendix and code release.
Analytic reference and decision-time evaluation. In the static Gaussian benchmark, the analytic decision-time reference under constant portfolios is available in closed form. We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint. For $\gamma>1$ we use the CRRA reference (38) and, for $\gamma=1$, the log-utility reference (31). We evaluate each method at $t=0$ on a fixed grid $\{(X_{0}^{(i)},T^{(i)})\}_{i=1}^{N_{\mathrm{eval}}}$ and report RMSE to the analytic reference:
$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T^{(i)})-\pi_{q,\gamma}^{\mathrm{const}}(T^{(i)})\big\|^{2}\bigg)^{1/2},$ (82)
where $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output). With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (79)–(81), the remaining subsections compare baseline Stage 1 PG–DPO, post-hoc Stage 2 P–PGDPO projection, and interactive distillation under matched simulation budgets.
4.2 High-dimensional CRRA benchmark: projection and amortization
Mixed-moment estimation and a decoupling approximation. A practical issue throughout our experiments (both aligned and misaligned) is the estimation of mixed moments across the latent parameter, such as $\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]$ (and analogous products entering $\widehat{G}_{t}^{\mathrm{mix}}$), because the costate $p_{t}^{\theta}(z)$ is $\theta$-dependent and high-dimensional, and finite-sample covariance between $p_{t}^{\theta}$ and $\theta$ can lead to large Monte-Carlo variance once the subsequent linear solve is applied. For numerical stability and a uniform protocol across geometries, we therefore use a simple decoupling (independence) approximation for these mixed moments,
$\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]\;\approx\;\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)]\,\mathbb{E}_{\theta\sim q}[\theta],$
(and similarly for other mixed products), which is exact when the relevant Pontryagin objects are effectively $\theta$-invariant and is accurate whenever $\mathrm{Cov}_{q}(p_{t}^{\theta}(z),\theta)$ is small relative to marginal scales. While this approximation is most valuable under misalignment—where direction mixing can amplify mixed-moment noise—it also performs well in aligned regimes (where mixed moments are typically easier to estimate), and in the CRRA benchmark below it does not alter the qualitative scaling conclusions: projection remains stable, and the observed misaligned degradation is consistent with residual growth and curvature mismatch rather than catastrophic mixed-moment blow-ups.  [^1] We note, however, that in extreme uncertainty/misalignment—where $\theta$–costate dependence becomes pronounced—the decoupling can break down, in which case one should revert to full mixed-moment estimation (possibly with larger budgets and/or regularized/certified projection).
Figure 1: Decision-time RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window (computed from the last evaluation snapshots). Rows: uncertainty magnitude $s\in\{10^{-3},10^{-2},10^{-1}\}$. Columns: aligned vs. misaligned geometry. Curves compare Stage 1 (deployable) and Stage 2 (post-hoc projection), with and without interactive distillation. Solid vs. dashed lines correspond to MC base ($100\cdot d$) vs. high ($400\cdot d$) budgets.
Protocol and summary statistic. We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (38). We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (82). Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window. Unless stated otherwise, the projection/teacher direction uses the mixed-moment ($p_{\theta}$) aggregation.
What is compared in Figure 1. Stage 1 (PG–DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients. Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section 3.3.1. Interactive distillation (Section 3.3.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
Thus Figure 1 separates projection quality (Stage 2: post-hoc projected, still $\theta$-blind) from amortized deployable quality (Stage 1 distilled: single forward pass).
Stage 2 projection versus amortization: scaling with dimension. Aligned geometry. For small and moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 delivers a sharp reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$. Interactive distillation consistently improves the deployable policy (Stage 1 (distill.) below Stage 1) while leaving Stage 2 essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation reduces the policy-class approximation/optimization gap by injecting that signal into $\pi_{\varphi}$.
Misaligned geometry. The picture becomes more heterogeneous. For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$. For the largest uncertainty scale ($s=10^{-1}$), Stage 1 becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed. Across settings, the base and high MC budgets tend to yield similar tail-median RMSE curves, suggesting that linear-in-$d$ scaling of simulation budgets is sufficient for stable comparisons in this benchmark.
Mechanism: why misalignment can reduce projection gains. To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix E; see Figures 36. The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting. These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix directions and can amplify Monte-Carlo errors in mixed-moment quantities (e.g., $\mathbb{E}[p_{1}\theta]$), especially as $d$ increases.
Figure 2: Pathwise sanity check at $d=100$ under common random numbers (same sampled $\theta$ and Brownian increments). Top: aligned geometry. Bottom: misaligned geometry. Each panel shows $\log X_{t}$ trajectories induced by the warm Stage 1 policy (PGDPO), the online Stage 2 P–PGDPO teacher (residual form), and the analytic $q$-reference.
Pathwise sanity check. Figure 2 complements the decision-time RMSE with a trajectory-level view under common random numbers. In the aligned case, the online Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy. In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 1 and motivating amortization/reliability mechanisms in interactive distillation.
Overall, the benchmark highlights a separation of roles. Stage 2 projection supplies a structured stationarity-correction signal that is particularly effective under aligned uncertainty, and interactive distillation amortizes this signal into a fast deployable Stage 1 policy. Under misalignment, projection can become more sensitive as $d$ and $s$ grow, consistent with diagnostic evidence of increased residuals and curvature mismatch; nevertheless, amortization remains a robust route to improving deployable policies under fixed simulation budgets.
4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
Why include PPO, and how we match the setting. Proximal Policy Optimization (PPO) is a widely used and robust model-free policy-gradient baseline for continuous control (schulman2017proximal). We include PPO to answer a concrete question: can a generic, well-tuned model-free RL method recover the decision-time $q$-optimal $\theta$-blind allocation in our high-dimensional drift-uncertainty benchmark under comparable simulation budgets? This comparison is especially informative in our static Gaussian benchmark because the target decision-time rule is structurally simple (constant and available in closed form), so performance gaps primarily reflect optimization difficulty and credit assignment rather than policy-class expressiveness.
Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section 2.4), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure. For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our PG–DPO pipeline, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective. We also use the same action cap $u_{\max}$ (with the same dimension-scaling convention) so that exploration ranges are comparable across $d$. Implementation details are deferred to the appendix and code release.
Empirical outcome. Table LABEL:tab:crra_stage12_distill_ppo_landscape shows that PPO remains far from the analytic decision-time $q$-reference across essentially all conditions, with RMSE typically on the order of $10^{-1}$. In contrast, the Pontryagin-based pipeline attains substantially smaller errors: in aligned regimes Stage 2 projection reaches the $10^{-5}$$10^{-4}$ range for small and moderate uncertainty, while in misaligned regimes the projection advantage narrows but remains systematic. Distillation improves the deployable Stage 1 policy relative to basic PG–DPO, but does not eliminate the remaining gap to the post-hoc projection, consistent with the amortization interpretation in Section 4.2.
Why PPO underperforms in this benchmark. The gap is not evidence that PPO is intrinsically weak; rather, it reflects that our benchmark stresses regimes where a generic likelihood-ratio policy gradient is statistically disadvantaged compared to pathwise/adjoint-based updates. With terminal utility as the only reward, PPO faces a long-horizon credit-assignment problem whose gradient variance grows with both horizon and action dimension. Sampling $\theta\sim q$ further creates episode-wise heterogeneity under a single $\theta$-blind policy, inducing additional variance and potential cancellation across parameter draws. In contrast, Stage 1 exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction that is tailored to the $\theta$-blind ex–ante objective.
Under matched simulation budgets in our latent-$\theta$, $\theta$-blind benchmark, a generic model-free PPO baseline does not reliably recover the decision-time $q$-optimal allocation (Table LABEL:tab:crra_stage12_distill_ppo_landscape), motivating structure-exploiting alternatives—pathwise gradients, costates, and the $q$-aggregated Pontryagin projection—as in PG–DPO and P–PGDPO.
5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
Sections 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic. Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (campbell2002strategic; xia2001learning).
We use the mean-reverting Gaussian premium benchmark of Section 2.3.2. Decision-time statistical uncertainty enters through the (uncertain) initial premium state $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$, while a nonzero return–factor correlation $\rho$ generates hedging demand through the cross term $M_{\mathrm{cross}}(T)$ in (51). Crucially, we enforce a deployable restriction aligned with Section 3: the policy is $Y$-blind and does not observe the realized $Y_{0}$ nor the path $(Y_{t})$.
We compare: (i) Stage 1 PG–DPO (deployable end-to-end learning; Section 3.1), (ii) Stage 2 $q$-aggregated Pontryagin projection (post-hoc correction in residual form; Sections 3.2 and 3.3.1), (iii) interactive distillation (amortized projection guidance; Section 3.3.2), and (iv) a model-free PPO baseline trained under the same deployable $Y$-blind observation restriction. Performance is measured by decision-time RMSE against the analytic constant-portfolio OU reference (51) (which reduces to the independence-case benchmark when $\rho=0$).
In addition to the full allocation error, this benchmark provides a natural myopic + hedging decomposition (driven by return–factor correlation). We therefore report (a) the RMSE of the full decision-time allocation for all methods (including PPO), and (b) component-wise diagnostics for the projected (Stage 2) rules: RMSE of the hedging component (Table 1) and, in the appendix, the RMSE of the myopic component (Table 2) and the cosine similarity of the hedging direction (Table 3). Since PPO does not expose a compatible myopic/hedging decomposition for these diagnostics, we include it only in the full RMSE table.
To keep the main text focused, we include the full RMSE table (Table LABEL:tab:ou_rmse_s0_sweep_landscape) and the hedging-RMSE table (Table 1) in Section 5.2; the remaining two diagnostic tables are deferred to the appendix (Section F).
5.1 Experimental setting
$Y$-blind deployability (and what uses $Y$). Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes the realized initial premium $Y_{0}$ nor the factor path $(Y_{t})$ as an input (including the PPO baseline). The latent premium factor is sampled and propagated only inside the simulator to generate trajectories and to form Monte Carlo averages used by the stage 2 projection (and by the teacher in distillation). Any $Y$-indexed quantities are used only for offline evaluation and diagnostics.
OU premium market with a hedging channel. We adopt the OU premium benchmark of Section 2.3.2. Let $Y_{t}\in\mathbb{R}^{m}$ be a mean-reverting premium factor and $R_{t}\in\mathbb{R}^{d}$ the risky excess returns:
$dY_{t}$ $=\kappa(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$
$dR_{t}$ $:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},$
$d\langle W,W^{Y}\rangle_{t}$ $=\rho\,dt.$
A nonzero $\rho$ induces intertemporal hedging demand and enters the CRRA decision-time reference through the cross-covariance term $M_{\mathrm{cross}}(T)$ in (51). When $\rho=0$ (independent return and factor shocks), the hedging channel vanishes ($M_{\mathrm{cross}}(T)=0$) and the reference reduces to the independence-case benchmark.
Decision-time uncertainty geometry for $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$. We control the magnitude of decision-time statistical uncertainty by a scalar $s_{0}>0$ and construct $P_{0}$ from an identification-motivated baseline
$\widetilde{P}_{0}:=(B^{\top}\Sigma^{-1}B)^{-1}\in\mathbb{R}^{m\times m}.$
We consider two geometries. In the aligned case, we keep the principal directions of $\widetilde{P}_{0}$ and rescale it so that the average marginal variance equals $s_{0}$:
$P_{0}^{\mathrm{aligned}}(s_{0}):=\frac{s_{0}\,m}{\mathrm{tr}(\widetilde{P}_{0})}\,\widetilde{P}_{0},$ (83)
so that $\mathrm{tr}(P_{0}^{\mathrm{aligned}})/m=s_{0}$. In the misaligned case, we preserve the eigenvalue spectrum of $\widetilde{P}_{0}$ but randomize its eigenvectors via an orthogonal rotation: letting $\widetilde{P}_{0}=U\mathrm{diag}(\lambda)U^{\top}$ be an eigen-decomposition and drawing an orthogonal matrix $R$ (e.g., Haar), we define
$P_{0}^{\mathrm{misaligned}}(s_{0}):=\frac{s_{0}\,m}{\mathrm{tr}(\widetilde{P}_{0})}\,UR\,\mathrm{diag}(\lambda)\,R^{\top}U^{\top},$ (84)
which matches the same trace normalization while rotating the uncertainty directions away from those of $\widetilde{P}_{0}$. We sweep $s_{0}\in\{10^{-3},10^{-2},10^{-1}\}$ under both aligned and misaligned $P_{0}$.
Two-stage solver, amortization, and evaluation protocol. We use the two-stage pipeline of Section 3. Stage 1 trains a deployable policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1). Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section 3.3.1). Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section 3.3.2). As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table LABEL:tab:ou_rmse_s0_sweep_landscape.
We sweep $d\in\{5,10,50,100\}$ (one fixed market instance per $d$), train for $5000$ epochs, and evaluate every $100$ epochs. Unless stated otherwise we set $\gamma=2$, $r=0.03$, $\kappa=1.0$, $\xi_{\mathrm{scale}}=0.25$, and $\rho=0.5$. We evaluate the decision-time action at $t=0$ and report RMSE to the analytic constant-portfolio OU reference (51).
In addition to the full allocation error, we use the natural myopic + hedging decomposition induced by the OU factor structure. We report the RMSE of the full decision-time allocation for all methods, and component-wise diagnostics for the projected (Stage 2) rules, including the RMSE of the hedging component and (in the appendix) the RMSE of the myopic component and cosine similarity of the hedging direction. To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
We report decision-time RMSE at $t=0$ against the analytic OU reference (51). To reduce noise from stochastic optimization, we summarize each condition by a tail median over the last six evaluation checkpoints. Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE for all deployable objects (Stage 1 and Stage 1+Stage 2, with and without distillation), and also includes a model-free PPO baseline trained under the same $Y$-blind deployability restriction. To isolate the economic channel of interest, Table 1 reports the RMSE of the hedging component for the post-hoc projected (Stage 2) rules. Two additional diagnostics—the myopic-component RMSE and the hedging-direction cosine similarity—are deferred to the appendix (Tables 2 and 3). Since PPO does not expose a compatible myopic/hedging decomposition in our diagnostic protocol, we report it only in the full-RMSE table.
Projection and economic hedging-demand recovery. Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (Stage 1+Stage 2) substantially reduces decision-time RMSE relative to the deployable Stage 1 policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, Stage 1 attains $3.54\text{\times}{10}^{-03}$ whereas Stage 1+Stage 2 achieves $1.56\text{\times}{10}^{-04}$. The component-wise diagnostics indicate that the remaining discrepancy is largely driven by the hedging channel: in the same setting, the hedging RMSE is $1.55\text{\times}{10}^{-04}$ (Basic) and $1.42\text{\times}{10}^{-04}$ (Distill.) (Table 1), while the myopic RMSE is an order of magnitude smaller (Appendix Table 2). This pattern is consistent with the economic mechanism in this benchmark: once the (mostly) myopic component is captured, the dominant remaining challenge is to recover the intertemporal hedge induced by correlated return–factor shocks.
Amortization, robustness, and the PPO baseline. Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This matches the intended division of labor in Section 3.3: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables LABEL:tab:ou_rmse_s0_sweep_landscape1). Misalignment has a limited effect for small and moderate uncertainty scales, but can induce noticeable deterioration in the hardest settings, where the direction-of-hedge diagnostic can also weaken (Appendix Table 3).
Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, with degradation that becomes especially pronounced at larger $d$ (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This is consistent with PPO facing a terminal-only credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline. Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
$s_{0}$ Method $d=5$ $10$ $50$ $100$
Aligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $4.59\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.39\text{\times}{10}^{-05}$ $5.25\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $98\text{\times}{10}^{-05} 4.4$$47\text{\times}{10}^{-05}$ $1.44\text{\times}{10}^{-04}$ $1.74\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.43\text{\times}{10}^{-05}$ $4.95\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $5.27\text{\times}{10}^{-05}$ $3.99\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.95\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.72\text{\times}{10}^{-05}$ $5.36\text{\times}{10}^{-05}$ $2.58\text{\times}{10}^{-04}$ $2.86\text{\times}{10}^{-04}$
Misaligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $5.00\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.34\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.30\text{\times}{10}^{-05}$ $5.33\text{\times}{10}^{-05}$ $1.40\text{\times}{10}^{-04}$ $1.43\text{\times}{10}^{-04}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $4.93\text{\times}{10}^{-05}$ $5.65\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.56\text{\times}{10}^{-05}$ $6.09\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $5.45\text{\times}{10}^{-05}$ $1.55\text{\times}{10}^{-04}$ $3.19\text{\times}{10}^{-04}$ $3.28\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-05}$ $1.57\text{\times}{10}^{-04}$ $3.13\text{\times}{10}^{-04}$ $3.20\text{\times}{10}^{-04}$
Table 1: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the post-hoc projected (Stage 2) policies (tail median over the last six evaluation checkpoints). Component-wise diagnostics are reported for Stage 2 since Stage 1 does not explicitly output a myopic/hedging decomposition.
In a factor-driven market where return–factor correlation induces intertemporal hedging demand, the proposed two-stage pipeline recovers the analytic OU decision-time reference with high accuracy: projection provides the dominant gains, and distillation improves deployable policies by amortizing projection guidance. In contrast, a model-free PPO baseline does not reliably match the analytic reference in this $Y$-blind setting.
6 Conclusion
We studied continuous-time portfolio choice in diffusion markets whose coefficients are estimated and therefore subject to statistical uncertainty (Section 2.1). We model this uncertainty by an exogenous law $q(d\theta)$ over a latent parameter $\theta$ that is drawn once at time 0 and remains fixed over the investment horizon, while the investor must deploy a single $\theta$-blind Markov feedback policy evaluated under an ex–ante CRRA objective (Remark 1, Section 2.1). This information structure shifts the relevant optimality notion from $\theta$-conditional (full-information) criticality to a $q$-aggregated Pontryagin first-order condition that is enforceable within the deployable $\theta$-blind policy class (Section 2.2, Theorem 1).
Methodologically, we extended Pontryagin–Guided Direct Policy Optimization (PG–DPO) to the latent-parameter setting by sampling $\theta$ only inside the simulator and computing exact discrete-time gradients via BPTT (Section 3.1), and we leveraged the BPTT–PMP correspondence to extract the costate objects needed for structured control updates (Theorem 2). Building on the $q$-aggregated stationarity, we proposed uncertainty-aware projected PG–DPO (P–PGDPO), which aggregates Monte Carlo Pontryagin quantities across $\theta\sim q$ and projects them onto the deployable first-order condition to obtain a single $\theta$-blind rule (Section 3.2). We established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3). In experiments with finite-sample uncertainty, projection improves stability and accuracy in high dimensions and exhibits a two-time-scale stabilization effect (costates versus policies), while interactive distillation amortizes the projection into a fast deployable network (Sections 4 and 5; Section 3.3).
Several extensions are natural. A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to belief-aware decision rules (Remark 2, Appendix A). A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section 2.4; Section 4.2; Appendix D.5). Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Purpose and scope. Sections 2.3.1 and 2.3.2 focus on decision-time benchmarks in which an uncertainty description $q$ is treated as given and the investor optimizes under the corresponding $\theta$-blind deployability constraint. In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (barberis2000investor; pastor2000portfolio; xia2001learning). This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., bensoussan1985optimal; pham2017dynamic), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$. We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization. In particular, our simulation-based Pontryagin-guided solvers developed later (Section 3) can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
A linear–Gaussian hidden-premium model (OU state, observed returns). We use a stylized linear–Gaussian counterpart of the mean-reverting premium setting, but now assume the premium factor is not directly observed. Let $Y_{t}\in\mathbb{R}^{m}$ be a latent premium factor following an OU dynamics
$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(\hat{y}_{0},P_{0}),$ (85)
where $K\in\mathbb{R}^{m\times m}$ is stable, $\bar{y}\in\mathbb{R}^{m}$, and $\Xi\in\mathbb{R}^{m\times m}$. Risky assets satisfy
$\frac{dS_{t}}{S_{t}}=r\mathbf{1}\,dt+BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},\hskip 18.49988pt\Sigma\in\mathbb{R}^{d\times d}\ \text{s.p.d.},$ (86)
with $B\in\mathbb{R}^{d\times m}$. Equivalently, the investor observes the excess-return signal
$dZ_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t}.$ (87)
We write $\mathbb{F}^{Z}=(\mathcal{F}_{t}^{Z})_{t\in[0,T]}$ for the filtration generated by $(Z_{s})_{s\leq t}$. For clarity, we present the independent-noise case $W\perp W^{Y}$; the correlated-noise extension remains linear–Gaussian but leads to more cumbersome gain formulas.
Kalman–Bucy posterior $q_{t}=\mathcal{L}(Y_{t}\mid\mathcal{F}_{t}^{Z})$. Under (85)–(87), the conditional law of the latent factor remains Gaussian:
$q_{t}(dy):=\mathcal{L}(Y_{t}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(\hat{Y}_{t},P_{t}),$ (88)
where $(\hat{Y}_{t},P_{t})$ satisfy the Kalman–Bucy equations
$d\hat{Y}_{t}$ $=K(\bar{y}-\hat{Y}_{t})\,dt+P_{t}B^{\top}\Sigma^{-1}\Big(dZ_{t}-B\hat{Y}_{t}\,dt\Big),$ (89)
$\dot{P}_{t}$ $=KP_{t}+P_{t}K^{\top}+\Xi\Xi^{\top}-P_{t}B^{\top}\Sigma^{-1}BP_{t},\hskip 18.49988ptP_{0}\ \text{given}.$ (90)
Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (90); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (bensoussan1985optimal; pham2017dynamic).
From a posterior on $Y_{t}$ to a Gaussian uncertainty description for decision-time optimization. To mirror the decision-time perspective of the OU benchmark, we consider the remaining-horizon time-averaged premium
$\bar{\theta}_{t,T}:=\frac{1}{T-t}\int_{t}^{T}BY_{s}\,ds\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (91)
For the OU dynamics (85), one has the decomposition
$\int_{t}^{T}Y_{s}\,ds=\tau\bar{y}+K^{-1}\big(I-e^{-K\tau}\big)(Y_{t}-\bar{y})+\int_{t}^{T}K^{-1}\big(I-e^{-K(T-u)}\big)\Xi\,dW_{u}^{Y}.$ (92)
Conditioning on $\mathcal{F}_{t}^{Z}$, the random variable $Y_{t}$ is distributed as $\mathcal{N}(\hat{Y}_{t},P_{t})$ by (88), while the future increments $(W_{u}^{Y}-W_{t}^{Y})_{u\geq t}$ are independent of $\mathcal{F}_{t}^{Z}$ in the independent-noise case. Hence $\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}$ is Gaussian:
$\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}\!\big(m_{t,T},\,P_{t,T}\big),\hskip 18.49988ptm_{t,T}:=\frac{B\,m_{I}(t,T)}{\tau},\hskip 18.49988ptP_{t,T}:=\frac{1}{\tau^{2}}\,B\,C_{I}(t,T)\,B^{\top},$ (93)
where
$m_{I}(t,T)$ $:=\tau\bar{y}+K^{-1}\big(I-e^{-K\tau}\big)(\hat{Y}_{t}-\bar{y}),$ (94)
$C_{I}(t,T)$ $:=K^{-1}\big(I-e^{-K\tau}\big)P_{t}\big(I-e^{-K\tau}\big)^{\top}K^{-\top}+\int_{0}^{\tau}K^{-1}\big(I-e^{-Ks}\big)\Xi\Xi^{\top}\big(I-e^{-Ks}\big)^{\top}K^{-\top}\,ds.$ (95)
Equation (93) provides a concrete example of an online-updated Gaussian uncertainty description $q_{t,T}:=\mathcal{L}(\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(m_{t,T},P_{t,T}).$
A plug-in decision-time benchmark (receding-horizon fixed-$q_{t,T}$). Given $(m_{t,T},P_{t,T})$ from (93), a simple decision-time rule is obtained by treating $q_{t,T}$ as fixed over the remaining horizon and applying the Gaussian constant-allocation benchmark of Section 2.3.1 with horizon $\tau$:
$\pi_{t}^{\mathrm{plug}}:=\Big(\gamma\,\Sigma+(\gamma-1)\tau\,P_{t,T}\Big)^{-1}m_{t,T}.$ (96)
One may interpret (96) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (barberis2000investor; pastor2000portfolio; xia2001learning).
Remarks (relation to belief-aware control). The plug-in rule (96) is intentionally decision-time: it conditions on the current uncertainty description and does not attempt to optimize over how the posterior will evolve. In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (barberis2000investor; pastor2000portfolio; xia2001learning). Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (bensoussan1985optimal; pham2017dynamic). Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (haussmann1987maximum; li1995general; baghery2007maximum). Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in $(x,y,\hat{Y},P)$-space (or its sufficient-statistic analogues) is an important direction that we defer to future work.
Appendix B Proof of Theorem 2
Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in huh2025breaking). Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.
Important remark (what this proof does not use). This proof concerns the $\theta$-conditional Pontryagin adjoint/costate for the fixed-$\theta$ control problem induced by (53)–(54). It does not use the $\theta$-blind $q$-aggregated stationarity condition (Theorem 1 in Section 3.2). Those constructions affect only the deployable aggregation/projection target in stage 2 and are irrelevant to the BPTT–PMP convergence itself.
Notation and filtration. Fix $\theta\in\Theta$. We work conditionally on this $\theta$ and consider the augmented (simulator) filtration
$\mathbb{G}^{\theta}:=(\mathcal{G}_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988pt\mathcal{G}_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
All conditional expectations and $L^{2}$ projections below are taken with respect to $\mathcal{G}_{t_{k}}^{\theta}$. This choice matches the information set used by the simulator and by automatic differentiation/BPTT (which differentiates through the full forward recursion).
Let $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, $N\Delta t=T$. For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind in the sense of (52) but evaluated along the $\theta$-conditional trajectory.
Step 1: Conditioning on $\theta$ and uniformity of bounds. Fix a compact set $K\subset\Theta$. Assume the coefficients in (53)–(54) satisfy the usual Lipschitz and linear-growth conditions uniformly over $\theta\in K$, and that the block covariance structure of $(W,W^{Y})$ (including instantaneous correlation) is uniformly nondegenerate on $K$. Then, for each fixed $\theta\in K$, the controlled SDE system (53)–(54) is well posed and admits uniform-in-time $L^{2}$ moment bounds. Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
$\sup_{t\in[0,T]}\mathbb{E}\big[\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\|^{2}\big]^{1/2}\;\leq\;C_{K}\,\Delta t^{1/2},$
with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are the deterministic assumptions used in huh2025breaking, now stated uniformly on $K$.
Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme for (53)–(54) on the grid $(t_{k})$:
$Y_{k+1}^{\theta}=Y_{k}^{\theta}+a(Y_{k}^{\theta},\theta)\Delta t+\beta(Y_{k}^{\theta},\theta)\Delta W_{k}^{Y},$
$X_{k+1}^{\theta}=X_{k}^{\theta}+X_{k}^{\theta}\Big(r+\pi_{k}^{\top}b(Y_{k}^{\theta},\theta)\Big)\Delta t+X_{k}^{\theta}\,\pi_{k}^{\top}\sigma(Y_{k}^{\theta},\theta)\Delta W_{k},$
with terminal reward $U(X_{N}^{\theta})$. Define the discrete (pathwise) wealth costate
$p_{k}^{\mathrm{pw},\theta}:=\frac{\partial}{\partial X_{k}^{\theta}}\,U(X_{N}^{\theta}),\hskip 18.49988ptk=0,\dots,N,$
which is the same object as (59) (episode indices suppressed and dependence on $\theta$ made explicit). For the projected-control constructions we also consider the additional pathwise objects
$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}},$
which correspond to (60). Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph.
The algebraic form of the one-step backward recursion coincides with the deterministic-parameter analysis in huh2025breaking, with the replacements
$\mu\mapsto b(\cdot,\theta),\hskip 18.49988pt\sigma\mapsto\sigma(\cdot,\theta),$
and with the factor block $(Y^{\theta},\Delta W^{Y})$ handled exactly as in the wealth–factor extension therein. All one-step remainder terms are controlled by standard Taylor/Euler estimates with constants uniform in $\theta\in K$.
Step 3: One-step $L^{2}$ projection and discrete BSDE form (fixed $\theta$). Fix $\theta\in K$. As in huh2025breaking, take the conditional $L^{2}$-projection of $p_{k+1}^{\mathrm{pw},\theta}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$:
$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ conditionally on $\mathcal{G}_{t_{k}}^{\theta}$. Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ yields unique projection coefficients $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$.
Substituting this projection into the BPTT backward recursion from Step 2 yields a canonical discrete BSDE representation for $(p_{k}^{\mathrm{pw},\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ whose drift matches the Euler discretization of the $\theta$-conditional Pontryagin adjoint BSDE associated with (53)–(54). The same argument applies to the derivatives $(p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})$: they satisfy linearized discrete backward recursions obtained by differentiating the discrete adjoint equations, hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$.
Step 4: Passage to continuous time and identification with the PMP costate. For each fixed $\theta\in K$, the forward SDE and the $\theta$-conditional adjoint BSDE form a standard FBSDE with coefficients parametrized by $\theta$. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under policy $\pi_{\varphi}$, so $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
Define the piecewise-constant interpolations
$p_{t}^{\Delta t,\theta}:=p_{k}^{\mathrm{pw},\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\mathrm{pw},\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\mathrm{pw},\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
By the same stability and convergence arguments as in huh2025breaking (Euler convergence for the forward equation plus discrete-BSDE convergence for the backward equation), we obtain, for each fixed $\theta\in K$,
$\sup_{t\in[0,T]}\mathbb{E}\big[\,|p_{t}^{\Delta t,\theta}-p_{t}^{\theta}|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{x,t}^{\Delta t,\theta}-p_{x,t}^{\theta}\|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{y,t}^{\Delta t,\theta}-p_{y,t}^{\theta}\|^{2}\,\big]\to 0,$
as $\Delta t\to 0$. Because all Lipschitz, growth, ellipticity, and covariance constants were assumed uniform on $K$, the convergence constants can be chosen independently of $\theta\in K$. This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$, completing the proof. ∎
Appendix C Auxiliary results for Theorem 3
C.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
Proposition 1 (Stability of the projection map $(A,G)\mapsto-A^{-1}G$).
Let $D$ be a measurable domain and let $\mu$ be a reference measure on $D$. Let $A,\widetilde{A}:D\to\mathbb{R}^{d\times d}$ and $G,\widetilde{G}:D\to\mathbb{R}^{d}$ be measurable. Assume:
  1. (i)
    $A(z)$ is invertible for $\mu$-a.e. $z\in D$ and $\|A^{-1}\|_{L^{\infty}(D)}\leq\kappa$ for some $\kappa>0$;
  2. (ii)
    $\|G\|_{L^{\infty}(D)}\leq M$ for some $M>0$;
  3. (iii)
    $\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq(2\kappa)^{-1}$.
Define $\pi:=-A^{-1}G$ and $\widetilde{\pi}:=-\widetilde{A}^{-1}\widetilde{G}$. Then $\widetilde{A}(z)$ is invertible for $\mu$-a.e. $z\in D$ with $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$, and
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\,\big(M+\|\widetilde{G}\|_{L^{\infty}(D)}\big)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (97)
Proof.
Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm. For a matrix-valued function $M:D\to\mathbb{R}^{d\times d}$, write
$\|M\|_{L^{\infty}(D)}:=\operatorname*{ess\,sup}_{z\in D}\|M(z)\|,\hskip 18.49988pt\|M\|_{L^{2}(\mu)}:=\Big(\int_{D}\|M(z)\|^{2}\,\mu(dz)\Big)^{1/2},$
and similarly for vector-valued functions.
Step 1: Invertibility and inverse bound for $\widetilde{A}$. Fix $z\in D$ such that $A(z)$ is invertible (this holds for $\mu$-a.e. $z$). Let $E(z):=\widetilde{A}(z)-A(z)$. By (i) and (iii),
$\|A(z)^{-1}E(z)\|\leq\|A(z)^{-1}\|\,\|E(z)\|\leq\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq\kappa\cdot\frac{1}{2\kappa}=\frac{1}{2}.$
Hence $I+A(z)^{-1}E(z)$ is invertible and admits the Neumann-series inverse. Therefore,
$\widetilde{A}(z)^{-1}=(A(z)+E(z))^{-1}=(I+A(z)^{-1}E(z))^{-1}A(z)^{-1},$
and
$\|\widetilde{A}(z)^{-1}\|\leq\frac{1}{1-\|A(z)^{-1}E(z)\|}\,\|A(z)^{-1}\|\leq\frac{1}{1-1/2}\,\kappa=2\kappa.$
Taking the essential supremum over $z\in D$ yields
$\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa.$
Step 2: A pointwise bound for $\widetilde{A}^{-1}-A^{-1}$. For $\mu$-a.e. $z\in D$ where both inverses exist,
$\widetilde{A}(z)^{-1}-A(z)^{-1}=\widetilde{A}(z)^{-1}\big(A(z)-\widetilde{A}(z)\big)A(z)^{-1}.$
Thus,
$\|\widetilde{A}(z)^{-1}-A(z)^{-1}\|\leq\|\widetilde{A}(z)^{-1}\|\,\|\widetilde{A}(z)-A(z)\|\,\|A(z)^{-1}\|\leq(2\kappa)\,\|\widetilde{A}(z)-A(z)\|\,\kappa=2\kappa^{2}\|\widetilde{A}(z)-A(z)\|.$
Consequently,
$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\leq 2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
Step 3: Control error bound. Recall $\pi=-A^{-1}G$ and $\widetilde{\pi}=-\widetilde{A}^{-1}\widetilde{G}$. Then
$\widetilde{\pi}-\pi=-\widetilde{A}^{-1}(\widetilde{G}-G)-(\widetilde{A}^{-1}-A^{-1})G.$
Taking $L^{2}(\mu)$ norms and using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) gives
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\,\|G\|_{L^{\infty}(D)}.$
Using $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$ (Step 1), $\|G\|_{L^{\infty}(D)}\leq M$ (assumption (ii)), and Step 2, we obtain
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
Finally, since $M\leq M+\|\widetilde{G}\|_{L^{\infty}(D)}$, this implies (97). ∎
C.2 Slab-wise small-gain for the $q$-aggregated projection inputs
Time-slab decomposition.
In the portfolio problem the working domain $D$ carries a time coordinate; for concreteness, assume
$D\subset[0,T]\times\mathcal{S},\hskip 18.49988pt\mu(dt,d\xi)=dt\otimes\nu(d\xi),$
for some reference measure $\nu$ on $\mathcal{S}$. Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
$D_{k}:=D\cap\big([t_{k-1},t_{k}]\times\mathcal{S}\big),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
Then $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{K}\|f\|_{k}^{2}$.
Proposition 2 (Short-time (slab) Lipschitz gain).
Let $\mathcal{U}$ be a neighborhood of $\pi^{\star}$ in the deployable $\theta$-blind policy class such that for all $\pi\in\mathcal{U}$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$
Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G_{\pi_{1}}^{\mathrm{mix}}-G_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (98)
for all $\pi_{1},\pi_{2}\in\mathcal{U}$ and constants $\bar{L}_{A},\bar{L}_{G}>0$ that depend only on band data. Define
$\rho(\tau):=\Big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\Big)\tau^{1/2}.$ (99)
Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\;\leq\;\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988ptT(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}.$ (100)
In particular, if the partition is chosen with $\max_{k}\tau_{k}\leq\tau^{\star}$ for some $\tau^{\star}>0$ such that $\rho(\tau^{\star})<1$, then
$\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1$
and $T$ is a contraction on every slab with constant at most $\rho_{*}$.
Proof.
Fix $k$ and $\pi_{1},\pi_{2}\in\mathcal{U}$. Write $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G_{\pi_{i}}^{\mathrm{mix}}$. Then
$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},$
and
$A_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$, we obtain
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\kappa\,\|G_{1}-G_{2}\|_{k}+\kappa^{2}M_{G}\,\|A_{1}-A_{2}\|_{k}.$
Applying (98) yields (100) with $\rho(\tau_{k})$ as in (99). ∎
Remark 3 (Verification of the $\tau^{1/2}$ gain and relation to prior slab analyses).
The $\tau^{1/2}$-gain in (98) is the same short-time parabolic smoothing effect used in our prior PGDPO analysis (see, e.g., huh2025breaking): one combines a Duhamel/semigroup representation of the relevant adjoint/costate objects with Young-type convolution bounds to obtain a factor $\tau^{1/2}$ on each short slab. In the present paper, the only additional bookkeeping is that $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ are $q$-aggregated (in particular, linear expectations over $\theta$), which does not alter the semigroup estimates; it only changes constants through coefficient bounds uniform in $\theta$ on the compact parameter set.
C.3 Proof of Theorem 3
Setup.
Let $D$ be the working domain with reference measure $\mu$ and slab decomposition $\{D_{k},\mu_{k},\|\cdot\|_{k}\}_{k=1}^{K}$ as in Appendix C.2. For a deployable $\theta$-blind policy $\pi$, write $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ for the $q$-aggregated projection inputs corresponding to the mixed-moment aggregation in (65)–(66). Define the projection map
$T(\pi)(z):=-A_{\pi}(z)^{-1}G_{\pi}^{\mathrm{mix}}(z),\hskip 18.49988ptz\in D,$
whenever $A_{\pi}(z)$ is invertible.
Let $\pi^{\mathrm{warm}}$ be the warm-up policy and set
$A_{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}},\hskip 18.49988ptG_{\mathrm{warm}}:=G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}},\hskip 18.49988pt\pi_{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
Let $\widehat{A}_{t}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ be the BPTT/Monte Carlo estimators constructed under $\pi^{\mathrm{warm}}$, and denote
$\widehat{\pi}^{\mathrm{agg,mix}}:=-\widehat{A}_{t}^{-1}\widehat{G}_{t}^{\mathrm{mix}}\hskip 18.49988pt\text{on }D.$
Proof.
Step 0 (Fixed-point form of the deployable optimum). Let $\pi^{\star}:=\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem. By the $q$-aggregated stationarity (Theorem 1), $\pi^{\star}$ satisfies
$A_{\pi^{\star}}(z)\,\pi^{\star}(z)=-G_{\pi^{\star}}^{\mathrm{mix}}(z)\qquad\text{for }\mu\text{-a.e.\ }z\in D,$
hence (under invertibility on $D$) it is a fixed point of $T$:
$\pi^{\star}(z)=T(\pi^{\star})(z)=-A_{\pi^{\star}}(z)^{-1}G_{\pi^{\star}}^{\mathrm{mix}}(z),\hskip 18.49988pt\mu\text{-a.e.\ }z\in D.$
Step 1 (Triangle decomposition). Add and subtract $\pi_{\mathrm{proj}}$:
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}.$ (101)
Step 2 (Estimation error via Proposition 1). Apply Proposition 1 with
$(A,G)=(A_{\mathrm{warm}},G_{\mathrm{warm}}),\hskip 18.49988pt(\widetilde{A},\widetilde{G})=(\widehat{A}_{t},\widehat{G}_{t}^{\mathrm{mix}}).$
Assume the perturbative regime on $D$:
$\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\qquad\|G_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq M_{G},\qquad\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq(2\kappa)^{-1},\qquad\|\widehat{G}_{t}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$ (102)
Then (97) yields
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}$ $\leq 2\kappa\,\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}+4\kappa^{2}M_{G}\,\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}.$ (103)
By the definition of $\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta})$ in Theorem 3,
$\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}+\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}\leq\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
hence
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),\hskip 18.49988ptC_{2}:=2\kappa+4\kappa^{2}M_{G}.$ (104)
Step 3 (Slab-wise warm-up bias bound). Assume $\pi^{\mathrm{warm}},\pi^{\star}\in\mathcal{U}$ and the slab-wise contraction $\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\|\pi_{1}-\pi_{2}\|_{k}$ from Proposition 2. Let $\rho_{*}:=\max_{k}\rho(\tau_{k})<1$. Since $\pi_{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}=T(\pi^{\star})$, for each slab $D_{k}$ we have
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$ (105)
Step 4 (Residual identity and slab-wise closure). Define the warm-up aggregated stationarity residual (mixed-moment) on $D$ by
$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}(z):=A_{\mathrm{warm}}(z)\,\pi^{\mathrm{warm}}(z)+G_{\mathrm{warm}}(z),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)}.$
Also define the slab-wise residual sizes
$\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k},\hskip 18.49988pt\text{so that }\ (\varepsilon_{\mathrm{warm}}^{\mathrm{mix}})^{2}=\sum_{k=1}^{K}(\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}})^{2}.$
By construction of $\pi_{\mathrm{proj}}$,
$\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}=A_{\mathrm{warm}}^{-1}r_{\mathrm{FOC,mix}}^{\mathrm{warm}},$
hence on each slab
$\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (106)
Now combine the triangle inequality on each slab with (105):
$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$
Since $\rho_{*}<1$, we close slab-wise:
$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\frac{\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (107)
Plugging into (105) gives
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (108)
Summing over slabs yields the global bias bound
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$ (109)
Step 5 (Finish). Combine (101), (104), and (109) to obtain
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}+C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
which is the slab-wise version of (71) (with $\rho_{*}$ in place of a global $C_{1}$). ∎
Remark 4 (Relation to prior PGDPO slab analyses).
The closure step above uses the same slab-wise small-gain philosophy as in huh2025breaking: short-time parabolic smoothing yields a contraction on each time slab, and the global bound follows by concatenation. The key difference here is that the contraction is applied to the $q$-aggregated projection map $T(\pi)=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$, hence the additional use of the algebraic projection stability (Proposition 1) for the estimator $\widehat{\pi}^{\mathrm{agg,mix}}$.
Appendix D Implementation details for Section 3
This appendix provides reproducible step-by-step templates for the methods in Section 3. The high-level pipeline is summarized in Figure LABEL:fig:sec3-pipeline-sideways. For compactness we present one template per subsection of Section 3.
D.1 Stage 1 (PG–DPO) template for Section 3.1
Stage 1 performs stochastic gradient ascent on the fixed-$q$ ex–ante objective (56), with latent $\theta\sim q$ sampled inside the simulator while the policy remains $\theta$-blind.
Inputs.
Policy parameters $\varphi$; sampler $\nu$ over initial states; prior $q$; time grid $(N,\Delta t)$; batch size $M$; optimizer and step size $\alpha$.
Template (one training iteration).
  1. 1.
    Sample initial states. Draw a mini-batch $\{z_{0}^{(i)}=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})\}_{i=1}^{M}\sim\nu$.
  2. 2.
    Sample latent environment parameter. Sample $\theta\sim q$ inside the simulator (unseen by $\pi_{\varphi}$). (Variant: sample $\theta^{(i)}\sim q$ independently per episode; both are unbiased for $\nabla_{\varphi}J(\varphi)$.)
  3. 3.
    Simulate Euler rollouts. For each episode $i$, simulate the Euler scheme in (53)–(54) under the $\theta$-blind policy $\pi_{\varphi}$ and collect terminal utilities $\{U(X_{T}^{(i)})\}_{i=1}^{M}$.
  4. 4.
    Backpropagation through time (BPTT). Compute the Monte Carlo gradient estimator
    $\widehat{g}\;\leftarrow\;\frac{1}{M}\sum_{i=1}^{M}\nabla_{\varphi}U(X_{T}^{(i)}).$
  5. 5.
[…]
  6. 6.
    Checkpoint. Periodically save a warm-up checkpoint $\varphi^{\mathrm{warm}}$ for stage 2 projection.
D.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
Stage 2 is a post-processing map: given a warm-up $\theta$-blind policy $\pi_{\varphi^{\mathrm{warm}}}$, it estimates Pontryagin sensitivity objects by Monte Carlo and constructs a deployable projected control on a working-domain sample $z\sim\mu$.
The aggregation used here matches the mixed-moment $q$-aggregation in (65)–(66), yielding the projected control (67).
Inputs.
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; working-domain sampler $\mu$ on $D$; budgets $(M_{z},M_{\theta},M_{\mathrm{MC}})$.
Template (constructing projection targets on a batch of query states).
  1. 1.
    Sample working-domain query states. Draw $\{z_{j}=(t_{j},x_{j},y_{j})\}_{j=1}^{M_{z}}\sim\mu$.
  2. 2.
    For each query state $z_{j}$, sample latent parameters. Sample $\{\theta_{\ell}\}_{\ell=1}^{M_{\theta}}\sim q$.
  3. 3.
    For each frozen $\theta_{\ell}$, estimate costates at $z_{j}$. For each $\ell=1,\dots,M_{\theta}$:
    1. (a)
      Simulate $M_{\mathrm{MC}}$ trajectories from $z_{j}$ under $\pi_{\varphi^{\mathrm{warm}}}$ with frozen $\theta_{\ell}$.
    2. (b)
      Compute pathwise sensitivities by autodiff/BPTT and average as in (62) to obtain $\widehat{p}_{t}^{\theta_{\ell}}(z_{j})$, $\widehat{p}_{x,t}^{\theta_{\ell}}(z_{j})$, $\widehat{p}_{y,t}^{\theta_{\ell}}(z_{j})$.
    3. (c)
      Form the $\theta$-conditional inputs (cf. (63)–(64)):
      $\widehat{A}_{t}^{\theta_{\ell}}(z_{j})\leftarrow x_{j}\,\widehat{p}_{x,t}^{\theta_{\ell}}(z_{j})\,\Sigma(y_{j},\theta_{\ell}),\hskip 18.49988pt\widehat{G}_{t}^{\theta_{\ell}}(z_{j})\leftarrow\widehat{p}_{t}^{\theta_{\ell}}(z_{j})\,b(y_{j},\theta_{\ell})+\Sigma_{SY}(y_{j},\theta_{\ell})\,\widehat{p}_{y,t}^{\theta_{\ell}}(z_{j}).$
  4. 4.
    Aggregate across $\theta\sim q$ (mixed-moment). Compute
    $\widehat{A}_{t}(z_{j})\;\leftarrow\;\frac{1}{M_{\theta}}\sum_{\ell=1}^{M_{\theta}}\widehat{A}_{t}^{\theta_{\ell}}(z_{j}),\hskip 18.49988pt\widehat{G}_{t}^{\mathrm{mix}}(z_{j})\;\leftarrow\;\frac{1}{M_{\theta}}\sum_{\ell=1}^{M_{\theta}}\widehat{G}_{t}^{\theta_{\ell}}(z_{j}),$
    consistent with (65)–(66).
  5. 5.
    Solve the projection (mixed-moment aggregation). Whenever $\widehat{A}_{t}(z_{j})$ is invertible and the solve is numerically stable, compute the deployable projected control
    $\widehat{\pi}^{\mathrm{agg,mix}}(z_{j})\;\leftarrow\;-\big(\widehat{A}_{t}(z_{j})\big)^{-1}\widehat{G}_{t}^{\mathrm{mix}}(z_{j}),$
    which matches (67).
D.3 Coupling I: residual/control-variate projection (Section 3.3.1)
This subsection records a variance-reduced implementation of the stage 2 map using the residual identity (74). The residual form is applied around the warm-up policy and uses the mixed-moment aggregated inputs $(\widehat{A}_{t},\widehat{G}_{t}^{\mathrm{mix}})$.
Inputs.
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section D.2.
Template (statewise residual projection; mixed-moment aggregation).
  1. 1.
    Evaluate warm-up control. Compute $\pi_{\varphi^{\mathrm{warm}}}(z)$.
  2. 2.
    Form the aggregated residual. Compute
    $\widehat{r}_{\mathrm{FOC}}(z)\;\leftarrow\;\widehat{A}_{t}(z)\,\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{G}_{t}^{\mathrm{mix}}(z).$
  3. 3.
    Apply the residual correction. Compute
    $\widehat{\pi}^{\mathrm{agg,mix}}(z)\;\leftarrow\;\pi_{\varphi^{\mathrm{warm}}}(z)-\big(\widehat{A}_{t}(z)\big)^{-1}\widehat{r}_{\mathrm{FOC}}(z).$
D.4 Coupling II: interactive distillation (Section 3.3.2)
This subsection records an implementation template for interactive distillation: the projected output from stage 2 is used as a teacher signal during stage 1 training via the mixed objective (76). The teacher is built from the mixed-moment projected rule (possibly evaluated in residual form for variance reduction).
Inputs.
Student parameters $\varphi$; teacher refresh interval $K$; distillation schedule $\lambda(n)$; working-domain sampler $\mu$.
Template (training loop with intermittent teacher refresh).
  1. 1.
    Initialize. Set $\varphi^{-}\leftarrow\varphi$ and initialize an empty teacher buffer $\mathcal{B}\leftarrow\emptyset$.
  2. 2.
    Repeat for iterations $n=1,2,\dots$:
    1. (a)
      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section D.1.
    2. (b)
      Teacher refresh (every $K$ steps). If $n\bmod K=0$:
      1. i.
        Set $\varphi^{-}\leftarrow\varphi$ (lagged copy).
      2. ii.
        Sample working-domain states $\{z_{j}\}_{j=1}^{M_{z}}\sim\mu$.
      3. iii.
        For each $z_{j}$, run stage 2 under $\pi_{\varphi^{-}}$ (mixed-moment aggregation) to compute a projected teacher $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j})$. (In practice we compute it in residual form around $\pi_{\varphi^{-}}$ as in (75).)
      4. iv.
        Optionally filter states using diagnostics (Section D.5) and update the buffer:
        $\mathcal{B}\ \leftarrow\ \{(z_{j},\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j}))\}_{j=1}^{M_{z}}\ \ \text{(after filtering)}.$
    3. (c)
      Distillation step (when enabled). If $\lambda(n)>0$ and $\mathcal{B}\neq\emptyset$:
      1. i.
        Sample $(z,\pi^{\mathrm{teach}})$ from $\mathcal{B}$.
      2. ii.
        Apply a gradient step to minimize the proximity term $\|\pi_{\varphi}(z)-\mathrm{stopgrad}(\pi^{\mathrm{teach}})\|^{2}$ with coefficient $\lambda(n)$, consistent with (76).
D.5 Engineering notes and stabilizers
This subsection collects practical stabilizers that we found helpful for reliable training and projection in high dimensions.
  • Antithetic sampling for $\theta$. When $q$ is symmetric (e.g. Gaussian in a latent normal parameterization), sample $\theta$ in antithetic pairs by drawing $z\sim\mathcal{N}(0,I)$ and using (z,z)(z,-z) to construct $(\theta^{+},\theta^{-})$. This reduces the variance of $q$-averaged quantities and typically improves the stability of stage 2 diagnostics on the working domain.
  • Blockwise Monte Carlo and robust aggregation. To control rare-tail domination, split Monte Carlo replications into $B$ blocks and compute blockwise averages of costate-driven ingredients (e.g. $\widehat{A}_{t}^{\theta}(z)$ and $\widehat{G}_{t}^{\theta}(z)$). Aggregate across blocks using a robust statistic such as the median or median-of-means, which makes the projection less sensitive to outlier trajectories.
  • Curvature/denominator stability checks. Because the projection map $(A,G)\mapsto-A^{-1}G$ can be sensitive to near-singularity of $A$, monitor the conditioning of $\widehat{A}_{t}$ (or failure rates of the linear solve). When diagnostics indicate ill-conditioning, skip projection-guided updates at that state or increase Monte Carlo budgets locally.
  • Residual magnitude as a reliability diagnostic. For the residual form, compute $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{G}_{t}^{\mathrm{mix}}(z)$. Small $\|\widehat{r}_{\mathrm{FOC}}(z)\|$ indicates approximate satisfaction of the mixed-moment aggregated first-order condition at $z$ and empirically correlates with more reliable teacher targets.
  • Diagnostics-based teacher selection on the working domain. Rather than applying distillation on all sampled $\{z_{j}\}\sim\mu$, keep only states that pass a reliability predicate. In practice, filter using residual-magnitude thresholds together with stable linear-solve diagnostics to prevent a small subset of pathological states from contaminating the teacher buffer.
  • $\lambda$ schedule and safeguards. Use a warm-up period with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after stage 2 diagnostics on the working domain are stable. To prevent the teacher term from dominating the ex–ante objective, cap the effective coefficient via
    $\lambda_{\mathrm{eff}}:=\min\Big\{\lambda,\;c\,\frac{|L_{\mathrm{main}}|}{L_{\mathrm{distill}}+\varepsilon}\Big\},$
    with $c\in(0,1)$ and $\varepsilon>0$.
  • Initialization and scale control in high dimensions. To avoid early-time numerical blow-ups (often through quadratic variation terms of the form $\pi^{\top}\Sigma\pi$), initialize the policy output near zero and/or scale the output by $d^{-1/2}$. As a last-resort safety net, a mild log-wealth clamp can prevent overflow, but it should be used conservatively and monitored, since frequent clamping may distort higher-order sensitivities.
Appendix E Stage 2 projection diagnostics
We report Stage 2 diagnostic statistics as a visual supplement to Section 4.2. Each figure summarizes the same tail-median protocol and layout; see captions for definitions and interpretation.
Figure 3: Stage 2 stationarity residual (q50). All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We plot the median (q50) of the estimated Hamiltonian first-order condition residual norm at the query states. Larger residual indicates the warm policy is farther from stationarity, implying a larger correction is required in the residual-form projection. Growth of this residual with $d$ (especially under misalignment) supports the mechanism that projection becomes more sensitive in high dimension due to larger correction magnitudes and amplified mixed-moment noise.
Figure 4: Stage 2 denominator magnitude (q50). All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We plot a typical (q50) magnitude of the projection denominator/curvature term used in the residual-form update. Values bounded away from zero indicate that projection is not operating in a near-singular regime at typical quantiles. This helps rule out “catastrophic inversion” as the primary driver of degradation; instead, residual growth and curvature mismatch (Fig. 5) provide a more consistent explanation in misaligned/high-$d$ regimes.
Figure 5: Stage 2 curvature-consistency statistic $\kappa$. All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We report the stabilized median-after-floor statistic $\kappa$ and compare it to the nominal reference $1/\gamma$ (horizontal dotted line). For CRRA, costate ratios imply a characteristic curvature scale; sustained deviations of $\kappa$ from $1/\gamma$ indicate costate inconsistency and/or bias in mixed-moment estimation, and are most visible in the hardest misaligned/high-uncertainty regime.
Figure 6: Stage 2 bad-sign fraction. All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We plot the fraction of samples in which the estimated curvature/denominator violates the expected sign condition (loss of local concavity on the sampled batch). Near-zero bad-sign fractions across most regimes suggest that the projection typically operates in a locally well-behaved region and that failures are not dominated by sign flips, supporting the main-text conclusion that misalignment primarily increases residual/costate mismatch rather than inducing widespread concavity violations.
Appendix F Supplementary decomposition diagnostics for Section 5
Tables 23 report Stage 2 decomposition diagnostics at $t=0$.
$s_{0}$ Method $d=5$ $10$ $50$ $100$
Aligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $7.17\text{\times}{10}^{-06}$ $1.51\text{\times}{10}^{-05}$ $1.60\text{\times}{10}^{-05}$ $1.23\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-06}$ $9.81\text{\times}{10}^{-06}$ $1.64\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $6.21\text{\times}{10}^{-06}$ $1.38\text{\times}{10}^{-05}$ $1.63\text{\times}{10}^{-05}$ $1.42\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $7.13\text{\times}{10}^{-06}$ $7.13\text{\times}{10}^{-06}$ $1.62\text{\times}{10}^{-05}$ $1.76\text{\times}{10}^{-05}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $8.18\text{\times}{10}^{-06}$ $1.41\text{\times}{10}^{-05}$ $3.82\text{\times}{10}^{-05}$ $2.42\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $6.72\text{\times}{10}^{-06}$ $9.21\text{\times}{10}^{-06}$ $3.67\text{\times}{10}^{-05}$ $3.16\text{\times}{10}^{-05}$
Misaligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $1.10\text{\times}{10}^{-05}$ $1.41\text{\times}{10}^{-05}$ $70\text{\times}{10}^{-05} 1.1$$18\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $7.71\text{\times}{10}^{-06}$ $5.94\text{\times}{10}^{-06}$ $1.84\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $7.90\text{\times}{10}^{-06}$ $2.02\text{\times}{10}^{-05}$ $1.46\text{\times}{10}^{-05}$ $1.20\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $6.00\text{\times}{10}^{-06}$ $1.71\text{\times}{10}^{-05}$ $2.21\text{\times}{10}^{-05}$ $1.43\text{\times}{10}^{-05}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $1.24\text{\times}{10}^{-05}$ $1.93\text{\times}{10}^{-04}$ $5.77\text{\times}{10}^{-05}$ $3.14\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $1.20\text{\times}{10}^{-05}$ $1.90\text{\times}{10}^{-04}$ $7.40\text{\times}{10}^{-05}$ $2.47\text{\times}{10}^{-05}$
Table 2: Myopic-component RMSE at $t=0$ (tail medians).
$s_{0}$ Method $d=5$ $10$ $50$ $100$
Aligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $0.994$ $0.988$ $0.991$ $0.990$
Stage 1+Stage 2 (Distill.) $0.995$ $0.986$ $0.990$ $0.987$
$10^{-2}$ Stage 1+Stage 2 (Basic) $0.993$ $0.989$ $0.992$ $0.988$
Stage 1+Stage 2 (Distill.) $0.992$ $0.994$ $0.990$ $0.987$
$10^{-1}$ Stage 1+Stage 2 (Basic) $0.988$ $0.990$ $0.936$ $0.932$
Stage 1+Stage 2 (Distill.) $0.996$ $0.990$ $0.949$ $0.922$
Misaligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $0.988$ $0.988$ $0.993$ $0.990$
Stage 1+Stage 2 (Distill.) $0.994$ $0.995$ $0.990$ $0.987$
$10^{-2}$ Stage 1+Stage 2 (Basic) $0.994$ $0.976$ $0.992$ $0.988$
Stage 1+Stage 2 (Distill.) $0.994$ $0.980$ $0.988$ $0.989$
$10^{-1}$ Stage 1+Stage 2 (Basic) $0.990$ $0.005$ $0.668$ $0.851$
Stage 1+Stage 2 (Distill.) $0.992$ 0.009-0.009 $0.642$ $0.871$
Table 3: Hedging-direction cosine similarity at $t=0$ (tail medians). Higher is better; negative indicates direction reversal.

### v2

@@
 Abstract
-We study continuous-time portfolio choice in diffusion markets with parameter $\theta\in\Theta$ and uncertainty law $q(d\theta)$.
-Nature draws latent $\theta\sim q$ at time 0; the investor cannot observe it and must deploy a single $\theta$-blind feedback policy maximizing an ex–ante CRRA objective averaged over diffusion noise and $\theta$.
-Our methods access $q$ only by sampling and assume no parametric form.
-We extend Pontryagin–Guided Direct Policy Optimization (PG–DPO) by sampling $\theta$ inside the simulator and computing discrete-time gradients via backpropagation through time (BPTT), and we propose projected PG–DPO (P–PGDPO) that projects costate estimates to satisfy the $q$-aggregated Pontryagin first-order condition, yielding a deployable rule.
-We prove a BPTT–PMP correspondence uniform on compacts and a residual-based $\theta$-blind policy-gap bound under local stability with explicit discretization/Monte Carlo errors; experiments show projection-driven stability and accurate decision-time benchmark recovery in high dimensions.
+We study continuous-time CRRA portfolio choice in diffusion markets with uncertain estimated coefficients.
+Nature draws a latent parameter from a given distribution and keeps it fixed; the investor cannot observe this parameter and must commit to a parameter-blind policy maximizing an ex-ante objective.
+We treat the uncertainty distribution as an inference-agnostic sampling input.
+We develop a simulation-only two-stage solver.
+Stage 1 extends Pontryagin-Guided Direct Policy Optimization (PG-DPO) by sampling parameters internally and computing gradients via backpropagation through time.
+Stage 2 performs an aggregated Pontryagin projection: it aggregates costates across the parameter distribution to enforce a deployable stationarity condition, yielding a structured correction amortized via interactive distillation.
+We prove a uniform conditional BPTT-PMP correspondence and a residual-based policy-gap bound with explicit error terms.
+Experiments on high-dimensional Gaussian drift and factor-driven benchmarks show that projection stabilizes learning and accurately recovers analytic references, while a model-free PPO baseline remains far from the targets.
+Keywords: Portfolio choice; Parameter uncertainty; Deep learning; Pontryagin principle; Stochastic control
 1 Introduction
 A central problem in quantitative finance is to allocate wealth dynamically across many risky assets in continuous time.
-In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and the investor solves a Hamilton–Jacobi–Bellman (HJB) equation to obtain closed-form optimal portfolios and value functions; see, for example, merton1969lifetime; merton1971optimum and the subsequent literature.
-In realistic markets, however, neither expected returns nor volatilities are known: they must be estimated from finite samples, often using many assets and predictors and under nontrivial model selection and regularization.
-Empirically, expected-return forecasting is fragile and return predictability is unstable across samples and over time, with many proposed predictors delivering limited out-of-sample gains (e.g., goyalwelch2008; campbellthompson2008; rapach2010forecast; danglhalling2012tvp; lettauvanNieuwerburgh2008shifts; pettenuzzo2014constraints).
-In such settings it is crucial to distinguish diffusion risk (Brownian noise conditional on parameters) from statistical parameter uncertainty (error in estimated coefficients).
-A long line of portfolio-choice work shows that return predictability, learning, and parameter uncertainty can induce substantial intertemporal hedging effects and more conservative allocations (e.g., kandelstambaugh1996predictability; barberis2000investor; campbell2002strategic; brandt2005portfolio; brennan1998investor; xia2001learning; maenhout2004robust).
-We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, where the estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$.
-We treat $q$ as an input object that encapsulates all statistical information available at time 0: it may be derived from resampling approximations (e.g.
-bootstrap) (e.g., efron1979bootstrap; efron1994bootstrap), model averaging or Bayesian model uncertainty pipelines (e.g., pastor2000portfolio; avramov2002stock; cremers2002stock), approximate posteriors, or other procedures.
+In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and closed-form optimal portfolios follow from HJB analysis; see, for example, Merton (1969, 1971).
+In realistic markets, however, key inputs—especially expected returns—must be estimated from finite samples, often in high dimensions and under regularization/model selection.
+Empirically, return predictability is unstable and many predictors deliver limited out-of-sample gains (e.g., Goyal and Welch, 2008; Campbell and Thompson, 2008).
+These considerations motivate separating diffusion risk (Brownian noise conditional on parameters) from decision-time estimation error (uncertainty in estimated coefficients), a theme that also underlies intertemporal hedging and more conservative allocations in models with predictability/learning (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Campbell and Viceira, 2002; Xia, 2001; Maenhout, 2004).
+We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, and an external estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$.
+We treat $q$ as an input object that summarizes decision-time estimation error about $\theta$.
+Concretely, $q$ may be obtained from resampling-based approximations (e.g., the bootstrap) (e.g., Efron, 1979; Efron and Tibshirani, 1994), from subsample aggregation and model-selection variability (e.g., bagging), or from asymptotic (sandwich/normal) approximations of an estimator’s sampling distribution (van der Vaart, 1998).
 Our goal is not to revisit inference, but to optimize decisions given this uncertainty description.
-Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside the simulator; we do not assume closed-form densities, conjugate updates, or any particular parametric form.
-Concretely, we seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$.
-This formulation also supports offline diagnostics that quantify how recommended allocations vary across the statistically plausible models encoded in $q$.
+Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside a simulator; we do not assume a closed-form density, conjugacy, or a particular parametric family.
+We seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$.
 A key modeling choice is that $\theta$ is latent: Nature draws a fixed but unobserved $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed on $[0,T]$.
-While the investor knows $q$, she does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy.
-We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a belief/posterior process.
-This fixed-$q$ commitment is intended as a decision-time benchmark: it targets a single deployable rule given an exogenous uncertainty law, cleanly decoupling how uncertainty is produced (any pipeline yielding $q$) from how decisions are optimized (our solver given $q$).
-We do not claim that belief-state control is conceptually inappropriate; rather, it defines a different (and typically far more demanding) problem than computing a single $\theta$-blind deployable feedback rule from a fixed uncertainty description (e.g., bensoussan1985optimal; pham2017dynamic).
-At the same time, fixed-$q$ optimization couples heterogeneous market models: gradient signals can vary substantially across $\theta$ draws and may partially cancel when learning a single global policy end-to-end.
-The $\theta$-blind constraint also changes what a first-order optimality condition means.
-If $\theta$ were observable, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition and an associated $\theta$-conditional full-information feedback map (infeasible under latent $\theta$).
-Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind.
-Taking the first variation of the ex–ante objective and using Fubini’s theorem shows that the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the state process, in the standard stochastic maximum principle framework (e.g., yong1999stochastic; fleming2006controlled; pham2009continuous).
-Because $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$ for our portfolio Hamiltonian, this aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projected portfolio rule.
-Notably, the condition and resulting projection are agnostic to the internal construction of $q$ and depend only on its role as the ex–ante mixing law.
-These features place the problem outside the practical reach of classical dynamic programming in the high-dimensional regime we target.
-In low-dimensional deterministic-parameter Markov models, DP/HJB is canonical; however, even with several factors it requires solving an HJB equation in the state $(t,X_{t},Y_{t})$, where grid-based PDE methods are quickly defeated by the curse of dimensionality (e.g., bellman1961adaptive; kushner2001numerical).
-Deep PDE surrogates such as PINNs (e.g., raissi2019physics; sirignano2018dgm) and deep BSDE methods (e.g., han2018solving; beck2019machine) alleviate the need for grids, but fully nonlinear portfolio HJBs with many assets and factors remain numerically delicate, especially when accurate mixed derivatives are required.
-If one further models parameter uncertainty via belief-state augmentation, the state becomes a posterior measure and the control problem becomes infinite-dimensional (e.g., bensoussan1985optimal; pham2017dynamic).
-Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (huh2025breaking; huh2025constraint).
-PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and employs backpropagation through time (BPTT) to compute exact gradients of terminal utility.
-Crucially, intermediate pathwise sensitivities computed by BPTT coincide with the stochastic costates (adjoints) in PMP, mirroring the classical duality between backpropagation and adjoint methods (see lecun1988optimal; yong1999stochastic).
-In the latent-parameter setting, we approximate the ex–ante objective by sampling $\theta\sim q$ inside the simulator and fixing it along each trajectory, while the policy depends only on observable states.
-To stabilize learning under heterogeneous $\theta$ draws, we extend the projected variant, P–PGDPO, to latent $\theta$: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated Pontryagin first-order condition.
-This reconstruction yields a robust deployable $\theta$-blind rule obtained from the $q$-aggregated criticality, and can be amortized into a fast-to-evaluate policy.
-In high-dimensional scaling experiments under static Gaussian drift uncertainty, the two-stage projected pipeline substantially improves decision-time accuracy relative to end-to-end learning, with clear stabilization effects in aligned regimes.
-In misaligned regimes, projection gains diminish with dimension; diagnostics indicate that deterioration is driven primarily by growth of aggregated first-order residuals and curvature mismatch rather than by catastrophic numerical inversion.
-In factor-driven markets with mean-reverting investment opportunities where return–factor correlation induces intertemporal hedging demand, the projected pipeline recovers the analytic decision-time benchmark under the same $\theta$-blind deployability restriction, while a model-free PPO baseline remains far from the reference in the regimes we test.
-Our main theoretical guarantee is a residual-based ex–ante $\theta$-blind policy-gap bound for the deployable fixed-$q$ commitment problem: under mild slab-wise local stability conditions for the $q$-aggregated projection map, a small warm-up aggregated first-order residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal interior deployable $\theta$-blind policy, up to discretization and Monte Carlo error.
+The investor knows $q$ but does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy.
+We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a filtering/belief process.
+This is a one-shot decision-time problem: $q$ is treated as fixed over the trading horizon, and we do not model or hedge subsequent updates of the uncertainty description.
+Belief-state control (optimal learning/hedging of parameter uncertainty) is conceptually important, but it defines a different and typically far more demanding problem than computing a single deployable $\theta$-blind rule from a fixed $q$ (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+The $\theta$-blind deployability constraint also changes what a first-order optimality condition means.
+Under full information, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition, hence a $\theta$-conditional feedback map that is infeasible when $\theta$ is latent.
+Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind, and the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the controlled state process (e.g., Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009).
+In our portfolio setting, $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$, so $q$-aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projection map.
+This condition is inference-agnostic: it depends on $q$ only through its role as the ex–ante mixing law.
+These features place the problem beyond the reach of classical dynamic programming in the targeted high-dimensional regime.
+Even in moderate dimensions, HJB solvers face the curse of dimensionality (e.g., Bellman, 1961; Kushner and Dupuis, 2001).
+Although deep PDE (e.g., Raissi et al., 2019; Sirignano and Spiliopoulos, 2018) and BSDE methods (e.g., Han et al., 2018; Beck et al., 2019) reduce grid reliance, nonlinear portfolio HJBs remain numerically delicate.
+Furthermore, belief augmentation renders the problem infinite-dimensional (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+Accordingly, we exclude value-function solvers as practical baselines under our $\theta$-blind interface, validating instead on closed-form benchmarks and simulation-only model-free RL (e.g., PPO).
+Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (Huh et al., 2025b, a).
+PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and uses backpropagation through time (BPTT) to compute exact discrete-time gradients of terminal utility.
+Pathwise sensitivities computed by BPTT coincide with stochastic costates in PMP, reflecting the classical adjoint/backpropagation correspondence (e.g., LeCun, 1988; Yong and Zhou, 1999).
+In the latent-parameter setting, we sample $\theta\sim q$ inside the simulator (fixing it along each trajectory), while the policy depends only on observable states.
+To stabilize learning under heterogeneous $\theta$ draws, we develop uncertainty-aware projected P–PGDPO: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated PMP first-order condition, producing a robust deployable $\theta$-blind rule that can be amortized into a fast policy.
+Empirically, in high-dimensional static drift experiments, the projected pipeline improves decision-time accuracy over end-to-end learning, stabilizing aligned regimes.
+In factor-driven markets where correlation induces hedging demand, it recovers analytic references under the $\theta$-blind restriction, whereas a PPO baseline remains far off.
+Theoretically, we establish a residual-based policy-gap bound: under mild stability, a small aggregated residual implies the policy is close to a locally optimal deployable rule, up to discretization and Monte Carlo errors.
 Our contributions are threefold.
-(i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated Pontryagin first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$.
+(i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated PMP first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$.
 (ii) We extend PG–DPO to this setting by sampling $\theta$ only inside the simulator and using BPTT to compute exact discrete-time gradients and pathwise sensitivities, and we establish a conditional BPTT–PMP correspondence uniform over $\theta$ on compact subsets of $\Theta$.
-(iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization and stability gains from projection.
+(iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization.
 The remainder of the paper is organized as follows.
-Section 2 formulates the fixed-$q$ ex–ante portfolio problem under a latent parameter and a deployable $\theta$-blind Markov feedback restriction, and derives the $\theta$-conditional versus $q$-aggregated Pontryagin first-order conditions together with Gaussian decision-time reference models.
-Section 3 develops PG–DPO and uncertainty-aware P–PGDPO for the latent-$\theta$ setting, establishes the conditional BPTT–PMP correspondence, and proves a residual-based ex–ante $\theta$-blind policy-gap bound under local stability of the aggregated projection map.
-Section 4 reports high-dimensional scaling experiments under static Gaussian drift uncertainty, and Section 5 studies hedging-demand recovery in factor-driven markets with mean-reverting investment opportunities.
-Technical proofs and implementation details are collected in the appendix.
+Section 2 formulates the fixed-$q$ ex–ante problem and derives $\theta$-conditional versus $q$-aggregated PMP conditions together with Gaussian decision-time reference models.
+Section 3 develops PG–DPO and P–PGDPO for latent $\theta$ and proves the policy-gap bound.
+Sections 4 and 5 report numerical experiments, and the appendix collects technical proofs and implementation details.
 2 Dynamic Portfolio Choice in Estimated Diffusion Markets with Latent Parameter Uncertainty
-In this section we formulate a continuous-time dynamic portfolio choice problem with CRRA preferences in a diffusion market whose coefficients are estimated from data and are therefore statistically uncertain.
-Rather than committing to a particular estimation architecture, we treat the market as belonging to a (possibly high-dimensional) parameterized family indexed by $\theta\in\Theta$, and we represent the uncertainty in the estimated parameter by an exogenously given probability law $q(d\theta)$ over $\Theta$.
-  • Nature draws a fixed but unobserved $\theta\sim q$ at time 0 and keeps it constant on $[0,T]$.
-  • The investor knows $q$ but does not observe the realized $\theta$ and must deploy a single $\theta$-blind portfolio policy.
-  • Performance is evaluated ex–ante by averaging terminal utility over both diffusion noise and $\theta\sim q$.
-  • We restrict to $\theta$-blind Markov feedback rules $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ and do not augment the state by a belief/posterior process.
-This fixed-$q$, $\theta$-blind formulation is intentionally algorithm-facing: we view the estimation procedure that produced $q(d\theta)$ as exogenous, and our goal is to compute stable, scalable portfolio rules given this uncertainty description.
-It is also a commitment model: the investor commits at time 0 to a single feedback map and does not update $q$ during trading.
-As a result, one must distinguish between (i) $\theta$-conditional (full-information) optimality conditions and objects that would be available if $\theta$ were observable (infeasible under latent $\theta$), and (ii) $q$-aggregated conditions that characterize optimality within the $\theta$-blind admissible class.
-Throughout, $\theta$-conditional objects are used only for offline diagnostics (e.g., heterogeneity inspection and infeasible upper bounds), whereas our algorithms target a single deployable $\theta$-blind rule; see Section 2.2.
+In this section we formalize a continuous-time CRRA portfolio problem in an estimated diffusion market with latent parameter uncertainty.
+Rather than fixing an inference architecture, we take as input an exogenous decision-time law $q(d\theta)$ over $\Theta$ and work with a latent, time-constant parameter $\theta\sim q$ (see Section 2.1 for the full model and objective).
+We adopt a $\theta$-blind commitment viewpoint: the investor commits at the decision time to a single Markov feedback rule $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ and does not update $q$ during trading.
+This restriction shifts the relevant optimality notion from infeasible $\theta$-conditional (full-information) conditions to a $q$-aggregated Pontryagin stationarity condition within the $\theta$-blind admissible class (Section 2.2).
+Closed-form Gaussian decision-time reference models used for validation are collected in Section 2.3.
 2.1 Model and ex–ante objective in estimated diffusion markets
-We work on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ supporting Brownian motions of appropriate dimension.
-Time is continuous and runs over a fixed finite horizon $[0,T]$.
-Deterministic-parameter reference (classical CRRA Merton).
-There is one risk-free asset (money market account) with price process $B$ satisfying
-$\frac{dB_{t}}{B_{t}}=r\,dt,\hskip 18.49988ptB_{0}=1,$ (1)
-where $r\in\mathbb{R}$ is a constant short rate. In the classical Merton model, the $d$ risky assets have prices $S_{t}=(S_{t}^{1},\dots,S_{t}^{d})^{\top}$ solving
-$\frac{dS_{t}}{S_{t}}:=\bigg(\frac{dS_{t}^{1}}{S_{t}^{1}},\dots,\frac{dS_{t}^{d}}{S_{t}^{d}}\bigg)^{\top}=r\,\mathbf{1}\,dt+\mu\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (2)
-with constant excess returns $\mu\in\mathbb{R}^{d}$, volatility matrix $\Sigma^{1/2}\in\mathbb{R}^{d\times d}$, and a $d$-dimensional Brownian motion $W$.
-An investor with CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$, chooses a progressively measurable portfolio fraction $\pi_{t}\in\mathbb{R}^{d}$; the wealth process satisfies
-$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\lambda\big)dt+\pi_{t}^{\top}\Sigma^{1/2}dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (3)
-where $\lambda:=\mu$ is the vector of risk premia. In this benchmark setting the optimal policy is constant:
-$\pi^{\star}=\frac{1}{\gamma}\,\Sigma^{-1}\lambda,$ (4)
-and the corresponding value function is given explicitly by
-$V^{\mathrm{Merton}}(t,x;\lambda)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)\Big(r+\tfrac{1}{2\gamma}\lambda^{\top}\Sigma^{-1}\lambda\Big)(T-t)\Big\},$ (5)
-see, for example, merton1969lifetime; merton1971optimum. We use this constant-coefficient model only as a deterministic-parameter reference.
+We interpret time 0 as the decision time and study portfolio choice on a fixed horizon $[0,T]$.
+The uncertainty law $q(d\theta)$ is an exogenous input available at the decision time and is treated as fixed over $[0,T]$ (no online belief updates during trading).
 Estimated diffusion market family (conditional on a latent parameter).
-In our main formulation, drift and volatility are not assumed known.
+Drift and volatility are not assumed known.
 Instead, we consider a general multi-asset, multi-factor diffusion family indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$, where $\theta$ represents the (possibly high-dimensional) parameter produced by an estimation procedure.
@@
 Uncertainty law $q(d\theta)$ and information structure.
-The parameter $\theta$ is estimated from finite samples and is uncertain. We summarize this uncertainty by a probability distribution
+We summarize the uncertainty of the parameter $\theta$, estimated from finite samples, by a probability distribution
 $q(d\theta).$ (4)
-We deliberately do not tie $q$ to any specific inference paradigm.
-Concretely, $q$ may represent an empirical/sampling distribution produced by resampling procedures such as the bootstrap (efron1979bootstrap; efron1994bootstrap), a distribution induced by model averaging or sub-sample aggregation procedures such as bagging (breiman1996bagging), an approximate Bayesian posterior (when a prior and likelihood/criterion are specified), or an asymptotic normal (or sandwich) approximation in parametric or semiparametric estimation.
-For our purposes, $q$ is an input object describing statistically plausible market parameters.
+We treat $q$ as inference-agnostic; it may represent resampling-based distributions (e.g., bootstrap (Efron, 1979; Efron and Tibshirani, 1994), bagging (Breiman, 1996)) or asymptotic approximations (van der Vaart, 1998).
+For our purposes, $q$ is a fixed decision-time input describing plausible market parameters over $[0,T]$.
 Remark 1 (Latent parameter, observability, and admissible controls).
-We interpret $\theta$ as a latent (unobserved) market parameter: Nature draws an $\mathcal{F}_{0}$-measurable random variable $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed over $[0,T]$.
+We interpret $\theta$ as a latent (unobserved) market parameter: at the (shifted) decision time 0, Nature draws an $\mathcal{F}_{0}$-measurable random variable $\theta\sim q$ (independent of the Brownian drivers) and keeps it fixed over $[0,T]$.
 The investor knows $q$ but does not observe the realized $\theta$, so deployable portfolio rules cannot take $\theta$ as an input.
@@
 where $\mathcal{A}^{\mathrm{obs}}$ is defined below.
-This restriction reflects a fixed-$q$ commitment model: the investor uses historical data to form $q$ prior to trading and does not perform online filtering/belief-state updates during $[0,T]$.
+This restriction reflects a fixed-$q$ commitment model: the investor forms $q$ at the decision time and does not perform online filtering/belief-state updates during $[0,T]$.
 Whenever we display $\theta$-conditional (full-information) controls or sensitivity objects, they are computed under frozen-$\theta$ simulations and are used only for offline diagnostics; the deployed policy class and the learned policy remain $\theta$-blind.
@@
 Ex–ante averaging in (8) can also create gradient cancellation across heterogeneous parameter draws when one attempts to learn a single global policy end-to-end.
-While an $\mathcal{F}_{t}^{\mathrm{obs}}$-adapted policy could, in principle, filter $\theta$ online and solve a belief-state control problem (see, e.g., bensoussan1985optimal; pham2017dynamic), we do not pursue that formulation here.
+While an $\mathcal{F}_{t}^{\mathrm{obs}}$-adapted policy could, in principle, filter $\theta$ online and solve a belief-state control problem (see, e.g., Bensoussan and van Schuppen (1985); Pham and Wei (2017)), we do not pursue that formulation here.
 Approximating the outer expectation in (8) amounts to sampling $\theta\sim q$ inside the simulator (once per trajectory or once per update), running (1)–(2) under that frozen draw, and updating a $\theta$-blind feedback policy to perform well on average over such draws.
 This is the setting targeted by the simulation-based PG–DPO and P–PGDPO methods developed in Section 3.
+Decision-time commitment and (optional) plug-in replanning.
+Throughout, we treat $q(d\theta)$ as an exogenous decision-time input and keep it fixed over the trading horizon $[0,T]$: within a single run, we do not update $q$ using newly observed returns or additional data, and we do not hedge against future changes of the uncertainty description.
+If the decision is revisited at a later calendar time $t_{0}\in(0,T)$, we interpret this as starting a new decision-time problem on the remaining horizon with an externally supplied law $q_{t_{0}}$.
+Such a law may coincide with a static input (Section 2.3.1), may arise as a model-implied predictive (prior) propagation of an initial uncertainty through factor dynamics without conditioning on new observations (Section 2.3.2), or may be produced by an external filtering/estimation routine that outputs a time-indexed uncertainty description (Appendix A).
+Our model and algorithms are conditional on whichever $q$ is supplied at the decision time and always freeze that input over the ensuing horizon.
 2.2 Pontryagin optimality under latent parameters: full-information vs. aggregated conditions
 This subsection records the Hamiltonian structure underlying our projection step and clarifies what “Pontryagin first-order conditions” mean when the market parameter $\theta$ is latent and admissible controls are $\theta$-blind.
+Throughout, $q(d\theta)$ denotes the decision-time input law (Section 2.1) and is treated as fixed over the trading horizon $[0,T]$.
+All expectations in this subsection are taken conditional on the decision-time information (suppressed in notation).
 In particular, we distinguish between (i) $\theta$-conditional (full-information) criticality conditions that would apply if $\theta$ were observable (and are therefore infeasible under latent $\theta$), and (ii) $q$-aggregated criticality conditions that characterize stationarity within the $\theta$-blind admissible class for the fixed-$q$ ex–ante objective.
 Our discussion follows standard stochastic control/PMP arguments for diffusion control (e.g.
-yong1999stochastic; fleming2006controlled; pham2009continuous).
+Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009).
 We also comment on the relationship to partial-information (belief-state) PMP, but we do not develop that formulation here.
@@
 Fix $\theta\in\Theta$ and suppose, for the moment, that $\theta$ were observable to the controller.
-In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (8):
-$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,V_{x},V_{xx},V_{xy}):=x\,\pi^{\top}b(y,\theta)\,V_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,V_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,V_{xy},$ (15)
-where $V_{x},V_{xx}$ are evaluated at $(t,x,y)$ and $V_{xy}(t,x,y)\in\mathbb{R}^{m}$. The last term in (15) is the return–factor hedging term induced by $d\langle W,W^{Y}\rangle\neq 0$.
-The pointwise first-order condition for an interior optimizer is
-$\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}=x\,V_{x}^{\star,\theta}\,b(y,\theta)+x^{2}\,V_{xx}^{\star,\theta}\,\Sigma(y,\theta)\,\pi+x\,\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}\;=\;0.$ (16)
+In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (3).
+For clarity, we define the control Hamiltonian map as a function of generic “sensitivity” inputs $(\varphi_{x},\varphi_{xx},\varphi_{xy})$:
+$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,\varphi_{x},\varphi_{xx},\varphi_{xy}):=x\,\pi^{\top}b(y,\theta)\,\varphi_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,\varphi_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,\varphi_{xy},$ (10)
+where $\varphi_{x},\varphi_{xx}\in\mathbb{R}$ and $\varphi_{xy}\in\mathbb{R}^{m}$.
+In the full-information HJB interpretation, we evaluate (10) at $(\varphi_{x},\varphi_{xx},\varphi_{xy})=(V_{x},V_{xx},V_{xy})$.
+In the Pontryagin interpretation used below, the same affine map is evaluated at the corresponding $\theta$-conditional adjoint/sensitivity objects (denoted $(p^{\theta},p_{x}^{\theta},p_{y}^{\theta})$ along a trajectory); in smooth Markov regimes these objects coincide with the indicated value derivatives.
+The pointwise first-order condition for an interior optimizer is therefore
+$\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}=x\,V_{x}^{\star,\theta}\,b(y,\theta)+x^{2}\,V_{xx}^{\star,\theta}\,\Sigma(y,\theta)\,\pi+x\,\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}\;=\;0,$ (11)
+where $V_{x}^{\star,\theta},V_{xx}^{\star,\theta},V_{xy}^{\star,\theta}$ are evaluated at $(t,x,y)$.
 Assuming $\Sigma(y,\theta)$ is invertible and $V_{xx}^{\star,\theta}<0$, this yields the closed-form $\theta$-conditional full-information portfolio rule
@@
 $\left.\frac{d}{d\varepsilon}J^{\theta}(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)^{\top}h_{t}\,dt\,\Big|\,\theta\right],$ (13)
-where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market.
-Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
+where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity/adjoint objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market.
+(Equivalently, (13) evaluates the same affine map $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ defined by (10), with $(V_{x},V_{xx},V_{xy})$ replaced by the corresponding Pontryagin objects along the controlled trajectory.) Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
 $\left.\frac{d}{d\varepsilon}J(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\mathbb{E}_{\theta\sim q}\!\Big[\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)\Big]^{\top}h_{t}\,dt\right].$ (14)
@@
 The conditional first-variation identity (13) is standard for diffusion control under a fixed parameter $\theta$ (e.g.
-yong1999stochastic; fleming2006controlled; pham2009continuous).
+Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009).
 Taking the outer expectation over $\theta\sim q$ yields (14).
@@
 If one allows history-dependent policies that explicitly infer $\theta$ from observed returns, a principled partial-information formulation introduces a time-varying posterior/belief state $q_{t}(\cdot)=\mathbb{P}(\theta\in\cdot\mid\mathcal{F}_{t}^{\mathrm{obs}})$.
-In such belief-state problems, the corresponding PMP/Hamiltonian criticality condition is expressed in terms of conditional expectations under $q_{t}$ (or, equivalently, conditional on $\mathcal{F}_{t}^{\mathrm{obs}}$); see, e.g., haussmann1987maximum; li1995general; baghery2007maximum.
+In such belief-state problems, the corresponding PMP/Hamiltonian criticality condition is expressed in terms of conditional expectations under $q_{t}$ (or, equivalently, conditional on $\mathcal{F}_{t}^{\mathrm{obs}}$); see, e.g., Haussmann (1987); Li and Tang (1995); Baghery and Øksendal (2007).
 We do not pursue that learning/belief-state route here.
-Our algorithms and theory target the fixed-$q$, $q$-aggregated projection (25) under the $\theta$-blind Markov feedback restriction (52).
+Our algorithms and theory target the fixed-$q$, $q$-aggregated projection (20) under the $\theta$-blind Markov feedback restriction.
 2.3 Gaussian references at a fixed decision time
-This subsection collects Gaussian benchmarks that isolate decision-time statistical uncertainty and yield closed-form reference allocations.
-We fix a calendar decision time $t_{0}$ at which an external estimation pipeline outputs an uncertainty law $q_{t_{0}}(d\theta)$ for a risk-premium parameter, and we treat this law as an $\mathcal{F}_{t_{0}}$-measurable input for portfolio choice over the remaining horizon.
-This interface accommodates both Bayesian posterior/prior-like uncertainty descriptions (e.g., barberis2000investor; pastor2000portfolio) and frequentist sampling/resampling laws conditional on $\mathcal{F}_{t_{0}}$ (e.g., bootstrap or bagging) (e.g., efron1979bootstrap; efron1994bootstrap; breiman1996bagging).
-Throughout this subsection we work conditionally on $\mathcal{F}_{t_{0}}$, suppress conditioning by writing $q(d\theta)$, and shift the trading clock so that the decision time becomes 0 and the remaining horizon is $T$.
-These references are used as analytic targets and sanity checks for our numerical sections (Sections 4 and 5), rather than as characterizations of the unrestricted optimum of (14) over the full feedback class.
+This subsection collects Gaussian benchmarks that yield closed-form decision-time reference allocations.
+Fix a decision time $t\in[0,T)$ and suppose an external procedure provides an $\mathcal{F}_{t}$-measurable input law $q_{t}$ for the uncertain premium parameter/state.
+We treat $q_{t}$ as a decision-time input (not a belief process updated within a run): it is held fixed over the remaining horizon $[t,T]$ and the resulting references are used only as validation targets and controlled diagnostics (not as characterizations of the unrestricted optimum).
+If the decision is revisited at a later time $t^{\prime}$, the reference can be recomputed by resolving the decision-time problem with the new input $q_{t^{\prime}}$ (plug-in replanning).
 We present two decision-time references.
-Section 2.3.1 considers static drift uncertainty: a latent premium is drawn from $q$ once at time 0 and kept fixed on $[0,T]$, providing the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4.
-Section 2.3.2 considers a mean-reverting (OU) premium with Gaussian initial uncertainty, which induces a horizon-dependent Gaussian law for the time-averaged premium and yields a tractable closed-form reference used in the hedging-demand recovery study of Section 5.
-For completeness, an online linear–Gaussian illustration that produces a time-varying uncertainty law $q_{t}$ via Kalman–Bucy filtering is deferred to Appendix A; it is included only to motivate a plug-in (receding-horizon) decision-time workflow in which $q_{t}$ is treated as an externally updated input at each decision time, rather than solving the fully optimal belief-state control problem (e.g., bensoussan1985optimal; pham2017dynamic).
+Section 2.3.1 considers static drift uncertainty and provides the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4.
+Section 2.3.2 considers a mean-reverting (OU) premium factor with Gaussian decision-time uncertainty for the current factor level, which induces a horizon-dependent effective Gaussian law for the time-averaged premium and yields a tractable reference used in the hedging-demand recovery study of Section 5.
+Appendix A gives a linear–Gaussian illustration of how an external filtering routine may produce a time-indexed input law $q_{t}$; it is included only as an input mechanism (plug-in replanning), not as a belief-state control formulation.
 2.3.1 Static Gaussian drift uncertainty
-We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter drawn at the decision time.
-The agent commits to a single $\theta$-blind policy, and all ex–ante uncertainty is summarized by the decision-time law $q$.
+We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter.
+This “unknown mean” specification is a standard device for isolating estimation risk/parameter uncertainty in portfolio choice (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Pástor, 2000).
+Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$.
+The agent chooses a $\theta$-blind policy on $[t,T]$, and all decision-time uncertainty about the (time-constant) premium is summarized by an $\mathcal{F}_{t}$-measurable law $q_{t}(d\theta)$, which we do not update over $[t,T]$.
 Market model (static latent drift).
-Let $d$ risky assets satisfy
-$\frac{dS_{t}}{S_{t}}=r\,\mathbf{1}\,dt+\theta\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (26)
-where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and the latent excess-return vector is drawn at time 0 as
-$\theta\sim q(d\theta).$ (27)
-A $\theta$-blind portfolio fraction process $\pi_{t}\in\mathbb{R}^{d}$ generates wealth
-$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\theta\big)\,dt+\pi_{t}^{\top}\Sigma^{1/2}\,dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (28)
-and we evaluate the ex–ante objective
-$J(\pi):=\mathbb{E}_{\theta\sim q}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta\big]\Big].$ (29)
-For reference, under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (merton1969lifetime; merton1971optimum), which is infeasible here because $\theta$ is latent.
-Analytic $q$-references via constant portfolios.
-To obtain a transparent closed-form benchmark that depends only on the decision-time law $q$, we temporarily restrict attention to constant portfolio fractions
-$\pi_{t}\equiv\pi\in\mathbb{R}^{d}.$ (30)
-This restriction is used solely to define an analytic $q$-reference; it is not imposed on the learning problem.
-Log utility ($\gamma=1$).
-Let $m_{\theta}:=\mathbb{E}_{\theta\sim q}[\theta]$. For constant $\pi$, the objective depends on $q$ only through $m_{\theta}$, and the optimal constant portfolio is
-$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\theta},$ (31)
-which is independent of $T$ in this static benchmark. In the one-asset case ($d=1$, $\Sigma=\sigma^{2}$),
-$\pi_{q,\log}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\sigma^{2}}.$ (32)
-CRRA ($\gamma\neq 1$): tilted optimality and Gaussian shrinkage.
-For $\gamma\neq 1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal, and
-$J(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)rT-\tfrac{1}{2}\gamma(1-\gamma)T\,\pi^{\top}\Sigma\pi\Big\}\,M_{q}\big((1-\gamma)T\pi\big),$ (33)
-where $M_{q}(u):=\mathbb{E}_{\theta\sim q}[\exp(u^{\top}\theta)]$ is the moment generating function of $q$.
-Any interior optimizer $\pi_{q,\gamma}^{\mathrm{const}}(T)$ satisfies the tilted first-order condition
-$\gamma\,\Sigma\,\pi_{q,\gamma}^{\mathrm{const}}(T)=\nabla_{u}\log M_{q}(u)\Big|_{u=(1-\gamma)T\,\pi_{q,\gamma}^{\mathrm{const}}(T)}.$ (34)
-If $q$ is Gaussian,
-$\theta\sim\mathcal{N}(m_{\theta},P),\hskip 18.49988ptP\succeq 0,$ (35)
-then $\nabla_{u}\log M_{q}(u)=m_{\theta}+Pu$ and the reference reduces to the linear system
-$\big(\gamma\Sigma-(1-\gamma)T\,P\big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=m_{\theta},$ (36)
+Let $d$ risky assets satisfy, for $s\in[t,T]$,
+$\frac{dS_{s}}{S_{s}}=r\,\mathbf{1}\,ds+\theta\,ds+\Sigma^{1/2}dW_{s},\hskip 18.49988ptS_{t}\in(0,\infty)^{d},$ (21)
+where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\theta\in\mathbb{R}^{d}$ is a latent constant excess-return vector.
+At decision time $t$, uncertainty about $\theta$ is represented by the input law
+$\theta\sim q_{t}(d\theta).$ (22)
+A $\theta$-blind portfolio fraction process $\pi_{s}\in\mathbb{R}^{d}$ generates wealth on $[t,T]$:
+$\frac{dX_{s}^{\pi}}{X_{s}^{\pi}}=\big(r+\pi_{s}^{\top}\theta\big)\,ds+\pi_{s}^{\top}\Sigma^{1/2}\,dW_{s},\hskip 18.49988ptX_{t}^{\pi}=x>0.$ (23)
+We evaluate the (conditional) decision-time objective by mixing over $q_{t}$:
+$J_{t}(\pi):=\mathbb{E}_{\theta\sim q_{t}}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\,\mathcal{F}_{t}\big]\Big].$ (24)
+Under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (Merton, 1969, 1971), which is infeasible here because $\theta$ is latent.
+Analytic $q_{t}$-references via portfolios constant over $[t,T]$.
+To obtain transparent closed-form targets, we temporarily restrict attention to portfolio fractions that are constant in trading time $s$ on $[t,T]$:
+$\pi_{s}\equiv\pi\in\mathbb{R}^{d},\hskip 18.49988pts\in[t,T].$ (25)
+The superscript $\mathrm{const}$” used below refers to this constancy in $s$.
+The resulting reference may still depend on the decision time $t$ through the input $q_{t}$ and on the remaining horizon $\tau=T-t$.
+(Decision-time constant-portfolio targets of this kind are closely connected to myopic/plug-in benchmarks used in the parameter-uncertainty literature; see, e.g., Barberis (2000); Pástor (2000).)
+CRRA ($\gamma>1$): tilted optimality and Gaussian shrinkage.
+For $\gamma>1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal and
+$J_{t}(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)r\tau-\tfrac{1}{2}\gamma(1-\gamma)\tau\,\pi^{\top}\Sigma\pi\Big\}\,M_{q_{t}}\big((1-\gamma)\tau\pi\big),$ (26)
+where $M_{q_{t}}(u):=\mathbb{E}_{\theta\sim q_{t}}[\exp(u^{\top}\theta)]$ is the moment generating function of $q_{t}$.
+Any interior optimizer $\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)$ satisfies the tilted first-order condition
+$\gamma\,\Sigma\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\nabla_{u}\log M_{q_{t}}(u)\Big|_{u=(1-\gamma)\tau\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)}.$ (27)
+If the decision-time input is Gaussian,
+$q_{t}=\mathcal{N}(m_{t},P_{t}),\hskip 18.49988ptP_{t}\succeq 0,$ (28)
+then $\nabla_{u}\log M_{q_{t}}(u)=m_{t}+P_{t}u$ and the reference reduces to the linear system
+$\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=m_{t},$ (29)
 hence
-$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma-(1-\gamma)T\,P\big)^{-1}m_{\theta}.$ (37)
+$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)^{-1}m_{t}.$ (30)
 For $\gamma>1$, this takes the familiar shrinkage form
-$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma+(\gamma-1)T\,P\big)^{-1}m_{\theta},\hskip 18.49988pt(\gamma>1),$ (38)
-and in one dimension ($d=1$, $\Sigma=\sigma^{2}$, $P=p$),
-$\pi_{q,\gamma}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\gamma\sigma^{2}+(\gamma-1)T\,p},\hskip 18.49988pt(\gamma>1).$ (39)
+$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt(\gamma>1),$ (31)
+which makes explicit how decision-time uncertainty $P_{t}$ dampens exposures as the remaining horizon grows (a qualitative effect widely emphasized in return-predictability/parameter-uncertainty studies; see, e.g., Barberis (2000); Pástor (2000); Kandel and Stambaugh (1996)).
 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
-We next replace the static premium by a mean-reverting Gaussian premium process, a standard reduced-form device for return predictability and intertemporal hedging (campbell2002strategic; xia2001learning).
-Our goal here is not to introduce additional information structure, but to obtain a closed-form, decision-time Gaussian reference for the time-averaged premium over the remaining horizon.
-This induces a horizon-dependent effective premium law that can be used as a controlled analytic input in numerical experiments.
-OU premium dynamics and decision-time uncertainty.
-Let the uncertain initial state be $\vartheta$ and set $Y_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0})$, so the decision-time law is $q=\mathcal{N}(m_{0},P_{0})$. The premium factor follows
-$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0}),$ (40)
-and risky excess returns satisfy
-$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},$ (41)
-allowing instantaneous correlation
-$d\langle W,W^{Y}\rangle_{t}=\rho\,dt,\hskip 18.49988pt\rho\in\mathbb{R}^{d\times m}.$ (42)
-Integrated premium and induced Gaussian law.
-Define the integrated premium
-$I_{T}:=\int_{0}^{T}Y_{s}\,ds\in\mathbb{R}^{m}.$ (43)
-Since (40) is linear-Gaussian with Gaussian initial condition, $I_{T}$ is Gaussian. Its mean and covariance are
-$m_{I}(T)=\mathbb{E}[I_{T}]=T\bar{y}+K^{-1}\big(I-e^{-KT}\big)\,(m_{0}-\bar{y}),$ (44)
-$C_{I}(T)=\mathrm{Cov}(I_{T})$ $=K^{-1}\big(I-e^{-KT}\big)\,P_{0}\,\big(I-e^{-KT}\big)^{\top}K^{-\top}$ (45)
-$\qquad+\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\Xi^{\top}\,\big(I-e^{-K(T-s)}\big)^{\top}K^{-\top}\,ds.$
-Notably, (44)–(45) depend only on the OU dynamics and the decision-time uncertainty $(m_{0},P_{0})$; they do not depend on the return–factor correlation $\rho$.
-Horizon-averaged premium and effective Gaussian law.
-Define the horizon-averaged effective premium
-$\bar{\theta}_{T}:=\frac{1}{T}\,B\,I_{T}\in\mathbb{R}^{d}.$ (46)
-Then
-$\bar{\theta}_{T}\sim\mathcal{N}(m_{\bar{\theta}}(T),P_{\bar{\theta}}(T)),\hskip 18.49988ptm_{\bar{\theta}}(T)=\frac{1}{T}B\,m_{I}(T),\hskip 18.49988ptP_{\bar{\theta}}(T)=\frac{1}{T^{2}}B\,C_{I}(T)\,B^{\top}.$ (47)
-When $\rho=0$, this induced law can be plugged directly into the static Gaussian reference of Section 2.3.1.
-When $\rho\neq 0$, the marginal law (47) remains valid, but constant-portfolio expected utility involves an additional cross-covariance term capturing the return–state shock linkage that generates hedging demand (campbell2002strategic; xia2001learning).
-Closed-form references under constant portfolios.
-Restricting to constant fractions $\pi_{t}\equiv\pi$ turns the problem into a transparent decision-time benchmark: only the integrated premium $I_{T}=\int_{0}^{T}Y_{s}\,ds$ enters the drift of $\log X_{T}^{\pi}$, while the risk term remains time-homogeneous.
-This yields a closed-form target that depends on the decision-time law $q=\mathcal{N}(m_{0},P_{0})$ only through the induced mean $m_{I}(T)=\mathbb{E}[I_{T}]$ (and, for CRRA, through covariances as well).
-Log utility ($\gamma=1$).
-With $\pi_{t}\equiv\pi$, the log-utility criterion reduces to a strictly concave quadratic in $\pi$ whose drift term depends on the OU factor only through the mean integrated premium $m_{I}(T)=\mathbb{E}\!\left[\int_{0}^{T}Y_{s}\,ds\right]$.
-Hence the decision-time reference depends on $q=\mathcal{N}(m_{0},P_{0})$ only through $m_{\bar{\theta}}(T)=(1/T)B\,m_{I}(T)$ (and, in particular, does not involve return–factor correlation), giving
-$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\bar{\theta}}(T)=\frac{1}{T}\,\Sigma^{-1}B\,m_{I}(T).$ (48)
+We next replace the static premium by a mean-reverting Gaussian premium factor, a standard reduced-form device for return predictability and intertemporal hedging (Campbell and Viceira, 2002; Xia, 2001).
+Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$.
+At time $t$ we take as input a Gaussian law for the current factor level $Y_{t}$ (rather than a point estimate), capturing decision-time statistical uncertainty (e.g., interval-type uncertainty quantification, resampling variability, or asymptotic approximations).
+Holding this input fixed over $[t,T]$, OU propagation and time-averaging induce a horizon-dependent effective premium law; when $\rho\neq 0$, return–factor shock correlation adds a further quadratic term relevant for hedging demand (Campbell and Viceira, 2002; Xia, 2001).
+OU premium dynamics.
+Let $Y_{s}\in\mathbb{R}^{m}$ be a premium factor following
+$dY_{s}=K(\bar{y}-Y_{s})\,ds+\Xi\,dW_{s}^{Y},\hskip 18.49988ptK\in\mathbb{R}^{m\times m}\ \text{Hurwitz and invertible},$ (32)
+where $\bar{y}\in\mathbb{R}^{m}$ and $\Xi\in\mathbb{R}^{m\times m}$. Risky excess returns satisfy
+$dR_{s}:=\frac{dS_{s}}{S_{s}}-r\mathbf{1}\,ds=BY_{s}\,ds+\Sigma^{1/2}\,dW_{s},$ (33)
+with instantaneous correlation
+$d\langle W,W^{Y}\rangle_{s}=\rho\,ds,\hskip 18.49988pt\rho\in\mathbb{R}^{d\times m}.$ (34)
+Decision-time input law for the factor.
+At decision time $t$, we take as input an $\mathcal{F}_{t}$-measurable Gaussian law for the current factor level,
+$Y_{t}\sim q_{t}(dy)=\mathcal{N}(m_{t},P_{t}),\hskip 18.49988ptP_{t}\succeq 0.$ (35)
+A common special case is purely model-implied predictive propagation from an initial uncertainty $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$ without conditioning on observations:
+$m_{t}=\bar{y}+e^{-Kt}(m_{0}-\bar{y}),$
+$P_{t}=e^{-Kt}P_{0}e^{-K^{\top}t}+\int_{0}^{t}e^{-K(t-s)}\,\Xi\Xi^{\top}\,e^{-K^{\top}(t-s)}\,ds,$ (36)
+which is an OU transition (prior-predictive) law and should not be interpreted as an information update.
+(References for return predictability and OU-style investment-opportunity dynamics include Campbell and Viceira (2002); Xia (2001).)
+Induced Gaussian law for the time-averaged premium over $[t,T]$.
+Define the remaining-horizon integrated factor and the associated horizon-averaged effective premium:
+$I_{t,T}:=\int_{t}^{T}Y_{s}\,ds\in\mathbb{R}^{m},\hskip 18.49988pt\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\,B\,I_{t,T}\in\mathbb{R}^{d}.$ (37)
+Let
+$A(\tau):=K^{-1}\big(I-e^{-K\tau}\big)\in\mathbb{R}^{m\times m}.$
+Under the linear–Gaussian OU model, $I_{t,T}$ is Gaussian under the input law (35):
+$I_{t,T}\sim\mathcal{N}\big(m_{I}(t,\tau),C_{I}(t,\tau)\big),$
+with
+$m_{I}(t,\tau)=\mathbb{E}[I_{t,T}]=\tau\bar{y}+A(\tau)\,(m_{t}-\bar{y}),$ (38)
+$C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})=A(\tau)\,P_{t}\,A(\tau)^{\top}+\int_{0}^{\tau}A(u)\,\Xi\Xi^{\top}\,A(u)^{\top}\,du.$ (39)
+Consequently, the horizon-averaged effective premium is Gaussian,
+$\bar{\theta}_{t,\tau}\sim\mathcal{N}\!\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big),\hskip 18.49988ptm_{\bar{\theta}}(t,\tau)=\frac{1}{\tau}B\,m_{I}(t,\tau),\hskip 18.49988ptP_{\bar{\theta}}(t,\tau)=\frac{1}{\tau^{2}}B\,C_{I}(t,\tau)\,B^{\top}.$ (40)
+The pair $\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big)$ is the horizon-dependent effective Gaussian law that we use as a controlled analytic input.
+Closed-form decision-time references under portfolios constant over $[t,T]$.
+As in Section 2.3.1, we define an analytic reference by restricting to portfolio fractions that are constant in trading time $s$ over $[t,T]$, i.e.
+$\pi_{s}\equiv\pi\in\mathbb{R}^{d}$.
+The superscript $\mathrm{const}$” below refers to this constancy in $s$; the resulting reference may still depend on the decision time $t$ through $(m_{t},P_{t})$ and on the remaining horizon $\tau$ through OU averaging.
 CRRA ($\gamma>1$).
-Define
-$C_{IW}(T):=\mathrm{Cov}(I_{T},W_{T})=\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\,\rho^{\top}\,ds\;\in\;\mathbb{R}^{m\times d},$ (49)
+When $\rho\neq 0$, return shocks and factor shocks are contemporaneously correlated, producing an additional quadratic term in the constant-portfolio objective (hedging-demand channel; see, e.g., Campbell and Viceira (2002); Xia (2001)).
+Define the cross-covariance between $I_{t,T}$ and the return Brownian increment $W_{T}-W_{t}$:
+$C_{IW}(\tau):=\mathrm{Cov}\!\big(I_{t,T},\,W_{T}-W_{t}\big)=\int_{0}^{\tau}A(u)\,\Xi\,\rho^{\top}\,du\;\in\;\mathbb{R}^{m\times d},$ (41)
 and the induced symmetric cross term
-$M_{\mathrm{cross}}(T):=B\,C_{IW}(T)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(T)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (50)
-Then the Gaussian-$q$ decision-time reference under constant portfolios is characterized by
-$\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=B\,m_{I}(T),\hskip 18.49988pt(\gamma>1),$ (51)
-equivalently
-$\pi_{q,\gamma}^{\mathrm{const}}(T)=\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)^{-1}B\,m_{I}(T).$
-When $\rho=0$, we have $C_{IW}(T)=0$ and $M_{\mathrm{cross}}(T)=0$, recovering the independence-case shrinkage reference.
-2.4 Why dynamic programming and deep PDE surrogates break down in high-dimensional uncertain markets
-This subsection explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted here.
-DP is conceptually sound in low-dimensional Markovian settings (fleming2006controlled; pham2009continuous), but two issues dominate in our regime: (i) numerically learning the value-function derivatives required for optimal policies becomes prohibitive as dimension and nonlinearity grow, and (ii) principled parameter uncertainty magnifies these difficulties.
-Classical HJB: curse of dimensionality and full nonlinearity.
-With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (fleming2006controlled).
-Grid-based solvers scale exponentially in the state dimension (bellman1961adaptive; kushner2001numerical).
-In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points.
-Moreover, constraints, transaction costs, and non-affine dynamics typically yield fully nonlinear HJBs, where stable monotone schemes are delicate even in moderate dimension and become impractical in the regime we target (kushner2001numerical).
-Deep PDE surrogates: fewer grids, same derivative bottleneck.
-PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (raissi2019physics; sirignano2018dgm; han2018solving; beck2019machine), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities (e.g.
-$V_{xy}$) that drive intertemporal hedging.
-In practice this induces nonconvex, ill-conditioned objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable hedging demands in high dimension.
-Latent parameter uncertainty: belief-state blowup and $\theta$-blind aggregation.
-A principled DP treatment augments the state with a posterior/belief over parameters, leading to a value function $V(t,x,y,\Pi)$ on a space of measures in general (bensoussan1985optimal; pham2017dynamic).
-Even when finite-dimensional conjugate summaries exist, the enlarged HJB is substantially harder.
-For deep surrogates, uncertainty either requires solving many $\theta$-conditional problems (expensive) or treating $\theta$ as an extra input (higher effective dimension, worse conditioning).
-In our deployable $\theta$-blind setting, a single policy must perform well under $\theta\sim q$, coupling heterogeneous models and potentially causing high-variance gradients and cancellation across parameter draws.
-We therefore avoid value-function PDE/BSDE baselines in this regime and instead work with a $q$-aggregated Pontryagin stationarity condition and projection map, estimating expectations over $\theta\sim q$ via Monte Carlo inside the simulator.
-While $\theta$-conditional PMP objects can still be computed under frozen-$\theta$ simulations for inspection, our deployable target and guarantees are expressed in terms of $q$-aggregated stationarity, motivating the simulation-based methods in Section 3.
+$M_{\mathrm{cross}}(\tau):=B\,C_{IW}(\tau)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(\tau)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (42)
+Then the Gaussian decision-time reference under constant portfolios is characterized by
+$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1).$ (43)
+Equivalently,
+$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)^{-1}B\,m_{I}(t,\tau).$
+When $\rho=0$, we have $C_{IW}(\tau)=0$ and $M_{\mathrm{cross}}(\tau)=0$, and the reference reduces to the independence-case shrinkage form obtained by plugging the effective Gaussian law (40) into the static Gaussian formula of Section 2.3.1:
+$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{\bar{\theta}}(t,\tau)\big)^{-1}m_{\bar{\theta}}(t,\tau),\hskip 18.49988pt(\rho=0,\ \gamma>1).$
 3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
-We study the fixed-$q$ ex–ante portfolio choice problem of Section 2 under latent parameter uncertainty $\theta\sim q$.
-The investor must deploy a $\theta$-blind policy (Remark 1), so the control can depend on observable states $(t,X_{t},Y_{t})$ but cannot take $\theta$ as an input.
-We restrict attention to Markov feedback policies parameterized by a neural network $\pi_{\varphi}(t,x,y)$.
-Our solution approach follows a two-stage pipeline:
-  • Stage 1 (PG–DPO).
-We perform stochastic gradient ascent on the ex–ante objective $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$, sampling $\theta$ only inside the simulator while keeping $\pi_{\varphi}$ deployable and $\theta$-blind.
-  • Stage 2 (Pontryagin projection).
-Under a warm-up policy, we estimate Pontryagin sensitivity objects by BPTT (conditionally on frozen $\theta$), aggregate them across $\theta\sim q$, and construct a single deployable portfolio by projecting onto the aggregated first-order condition (20).
-A practical subtlety is that the $q$-aggregated Pontryagin condition involves mixed moments across $\theta$ (products of $\theta$-dependent costates and $\theta$-dependent coefficients).
-In moderate to high dimensions, these quantities can be statistically noisy under finite Monte Carlo budgets.
-In our implementation, the main stabilization mechanisms are (i) estimating stage 2 objects under a warm-up policy (two-time-scale stabilization), (ii) computing the same projection in a residual/control-variate form (Section 3.3.1), and (iii) amortizing projection via interactive distillation (Section 3.3.2).
-Section 3.1 reviews baseline PG–DPO and the conditional BPTT–PMP correspondence.
-Section 3.2 develops the stage 2 $q$-aggregated projection under latent $\theta$, together with a residual-based policy-gap guarantee.
-Section 3.3 records two practical couplings between stage 1 and stage 2 (residual form and interactive distillation).
-3.1 PG–DPO as stochastic gradient ascent and BPTT–PMP correspondence
+We solve the fixed-$q$ ex–ante portfolio problem of Section 2 under a latent, time-constant parameter $\theta\sim q$.
+The deployed controller is $\theta$-blind: it depends only on observable states $(t,X_{t},Y_{t})$ and never conditions on the realized $\theta$.
+Uncertainty enters only through sampling $\theta$ inside the simulator.
+Our solver has two components.
+Stage 1 (PG–DPO) performs stochastic gradient ascent on $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$ by differentiating through an Euler simulator via BPTT.
+Stage 2 (P–PGDPO) aggregates the BPTT costate blocks across $\theta\sim q$ and enforces the $q$-aggregated Pontryagin first-order condition, yielding a deployable projected control.
+Section 3.1 establishes the conditional BPTT–PMP correspondence and specifies the adjoint blocks needed for projection.
+Section 3.2 develops the $q$-aggregated projection and a residual-based policy-gap bound.
+Practical coupling mechanisms (residual-form projection and interactive distillation) are deferred to Appendix F.
+3.1 PG–DPO as stochastic gradient ascent and a conditional BPTT–PMP correspondence
 Setup and objectives (frozen $\theta$, deployable $\theta$-blind feedback).
@@
 $\pi_{\varphi}:\ [0,T]\times(0,\infty)\times\mathbb{R}^{m}\to\mathbb{R}^{d},\hskip 18.49988pt(t,x,y)\mapsto\pi_{\varphi}(t,x,y),\hskip 18.49988pt\varphi\in\mathbb{R}^{p},$ (44)
-which does not take $\theta$ as an input. For a fixed frozen $\theta$, the $\theta$-conditional state is $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ and evolves as
+which does not take $\theta$ as an input. Conditional on a fixed frozen $\theta$, the controlled state $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ evolves as
 $\frac{dX_{t}^{\pi,\theta}}{X_{t}^{\pi,\theta}}$ $=\Big(r+\pi_{t}^{\top}b\big(Y_{t}^{\theta},\theta\big)\Big)\,dt+\pi_{t}^{\top}\sigma\big(Y_{t}^{\theta},\theta\big)\,dW_{t},\hskip 18.49988ptX_{0}=x>0,$ (45)
-$dY_{t}^{\theta}$ $=a\big(Y_{t}^{\theta},\theta\big)\,dt+\beta\big(Y_{t}^{\theta},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m}.$ (54)
-For each fixed $\theta$ we evaluate $\pi_{\varphi}$ by the conditional objective
+$dY_{t}^{\theta}$ $=a\big(Y_{t}^{\theta},\theta\big)\,dt+\beta\big(Y_{t}^{\theta},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m},$ (46)
+where $(W,W^{Y})$ may be instantaneously correlated as in Section 2. For each fixed $\theta$ we evaluate $\pi_{\varphi}$ by the conditional objective
 $J^{\theta}(\varphi):=\mathbb{E}\big[U\big(X_{T}^{\pi_{\varphi},\theta}\big)\,\big|\,\theta\big],$ (47)
@@
 where the last expectation is joint over $\theta\sim q$ and $(W,W^{Y})$.
-Thus $\sup_{\varphi}J(\varphi)$ is a standard stochastic optimization problem: $\theta$ is sampled inside the simulator while the policy remains $\theta$-blind.
+Thus $\sup_{\varphi}J(\varphi)$ is a stochastic optimization problem in which $\theta$ is sampled inside the simulator while the policy remains $\theta$-blind.
 Discretization, sampling over $\theta$, and baseline PG–DPO update.
-We discretize $[0,T]$ into $N$ steps of length $\Delta t$ and approximate (53)–(54) by an Euler scheme.
-For episode $i$ we denote the discrete state by $(X_{k}^{(i)},Y_{k}^{(i)})_{k=0,\dots,N}$ and write $\theta^{(i)}$ for the frozen parameter used to generate that simulated environment.
+We discretize the horizon using an Euler scheme.
+In training we may sample initial states $z_{0}^{(i)}=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$ from a user-chosen distribution $\nu$ on $0,T)\times(0,\infty)\times\mathbb{R}^{m}$.
+Given $t_{0}^{(i)}$, we discretize $t_{0}^{(i)},T]$ into $N$ steps with step size
+$\Delta t^{(i)}:=\frac{T-t_{0}^{(i)}}{N},\hskip 18.49988ptt_{k}^{(i)}:=t_{0}^{(i)}+k\Delta t^{(i)},\qquad k=0,\dots,N.$
+For episode $i$, we denote the discrete state by $(X_{k}^{(i)},Y_{k}^{(i)})_{k=0,\dots,N}$ and write $\theta^{(i)}$ for the frozen parameter used to generate that simulated environment.
 Given $\pi_{\varphi}$ and Brownian increments, the mapping
 $\big(x^{(i)},y^{(i)},\theta^{(i)},\{\Delta W_{k}^{(i)},\Delta W_{k}^{Y,(i)}\}_{k=0}^{N-1},\varphi\big)\longmapsto U\big(X_{N}^{(i)}\big)$
-is a finite computational graph, so automatic differentiation computes exact discrete gradients $\nabla_{\varphi}U(X_{N}^{(i)})$.
-A typical PG–DPO update samples a mini-batch of initial states $\{(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})\}_{i=1}^{M}$ from a user-chosen training distribution $\nu$ on $0,T)\times(0,\infty)\times\mathbb{R}^{m}$, samples $\theta\sim q$ inside the simulator (unseen by the policy) and holds it frozen for the update, and simulates forward:
-$Y_{k+1}^{(i)}$ $=Y_{k}^{(i)}+a\big(Y_{k}^{(i)},\theta\big)\Delta t^{(i)}+\beta\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{Y,(i)},$
-$X_{k+1}^{(i)}$ $=X_{k}^{(i)}+X_{k}^{(i)}\Big(r+\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}b\big(Y_{k}^{(i)},\theta\big)\Big)\Delta t^{(i)}$
-$\hphantom{={}}{}+X_{k}^{(i)}\,\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}\sigma\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{(i)},$
-starting from $X_{0}^{(i)}=x_{0}^{(i)}$, $Y_{0}^{(i)}=y_{0}^{(i)}$. The episode reward is
+is a finite computational graph; automatic differentiation thus computes exact discrete-time gradients $\nabla_{\varphi}U(X_{N}^{(i)})$ for the chosen discretization.
+A typical PG–DPO update samples a mini-batch $\{z_{0}^{(i)}\}_{i=1}^{M}\sim\nu$, samples latent parameters either independently per episode ($\theta^{(i)}\sim q$) or shared within the batch (one $\theta\sim q$ reused across $i$), simulates Euler rollouts, and performs stochastic gradient ascent.
+Suppressing episode indices and writing $\theta$ for the frozen draw used in the rollout, the Euler recursion is
+$Y_{k+1}$ $=Y_{k}+a\big(Y_{k},\theta\big)\Delta t+\beta\big(Y_{k},\theta\big)\Delta W_{k}^{Y},$
+$X_{k+1}$ $=X_{k}+X_{k}\Big(r+\pi_{\varphi}\big(t_{k},X_{k},Y_{k}\big)^{\top}b\big(Y_{k},\theta\big)\Big)\Delta t+X_{k}\,\pi_{\varphi}\big(t_{k},X_{k},Y_{k}\big)^{\top}\sigma\big(Y_{k},\theta\big)\Delta W_{k},$
+starting from $(t_{0},X_{0},Y_{0})=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$. The episode reward is
 $J^{(i)}(\varphi):=U\big(X_{N}^{(i)}\big),$ (49)
@@
 $\varphi\leftarrow\varphi+\alpha\,\frac{1}{M}\sum_{i=1}^{M}\nabla_{\varphi}J^{(i)}(\varphi).$ (50)
-Sampling $\theta$ independently per episode (i.e.
-$\theta^{(i)}\sim q$) or sampling one $\theta\sim q$ per update and reusing it across the batch both yield unbiased stochastic gradients for $J(\varphi)$.
-Pathwise costates from BPTT and the (conditional) BPTT–PMP correspondence.
-BPTT returns not only $\nabla_{\varphi}J^{(i)}(\varphi)$ but also pathwise sensitivities with respect to intermediate state variables, which coincide with discrete-time adjoint variables (costates) in the sense of Pontryagin.
-For a single episode (suppressing $i$ and $\theta$ in notation), define the pathwise wealth costate
+Under standard interchange conditions (dominated convergence / uniform integrability for the discretized simulator), this is an unbiased stochastic gradient for the discretized ex–ante objective; as $\Delta t\to 0$ it approximates the continuous-time objective.
+BPTT adjoints and their interpretation: first adjoint and second-adjoint blocks.
+BPTT returns not only $\nabla_{\varphi}J^{(i)}(\varphi)$ but also adjoint variables with respect to intermediate state variables.
+For a single episode (suppressing $i$ and $\theta$ in notation), define the pathwise wealth adjoint
 $p_{k}:=\frac{\partial U(X_{N})}{\partial X_{k}},\hskip 18.49988ptk=0,\dots,N,$ (51)
-and the additional pathwise sensitivity objects used in projected controls:
-$p_{x,k}:=\frac{\partial p_{k}}{\partial X_{k}},\hskip 18.49988ptp_{y,k}:=\frac{\partial p_{k}}{\partial Y_{k}},\hskip 18.49988ptk=0,\dots,N.$ (60)
-Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, uniform on compacts)).
-Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (53)–(54) under the $\theta$-blind policy $\pi_{\varphi}$ and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle (adjoint) system.
-Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the resulting continuous-time Pontryagin objects under $\pi_{\varphi}$ (and, in smooth Markov regimes, the corresponding spatial derivatives of the decoupling field).
+which is the discrete-time analogue of the wealth-component first adjoint.
+For the Pontryagin projection in Section 3.2, we also require curvature-type objects.
+In smooth Markov regimes these correspond to $(V_{xx},V_{xy})$; in the stochastic maximum principle they are naturally interpreted as blocks of the second adjoint.
+Accordingly, we track the following discrete-time blocks (computed by higher-order automatic differentiation):
+$p_{x,k}:=\frac{\partial p_{k}}{\partial X_{k}}\;=\;\frac{\partial^{2}U(X_{N})}{\partial X_{k}^{2}},\hskip 18.49988ptp_{y,k}:=\frac{\partial p_{k}}{\partial Y_{k}}\;=\;\frac{\partial^{2}U(X_{N})}{\partial Y_{k}\,\partial X_{k}},\hskip 18.49988ptk=0,\dots,N.$ (52)
+(Here $p_{y,k}\in\mathbb{R}^{m}$ is the wealth–factor cross block.) These are the discrete-time objects that enter the affine-in-control Hamiltonian gradient and hence the projection step.
+Pathwise vs.
+SMP-adapted adjoints.
+The quantities in (51)–(52) are pathwise (sample-wise) derivatives on the simulated computation graph and, for $k<N$, they need not be adapted to time-$t_{k}$ information (they can depend on future increments).
+To connect them to the stochastic maximum principle (which yields adapted adjoint processes and associated martingale terms), one performs the standard conditional $L^{2}$ projection onto the span of the driving increments on each step, yielding a discrete-time BSDE representation.
+Concretely, for a fixed frozen $\theta$ and the simulator filtration
+$G_{t_{k}}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t_{k}\}\big)\qquad\text{(with the usual augmentation)},$
+we define the adapted (projected) discrete adjoints
+$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid G_{t_{k}}^{\theta}].$
+Theorem 2 concerns the convergence of these adapted discrete objects (together with the associated one-step BSDE martingale coefficients) to the continuous-time Pontryagin adjoints.
+Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, including second-adjoint blocks; uniform on compacts)).
+Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (45)–(46) under the $\theta$-blind policy $\pi_{\varphi}$, and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle, including the wealth-component first adjoint and the relevant blocks of the second adjoint needed for the portfolio Hamiltonian gradient.
+Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote these continuous-time Pontryagin objects under $\pi_{\varphi}$.
+In smooth Markov regimes, they coincide (along the controlled state process) with $(V_{x},V_{xx},V_{xy})$ for the $\theta$-conditional value function.
 Let $(p_{k},p_{x,k},p_{y,k})$ be the discrete pathwise quantities computed by BPTT for the Euler discretization with step $\Delta t$, as defined in (51)–(52).
-Then, as $\Delta t\to 0$, the BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories).
+Let $(p_{k}^{\Delta t,\theta},p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$ be their adapted projections onto $G_{t_{k}}^{\theta}$ as defined above, and let the associated discrete-time BSDE martingale coefficients be defined via the one-step conditional $L^{2}$ projection on each step.
+Then, as $\Delta t\to 0$, the projected (adapted) BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories).
 Moreover, for any compact set $K\subset\Theta$, the constants in the convergence bounds can be chosen uniformly for all $\theta\in K$.
 Proof.
-See Appendix B. ∎
+See Appendix C. ∎
 Across $\theta\sim q$, these Pontryagin objects form a $\theta$-indexed family.
 Baseline PG–DPO trains against the ex–ante objective (48) by repeatedly sampling $\theta\sim q$ inside the simulator, while the deployable policy remains $\theta$-blind.
-3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection and a residual-based policy-gap bound
-Stage 2 is a projection step: given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin sensitivity objects by BPTT/Monte Carlo under frozen $\theta\sim q$, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated Pontryagin stationarity condition derived in Section 2.2.
-The main point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from estimated Pontryagin objects to a portfolio rule.
+3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection, diagnostics, and a residual-based policy-gap bound
+Stage 2 is a projection step.
+Given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin adjoint objects under frozen $\theta\sim q$ by BPTT/Monte Carlo, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated stationarity condition derived in Section 2.2.
+Equivalently, Stage 2 can be viewed as a plug-in approximation of one application of the population projection map $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ (defined below) evaluated at $\pi=\pi^{\mathrm{warm}}$ on the working domain.
+The key point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from (estimated) adjoint objects to a portfolio rule.
+Throughout, the adjoint objects used in projection consist of the wealth-component first adjoint together with the relevant blocks of the second adjoint (cf.
+Section 3.1): in smooth Markov regimes these correspond to $(V_{x},V_{xx},V_{xy})$ evaluated along the controlled state.
 Working domain and norms.
@@
 Fix a query state $z=(t,x,y)\in D$ and a frozen parameter $\theta$.
-We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute pathwise Pontryagin sensitivity objects by autodiff/BPTT; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
-$\widehat{p}_{t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{x,t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{y,t}^{\theta}(z).$ (62)
+We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute discrete-time adjoint blocks by autodiff/BPTT on the Euler simulator; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
+$\widehat{p}_{t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{x,t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{y,t}^{\theta}(z),$ (54)
+where $\widehat{p}_{t}^{\theta}$ corresponds to the wealth-component first adjoint and $(\widehat{p}_{x,t}^{\theta},\widehat{p}_{y,t}^{\theta})$ correspond to the blocks of the second adjoint relevant for the Hamiltonian gradient.
+In the analysis, these are interpreted as estimators of the corresponding $\theta$-conditional Pontryagin objects (cf.
+the conditional BPTT–PMP correspondence in Section 3.1).
 Using these, define the $\theta$-conditional estimated projection inputs
 $\widehat{A}_{t}^{\theta}(t,x,y)$ $:=x\,\widehat{p}_{x,t}^{\theta}(t,x,y)\,\Sigma(y,\theta)\in\mathbb{R}^{d\times d},$ (55)
-$\widehat{G}_{t}^{\theta}(t,x,y)$ $:=\widehat{p}_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,\widehat{p}_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d}.$ (64)
-Aggregating across $\theta\sim q$ (approximated in practice by sampling $M_{\theta}$ frozen parameters) gives
+$\widehat{G}_{t}^{\theta}(t,x,y)$ $:=\widehat{p}_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,\widehat{p}_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d},$ (56)
+where $\Sigma(\cdot,\theta)$ and $\Sigma_{SY}(\cdot,\theta)$ are the instantaneous covariance objects appearing in the $\theta$-conditional Hamiltonian (as in Section 2).
+Aggregating across $\theta\sim q$ (approximated in practice by sampling $M_{\theta}$ frozen parameters) gives the mixed-moment estimators
 $\widehat{A}_{t}(t,x,y)$ $:=\mathbb{E}_{\theta\sim q}\Big[\widehat{A}_{t}^{\theta}(t,x,y)\Big],$ (57)
@@
 $\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y):=-\,\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{G}_{t}^{\mathrm{mix}}(t,x,y).$ (59)
+Stabilized inversion and projection diagnostics (implementation).
+In practice, finite-sample estimates can make $\widehat{A}_{t}$ ill-conditioned or even singular on parts of $D$.
+Our default implementation therefore uses standard stabilizers (Appendix E.5), including: (i) ridge regularization $\widehat{A}_{t,\lambda}:=\widehat{A}_{t}+\lambda I_{d}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\lambda}:=-\widehat{A}_{t,\lambda}^{-1}\widehat{G}_{t}^{\mathrm{mix}}$, (ii) condition-number or eigenvalue diagnostics to skip/clip projection on unreliable states (falling back to $\pi^{\mathrm{warm}}$), and (iii) residual-form projection (Section F.1) to reduce variance.
+The analysis below is stated for the population projection $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ and for small estimator perturbations, which is precisely the regime enforced by these stabilizers.
+Remark (decoupled aggregation heuristic).
+The mixed-moment aggregation (57)–(58) is the object appearing in the exact $q$-aggregated first-order condition.
+In some high-dimensional experiments, we additionally consider a variance-reduction heuristic that replaces mixed moments by products of empirical means, e.g.
+$\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}\,b(y,\theta)]\approx\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}]\cdot\mathbb{E}_{\theta}[b(y,\theta)]$, and similarly for the other terms.
+All theoretical statements in this subsection are for the mixed-moment projection (59).
 Residual diagnostic and a slab-wise small-gain policy-gap bound.
-To connect the projected policy (67) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity.
+To connect (59) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity.
 Let $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ denote the mixed-moment $q$-aggregated projection inputs induced by a policy $\pi$ (i.e.
-the objects in (23) evaluated using the $\theta$-conditional Pontryagin objects generated by $\pi$).
+the objects in (18) evaluated using the $\theta$-conditional Pontryagin adjoint blocks generated by $\pi$).
 Define the warm-up aggregated stationarity residual on $D$ by
 $r^{\mathrm{warm}}_{\mathrm{FOC,mix}}(t,x,y):=A_{\pi^{\mathrm{warm}}}(t,x,y)\,\pi^{\mathrm{warm}}(t,x,y)+G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}(t,x,y),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\big\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\big\|_{L^{2}(\mu)}.$ (60)
-In practice we monitor the estimator $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$ computed from the same BPTT/Monte Carlo pipeline.
+Note that $r^{\mathrm{warm}}_{\mathrm{FOC,mix}}=0$ is equivalent to $\pi^{\mathrm{warm}}$ being a fixed point of $T$ on $D$. More generally, the population projection admits the residual form
+$T(\pi^{\mathrm{warm}})\;=\;\pi^{\mathrm{warm}}\;-\;A_{\pi^{\mathrm{warm}}}^{-1}\,r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$ (61)
+In practice we monitor the empirical analogue $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$, and use it as a diagnostic to flag states where projection is unreliable (cf.
+Appendix E.5).
+(Equivalently, $\widehat{\pi}^{\mathrm{agg,mix}}=\pi^{\mathrm{warm}}-\widehat{A}_{t}^{-1}\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}$, or its stabilized variant, which motivates the residual-form implementation.)
+BPTT/Monte Carlo error term (one explicit choice). To make the discretization/Monte Carlo dependence in (65) explicit, one may take for example
+$\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}):=\big\|\widehat{A}_{t}-A_{\pi^{\mathrm{warm}}}\big\|_{L^{2}(\mu)}+\big\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}\big\|_{L^{2}(\mu)}.$ (62)
+Under standard Lipschitz/bounded-moment assumptions (Euler strong error and Monte Carlo sampling error), one typically has a scaling of the form $\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta})\lesssim\Delta t^{1/2}+M_{\mathrm{MC}}^{-1/2}+M_{\theta}^{-1/2}$, up to constants depending on the working domain and stability bounds.
 A technical point is that a global small-gain condition of the form $C_{1}<1$ can be overly restrictive.
 Following the slab-wise philosophy in our prior PGDPO analysis (e.g.
-huh2025breaking), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab.
+Huh et al.
+(2025b, Appendix B)), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab.
 Concretely, assume $D$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$.
@@
 a short-time contraction of $T$ on each $D_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error.
-The proof combines a projection-map stability bound (Appendix C.1) with a slab-wise closure (Appendix C.2), in the same spirit as the slab analyses used in huh2025breaking.
+The proof combines a projection-map stability bound (Appendix D.1) with a slab-wise closure (Appendix D.2), in the same spirit as the slab analyses used in Huh et al.
+(2025b).
+Stage 1: PG–DPO
+Sample $\theta\sim q$ in the simulator; Euler rollout; BPTT; update $\varphi$.
+Output: $\varphi^{\mathrm{warm}}$.
+Stage 2: $q$-aggregated projection
+Under $\pi_{\varphi^{\mathrm{warm}}}$, estimate adjoint blocks; aggregate over $\theta$; solve $A\pi=-G$.
+Output: $\widehat{\pi}^{\mathrm{agg,mix}}$.
+Optional couplings
+Residual-form projection; interactive distillation (teacher refresh).
+Figure 1: Two-stage pipeline of Section 3. A detailed schematic (including the coupling mechanisms and implementation notes) is provided in Appendix F, Fig. LABEL:fig:sec3-pipeline-sideways.
 Theorem 3 (Residual-based ex–ante $\theta$-blind policy-gap bound for P–PGDPO (mixed-moment, deployable, slab-wise local)).
-Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix C.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$.
+Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix D.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$ (in particular, that $\widehat{A}_{t}$ remains a small perturbation of $A_{\pi^{\mathrm{warm}}}$ on $D$ under the adopted stabilizers).
 Let $\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem.
@@
 $\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G},$
-and assume the slab-wise Lipschitz gain of Appendix C.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
+and assume the slab-wise Lipschitz gain of Appendix D.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
 $\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G_{\pi_{1}}^{\mathrm{mix}}-G_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k}.$ (63)
@@
 Assume the slab partition is chosen so that $\rho_{*}<1$.
-Let $\widehat{\pi}^{\mathrm{agg,mix}}$ be the mixed-moment projected policy (67) computed from BPTT/Monte Carlo estimates under $\pi^{\mathrm{warm}}$, and let $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ be the population residual (68).
+Let $\widehat{\pi}^{\mathrm{agg,mix}}$ be the mixed-moment projected policy (59) computed from BPTT/Monte Carlo estimates under $\pi^{\mathrm{warm}}$ (with stabilized inversion as needed), and let $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ be the population residual (60).
 Then there exists $C_{2}>0$ such that
-$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}).$ (71)
-Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
+$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$ (65)
+where one admissible explicit choice of $\delta_{\mathrm{BPTT}}$ is given in (62). Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
 Proof.
-See Appendix C.3. ∎
-3.3 Coupling stage 1 and stage 2: residual projection and interactive distillation
-We keep the ex–ante objective (56) and the $\theta$-blind deployability constraint throughout.
-Stage 2 is not a separate optimization problem: it reuses the current stage 1 policy as a warm-up control under which the (costate-based) projection ingredients are estimated, and then applies a $q$-aggregated Pontryagin projection as a post-processing map.
-This subsection records two couplings between the two stages, each with a distinct role.
-First, we implement the projected rule in a residual (control-variate) form, which is algebraically equivalent to the direct projection but typically reduces Monte Carlo variance and improves numerical stability in high dimensions.
-Second, we use the projected output as a teacher signal via interactive distillation.
-Beyond acting as an optimization aid, distillation serves an amortization purpose: stage 2 projection can be accurate but Monte-Carlo intensive, whereas a distilled student policy can approximate the projected rule with a single forward pass at stage 1 inference cost.
-3.3.1 Control-variate (residual) form of the projected rule
-Recall the mixed-moment projected rule (67).
-In high dimensions, Monte Carlo noise in the projection inputs can be non-negligible, and solving a linear system with $\widehat{A}_{t}$ can amplify this noise.
-A convenient stabilization is to compute the same projected rule in a residual form around the warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$.
-Define the $\theta$-conditional residual (under frozen-$\theta$ simulations)
-$\widehat{r}_{\mathrm{FOC}}^{\theta}(t,x,y):=\widehat{A}_{t}^{\theta}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\theta}(t,x,y),$ (72)
-and the aggregated residual (the quantity we actually solve against)
-$\widehat{r}_{\mathrm{FOC}}(t,x,y):=\widehat{A}_{t}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\mathrm{mix}}(t,x,y).$ (73)
-Whenever $\widehat{A}_{t}(t,x,y)$ is invertible, the projected rule admits the identity
-$\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y)=\pi_{\varphi^{\mathrm{warm}}}(t,x,y)-\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{r}_{\mathrm{FOC}}(t,x,y),$ (74)
-which is an algebraic rewriting of (67) (hence it does not change the target).
-Its practical value is variance reduction: when the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small, and it often concentrates faster because the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and partially cancel.
-3.3.2 Interactive distillation: projection-guided training and amortized deployment
-Let $\pi_{\varphi}$ be the trainable stage 1 policy network.
-At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher.
-This coupling serves two purposes.
-During training it provides projection-guided targets that can stabilize and accelerate stage 1 optimization; after training it amortizes the expensive projection by distilling it into a fast deployable policy network.
-In residual form (74), the teacher is the $\theta$-blind map
-$\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y):=\pi_{\varphi^{-}}(t,x,y)-\big(\widehat{A}_{t}^{\varphi^{-}}(t,x,y)\big)^{-1}\,\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}(t,x,y),$ (75)
-where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy.
-We then train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to this teacher on the working domain:
-$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(t,x,y)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y)\big)\big\|^{2}\Big],$ (76)
-where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance.
-The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2; once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
-In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps: we hold $\varphi^{-}$ fixed for several updates of $\varphi$ under (76), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher.
-A practical schedule is to start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain (e.g., residual magnitudes and curvature/denominator stability) indicate that the stage 2 map has become reliable.
-Moreover, to avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we may apply projection guidance only on states where the projection checks certify reliability (an “adaptive teacher selection”); implementation details are deferred to the appendix (Appendix D).
+See Appendix D.3. ∎
+Coupling and amortization.
+Figure 1 summarizes the overall two-stage design.
+Stage 2 is used purely as a warm-started post-processing map: it reuses the current stage 1 policy, estimates the adjoint blocks under frozen $\theta\sim q$, and outputs a single deployable $\theta$-blind projected rule.
+For numerical stability and amortized deployment, we additionally consider (i) a residual/control-variate form of the projection and (ii) interactive distillation of the projected rule into the policy network.
+These practical coupling mechanisms are described in Appendix F.
 4 Breaking the Dimensional Barrier under Drift Uncertainty
-This section instantiates the decision-time static Gaussian drift-uncertainty benchmark in Section 2.3.1 and uses its closed-form constant-portfolio $q$-reference as an analytic target.
-Nature draws a fixed latent drift $\theta\sim q$ at $t=0$ and keeps it constant over $[0,T]$, while the investor cannot observe $\theta$ and must deploy a single $\theta$-blind policy under an ex–ante CRRA objective.
-Because the benchmark admits a transparent decision-time reference, we can measure accuracy directly via decision-time RMSE, rather than relying only on realized utility.
+This section instantiates the decision-time static Gaussian drift-uncertainty benchmark from Section 2.3.1 and uses its closed-form constant-portfolio $q$-reference as an analytic target.
+Nature draws a latent drift $\theta\sim q$ at $t=0$ and keeps it constant over $[0,T]$, while the investor cannot observe $\theta$ and must deploy a single $\theta$-blind policy under an ex–ante CRRA objective.
+Because the benchmark admits a transparent decision-time reference, we can measure accuracy directly via a decision-time Euclidean RMSE to the analytic reference, rather than relying only on realized utility.
 Our goal is to test whether Pontryagin-guided learning and projection remain stable as the number of assets grows.
 We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions.
-We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection) (and, when applicable, amortized variants via interactive distillation) under matched simulation budgets that scale linearly with $d$.
+We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection), and (when applicable) amortized variants via interactive distillation, under Monte Carlo budgets that scale linearly with $d$.
 4.1 Benchmark market and evaluation protocol
@@
 $\theta$-blind deployability (and what uses $\theta$).
-Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $t=0$ and $X_{0}$), and never takes the realized latent premium $\theta$ as an input.
+Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $(t,X_{0})$ with fixed horizon $T$), and never takes the realized latent premium $\theta$ as an input.
 The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (notably in Stage 2 projection).
@@
 APT-style factor construction of $(m,\Sigma)$.
-We construct the mean premium and covariance via a low-dimensional factor representation.
+We construct the mean premium and covariance via a low-dimensional factor representation in the spirit of arbitrage pricing theory (APT) Ross (1976).
 Let $W^{f}$ be a $k_{\Sigma}$-dimensional Brownian motion (factor shocks) and $W^{\varepsilon}$ a $d$-dimensional Brownian motion (idiosyncratic shocks), independent of $W^{f}$.
@@
 $P=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}+s\,\mathrm{diag}(D),$ (70)
-where $\widetilde{B}$ is generated independently of $B$ (or explicitly orthogonalized against the span of $B$ to enforce large principal angles).
-The factor term is rescaled so that its overall magnitude matches the aligned case under the same $s$ (e.g., by matching $\mathrm{tr}(P)$ or $\|P\|_{F}$ up to the shared diagonal component).
+where $\widetilde{B}$ is generated independently of $B$ (and can be orthogonalized against the span of $B$ to enforce large principal angles).
+The factor term is rescaled to match the aligned factor magnitude under the same $s$; concretely, writing $\Sigma_{\mathrm{fac}}:=B\Sigma_{f}B^{\top}$ and $\widetilde{\Sigma}_{\mathrm{fac}}:=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}$, we rescale $\widetilde{\Sigma}_{\mathrm{fac}}$ so that $\mathrm{tr}(\widetilde{\Sigma}_{\mathrm{fac}})=\mathrm{tr}(s\,\Sigma_{\mathrm{fac}})$ while keeping the diagonal term fixed as $s\,\mathrm{diag}(D)$.
 This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
 Experiment grid and simulation budgets.
-We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$, for both aligned and misaligned geometries.
-To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ sampling budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ paths and a high regime with $N_{\mathrm{MC}}=400\cdot d$ (where $N_{\mathrm{MC}}$ denotes the per-update or per-estimator path budget, depending on the stage).
-All methods share the same discretization scheme (Euler) and action constraints; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs.
+We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$ for both aligned and misaligned geometries.
+To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ trajectory budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ rollouts and a high regime with $N_{\mathrm{MC}}=400\cdot d$ rollouts.
+Here $N_{\mathrm{MC}}$ denotes the number of independent simulated trajectories used to approximate the relevant $q$-expectations (in Stage 1 for stochastic gradient estimates, and in Stage 2 for costate/mixed-moment estimators at queried states).
+Each trajectory draws its own latent $\theta\sim q$ at time 0 and its own Brownian path thereafter; we do not use a separate inner loop of $\theta$-samples per trajectory.
+All methods share the same discretization scheme (Euler) and the same market instances within each $d$; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs.
 Stage 2) are reported in the implementation appendix and code release.
@@
 We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint.
-For $\gamma>1$ we use the CRRA reference (38) and, for $\gamma=1$, the log-utility reference (31).
-We evaluate each method at $t=0$ on a fixed grid $\{(X_{0}^{(i)},T^{(i)})\}_{i=1}^{N_{\mathrm{eval}}}$ and report RMSE to the analytic reference:
-$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T^{(i)})-\pi_{q,\gamma}^{\mathrm{const}}(T^{(i)})\big\|^{2}\bigg)^{1/2},$ (82)
-where $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output).
+For $\gamma>1$ we use the CRRA reference (31).
+We evaluate each method at $t=0$ with a fixed horizon $T=1.5$ on a fixed grid $\{X_{0}^{(i)}\}_{i=1}^{N_{\mathrm{eval}}}$ and report Euclidean RMSE to the analytic reference:
+$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T)-\pi_{q,\gamma}^{\mathrm{const}}(T)\big\|^{2}\bigg)^{1/2},$ (71)
+where $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$ and $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output).
 With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (68)–(70), the remaining subsections compare baseline Stage 1 PG–DPO, post-hoc Stage 2 P–PGDPO projection, and interactive distillation under matched simulation budgets.
@@
 [^1] We note, however, that in extreme uncertainty/misalignment—where $\theta$–costate dependence becomes pronounced—the decoupling can break down, in which case one should revert to full mixed-moment estimation (possibly with larger budgets and/or regularized/certified projection).
-Figure 1: Decision-time RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window (computed from the last evaluation snapshots).
+Figure 2: Decision-time Euclidean RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window.
+Concretely, for each condition we take the median RMSE over the last evaluation snapshots recorded during training.
 Rows: uncertainty magnitude $s\in\{10^{-3},10^{-2},10^{-1}\}$.
@@
 dashed lines correspond to MC base ($100\cdot d$) vs.
-high ($400\cdot d$) budgets.
+high ($400\cdot d$) trajectory budgets.
 Protocol and summary statistic.
 We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (31).
-We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (82).
+We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (71) with fixed horizon $T=1.5$.
 Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window.
-Unless stated otherwise, the projection/teacher direction uses the mixed-moment ($p_{\theta}$) aggregation.
+Unless stated otherwise, the projection/teacher direction uses the $p_{\theta}$ aggregation (Section 3.2).
 What is compared in Figure 2.
 Stage 1 (PG–DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients.
-Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section 3.3.1.
-Interactive distillation (Section 3.3.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
+Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section F.1.
+Interactive distillation (Section F.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
 Thus Figure 2 separates projection quality (Stage 2: post-hoc projected, still $\theta$-blind) from amortized deployable quality (Stage 1 distilled: single forward pass).
+Figure 3: Pathwise sanity check at $d=100$ under common random numbers.
+Black: analytic Gaussian constant-fraction benchmark evaluated with the remaining horizon $\tau=T-t$ (fixed $q$).
+Blue: Stage 1 PG–DPO.
+Orange: on-the-fly Stage 2 P–PGDPO teacher (computed via Monte Carlo projection at visited states).
+Top/bottom: aligned/misaligned.
 Stage 2 projection versus amortization: scaling with dimension.
-Aligned geometry.
-For small and moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 delivers a sharp reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$.
+For the aligned geometry, small and moderate uncertainty ($s=10^{-3},10^{-2}$) yield a sharp Stage 2 reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$.
 Interactive distillation consistently improves the deployable policy (Stage 1 (distill.) below Stage 1) while leaving Stage 2 essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation reduces the policy-class approximation/optimization gap by injecting that signal into $\pi_{\varphi}$.
-Misaligned geometry.
-The picture becomes more heterogeneous.
+In contrast, under the misaligned geometry the picture becomes more heterogeneous.
 For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$.
@@
 Mechanism: why misalignment can reduce projection gains.
-To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix E; see Figures 36.
+To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix G; see Figures 47.
 The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting.
-These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix directions and can amplify Monte-Carlo errors in mixed-moment quantities (e.g., $\mathbb{E}[p_{1}\theta]$), especially as $d$ increases.
-Figure 2: Pathwise sanity check at $d=100$ under common random numbers (same sampled $\theta$ and Brownian increments).
-Top: aligned geometry.
-Bottom: misaligned geometry.
-Each panel shows $\log X_{t}$ trajectories induced by the warm Stage 1 policy (PGDPO), the online Stage 2 P–PGDPO teacher (residual form), and the analytic $q$-reference.
+These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix uncertainty directions that are poorly aligned with risk directions, which can amplify estimation noise in the projected correction as $d$ grows.
 Pathwise sanity check.
 Figure 3 complements the decision-time RMSE with a trajectory-level view under common random numbers.
-In the aligned case, the online Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy.
+In the aligned case, the on-the-fly Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy.
 In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 2 and motivating amortization/reliability mechanisms in interactive distillation.
-Overall, the benchmark highlights a separation of roles.
-Stage 2 projection supplies a structured stationarity-correction signal that is particularly effective under aligned uncertainty, and interactive distillation amortizes this signal into a fast deployable Stage 1 policy.
-Under misalignment, projection can become more sensitive as $d$ and $s$ grow, consistent with diagnostic evidence of increased residuals and curvature mismatch; nevertheless, amortization remains a robust route to improving deployable policies under fixed simulation budgets.
 4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
 Why include PPO, and how we match the setting.
-Proximal Policy Optimization (PPO) is a widely used and robust model-free policy-gradient baseline for continuous control (schulman2017proximal).
+Proximal Policy Optimization (PPO) is a widely used and robust model-free policy-gradient baseline for continuous control (Schulman et al., 2017).
 We include PPO to answer a concrete question: can a generic, well-tuned model-free RL method recover the decision-time $q$-optimal $\theta$-blind allocation in our high-dimensional drift-uncertainty benchmark under comparable simulation budgets?
 This comparison is especially informative in our static Gaussian benchmark because the target decision-time rule is structurally simple (constant and available in closed form), so performance gaps primarily reflect optimization difficulty and credit assignment rather than policy-class expressiveness.
-Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section 2.4), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure.
+Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section B), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure.
 For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our PG–DPO pipeline, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective.
-We also use the same action cap $u_{\max}$ (with the same dimension-scaling convention) so that exploration ranges are comparable across $d$.
-Implementation details are deferred to the appendix and code release.
+We match simulation budgets at the level of total environment interactions (simulated trajectories $\times$ time steps); implementation and PPO tuning details are deferred to the appendix and code release.
 Empirical outcome.
@@
 The gap is not evidence that PPO is intrinsically weak; rather, it reflects that our benchmark stresses regimes where a generic likelihood-ratio policy gradient is statistically disadvantaged compared to pathwise/adjoint-based updates.
-With terminal utility as the only reward, PPO faces a long-horizon credit-assignment problem whose gradient variance grows with both horizon and action dimension.
-Sampling $\theta\sim q$ further creates episode-wise heterogeneity under a single $\theta$-blind policy, inducing additional variance and potential cancellation across parameter draws.
+Terminal-reward credit assignment.
+With terminal utility as the only reward signal, PPO faces a long-horizon credit-assignment problem: most intermediate transitions carry zero reward, so advantage estimates inherit high variance and weak learning signals, especially in higher dimensions.
+One contributor is terminal-reward credit assignment: with terminal utility as the only reward signal, most intermediate transitions carry no direct learning signal, so advantage estimates become noisy and the gradient signal weakens as the horizon and dimension grow.
+A second contributor is high-dimensional action noise and variance scaling: in continuous control, exploration is typically implemented through stochastic policies (e.g., diagonal Gaussians), and as $d$ increases the effective exploration noise and likelihood-ratio gradient variance can inflate rapidly, making it difficult to consistently recover the precise constant-fraction structure of the $q$-optimal $\theta$-blind rule under fixed simulation budgets.
 In contrast, Stage 1 exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction that is tailored to the $\theta$-blind ex–ante objective.
-Under matched simulation budgets in our latent-$\theta$, $\theta$-blind benchmark, a generic model-free PPO baseline does not reliably recover the decision-time $q$-optimal allocation (Table LABEL:tab:crra_stage12_distill_ppo_landscape), motivating structure-exploiting alternatives—pathwise gradients, costates, and the $q$-aggregated Pontryagin projection—as in PG–DPO and P–PGDPO.
 5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
 Sections 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic.
-Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (campbell2002strategic; xia2001learning).
-We use the mean-reverting Gaussian premium benchmark of Section 2.3.2.
-Decision-time statistical uncertainty enters through the (uncertain) initial premium state $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$, while a nonzero return–factor correlation $\rho$ generates hedging demand through the cross term $M_{\mathrm{cross}}(T)$ in (51).
-Crucially, we enforce a deployable restriction aligned with Section 3: the policy is $Y$-blind and does not observe the realized $Y_{0}$ nor the path $(Y_{t})$.
-We compare: (i) Stage 1 PG–DPO (deployable end-to-end learning; Section 3.1), (ii) Stage 2 $q$-aggregated Pontryagin projection (post-hoc correction in residual form; Sections 3.2 and 3.3.1), (iii) interactive distillation (amortized projection guidance; Section 3.3.2), and (iv) a model-free PPO baseline trained under the same deployable $Y$-blind observation restriction.
-Performance is measured by decision-time RMSE against the analytic constant-portfolio OU reference (51) (which reduces to the independence-case benchmark when $\rho=0$).
-In addition to the full allocation error, this benchmark provides a natural myopic + hedging decomposition (driven by return–factor correlation).
-We therefore report (a) the RMSE of the full decision-time allocation for all methods (including PPO), and (b) component-wise diagnostics for the projected (Stage 2) rules: RMSE of the hedging component (Table 1) and, in the appendix, the RMSE of the myopic component (Table 2) and the cosine similarity of the hedging direction (Table 3).
-Since PPO does not expose a compatible myopic/hedging decomposition for these diagnostics, we include it only in the full RMSE table.
-To keep the main text focused, we include the full RMSE table (Table LABEL:tab:ou_rmse_s0_sweep_landscape) and the hedging-RMSE table (Table 1) in Section 5.2; the remaining two diagnostic tables are deferred to the appendix (Section F).
+Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (Campbell and Viceira, 2002; Xia, 2001).
+We adopt the mean-reverting Gaussian premium benchmark of Section 2.3.2.
+At decision time ($t=0$), the investor is assumed to have access only to an interval/posterior uncertainty description $q(dy)$ for the initial premium state $Y_{0}$ (e.g., $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$), and we restrict attention to deployable $Y$-blind policies that depend only on observable wealth and time-to-go.
+When return and factor shocks are correlated ($\rho\neq 0$), the analytic OU decision-time reference (43) contains the cross term $M_{\mathrm{cross}}(\tau)$ (with $\tau:=T-t$) that encodes intertemporal hedging demand.
+We evaluate the two-stage PG–DPO / $q$-aggregated Pontryagin projection pipeline (and interactive distillation) against a model-free PPO baseline, all under the same $Y$-blind deployability restriction.
+Performance is measured by decision-time RMSE at $t=0$ relative to the analytic constant-portfolio OU reference (43).
 5.1 Experimental setting
-$Y$-blind deployability (and what uses $Y$).
-Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes the realized initial premium $Y_{0}$ nor the factor path $(Y_{t})$ as an input (including the PPO baseline).
+Decision-time interval/posterior estimate for $Y_{0}$ and deployability.
+Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes as input (i) a point estimate of $Y_{0}$, (ii) the realized initial premium $Y_{0}$, or (iii) the factor path $(Y_{t})$ (including the PPO baseline).
+At deployment, the investor is assumed to have access only to an uncertainty description $q(dy)$ for $Y_{0}$ (interval/posterior output of an external estimation pipeline), and commits to a single policy optimized ex ante under $Y_{0}\sim q$ (without filtering/updating over $[0,T]$).
 The latent premium factor is sampled and propagated only inside the simulator to generate trajectories and to form Monte Carlo averages used by the stage 2 projection (and by the teacher in distillation).
@@
 We adopt the OU premium benchmark of Section 2.3.2.
+To avoid a notational clash with other uses of $\kappa$ elsewhere in the paper, we denote the OU mean-reversion rate by $\kappa_{Y}$ in this section.
 Let $Y_{t}\in\mathbb{R}^{m}$ be a mean-reverting premium factor and $R_{t}\in\mathbb{R}^{d}$ the risky excess returns:
-$dY_{t}$ $=\kappa(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$
+$dY_{t}$ $=\kappa_{Y}(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$
 $dR_{t}$ $:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},$
 $d\langle W,W^{Y}\rangle_{t}$ $=\rho\,dt.$
-A nonzero $\rho$ induces intertemporal hedging demand and enters the CRRA decision-time reference through the cross-covariance term $M_{\mathrm{cross}}(T)$ in (51).
-When $\rho=0$ (independent return and factor shocks), the hedging channel vanishes ($M_{\mathrm{cross}}(T)=0$) and the reference reduces to the independence-case benchmark.
+Here $\rho\in\mathbb{R}^{d\times m}$ denotes the instantaneous return–factor shock correlation matrix, satisfying the feasibility condition $\rho^{\top}\rho\preceq I_{m}$.
+A nonzero $\rho$ induces intertemporal hedging demand and enters the CRRA decision-time reference through the cross-covariance term $M_{\mathrm{cross}}(\tau)$ in (43), where $\tau:=T-t$ is time-to-go (so at decision time $t=0$, $\tau=T$).
+When $\rho=0$ (independent return and factor shocks), the hedging channel vanishes ($M_{\mathrm{cross}}(\tau)=0$) and the reference reduces to the independence-case benchmark.
 Decision-time uncertainty geometry for $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$.
-We control the magnitude of decision-time statistical uncertainty by a scalar $s_{0}>0$ and construct $P_{0}$ from an identification-motivated baseline
+We control the magnitude of decision-time estimation uncertainty by a scalar $s_{0}>0$ and construct $P_{0}$ from an identification-motivated baseline
 $\widetilde{P}_{0}:=(B^{\top}\Sigma^{-1}B)^{-1}\in\mathbb{R}^{m\times m}.$
@@
 Stage 1 trains a deployable policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1).
-Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section 3.3.1).
-Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section 3.3.2).
+Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section F.1).
+Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section F.2).
 As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table LABEL:tab:ou_rmse_s0_sweep_landscape.
-We sweep $d\in\{5,10,50,100\}$ (one fixed market instance per $d$), train for $5000$ epochs, and evaluate every $100$ epochs.
-Unless stated otherwise we set $\gamma=2$, $r=0.03$, $\kappa=1.0$, $\xi_{\mathrm{scale}}=0.25$, and $\rho=0.5$.
-We evaluate the decision-time action at $t=0$ and report RMSE to the analytic constant-portfolio OU reference (51).
-In addition to the full allocation error, we use the natural myopic + hedging decomposition induced by the OU factor structure.
-We report the RMSE of the full decision-time allocation for all methods, and component-wise diagnostics for the projected (Stage 2) rules, including the RMSE of the hedging component and (in the appendix) the RMSE of the myopic component and cosine similarity of the hedging direction.
+We sweep $d\in\{5,10,50,100\}$ (one fixed market instance per $d$, following the same market-generation protocol used in Section 4), train for $5000$ epochs, and evaluate every $100$ epochs.
+We evaluate the decision-time action at $t=0$ (i.e., $\tau=T$) and report RMSE to the analytic constant-portfolio OU reference (43).
+Unless stated otherwise we set $\gamma=2$, $r=0.03$, $\kappa_{Y}=1.0$, $\xi_{\mathrm{scale}}=0.25$, and a correlation magnitude $\rho_{0}=0.5$.
+We instantiate the model parameters as follows: (i) we take $\Xi=\xi_{\mathrm{scale}}I_{m}$, (ii) we implement the return–factor correlation by choosing a matrix $\rho=\rho_{0}Q$ with $Q\in\mathbb{R}^{d\times m}$ having orthonormal columns ($Q^{\top}Q=I_{m}$), ensuring $\rho^{\top}\rho=\rho_{0}^{2}I_{m}\preceq I_{m}$.
+(A simple choice is $Q=[I_{m};0]\in\mathbb{R}^{d\times m}$ when $m\leq d$; any fixed orthonormal-column $Q$ yields an equivalent correlation magnitude.) We fix the horizon to $T=1.5$ in this section.
+In addition to the full allocation error, we use the myopic+hedging decomposition induced by the OU structure, and report component-wise diagnostics for the projected (Stage 2) rules.
 To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
-We report decision-time RMSE at $t=0$ against the analytic OU reference (51).
-To reduce noise from stochastic optimization, we summarize each condition by a tail median over the last six evaluation checkpoints.
-Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE for all deployable objects (Stage 1 and Stage 1+Stage 2, with and without distillation), and also includes a model-free PPO baseline trained under the same $Y$-blind deployability restriction.
-To isolate the economic channel of interest, Table 1 reports the RMSE of the hedging component for the post-hoc projected (Stage 2) rules.
-Two additional diagnostics—the myopic-component RMSE and the hedging-direction cosine similarity—are deferred to the appendix (Tables 2 and 3).
-Since PPO does not expose a compatible myopic/hedging decomposition in our diagnostic protocol, we report it only in the full-RMSE table.
-Projection and economic hedging-demand recovery.
-Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (Stage 1+Stage 2) substantially reduces decision-time RMSE relative to the deployable Stage 1 policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
-For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, Stage 1 attains $3.54\text{\times}{10}^{-03}$ whereas Stage 1+Stage 2 achieves $1.56\text{\times}{10}^{-04}$.
-The component-wise diagnostics indicate that the remaining discrepancy is largely driven by the hedging channel: in the same setting, the hedging RMSE is $1.55\text{\times}{10}^{-04}$ (Basic) and $1.42\text{\times}{10}^{-04}$ (Distill.) (Table 1), while the myopic RMSE is an order of magnitude smaller (Appendix Table 2).
-This pattern is consistent with the economic mechanism in this benchmark: once the (mostly) myopic component is captured, the dominant remaining challenge is to recover the intertemporal hedge induced by correlated return–factor shocks.
-Amortization, robustness, and the PPO baseline.
-Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
-This matches the intended division of labor in Section 3.3: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
-As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables LABEL:tab:ou_rmse_s0_sweep_landscape1).
-Misalignment has a limited effect for small and moderate uncertainty scales, but can induce noticeable deterioration in the hardest settings, where the direction-of-hedge diagnostic can also weaken (Appendix Table 3).
-Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, with degradation that becomes especially pronounced at larger $d$ (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
-This is consistent with PPO facing a terminal-only credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline.
-Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
+Metrics and diagnostics.
+Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE at $t=0$ for all deployable policies, including the PPO baseline.
+To isolate the economic hedging channel, we use the OU-induced decomposition of the analytic reference
+$\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau),$
+where $\pi^{\mathrm{myo}}$ is the independence-case reference obtained by setting $\rho=0$ (equivalently, dropping $M_{\mathrm{cross}}(\tau)$ in (43)), and $\pi^{\mathrm{hedge}}:=\pi^{\mathrm{ref}}-\pi^{\mathrm{myo}}$ is the residual $\rho$-dependent term.
+We report the hedging-component RMSE for the projected (Stage 2) rules in Table 1; the myopic-component RMSE and hedging-direction cosine similarity are deferred to Appendix H.
+Since Stage 1 and PPO do not expose a compatible myopic/hedging split under our diagnostic protocol, component-wise diagnostics are reported only for Stage 2.
 $s_{0}$ Method $d=5$ $10$ $50$ $100$
@@
 Table 1: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the post-hoc projected (Stage 2) policies (tail median over the last six evaluation checkpoints).
-Component-wise diagnostics are reported for Stage 2 since Stage 1 does not explicitly output a myopic/hedging decomposition.
-In a factor-driven market where return–factor correlation induces intertemporal hedging demand, the proposed two-stage pipeline recovers the analytic OU decision-time reference with high accuracy: projection provides the dominant gains, and distillation improves deployable policies by amortizing projection guidance.
-In contrast, a model-free PPO baseline does not reliably match the analytic reference in this $Y$-blind setting.
+We use the decomposition $\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau)$ where $\pi^{\mathrm{myo}}$ is the $\rho=0$ reference and $\pi^{\mathrm{hedge}}$ is the residual $\rho$-dependent term (cf.
+(43)).
+Component-wise diagnostics are reported for Stage 2 since Stage 1 and PPO do not explicitly output a compatible myopic/hedging split under our protocol.
+Projection and economic hedging-demand recovery.
+Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (Stage 1+Stage 2) substantially reduces decision-time RMSE relative to the deployable Stage 1 policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
+For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, Stage 1 attains $3.54\text{\times}{10}^{-03}$ whereas Stage 1+Stage 2 achieves $1.56\text{\times}{10}^{-04}$.
+The component-wise diagnostics indicate that the remaining discrepancy is largely driven by the hedging channel: in the same setting, the hedging RMSE is $1.55\text{\times}{10}^{-04}$ (Basic) and $1.42\text{\times}{10}^{-04}$ (Distill.) (Table 1), while the myopic RMSE is an order of magnitude smaller (Appendix Table 2).
+This pattern is consistent with the economic mechanism in this benchmark: once the (mostly) myopic component is captured, the dominant remaining challenge is to recover the intertemporal hedge induced by correlated return–factor shocks.
+Amortization, robustness, and the PPO baseline.
+Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
+This matches the intended division of labor in Section F: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
+As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables LABEL:tab:ou_rmse_s0_sweep_landscape1).
+Misalignment has a limited effect for small and moderate uncertainty scales, but can induce noticeable deterioration in the hardest settings, where the direction-of-hedge diagnostic can also weaken (Appendix Table 3).
+Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, and can be unstable in this terminal-only setting (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
+This is consistent with PPO facing a difficult credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline.
+Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
 6 Conclusion
@@
 We established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3).
-In experiments with finite-sample uncertainty, projection improves stability and accuracy in high dimensions and exhibits a two-time-scale stabilization effect (costates versus policies), while interactive distillation amortizes the projection into a fast deployable network (Sections 4 and 5; Section 3.3).
+In experiments, projection provides the dominant gains—improving stability and decision-time accuracy in high-dimensional drift-uncertainty scaling tests and recovering intertemporal hedging demand in factor-driven markets—while interactive distillation amortizes projection into a fast deployable network (Sections 4 and 5; Section F).
 Several extensions are natural.
 A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to belief-aware decision rules (Remark 2, Appendix A).
-A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section 2.4; Section 4.2; Appendix D.5).
+A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section B; Section 4.2; Appendix E.5).
 Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
@@
 Sections 2.3.1 and 2.3.2 focus on decision-time benchmarks in which an uncertainty description $q$ is treated as given and the investor optimizes under the corresponding $\theta$-blind deployability constraint.
-In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (barberis2000investor; pastor2000portfolio; xia2001learning).
-This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., bensoussan1985optimal; pham2017dynamic), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$.
+In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (Barberis, 2000; Pástor, 2000; Xia, 2001).
+This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., Bensoussan and van Schuppen (1985); Pham and Wei (2017)), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$.
 We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization.
@@
 $\dot{P}_{t}$ $=KP_{t}+P_{t}K^{\top}+\Xi\Xi^{\top}-P_{t}B^{\top}\Sigma^{-1}BP_{t},\hskip 18.49988ptP_{0}\ \text{given}.$ (79)
-Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (90); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (bensoussan1985optimal; pham2017dynamic).
+Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (79); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
 From a posterior on $Y_{t}$ to a Gaussian uncertainty description for decision-time optimization.
@@
 $\pi_{t}^{\mathrm{plug}}:=\Big(\gamma\,\Sigma+(\gamma-1)\tau\,P_{t,T}\Big)^{-1}m_{t,T}.$ (85)
-One may interpret (96) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (barberis2000investor; pastor2000portfolio; xia2001learning).
+One may interpret (85) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (Barberis, 2000; Pástor, 2000; Xia, 2001).
 Remarks (relation to belief-aware control).
 The plug-in rule (85) is intentionally decision-time: it conditions on the current uncertainty description and does not attempt to optimize over how the posterior will evolve.
-In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (barberis2000investor; pastor2000portfolio; xia2001learning).
-Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (bensoussan1985optimal; pham2017dynamic).
-Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (haussmann1987maximum; li1995general; baghery2007maximum).
+In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (Barberis, 2000; Pástor, 2000; Xia, 2001).
+Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (Haussmann, 1987; Li and Tang, 1995; Baghery and Øksendal, 2007).
 Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in $(x,y,\hat{Y},P)$-space (or its sufficient-statistic analogues) is an important direction that we defer to future work.
-Appendix B Proof of Theorem 2
-Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in huh2025breaking).
+Appendix B Why dynamic programming and deep PDE surrogates are not primary baselines under the $\theta$-blind fixed-$q$ interface
+This appendix explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted in this paper.
+DP is conceptually sound in low-dimensional Markovian settings (Fleming and Soner, 2006; Pham, 2009), but two obstructions dominate in our regime: (i) value-function approaches require accurate derivatives (often including mixed sensitivities) to recover optimal portfolios, and this becomes prohibitive as dimension and nonlinearity grow; (ii) principled parameter uncertainty magnifies both the state dimension and the aggregation difficulties.
+These points are standard for grid-based HJB solvers (Bellman, 1961; Kushner and Dupuis, 2001) and remain relevant for modern deep PDE surrogates (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), especially under partial information/belief-state formulations (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+B.1 The value-function target for latent $\theta$ is a belief-state functional
+Fixed-$q$ ex–ante objective and $\theta$-blind deployability.
+Our decision-time interface is
+$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}J(\pi)\;:=\;\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right]\right],$ (86)
+where $\theta\sim q$ is latent, fixed along each episode, and $\mathcal{A}^{\mathrm{blind}}$ denotes admissible controls measurable with respect to the observable filtration (so the deployed policy cannot take $\theta$ as an input).
+Why $(t,X_{t},Y_{t})$ is not a closed DP state.
+In general, the observable history carries information about $\theta$, so the continuation value at time $t$ depends on the conditional law
+$q_{t}(\cdot):=\mathcal{L}\!\big(\theta\mid\mathcal{F}_{t}^{\mathrm{obs}}\big),$ (87)
+not only on $(t,X_{t},Y_{t})$.
+This is the standard partially observed control phenomenon: DP becomes Markov only after augmenting the state by a belief (or a finite-dimensional sufficient statistic when available) (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+Belief-state DP is infinite-dimensional under our interface.
+The principled DP value function is therefore a functional of a probability measure:
+$V(t,x,y,\mu):=\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}^{t,x,y,\mu}\!\left[U(X_{T}^{\pi})\right],\hskip 18.49988pt\mu\in\mathcal{P}(\Theta),$ (88)
+so the associated HJB equation is posed on a space of measures in general (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+Under our inference-agnostic interface, $q$ need not belong to a conjugate family and we do not specify a filtering model that would keep $q_{t}$ in a tractable finite-dimensional class.
+Consequently, a DP/PDE baseline would be forced to approximate $V$ over an infinite-dimensional belief state, which is not the computational target of this paper.
+B.2 Even with deterministic parameters, HJB solvers face a derivative bottleneck in high dimension
+Classical HJB: curse of dimensionality and full nonlinearity.
+With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (Fleming and Soner, 2006; Pham, 2009).
+Grid-based solvers scale exponentially in the state dimension (Bellman, 1961; Kushner and Dupuis, 2001).
+In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points.
+Moreover, realistic features (constraints, transaction costs, non-affine dynamics) typically yield fully nonlinear HJBs.
+Recovering the optimal portfolio from a value function also requires derivatives (schematically, in Merton-type problems)
+$\pi^{\star}(t,x,y)\;\propto\;-\frac{V_{x}(t,x,y)}{V_{xx}(t,x,y)}\times(\text{risk model})^{-1}\times(\text{drift model}),$ (89)
+and in factor-driven settings intertemporal hedging demands are driven by mixed sensitivities (e.g.
+$V_{xy}$).
+Thus the numerical burden is not only approximating $V$ but also obtaining stable first/second/mixed derivatives in high dimension.
+B.3 Deep PDE surrogates remove grids but not the value-function derivative bottleneck
+PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities that drive intertemporal hedging.
+In practice this can induce nonconvex and ill-conditioned training objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable high-dimensional portfolio policies.
+This is not a claim that deep PDE methods are “invalid”; rather, under our high-dimensional and $\theta$-blind interface, they are not practical baselines relative to direct, simulation-only policy optimization methods.
+B.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
+A natural alternative is to condition on $\theta$ and solve a family of full-information problems.
+$\theta$-conditional HJB.
+If $\theta$ were observed, the full-information value function $V^{\theta}$ solves
+$0=\partial_{t}V^{\theta}(t,x,y)+\sup_{\pi\in\mathbb{R}^{d}}\Big\{\mathcal{L}^{\pi,\theta}V^{\theta}(t,x,y)\Big\},\hskip 18.49988ptV^{\theta}(T,x,y)=U(x),$ (90)
+for the controlled generator $\mathcal{L}^{\pi,\theta}$ (Fleming and Soner, 2006; Pham, 2009).
+This yields a $\theta$-dependent optimizer $\pi^{\star}(t,x,y,\theta)$, which is infeasible in our deployable $\theta$-blind setting.
+Supremum–expectation non-commutativity.
+More fundamentally, even perfect knowledge of the $\theta$-conditional optimizer does not solve the $\theta$-blind objective, because
+$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[J(\pi,\theta)\right]\;\leq\;\mathbb{E}_{\theta\sim q}\!\left[\sup_{\pi}J(\pi,\theta)\right],\hskip 18.49988ptJ(\pi,\theta):=\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right],$ (91)
+and the inequality is typically strict. The right-hand side corresponds to the (infeasible) $\theta$-adaptive controller.
+B.5 Why “sample $\theta$ only in the PDE loss” is mismatched to the ex–ante objective
+One might try to train a single value network $V_{\varphi}(t,x,y)$ that does not take $\theta$ as input, and sample $\theta\sim q$ only inside a collocation/pathwise residual loss:
+$\min_{\varphi}\;\mathbb{E}_{\theta\sim q}\!\Big[\big\|\mathrm{HJB}_{\theta}\!\big(V_{\varphi}\big)(t,x,y)\big\|^{2}\Big],\hskip 18.49988pt\text{(schematic)}$ (92)
+where $\mathrm{HJB}_{\theta}(\cdot)$ is the full-information HJB operator for parameter $\theta$. This is appealing but generally mismatched to (86):
+  • Different $\theta$ impose different (often conflicting) optimality conditions at the same $(t,x,y)$, so there is generally no single $V$ satisfying $\mathrm{HJB}_{\theta}(V)=0$ for many $\theta$ simultaneously.
+  • Minimizing an expected squared PDE residual (92) is not equivalent to maximizing expected terminal utility under $\theta$-blind controls.
+  • Portfolio feedback rules are typically extracted through nonlinear operations on derivatives (ratios, inverses), so small value-function approximation errors can induce large control errors.
+B.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
+The aggregation issue is visible even in a one-asset Merton model with static latent drift. Let
+$\frac{dS_{t}}{S_{t}}=r\,dt+\theta\,dt+\sigma\,dW_{t},\hskip 18.49988pt\theta\sim\mathcal{N}(m,p),$
+and restrict to constant fractions $\pi$ over a remaining horizon $\tau$.
+For CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ with $\gamma>1$, the $\theta$-conditional full-information constant Merton rule is
+$\pi^{\star}(\theta)=\frac{\theta}{\gamma\sigma^{2}}.$ (93)
+Averaging this infeasible rule gives $\mathbb{E}[\pi^{\star}(\theta)]=m/(\gamma\sigma^{2})$. By contrast, the deployable $\theta$-blind ex–ante optimizer within the constant-fraction class is
+$\pi_{q}^{\mathrm{const}}(\tau)=\frac{m}{\gamma\sigma^{2}+(\gamma-1)\tau\,p},\hskip 18.49988pt(\gamma>1),$ (94)
+which exhibits shrinkage due to estimation uncertainty ($p>0$) and depends on the horizon $\tau$. Thus $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\mathrm{const}}(\tau)$ whenever $p>0$.
+B.7 Implication for baselines in this paper
+The discussion above motivates our baseline choices in the main text:
+  • We validate on controlled drift/premium benchmarks where closed-form decision-time Gaussian targets are available and match the fixed-$q$ deployability restriction.
+  • For learning-based comparisons, we emphasize simulation-only direct policy search baselines (e.g., PPO) that operate under the same observation restriction (no access to $\theta$; $\theta$ sampled only inside the simulator), rather than value-function PDE/BSDE solvers whose natural DP target is belief-augmented.
+Appendix C Proof of Theorem 2
+Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in Huh et al.
+(2025b)).
 Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.
+(This appendix replaces the current Appendix C in the draft.)
 Important remark (what this proof does not use).
@@
 It does not use the $\theta$-blind $q$-aggregated stationarity condition (Theorem 1 in Section 3.2).
-Those constructions affect only the deployable aggregation/projection target in stage 2 and are irrelevant to the BPTT–PMP convergence itself.
-Notation and filtration. Fix $\theta\in\Theta$. We work conditionally on this $\theta$ and consider the augmented (simulator) filtration
-$\mathbb{G}^{\theta}:=(\mathcal{G}_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988pt\mathcal{G}_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
-All conditional expectations and $L^{2}$ projections below are taken with respect to $\mathcal{G}_{t_{k}}^{\theta}$.
-This choice matches the information set used by the simulator and by automatic differentiation/BPTT (which differentiates through the full forward recursion).
+Those constructions affect only stage 2 projection targets and are irrelevant to the BPTT–PMP convergence itself.
+Notation, filtration, and the pathwise-vs-adapted issue. Fix $\theta\in\Theta$. We work conditionally on this $\theta$ and consider the augmented (simulator) filtration
+$\mathbb{G}^{\theta}:=(G_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988ptG_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
+All conditional expectations and $L^{2}$ projections below are taken with respect to $G_{t_{k}}^{\theta}$.
+Key point.
+The raw BPTT “adjoints” are pathwise derivatives of the terminal utility along the full discrete computation graph.
+Hence, for $k<N$, they are typically $G_{T}^{\theta}$-measurable and need not be $G_{t_{k}}^{\theta}$-measurable (they can depend on future increments).
+In contrast, the stochastic maximum principle (SMP/PMP) costate is $\mathbb{G}^{\theta}$-adapted.
+Therefore, the correspondence is formulated after a standard one-step conditional $L^{2}$ projection (equivalently, a discrete-time BSDE representation), which turns pathwise objects into adapted discrete adjoints.
 Let $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, $N\Delta t=T$.
-For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind in the sense of (52) but evaluated along the $\theta$-conditional trajectory.
+For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind but evaluated along the $\theta$-conditional trajectory.
 Step 1: Conditioning on $\theta$ and uniformity of bounds.
@@
 Assume the coefficients in (45)–(46) satisfy the usual Lipschitz and linear-growth conditions uniformly over $\theta\in K$, and that the block covariance structure of $(W,W^{Y})$ (including instantaneous correlation) is uniformly nondegenerate on $K$.
-Then, for each fixed $\theta\in K$, the controlled SDE system (53)–(54) is well posed and admits uniform-in-time $L^{2}$ moment bounds.
+Assume also the regularity/integrability needed for the $\theta$-conditional SMP (including the blocks relevant for the portfolio Hamiltonian gradient) to be well posed.
+Then, for each fixed $\theta\in K$, the controlled SDE system is well posed and admits uniform-in-time $L^{2}$ moment bounds.
 Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
-$\sup_{t\in[0,T]}\mathbb{E}\big[\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\|^{2}\big]^{1/2}\;\leq\;C_{K}\,\Delta t^{1/2},$
-with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are the deterministic assumptions used in huh2025breaking, now stated uniformly on $K$.
-Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme for (53)–(54) on the grid $(t_{k})$:
+$\sup_{t\in[0,T]}\mathbb{E}\!\left[\left\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\right\|^{2}\right]^{1/2}\leq C_{K}\,\Delta t^{1/2},$
+with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are exactly the deterministic assumptions used in Huh et al. (2025b), now stated uniformly on $K$.
+Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme on the grid $(t_{k})$:
 $Y_{k+1}^{\theta}=Y_{k}^{\theta}+a(Y_{k}^{\theta},\theta)\Delta t+\beta(Y_{k}^{\theta},\theta)\Delta W_{k}^{Y},$
 $X_{k+1}^{\theta}=X_{k}^{\theta}+X_{k}^{\theta}\Big(r+\pi_{k}^{\top}b(Y_{k}^{\theta},\theta)\Big)\Delta t+X_{k}^{\theta}\,\pi_{k}^{\top}\sigma(Y_{k}^{\theta},\theta)\Delta W_{k},$
-with terminal reward $U(X_{N}^{\theta})$. Define the discrete (pathwise) wealth costate
+with terminal reward $U(X_{N}^{\theta})$. Define the discrete pathwise wealth costate
 $p_{k}^{\mathrm{pw},\theta}:=\frac{\partial}{\partial X_{k}^{\theta}}\,U(X_{N}^{\theta}),\hskip 18.49988ptk=0,\dots,N,$
-which is the same object as (59) (episode indices suppressed and dependence on $\theta$ made explicit). For the projected-control constructions we also consider the additional pathwise objects
-$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}},$
-which correspond to (60).
+and the additional pathwise blocks
+$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}}.$
 Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph.
-The algebraic form of the one-step backward recursion coincides with the deterministic-parameter analysis in huh2025breaking, with the replacements
+The one-step algebraic form of the backward recursion coincides with the deterministic-parameter analysis in Huh et al.
+(2025b), with the replacements
 $\mu\mapsto b(\cdot,\theta),\hskip 18.49988pt\sigma\mapsto\sigma(\cdot,\theta),$
-and with the factor block $(Y^{\theta},\Delta W^{Y})$ handled exactly as in the wealth–factor extension therein.
+and with the factor block handled exactly as in the wealth–factor extension therein.
 All one-step remainder terms are controlled by standard Taylor/Euler estimates with constants uniform in $\theta\in K$.
-Step 3: One-step $L^{2}$ projection and discrete BSDE form (fixed $\theta$).
+Step 3: Conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables (fixed $\theta$).
 Fix $\theta\in K$.
-As in huh2025breaking, take the conditional $L^{2}$-projection of $p_{k+1}^{\mathrm{pw},\theta}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$:
-$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
-where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ conditionally on $\mathcal{G}_{t_{k}}^{\theta}$.
-Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ yields unique projection coefficients $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$.
-Substituting this projection into the BPTT backward recursion from Step 2 yields a canonical discrete BSDE representation for $(p_{k}^{\mathrm{pw},\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ whose drift matches the Euler discretization of the $\theta$-conditional Pontryagin adjoint BSDE associated with (53)–(54).
-The same argument applies to the derivatives $(p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})$: they satisfy linearized discrete backward recursions obtained by differentiating the discrete adjoint equations, hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$.
-Step 4: Passage to continuous time and identification with the PMP costate.
+Let $\Delta W_{k}:=W_{t_{k+1}}-W_{t_{k}}$ and $\Delta W_{k}^{Y}:=W^{Y}_{t_{k+1}}-W^{Y}_{t_{k}}$.
+Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ implies that the conditional $L^{2}$-projection onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ is well defined (and yields unique coefficients).
+(a) One-step projection (defining the discrete martingale coefficients).
+Define $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ by the conditional orthogonality relations (equivalently, the $L^{2}$-projection) such that
+$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
+where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $G_{t_{k}}^{\theta}$.
+(b) Adapted discrete adjoints (the objects compared to PMP). Define the adapted discrete-time costates by conditional expectation:
+$p_{k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{x,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{y,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right].$
+These are $G_{t_{k}}^{\theta}$-measurable and are the discrete-time counterparts of the $\theta$-conditional Pontryagin objects.
+(c) Discrete BSDE form and identification of the driver.
+Substituting the projection in (a) into the one-step BPTT recursion from Step 2 and then taking conditional expectations with respect to $G_{t_{k}}^{\theta}$ yields a canonical discrete-time BSDE representation for the adapted triple $(p_{k}^{\Delta t,\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$.
+The resulting drift coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE associated with (45)–(46) under the fixed policy $\pi_{\varphi}$, up to standard one-step remainder terms of higher order in $\Delta t$.
+The same argument applies to the blocks $(p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$: they satisfy linearized discrete backward recursions (obtained by differentiating the discrete adjoint equations) and hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$.
+All estimates and remainder controls are identical to Huh et al.
+(2025b) once we condition on $\theta$, and the constants can be chosen uniformly on $K$ by the uniform-in-$\theta$ assumptions from Step 1.
+Step 4: Passage to continuous time and convergence to the $\theta$-conditional PMP costate.
 For each fixed $\theta\in K$, the forward SDE and the $\theta$-conditional adjoint BSDE form a standard FBSDE with coefficients parametrized by $\theta$.
-Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under policy $\pi_{\varphi}$, so $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
-Define the piecewise-constant interpolations
-$p_{t}^{\Delta t,\theta}:=p_{k}^{\mathrm{pw},\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\mathrm{pw},\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\mathrm{pw},\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
-By the same stability and convergence arguments as in huh2025breaking (Euler convergence for the forward equation plus discrete-BSDE convergence for the backward equation), we obtain, for each fixed $\theta\in K$,
+Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under $\pi_{\varphi}$, with terminal condition $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
+Define the piecewise-constant interpolations of the adapted discrete variables:
+$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
+Then, by the same stability and convergence arguments as in Huh et al.
+(2025b) (Euler convergence for the forward equation plus convergence of the discrete-BSDE scheme for the backward equation), we obtain, for each fixed $\theta\in K$,
 $\sup_{t\in[0,T]}\mathbb{E}\big[\,|p_{t}^{\Delta t,\theta}-p_{t}^{\theta}|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{x,t}^{\Delta t,\theta}-p_{x,t}^{\theta}\|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{y,t}^{\Delta t,\theta}-p_{y,t}^{\theta}\|^{2}\,\big]\to 0,$
 as $\Delta t\to 0$.
-Because all Lipschitz, growth, ellipticity, and covariance constants were assumed uniform on $K$, the convergence constants can be chosen independently of $\theta\in K$.
+Because all Lipschitz, growth, ellipticity, and block-covariance constants were assumed uniform on $K$, the constants in these convergence bounds can be chosen independently of $\theta\in K$.
 This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$, completing the proof.
 ∎
-Appendix C Auxiliary results for Theorem 3
-C.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
+Appendix D Auxiliary results for Theorem 3
+D.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
 Proposition 1 (Stability of the projection map $(A,G)\mapsto-A^{-1}G$).
@@
 $z\in D$ with $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$, and
-$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\,\big(M+\|\widetilde{G}\|_{L^{\infty}(D)}\big)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (97)
+$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (95)
 Proof.
-Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm. For a matrix-valued function $M:D\to\mathbb{R}^{d\times d}$, write
-$\|M\|_{L^{\infty}(D)}:=\operatorname*{ess\,sup}_{z\in D}\|M(z)\|,\hskip 18.49988pt\|M\|_{L^{2}(\mu)}:=\Big(\int_{D}\|M(z)\|^{2}\,\mu(dz)\Big)^{1/2},$
-and similarly for vector-valued functions.
-Step 1: Invertibility and inverse bound for $\widetilde{A}$. Fix $z\in D$ such that $A(z)$ is invertible (this holds for $\mu$-a.e. $z$). Let $E(z):=\widetilde{A}(z)-A(z)$. By (i) and (iii),
-$\|A(z)^{-1}E(z)\|\leq\|A(z)^{-1}\|\,\|E(z)\|\leq\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq\kappa\cdot\frac{1}{2\kappa}=\frac{1}{2}.$
-Hence $I+A(z)^{-1}E(z)$ is invertible and admits the Neumann-series inverse. Therefore,
-$\widetilde{A}(z)^{-1}=(A(z)+E(z))^{-1}=(I+A(z)^{-1}E(z))^{-1}A(z)^{-1},$
-and
-$\|\widetilde{A}(z)^{-1}\|\leq\frac{1}{1-\|A(z)^{-1}E(z)\|}\,\|A(z)^{-1}\|\leq\frac{1}{1-1/2}\,\kappa=2\kappa.$
-Taking the essential supremum over $z\in D$ yields
-$\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa.$
-Step 2: A pointwise bound for $\widetilde{A}^{-1}-A^{-1}$. For $\mu$-a.e. $z\in D$ where both inverses exist,
-$\widetilde{A}(z)^{-1}-A(z)^{-1}=\widetilde{A}(z)^{-1}\big(A(z)-\widetilde{A}(z)\big)A(z)^{-1}.$
-Thus,
-$\|\widetilde{A}(z)^{-1}-A(z)^{-1}\|\leq\|\widetilde{A}(z)^{-1}\|\,\|\widetilde{A}(z)-A(z)\|\,\|A(z)^{-1}\|\leq(2\kappa)\,\|\widetilde{A}(z)-A(z)\|\,\kappa=2\kappa^{2}\|\widetilde{A}(z)-A(z)\|.$
-Consequently,
-$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\leq 2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
-Step 3: Control error bound. Recall $\pi=-A^{-1}G$ and $\widetilde{\pi}=-\widetilde{A}^{-1}\widetilde{G}$. Then
+Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm.
+Step 1: invertibility and uniform bound for $\widetilde{A}^{-1}$. For $\mu$-a.e. $z$, write
+$\widetilde{A}(z)\;=\;A(z)\Bigl(I+A(z)^{-1}\bigl(\widetilde{A}(z)-A(z)\bigr)\Bigr).$
+By (i) and (iii),
+$\bigl\|A^{-1}(z)\bigl(\widetilde{A}(z)-A(z)\bigr)\bigr\|\;\leq\;\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\;\leq\;\kappa\cdot(2\kappa)^{-1}\;=\;\tfrac{1}{2}.$
+Hence $I+A^{-1}(\widetilde{A}-A)$ is invertible and $\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}\bigr\|\leq(1-\tfrac{1}{2})^{-1}=2$. Therefore $\widetilde{A}(z)$ is invertible and
+$\|\widetilde{A}^{-1}(z)\|\;=\;\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}A^{-1}(z)\bigr\|\;\leq\;2\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa,$
+so $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$.
+Step 2: difference of inverses. Using the identity $\widetilde{A}^{-1}-A^{-1}=\widetilde{A}^{-1}(A-\widetilde{A})A^{-1}$ and Hölder ($L^{\infty}\times L^{2}\times L^{\infty}\to L^{2}$),
+$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\;\leq\;\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\,\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
+Step 3: control error bound. Since $\pi=-A^{-1}G$ and $\widetilde{\pi}=-\widetilde{A}^{-1}\widetilde{G}$,
 $\widetilde{\pi}-\pi=-\widetilde{A}^{-1}(\widetilde{G}-G)-(\widetilde{A}^{-1}-A^{-1})G.$
-Taking $L^{2}(\mu)$ norms and using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) gives
+Taking $L^{2}(\mu)$ norms and applying Hölder ($L^{\infty}\times L^{2}\to L^{2}$) yields
 $\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\,\|G\|_{L^{\infty}(D)}.$
-Using $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$ (Step 1), $\|G\|_{L^{\infty}(D)}\leq M$ (assumption (ii)), and Step 2, we obtain
-$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
-Finally, since $M\leq M+\|\widetilde{G}\|_{L^{\infty}(D)}$, this implies (97). ∎
-C.2 Slab-wise small-gain for the $q$-aggregated projection inputs
+Using Step 1, Step 2, and (ii),
+$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\|\widetilde{A}-A\|_{L^{2}(\mu)},$
+which is (95). ∎
+D.2 Slab-wise small-gain for the $q$-aggregated projection inputs
 Time-slab decomposition.
-In the portfolio problem the working domain $D$ carries a time coordinate; for concreteness, assume
-$D\subset[0,T]\times\mathcal{S},\hskip 18.49988pt\mu(dt,d\xi)=dt\otimes\nu(d\xi),$
-for some reference measure $\nu$ on $\mathcal{S}$. Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
-$D_{k}:=D\cap\big([t_{k-1},t_{k}]\times\mathcal{S}\big),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
+Assume the working domain carries a time coordinate and, for concreteness, $D\subset[0,T]\times S$ and $\mu(dt,d\xi)=dt\otimes\nu(d\xi)$ for some reference measure $\nu$ on $S$.
+Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
+$D_{k}:=D\cap([t_{k-1},t_{k}]\times S),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
 Then $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{K}\|f\|_{k}^{2}$.
 Proposition 2 (Short-time (slab) Lipschitz gain).
-Let $\mathcal{U}$ be a neighborhood of $\pi^{\star}$ in the deployable $\theta$-blind policy class such that for all $\pi\in\mathcal{U}$,
-$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$
-Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
-$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G_{\pi_{1}}^{\mathrm{mix}}-G_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (98)
-for all $\pi_{1},\pi_{2}\in\mathcal{U}$ and constants $\bar{L}_{A},\bar{L}_{G}>0$ that depend only on band data. Define
-$\rho(\tau):=\Big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\Big)\tau^{1/2}.$ (99)
-Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
-$\|T(\pi_{1})-T(\pi_{2})\|_{k}\;\leq\;\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988ptT(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}.$ (100)
-In particular, if the partition is chosen with $\max_{k}\tau_{k}\leq\tau^{\star}$ for some $\tau^{\star}>0$ such that $\rho(\tau^{\star})<1$, then
-$\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1$
-and $T$ is a contraction on every slab with constant at most $\rho_{*}$.
+Let $U$ be a neighborhood of a locally optimal deployable $\theta$-blind policy $\pi^{\star}_{\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in U$,
+$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi}\|_{L^{\infty}(D)}\leq M_{G}$
+for some constants $\kappa,M_{G}>0$. Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
+$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\overline{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi_{1}}-G^{\mathrm{mix}}_{\pi_{2}}\|_{k}\leq\overline{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (96)
+for all $\pi_{1},\pi_{2}\in U$, with constants $\overline{L}_{A},\overline{L}_{G}>0$ depending only on band data.
+Define the (population) projection map $T(\pi):=-A_{\pi}^{-1}G^{\mathrm{mix}}_{\pi}$ and
+$\rho(\tau):=\Bigl(\kappa\,\overline{L}_{G}+\kappa^{2}M_{G}\,\overline{L}_{A}\Bigr)\tau^{1/2}.$
+Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in U$,
+$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}.$
+In particular, if the partition is chosen so that
+$\rho^{\ast}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1,$
+then $T$ is a strict contraction on every slab with constant at most $\rho^{\ast}$.
 Proof.
-Fix $k$ and $\pi_{1},\pi_{2}\in\mathcal{U}$. Write $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G_{\pi_{i}}^{\mathrm{mix}}$. Then
-$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},$
-and
-$A_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
-Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$, we obtain
+Fix $k$ and $\pi_{1},\pi_{2}\in U$. Set $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G^{\mathrm{mix}}_{\pi_{i}}$. Then
+$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},\hskip 18.49988ptA_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
+Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$ yields
 $\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\kappa\|G_{1}-G_{2}\|_{k}+\kappa^{2}M_{G}\|A_{1}-A_{2}\|_{k}.$
-Applying (98) yields (100) with $\rho(\tau_{k})$ as in (99). ∎
-Remark 3 (Verification of the $\tau^{1/2}$ gain and relation to prior slab analyses).
-The $\tau^{1/2}$-gain in (98) is the same short-time parabolic smoothing effect used in our prior PGDPO analysis (see, e.g., huh2025breaking): one combines a Duhamel/semigroup representation of the relevant adjoint/costate objects with Young-type convolution bounds to obtain a factor $\tau^{1/2}$ on each short slab.
-In the present paper, the only additional bookkeeping is that $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ are $q$-aggregated (in particular, linear expectations over $\theta$), which does not alter the semigroup estimates; it only changes constants through coefficient bounds uniform in $\theta$ on the compact parameter set.
-C.3 Proof of Theorem 3
-Setup.
-Let $D$ be the working domain with reference measure $\mu$ and slab decomposition $\{D_{k},\mu_{k},\|\cdot\|_{k}\}_{k=1}^{K}$ as in Appendix C.2.
-For a deployable $\theta$-blind policy $\pi$, write $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ for the $q$-aggregated projection inputs corresponding to the mixed-moment aggregation in (65)–(66).
-Define the projection map
-$T(\pi)(z):=-A_{\pi}(z)^{-1}G_{\pi}^{\mathrm{mix}}(z),\hskip 18.49988ptz\in D,$
-whenever $A_{\pi}(z)$ is invertible.
-Let $\pi^{\mathrm{warm}}$ be the warm-up policy and set
-$A_{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}},\hskip 18.49988ptG_{\mathrm{warm}}:=G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}},\hskip 18.49988pt\pi_{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
-Let $\widehat{A}_{t}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ be the BPTT/Monte Carlo estimators constructed under $\pi^{\mathrm{warm}}$, and denote
-$\widehat{\pi}^{\mathrm{agg,mix}}:=-\widehat{A}_{t}^{-1}\widehat{G}_{t}^{\mathrm{mix}}\hskip 18.49988pt\text{on }D.$
-Proof.
-Step 0 (Fixed-point form of the deployable optimum).
-Let $\pi^{\star}:=\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem.
-By the $q$-aggregated stationarity (Theorem 1), $\pi^{\star}$ satisfies
-$A_{\pi^{\star}}(z)\,\pi^{\star}(z)=-G_{\pi^{\star}}^{\mathrm{mix}}(z)\qquad\text{for }\mu\text{-a.e.\ }z\in D,$
-hence (under invertibility on $D$) it is a fixed point of $T$:
-$\pi^{\star}(z)=T(\pi^{\star})(z)=-A_{\pi^{\star}}(z)^{-1}G_{\pi^{\star}}^{\mathrm{mix}}(z),\hskip 18.49988pt\mu\text{-a.e.\ }z\in D.$
-Step 1 (Triangle decomposition). Add and subtract $\pi_{\mathrm{proj}}$:
-$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}.$ (101)
-Step 2 (Estimation error via Proposition 1). Apply Proposition 1 with
-$(A,G)=(A_{\mathrm{warm}},G_{\mathrm{warm}}),\hskip 18.49988pt(\widetilde{A},\widetilde{G})=(\widehat{A}_{t},\widehat{G}_{t}^{\mathrm{mix}}).$
-Assume the perturbative regime on $D$:
-$\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\qquad\|G_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq M_{G},\qquad\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq(2\kappa)^{-1},\qquad\|\widehat{G}_{t}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$ (102)
-Then (97) yields
-$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}$ $\leq 2\kappa\,\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}+4\kappa^{2}M_{G}\,\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}.$ (103)
-By the definition of $\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta})$ in Theorem 3,
-$\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}+\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}\leq\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
-hence
-$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),\hskip 18.49988ptC_{2}:=2\kappa+4\kappa^{2}M_{G}.$ (104)
-Step 3 (Slab-wise warm-up bias bound).
-Assume $\pi^{\mathrm{warm}},\pi^{\star}\in\mathcal{U}$ and the slab-wise contraction $\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\|\pi_{1}-\pi_{2}\|_{k}$ from Proposition 2.
-Let $\rho_{*}:=\max_{k}\rho(\tau_{k})<1$.
-Since $\pi_{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}=T(\pi^{\star})$, for each slab $D_{k}$ we have
-$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$ (105)
-Step 4 (Residual identity and slab-wise closure). Define the warm-up aggregated stationarity residual (mixed-moment) on $D$ by
-$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}(z):=A_{\mathrm{warm}}(z)\,\pi^{\mathrm{warm}}(z)+G_{\mathrm{warm}}(z),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)}.$
-Also define the slab-wise residual sizes
-$\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k},\hskip 18.49988pt\text{so that }\ (\varepsilon_{\mathrm{warm}}^{\mathrm{mix}})^{2}=\sum_{k=1}^{K}(\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}})^{2}.$
-By construction of $\pi_{\mathrm{proj}}$,
-$\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}=A_{\mathrm{warm}}^{-1}r_{\mathrm{FOC,mix}}^{\mathrm{warm}},$
-hence on each slab
-$\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (106)
-Now combine the triangle inequality on each slab with (105):
-$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$
-Since $\rho_{*}<1$, we close slab-wise:
-$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\frac{\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (107)
-Plugging into (105) gives
-$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (108)
-Summing over slabs yields the global bias bound
-$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$ (109)
-Step 5 (Finish). Combine (101), (104), and (109) to obtain
-$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}+C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
-which is the slab-wise version of (71) (with $\rho_{*}$ in place of a global $C_{1}$). ∎
-Remark 4 (Relation to prior PGDPO slab analyses).
-The closure step above uses the same slab-wise small-gain philosophy as in huh2025breaking: short-time parabolic smoothing yields a contraction on each time slab, and the global bound follows by concatenation.
-The key difference here is that the contraction is applied to the $q$-aggregated projection map $T(\pi)=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$, hence the additional use of the algebraic projection stability (Proposition 1) for the estimator $\widehat{\pi}^{\mathrm{agg,mix}}$.
-Appendix D Implementation details for Section 3
+Applying (96) gives the claim with $\rho(\tau_{k})$. ∎
+Remark 3 (Where the $\tau^{1/2}$ gain comes from; relation to huh2025a).
+The short-time factor $\tau^{1/2}$ in (96) is the same parabolic smoothing mechanism used in our earlier slab-wise PGDPO analysis:
+  • One represents the relevant value-gap / adjoint differences on a short slab via a Duhamel (semigroup) formula, differentiates in space, and applies Young-type convolution bounds.
+  • In mixed norms, the gradient smoothing bound contributes a factor $\tau^{1/2-1/q}$ and the trivial embedding $L^{\infty}_{t}\hookrightarrow L^{q}_{t}$ contributes $\tau^{1/q}$, yielding net $\tau^{1/2}$.
+A full PDE derivation of this short-time contraction philosophy is given in Huh et al.
+(2025b, Appendix B); see also the companion supplementary note to Huh et al.
+(2025b) (“On Proof of Theorem 4 (Policy Gap Bound)”) for a step-by-step slab contraction argument in a closely related setting.
+In the present paper the same smoothing principle is applied not to the costate update map directly, but to the induced $q$-aggregated projection inputs $(A_{\pi},G^{\mathrm{mix}}_{\pi})$.
+D.3 Proof of Theorem 3
+Proof of Theorem 3.
+We keep the notation of the main text.
+Let $\pi^{\mathrm{warm}}$ be the warm-up (deployable, $\theta$-blind) policy and define the population $q$-aggregated projection inputs $(A_{\mathrm{warm}},G_{\mathrm{warm}})$ by
+$A_{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}},\hskip 18.49988ptG_{\mathrm{warm}}:=G^{\mathrm{mix}}_{\pi^{\mathrm{warm}}}.$
+Define the population projected policy
+$\pi^{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
+Let $\widehat{A}$ and $\widehat{G}$ be the stage-2 (BPTT/Monte Carlo) estimators of $A_{\mathrm{warm}}$ and $G_{\mathrm{warm}}$ used to form the deployable stage-2 control
+$\widehat{\pi}_{\mathrm{agg,mix}}:=-\widehat{A}^{-1}\widehat{G}$.
+Let $\pi^{\star}_{\mathrm{blind}}$ denote a locally optimal interior deployable $\theta$-blind policy, so in particular it satisfies the fixed-point (stationarity) identity
+$\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}}).$
+Step 1 (Estimator-to-population projection stability).
+Assume we are in the stable inversion regime (as enforced in practice by ridge/diagnostics) so that the hypotheses of Proposition 1 apply on $D$ with $A=A_{\mathrm{warm}}$ and $G=G_{\mathrm{warm}}$.
+Then there exists a constant $C_{2}>0$ (depending only on the uniform inverse bound $\kappa$ and uniform $L^{\infty}$ bounds on the relevant $G$ terms) such that
+$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}},$ (97)
+where one convenient choice is
+$\delta_{\mathrm{BPTT}}:=\|\widehat{A}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}+\|\widehat{G}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}.$
+[…]
+Step 2 (Slab-wise contraction for the population projection map).
+Assume $\pi^{\mathrm{warm}},\pi^{\star}_{\mathrm{blind}}\in U$ and let $\rho^{\ast}<1$ be the slab contraction constant of Proposition 2.
+Since $\pi^{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}})$, for each slab $D_{k}$ we have
+$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star}_{\mathrm{blind}})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$ (98)
+Step 3 (Residual identity and slab-wise warm-up deviation). Define the population mixed-moment aggregated stationarity residual of the warm-up policy
+$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}.$
+By definition of $\pi^{\mathrm{proj}}$,
+$\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}=\pi^{\mathrm{warm}}-(-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}})=A_{\mathrm{warm}}^{-1}\bigl(A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}}\bigr)=A_{\mathrm{warm}}^{-1}r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$
+Hence, on each slab,
+$\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\leq\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\,\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (99)
+Step 4 (Slab-wise closure). Combine the triangle inequality with (98) and (99):
+$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}+\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$
+Since $\rho^{\ast}<1$, we can close slab-wise:
+$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (100)
+Plugging (100) into (98) yields
+$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (101)
+Step 5 (Global bound and completion). Sum over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k}\|f\|_{k}^{2}$:
+$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}.$
+Finally, combine with (97) and the triangle inequality:
+$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}+\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}},$
+which is the claimed slab-wise residual-based $\theta$-blind policy-gap bound. ∎
+Appendix E Implementation details for Section 3
 This appendix provides reproducible step-by-step templates for the methods in Section 3.
@@
 For compactness we present one template per subsection of Section 3.
-D.1 Stage 1 (PG–DPO) template for Section 3.1
+E.1 Stage 1 (PG–DPO) template for Section 3.1
 Stage 1 performs stochastic gradient ascent on the fixed-$q$ ex–ante objective (48), with latent $\theta\sim q$ sampled inside the simulator while the policy remains $\theta$-blind.
@@
     Checkpoint. Periodically save a warm-up checkpoint $\varphi^{\mathrm{warm}}$ for stage 2 projection.
-D.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
+E.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
 Stage 2 is a post-processing map: given a warm-up $\theta$-blind policy $\pi_{\varphi^{\mathrm{warm}}}$, it estimates Pontryagin sensitivity objects by Monte Carlo and constructs a deployable projected control on a working-domain sample $z\sim\mu$.
@@
     which matches (59).
-D.3 Coupling I: residual/control-variate projection (Section 3.3.1)
+E.3 Coupling I: residual/control-variate projection (Section F.1)
 This subsection records a variance-reduced implementation of the stage 2 map using the residual identity (104).
@@
 Inputs.
-Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section D.2.
+Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section E.2.
 Template (statewise residual projection; mixed-moment aggregation).
@@
     $\widehat{\pi}^{\mathrm{agg,mix}}(z)\;\leftarrow\;\pi_{\varphi^{\mathrm{warm}}}(z)-\big(\widehat{A}_{t}(z)\big)^{-1}\widehat{r}_{\mathrm{FOC}}(z).$
-D.4 Coupling II: interactive distillation (Section 3.3.2)
+E.4 Coupling II: interactive distillation (Section F.2)
 This subsection records an implementation template for interactive distillation: the projected output from stage 2 is used as a teacher signal during stage 1 training via the mixed objective (106).
@@
     1. (a)
-      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section D.1.
+      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section E.1.
     2. (b)
@@
       4. iv.
-        Optionally filter states using diagnostics (Section D.5) and update the buffer:
+        Optionally filter states using diagnostics (Section E.5) and update the buffer:
         $\mathcal{B}\ \leftarrow\ \{(z_{j},\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j}))\}_{j=1}^{M_{z}}\ \ \text{(after filtering)}.$
@@
         Apply a gradient step to minimize the proximity term $\|\pi_{\varphi}(z)-\mathrm{stopgrad}(\pi^{\mathrm{teach}})\|^{2}$ with coefficient $\lambda(n)$, consistent with (106).
-D.5 Engineering notes and stabilizers
+E.5 Engineering notes and stabilizers
 This subsection collects practical stabilizers that we found helpful for reliable training and projection in high dimensions.
@@
 As a last-resort safety net, a mild log-wealth clamp can prevent overflow, but it should be used conservatively and monitored, since frequent clamping may distort higher-order sensitivities.
-Appendix E Stage 2 projection diagnostics
+Appendix F Coupling stage 1 and stage 2: residual projection and interactive distillation
+We keep the ex–ante objective (48) and the $\theta$-blind deployability constraint throughout.
+Stage 2 is not a separate optimization problem: it is a (warm-started) post-processing map that reuses the current stage 1 policy as a warm-up control, estimates the projection ingredients under this warm-up policy, and then applies the $q$-aggregated Pontryagin projection as a deterministic transformation of the policy.
+Concretely, stage 2 takes $\pi^{\mathrm{warm}}$ as input, runs BPTT/Monte Carlo under frozen $\theta\sim q$ to estimate the adjoint blocks required for the Hamiltonian gradient, aggregates them across $\theta$, and outputs a single deployable $\theta$-blind rule $\widehat{\pi}^{\mathrm{agg,mix}}$ (Section 3.2).
+This appendix records two couplings between the two stages, each with a distinct role:
+  • Residual (control-variate) projection.
+We compute the same projected rule in a residual form around the warm-up policy.
+This is algebraically equivalent to the direct projection, but typically reduces Monte Carlo variance and improves numerical stability in high dimensions.
+  • Interactive distillation (amortized projection).
+We use the projected rule as a teacher signal at intermittent refresh times.
+Beyond accelerating/stabilizing training, distillation amortizes the cost of projection: stage 2 can be accurate but Monte-Carlo intensive, whereas a distilled student policy approximates the projected rule with a single forward pass at stage 1 inference cost.
+F.1 Control-variate (residual) form of the projected rule
+Recall the mixed-moment projected rule (59).
+In high dimensions, Monte Carlo noise in the projection inputs can be non-negligible, and solving a linear system with $\widehat{A}_{t}$ can amplify this noise.
+A convenient stabilization is to compute the same projected rule in a residual (control-variate) form around the warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$.
+Define the $\theta$-conditional residual (under frozen-$\theta$ simulations)
+$\widehat{r}_{\mathrm{FOC}}^{\theta}(t,x,y):=\widehat{A}_{t}^{\theta}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\theta}(t,x,y),$ (102)
+and the aggregated residual (the quantity we actually solve against)
+$\widehat{r}_{\mathrm{FOC}}(t,x,y):=\widehat{A}_{t}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\mathrm{mix}}(t,x,y).$ (103)
+Whenever $\widehat{A}_{t}(t,x,y)$ is (numerically) invertible, the projected rule admits the identity
+$\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y)=\pi_{\varphi^{\mathrm{warm}}}(t,x,y)-\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{r}_{\mathrm{FOC}}(t,x,y),$ (104)
+which is an algebraic rewriting of (59) (hence it does not change the target).
+Practical value (variance reduction and stability).
+When the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small.
+Moreover, the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and often partially cancel, improving concentration of $\widehat{r}_{\mathrm{FOC}}$.
+In implementation, we combine (104) with the inversion stabilizers of Appendix E.5 (e.g.
+ridge regularization and diagnostic-based fallback to $\pi_{\varphi^{\mathrm{warm}}}$ on unreliable states).
+F.2 Interactive distillation: projection-guided training and amortized deployment
+Let $\pi_{\varphi}$ be the trainable stage 1 policy network.
+At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher.
+This coupling serves two purposes:
+  • Training aid. Projection-guided targets can stabilize and accelerate stage 1 optimization, especially once the warm-up policy enters a locally stable regime.
+  • Amortization. Stage 2 projection can be accurate but Monte-Carlo intensive; distillation compresses it into a single forward pass for deployment.
+Teacher definition (computed by stage 2, treated as fixed).
+In residual form (104), the teacher is the $\theta$-blind map
+$\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y):=\pi_{\varphi^{-}}(t,x,y)-\big(\widehat{A}_{t}^{\varphi^{-}}(t,x,y)\big)^{-1}\,\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}(t,x,y),$ (105)
+where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy.
+In practice, the inversion in (105) is implemented with the same stabilizers as in Section 3.2 (ridge/diagnostics/residual form); for notational simplicity we keep the formal inverse.
+Student objective (hybrid PG–DPO + distillation loss).
+We train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to the teacher on the working domain:
+$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(t,x,y)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y)\big)\big\|^{2}\Big],$ (106)
+where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance.
+The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2: once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
+Refresh schedule, annealing, and diagnostic gating (implementation).
+In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps. A typical and robust configuration is:
+  • Two-time-scale refresh. Hold $\varphi^{-}$ fixed for $K_{\mathrm{refresh}}$ stage 1 updates of $\varphi$ under (106), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher via stage 2.
+  • Anneal $\lambda$.
+Start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain indicate that the stage 2 map is reliable.
+Concretely, we anneal $\lambda$ toward a target value $\lambda_{\max}$ once the warm-up residual $\|\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)}$ and curvature diagnostics remain stable over several refresh cycles.
+  • Adaptive teacher selection (gating).
+To avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we optionally apply the teacher loss only on states where projection diagnostics certify reliability (e.g.
+bounded condition number of $\widehat{A}_{t}$, eigenvalue/denominator lower bounds, and moderate residual magnitude), falling back to pure PG–DPO elsewhere.
+Implementation details, including the precise diagnostics and stabilizers, are deferred to Appendix E.
+Deployment.
+After training, the deployed controller is simply the student network $\pi_{\varphi}$, which approximates the stage 2 projected rule with a single forward pass, without repeatedly running BPTT/Monte Carlo projection at inference time.
+Appendix G Stage 2 projection diagnostics
 We report Stage 2 diagnostic statistics as a visual supplement to Section 4.2. Each figure summarizes the same tail-median protocol and layout; see captions for definitions and interpretation.
@@
 This helps rule out “catastrophic inversion” as the primary driver of degradation; instead, residual growth and curvature mismatch (Fig.
-5) provide a more consistent explanation in misaligned/high-$d$ regimes.
+6) provide a more consistent explanation in misaligned/high-$d$ regimes.
 Figure 6: Stage 2 curvature-consistency statistic $\kappa$.
@@
 Near-zero bad-sign fractions across most regimes suggest that the projection typically operates in a locally well-behaved region and that failures are not dominated by sign flips, supporting the main-text conclusion that misalignment primarily increases residual/costate mismatch rather than inducing widespread concavity violations.
-Appendix F Supplementary decomposition diagnostics for Section 5
+Appendix H Supplementary decomposition diagnostics for Section 5
 Tables 23 report Stage 2 decomposition diagnostics at $t=0$.

### v3

@@
 Abstract
-We study continuous-time CRRA portfolio choice in diffusion markets with uncertain estimated coefficients.
-Nature draws a latent parameter from a given distribution and keeps it fixed; the investor cannot observe this parameter and must commit to a parameter-blind policy maximizing an ex-ante objective.
-We treat the uncertainty distribution as an inference-agnostic sampling input.
-We develop a simulation-only two-stage solver.
-Stage 1 extends Pontryagin-Guided Direct Policy Optimization (PG-DPO) by sampling parameters internally and computing gradients via backpropagation through time.
-Stage 2 performs an aggregated Pontryagin projection: it aggregates costates across the parameter distribution to enforce a deployable stationarity condition, yielding a structured correction amortized via interactive distillation.
-We prove a uniform conditional BPTT-PMP correspondence and a residual-based policy-gap bound with explicit error terms.
-Experiments on high-dimensional Gaussian drift and factor-driven benchmarks show that projection stabilizes learning and accurately recovers analytic references, while a model-free PPO baseline remains far from the targets.
+We study continuous-time CRRA portfolio choice in diffusion markets with estimated and hence uncertain coefficients.
+Nature draws a latent parameter $\theta\sim q$ at time 0 and keeps it fixed; the investor never observes $\theta$ and must commit to a single $\theta$-blind policy maximizing an ex–ante objective, treating $q$ as a decision-time input.
+We propose a simulation-only two-stage solver.
+Stage 1 (DPO) performs BPTT-based stochastic gradient ascent through an Euler simulator while sampling $\theta$ only inside the simulator.
+Stage 2 (Pontryagin projection) aggregates costate blocks across $\theta\sim q$ and enforces the $q$-aggregated stationarity condition within the deployable class; the resulting correction can be amortized via interactive distillation.
+We refer to the full Stage 1+Stage 2 pipeline as PG–DPO.
+We prove a uniform conditional BPTT–PMP correspondence and a residual-based policy-gap bound with explicit discretization and Monte Carlo error terms.
+Experiments on high-dimensional Gaussian drift-uncertainty and factor-driven benchmarks show that projection stabilizes learning and accurately recovers analytic decision-time references, while a model-free PPO baseline remains far from the targets.
 Keywords: Portfolio choice; Parameter uncertainty; Deep learning; Pontryagin principle; Stochastic control
 1 Introduction
-A central problem in quantitative finance is to allocate wealth dynamically across many risky assets in continuous time.
-In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and closed-form optimal portfolios follow from HJB analysis; see, for example, Merton (1969, 1971).
-In realistic markets, however, key inputs—especially expected returns—must be estimated from finite samples, often in high dimensions and under regularization/model selection.
-Empirically, return predictability is unstable and many predictors deliver limited out-of-sample gains (e.g., Goyal and Welch, 2008; Campbell and Thompson, 2008).
-These considerations motivate separating diffusion risk (Brownian noise conditional on parameters) from decision-time estimation error (uncertainty in estimated coefficients), a theme that also underlies intertemporal hedging and more conservative allocations in models with predictability/learning (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Campbell and Viceira, 2002; Xia, 2001; Maenhout, 2004).
-We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, and an external estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$.
-We treat $q$ as an input object that summarizes decision-time estimation error about $\theta$.
-Concretely, $q$ may be obtained from resampling-based approximations (e.g., the bootstrap) (e.g., Efron, 1979; Efron and Tibshirani, 1994), from subsample aggregation and model-selection variability (e.g., bagging), or from asymptotic (sandwich/normal) approximations of an estimator’s sampling distribution (van der Vaart, 1998).
-Our goal is not to revisit inference, but to optimize decisions given this uncertainty description.
-Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside a simulator; we do not assume a closed-form density, conjugacy, or a particular parametric family.
-We seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$.
-A key modeling choice is that $\theta$ is latent: Nature draws a fixed but unobserved $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed on $[0,T]$.
-The investor knows $q$ but does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy.
-We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a filtering/belief process.
-This is a one-shot decision-time problem: $q$ is treated as fixed over the trading horizon, and we do not model or hedge subsequent updates of the uncertainty description.
-Belief-state control (optimal learning/hedging of parameter uncertainty) is conceptually important, but it defines a different and typically far more demanding problem than computing a single deployable $\theta$-blind rule from a fixed $q$ (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
-The $\theta$-blind deployability constraint also changes what a first-order optimality condition means.
-Under full information, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition, hence a $\theta$-conditional feedback map that is infeasible when $\theta$ is latent.
-Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind, and the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the controlled state process (e.g., Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009).
-In our portfolio setting, $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$, so $q$-aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projection map.
-This condition is inference-agnostic: it depends on $q$ only through its role as the ex–ante mixing law.
-These features place the problem beyond the reach of classical dynamic programming in the targeted high-dimensional regime.
-Even in moderate dimensions, HJB solvers face the curse of dimensionality (e.g., Bellman, 1961; Kushner and Dupuis, 2001).
-Although deep PDE (e.g., Raissi et al., 2019; Sirignano and Spiliopoulos, 2018) and BSDE methods (e.g., Han et al., 2018; Beck et al., 2019) reduce grid reliance, nonlinear portfolio HJBs remain numerically delicate.
-Furthermore, belief augmentation renders the problem infinite-dimensional (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
-Accordingly, we exclude value-function solvers as practical baselines under our $\theta$-blind interface, validating instead on closed-form benchmarks and simulation-only model-free RL (e.g., PPO).
-Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (Huh et al., 2025b, a).
-PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and uses backpropagation through time (BPTT) to compute exact discrete-time gradients of terminal utility.
-Pathwise sensitivities computed by BPTT coincide with stochastic costates in PMP, reflecting the classical adjoint/backpropagation correspondence (e.g., LeCun, 1988; Yong and Zhou, 1999).
-In the latent-parameter setting, we sample $\theta\sim q$ inside the simulator (fixing it along each trajectory), while the policy depends only on observable states.
-To stabilize learning under heterogeneous $\theta$ draws, we develop uncertainty-aware projected P–PGDPO: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated PMP first-order condition, producing a robust deployable $\theta$-blind rule that can be amortized into a fast policy.
-Empirically, in high-dimensional static drift experiments, the projected pipeline improves decision-time accuracy over end-to-end learning, stabilizing aligned regimes.
-In factor-driven markets where correlation induces hedging demand, it recovers analytic references under the $\theta$-blind restriction, whereas a PPO baseline remains far off.
-Theoretically, we establish a residual-based policy-gap bound: under mild stability, a small aggregated residual implies the policy is close to a locally optimal deployable rule, up to discretization and Monte Carlo errors.
-Our contributions are threefold.
-(i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated PMP first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$.
-(ii) We extend PG–DPO to this setting by sampling $\theta$ only inside the simulator and using BPTT to compute exact discrete-time gradients and pathwise sensitivities, and we establish a conditional BPTT–PMP correspondence uniform over $\theta$ on compact subsets of $\Theta$.
-(iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization.
-The remainder of the paper is organized as follows.
-Section 2 formulates the fixed-$q$ ex–ante problem and derives $\theta$-conditional versus $q$-aggregated PMP conditions together with Gaussian decision-time reference models.
-Section 3 develops PG–DPO and P–PGDPO for latent $\theta$ and proves the policy-gap bound.
-Sections 4 and 5 report numerical experiments, and the appendix collects technical proofs and implementation details.
+Dynamic portfolio choice is typically studied under the assumption that investment opportunities are known, yielding Merton-type closed-form rules and their intertemporal extensions (Markowitz, 1952; Samuelson, 1969; Merton, 1969, 1971; Karatzas and Shreve, 1998; Campbell and Viceira, 2002).
+In practice, expected returns and other inputs are estimated from finite samples, are unstable across specifications, and can be highly uncertain at the time a decision is made (Goyal and Welch, 2008; Campbell and Thompson, 2008; Lettau and Van Nieuwerburgh, 2008; Rapach et al., 2010).
+A classical response is learning and belief-state control: investors update posteriors over time and hedge against future revisions in investment opportunities (Kandel and Stambaugh, 1996; Barberis, 2000; Xia, 2001; Brandt et al., 2005).
+While conceptually appealing, belief-state formulations lead to high-dimensional (sometimes infinite-dimensional) state spaces and are difficult to implement in realistic multi-asset settings (Lakner, 1995; Bensoussan and van Schuppen, 1985; Björk et al., 2010; Pham, 2009).
+This paper takes a complementary decision-time perspective centered on deployability:
+Given an externally supplied description of parameter uncertainty at the decision time, what constitutes an optimal deployable portfolio rule when the true parameter is latent and cannot be conditioned on during trading?
+We treat estimation risk as an exogenous decision-time uncertainty object—a fixed law $q(d\theta)$ over market parameters produced by an external pipeline (e.g., bootstrap, subsampling, or asymptotic approximations) (Efron, 1979; Efron and Tibshirani, 1994; Breiman, 1996; van der Vaart, 1998).
+Nature draws $\theta\sim q$ at time 0 and keeps it fixed over the horizon, but the investor never observes $\theta$ and must commit to a single $\theta$-blind policy.
+This information structure changes the relevant first-order condition.
+Full-information policies such as the Merton rule are defined conditional on the realized parameter and are infeasible when $\theta$ is latent (Merton, 1969, 1971; Campbell and Viceira, 2002).
+More fundamentally, optimality must be enforced within the deployable class: admissible perturbations are $\theta$-blind, so stationarity requires the $q$-expectation of the Hamiltonian gradient to vanish along the controlled state process.
+In portfolio problems, this $q$-aggregated Pontryagin condition reduces to a statewise linear system whose solution defines a canonical deployable projected rule.
+Averaging full-information optima does not, in general, produce such a rule; optimality does not commute with averaging (Pástor, 2000; Avramov, 2002; Cremers, 2002; Brown, 1978).
+These observations lead to three contributions.
+First, we derive the $q$-aggregated Pontryagin first-order condition for the fixed-$q$, $\theta$-blind problem and show how it induces a deployable projection map in portfolio settings (Pontryagin, 2018; Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009).
+Second, we propose a simulation-only two-stage solver that avoids dynamic programming, which is infeasible in the high-dimensional uncertain regimes of interest (Bellman, 1961; Kushner and Dupuis, 2001; Han et al., 2018; Beck et al., 2019).
+Stage 1 (DPO) trains a $\theta$-blind policy by stochastic gradient ascent using exact discrete-time gradients computed by backpropagation through time through an Euler simulator, while sampling $\theta$ only inside the simulator.
+Stage 2 applies a Pontryagin projection by aggregating costate objects across $\theta\sim q$ and projecting the warm policy onto the deployable $q$-aggregated stationarity condition, yielding a structured correction that can be amortized via interactive distillation (Werbos, 1990; Williams and Zipser, 1989; Pardoux and Peng, 1990; Ma and Yong, 1999; Hinton et al., 2015; Jin et al., 2020).
+We refer to the full Stage 1+Stage 2 pipeline as PG–DPO.
+Third, we provide a theory that matches this computational design: we prove a conditional BPTT–PMP correspondence that is uniform over compact parameter sets and includes the costate blocks required by the projection step, and we establish a residual-based policy-gap bound with explicit discretization and Monte Carlo error terms.
+In Gaussian benchmarks, the projected rule recovers familiar shrinkage and horizon-dependent hedging effects connected to classical return-predictability insights (Kandel and Stambaugh, 1996; Campbell and Viceira, 1999; Kim and Omberg, 1996; Lynch, 2001).
+Empirically, in high-dimensional settings the Stage 2 projection stabilizes learning and accurately recovers analytic decision-time references, while a model-free PPO baseline remains far from the targets under the same deployability restriction (Schulman et al., 2017; Wang et al., 2020; Jia and Zhou, 2022).
+Overall, the paper reframes portfolio choice under parameter uncertainty as a problem of commitment under imperfect information at the decision time, rather than explicit belief-state learning.
+Section 2 formulates the fixed-$q$ ex–ante problem, contrasts $\theta$-conditional and $q$-aggregated PMP conditions, and presents Gaussian decision-time reference models.
+Section 3 develops Stage 1 DPO and Stage 2 Pontryagin projection (PG–DPO) for latent $\theta$ and proves the policy-gap bound.
+Sections 4 and 5 report numerical experiments, and the appendix collects proofs and implementation details.
 2 Dynamic Portfolio Choice in Estimated Diffusion Markets with Latent Parameter Uncertainty
@@
 $dY_{t}$ $=a\big(Y_{t},\theta\big)\,dt+\beta\big(Y_{t},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m},$ (2)
-where $W$ and $W^{Y}$ are Brownian motions (possibly of different dimension) that may be instantaneously correlated. We write the instantaneous covariance and return–factor cross-covariance as
+where $W$ and $W^{Y}$ are Brownian motions (possibly of different dimension) that may be instantaneously correlated.
+For bookkeeping, let $W$ be $d_{W}$-dimensional, $W^{Y}$ be $d_{Y}$-dimensional, and write
+$d\langle W,W^{Y}\rangle_{t}=\rho\,dt,\hskip 18.49988pt\rho\in\mathbb{R}^{d_{W}\times d_{Y}}.$ (3)
+We write the instantaneous covariance and return–factor cross-covariance as
 $\Sigma(y,\theta):=\sigma(y,\theta)\sigma(y,\theta)^{\top},\hskip 18.49988pt\Sigma_{SY}(y,\theta):=\sigma(y,\theta)\,\rho\,\beta(y,\theta)^{\top},$ (4)
-where $\rho$ is defined by $d\langle W,W^{Y}\rangle_{t}=\rho\,dt$. Thus $\Sigma(y,\theta)\in\mathbb{R}^{d\times d}$ and $\Sigma_{SY}(y,\theta)\in\mathbb{R}^{d\times m}$.
+so that $\Sigma(y,\theta)\in\mathbb{R}^{d\times d}$ and $\Sigma_{SY}(y,\theta)\in\mathbb{R}^{d\times m}$.
 Uncertainty law $q(d\theta)$ and information structure.
@@
 Remark 1 (Latent parameter, observability, and admissible controls).
+All processes are defined on a filtered probability space $(\Omega,\mathcal{F},\mathbb{F}=(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$ satisfying the usual conditions and carrying $(W,W^{Y},\theta)$.
 We interpret $\theta$ as a latent (unobserved) market parameter: at the (shifted) decision time 0, Nature draws an $\mathcal{F}_{0}$-measurable random variable $\theta\sim q$ (independent of the Brownian drivers) and keeps it fixed over $[0,T]$.
 The investor knows $q$ but does not observe the realized $\theta$, so deployable portfolio rules cannot take $\theta$ as an input.
-We consider the observable market filtration
+We consider the observable market filtration (a subfiltration of $\mathbb{F}$)
 $\mathcal{F}_{t}^{\mathrm{obs}}:=\sigma\{(S_{s},Y_{s}):0\leq s\leq t\},\hskip 18.49988pt0\leq t\leq T,$ (6)
@@
 Whenever we display $\theta$-conditional (full-information) controls or sensitivity objects, they are computed under frozen-$\theta$ simulations and are used only for offline diagnostics; the deployed policy class and the learned policy remain $\theta$-blind.
+Notation.
+Throughout Section 2, $\rho$ denotes the Brownian correlation matrix in (3) (and (33)); $T$ denotes the terminal time/horizon; and $K$ denotes the OU mean-reversion matrix in (31).
+We do not reuse $\rho$, $T$, or $K$ for unrelated quantities elsewhere.
 Wealth dynamics and admissibility (given $\theta$).
@@
 Approximating the outer expectation in (9) amounts to sampling $\theta\sim q$ inside the simulator (once per trajectory or once per update), running (1)–(2) under that frozen draw, and updating a $\theta$-blind feedback policy to perform well on average over such draws.
-This is the setting targeted by the simulation-based PG–DPO and P–PGDPO methods developed in Section 3.
+This is the setting targeted by the simulation-based DPO and PG–DPO methods developed in Section 3.
 Decision-time commitment and (optional) plug-in replanning.
@@
 If the decision is revisited at a later calendar time $t_{0}\in(0,T)$, we interpret this as starting a new decision-time problem on the remaining horizon with an externally supplied law $q_{t_{0}}$.
-Such a law may coincide with a static input (Section 2.3.1), may arise as a model-implied predictive (prior) propagation of an initial uncertainty through factor dynamics without conditioning on new observations (Section 2.3.2), or may be produced by an external filtering/estimation routine that outputs a time-indexed uncertainty description (Appendix A).
+Such a law may coincide with a static input (Section 2.3.1), may arise as a model-implied predictive (prior) propagation of an initial uncertainty through factor dynamics without conditioning on new observations (Section 2.3.2), or may be produced by an external filtering/estimation routine that outputs a time-indexed uncertainty description (Appendix B).
 Our model and algorithms are conditional on whichever $q$ is supplied at the decision time and always freeze that input over the ensuing horizon.
@@
 In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (4).
-For clarity, we define the control Hamiltonian map as a function of generic “sensitivity” inputs $(\varphi_{x},\varphi_{xx},\varphi_{xy})$:
-$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,\varphi_{x},\varphi_{xx},\varphi_{xy}):=x\,\pi^{\top}b(y,\theta)\,\varphi_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,\varphi_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,\varphi_{xy},$ (10)
-where $\varphi_{x},\varphi_{xx}\in\mathbb{R}$ and $\varphi_{xy}\in\mathbb{R}^{m}$.
-In the full-information HJB interpretation, we evaluate (10) at $(\varphi_{x},\varphi_{xx},\varphi_{xy})=(V_{x},V_{xx},V_{xy})$.
+For clarity, we define the control Hamiltonian map as a function of generic “sensitivity” inputs $(\psi_{x},\psi_{xx},\psi_{xy})$:
+$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,\psi_{x},\psi_{xx},\psi_{xy}):=x\,\pi^{\top}b(y,\theta)\,\psi_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,\psi_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,\psi_{xy}.$ (11)
+In the full-information HJB interpretation, we evaluate (11) at $(\psi_{x},\psi_{xx},\psi_{xy})=(V_{x},V_{xx},V_{xy})$.
 In the Pontryagin interpretation used below, the same affine map is evaluated at the corresponding $\theta$-conditional adjoint/sensitivity objects (denoted $(p^{\theta},p_{x}^{\theta},p_{y}^{\theta})$ along a trajectory); in smooth Markov regimes these objects coincide with the indicated value derivatives.
@@
 To see why ex–ante aggregation enters the first-order condition, take any $\theta$-blind admissible perturbation $h=\{h_{t}\}_{t\in[0,T]}$ that is progressively measurable with respect to the observation filtration $(\mathcal{F}_{t}^{\mathrm{obs}})$ and square-integrable, and define $\pi^{\varepsilon}:=\pi+\varepsilon h$ for small $\varepsilon$.
-For each fixed $\theta$, the stochastic maximum principle yields the first-variation identity
+For each fixed $\theta$, define $J^{\theta}(\pi):=\mathbb{E}[U(X_{T}^{\pi})\mid\theta]$.
+The stochastic maximum principle yields the first-variation identity
 $\left.\frac{d}{d\varepsilon}J^{\theta}(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)^{\top}h_{t}\,dt\,\Big|\,\theta\right],$ (14)
 where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity/adjoint objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market.
-(Equivalently, (13) evaluates the same affine map $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ defined by (10), with $(V_{x},V_{xx},V_{xy})$ replaced by the corresponding Pontryagin objects along the controlled trajectory.) Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
+(Equivalently, (14) evaluates the same affine map $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ defined by (11), with $(V_{x},V_{xx},V_{xy})$ replaced by the corresponding Pontryagin objects along the controlled trajectory.)
+Frozen-$\theta$ adjoint BSDE and the blocks used in projection.
+To make the $\theta$-conditional Pontryagin objects in (14) concrete, we briefly record the backward (stochastic) equation they satisfy for a fixed policy in a frozen-$\theta$ market.
+Fix a $\theta$-blind feedback policy $\pi\in\mathcal{A}^{\mathrm{fb}}$ and condition on a frozen parameter $\theta$.
+Let $\mathcal{G}_{t}^{\theta}:=\sigma\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)$ denote the simulator filtration (with the usual augmentation).
+Under standard hypotheses ensuring well-posedness of the stochastic maximum principle (summarized in Assumption 1), there exists an adapted triple $(p^{\theta},z^{\theta},\tilde{z}^{\theta})$ solving the $\theta$-conditional adjoint BSDE
+$p_{t}^{\theta}=U^{\prime}\!\big(X_{T}^{\pi,\theta}\big)+\int_{t}^{T}f_{\theta}\!\big(s,X_{s}^{\pi,\theta},Y_{s}^{\theta},\pi_{s};\,p_{s}^{\theta},z_{s}^{\theta},\tilde{z}_{s}^{\theta}\big)\,ds-\int_{t}^{T}z_{s}^{\theta}\cdot dW_{s}-\int_{t}^{T}\tilde{z}_{s}^{\theta}\cdot dW_{s}^{Y},$ (15)
+for a driver $f_{\theta}$ determined by the frozen-$\theta$ dynamics and the fixed policy $\pi$ (e.g.
+Yong and Zhou, 1999; Pham, 2009).
+In a smooth Markov regime one may identify $p_{t}^{\theta}=V_{x}^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$ along the controlled state and interpret the blocks entering the portfolio Hamiltonian gradient (11) as
+$p_{x,t}^{\theta}:=V_{xx}^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta}),\hskip 18.49988ptp_{y,t}^{\theta}:=V_{xy}^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta}).$ (16)
+Equivalently, one may work directly with the second-order stochastic maximum principle, where $(p_{x,t}^{\theta},p_{y,t}^{\theta})$ coincide with the relevant blocks of the second adjoint.
+Assumption 1 (Frozen-$\theta$ Markov/BSDE regime (used in Theorems 1 and 2)).
+Fix a compact parameter set $\Theta_{0}\subset\Theta$ and a working domain $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$.
+For every $\theta\in\Theta_{0}$ and every $\theta$-blind feedback policy $\pi$ considered in the paper, assume:
+  1. (A1)
+    Forward well-posedness and moments. The frozen-$\theta$ forward system (8)–(2) admits a unique strong solution under $\pi$, and for some $p\geq 2$,
+    $\sup_{\theta\in\Theta_{0}}\mathbb{E}\Big[\sup_{t\in[0,T]}\big(|X_{t}^{\pi,\theta}|^{p}+\|Y_{t}^{\theta}\|^{p}\big)\Big]<\infty.$
+  2. (A2)
+    Positivity and utility integrability. Wealth stays strictly positive a.s., and $U$ is $C^{3}$ on $(0,\infty)$ with
+    $\sup_{\theta\in\Theta_{0}}\mathbb{E}\Big[|U^{\prime}(X_{T}^{\pi,\theta})|^{2}+|U^{\prime\prime}(X_{T}^{\pi,\theta})|^{2}\Big]<\infty.$
+    (For CRRA, this entails suitable negative-moment bounds for $X_{T}^{\pi,\theta}$.)
+  3. (A3)
+    Adjoint BSDE and smooth blocks.
+For each $\theta\in\Theta_{0}$, the adjoint BSDE (15) is well posed in $L^{2}$ and admits a Markov decoupling field $u^{\theta}$ on $\mathcal{D}$ of class $C^{1,2}$ in $(t,x,y)$ such that $p_{t}^{\theta}=u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, $p_{x,t}^{\theta}=\partial_{x}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, and $p_{y,t}^{\theta}=\partial_{y}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, with square-integrable bounds.
+  4. (A4)
+    Nondegenerate driving noise for conditional projections.
+The block covariance matrix $\Gamma:=\begin{pmatrix}I_{d_{W}}&\rho\\ \rho^{\top}&I_{d_{Y}}\end{pmatrix}$ is positive definite (equivalently, the conditional $L^{2}$ projections used below are taken onto a nondegenerate basis of the driving increments).
+  5. (A5)
+    Uniformity on compacts. All Lipschitz/growth/nondegeneracy/smoothness constants in (A1)–(A4) can be chosen uniformly for $\theta\in\Theta_{0}$.
+Remark 2 (What we assume vs. what we prove).
+Assumption 1 collects the analytic conditions required to define the Pontryagin/BSDE objects and to justify the interchanges of limits and conditional expectations used later (notably in the BPTT–BSDE correspondence of Theorem 2).
+Establishing such smooth decoupling fields and uniform bounds for general high-dimensional coefficients is nontrivial; we do not attempt to do so here.
+Instead, our contributions are conditional: given a regime where the frozen-$\theta$ SMP/BSDE objects are well behaved (uniformly on compacts), we show that BPTT recovers them in the small-step limit and that the resulting projection yields quantitative residual-based policy-gap control.
+Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
 $\left.\frac{d}{d\varepsilon}J(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\mathbb{E}_{\theta\sim q}\!\Big[\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)\Big]^{\top}h_{t}\,dt\right].$ (17)
@@
 Consider the fixed-$q$ ex–ante objective (9) over the $\theta$-blind Markov feedback class $\mathcal{A}^{\mathrm{fb}}$.
-Assume standard smoothness/integrability conditions ensuring validity of first variations within $\mathcal{A}^{\mathrm{fb}}$ and existence of the associated $\theta$-conditional Pontryagin objects.
+Assume Assumption 1 holds (in particular, the first-variation identity (14) is valid within $\mathcal{A}^{\mathrm{fb}}$ and the associated $\theta$-conditional Pontryagin/BSDE objects exist).
 If $\pi^{\star,\mathrm{blind}}$ is a locally optimal interior policy in $\mathcal{A}^{\mathrm{fb}}$, then (18) holds.
@@
 Proof sketch.
-The conditional first-variation identity (13) is standard for diffusion control under a fixed parameter $\theta$ (e.g.
+Fix an interior locally optimal $\theta$-blind feedback $\pi^{\star,\mathrm{blind}}\in\mathcal{A}^{\mathrm{fb}}$.
+For any $\theta$-blind admissible perturbation $h=\{h_{t}\}_{t\in[0,T]}$ (progressively measurable w.r.t.
+$(\mathcal{F}_{t}^{\mathrm{obs}})$ and square-integrable), set $\pi^{\varepsilon}:=\pi^{\star,\mathrm{blind}}+\varepsilon h$.
+  1. 1.
+    First variation and $q$-aggregation.
+For each fixed $\theta$, the $\theta$-conditional first-variation identity (14) is standard for diffusion control under a frozen parameter (e.g.
 Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009).
-Taking the outer expectation over $\theta\sim q$ yields (14).
-Since $h$ is an arbitrary $\theta$-blind admissible perturbation, vanishing of the first variation at an interior optimum implies (15).
-For the quadratic portfolio Hamiltonian (10), substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ and introducing (16)–(18) yields the linear system (19) and the projected form (20) whenever $A_{t}$ is invertible.
-∎
+Taking the outer expectation over $\theta\sim q$ and using Fubini/Tonelli under the standing integrability assumptions yields (17).
+Remark.
+This step implements the fixed-$q$ decision-time commitment model: $\theta$ is sampled only inside the simulator to form the ex–ante objective $J(\cdot)$, while admissible perturbations remain $\theta$-blind.
+We do not formulate the belief-state/partial-information PMP, where stationarity would be expressed in terms of time-varying posteriors $q_{t}$ (cf.
+Remark 3).
+  2. 2.
+    Vanishing first variation $\Rightarrow$ aggregated stationarity. Since $\pi^{\star,\mathrm{blind}}$ is an (interior) local maximizer, the first variation must vanish:
+    $\left.\frac{d}{d\varepsilon}J(\pi^{\varepsilon})\right|_{\varepsilon=0}=0\hskip 18.49988pt\text{for all such $\theta$-blind perturbations $h$.}$
+    Using standard spike-variation/density arguments within the $\theta$-blind admissible class (i.e., choosing $h_{t}$ supported on arbitrarily small time intervals with arbitrary directions), we obtain the aggregated Pontryagin condition (18), a.s.
+for a.e.
+$t\in[0,T]$.
+  3. 3.
+    Affine-in-control Hamiltonian $\Rightarrow$ linear system and projected form.
+For the quadratic portfolio Hamiltonian (11), $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$.
+Substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ into (18) and introducing the projection inputs (19)–(21) (and using $x>0$ to divide out a common factor) yields the statewise linear system (22).
+Whenever $A_{t}(t,x,y)$ is invertible on the working domain, this is equivalently expressed as the projected rule $\pi^{\mathrm{agg}}(t,x,y)=-A_{t}(t,x,y)^{-1}g_{t}(t,x,y)$, i.e.
+(23).
 Note that $\pi^{\mathrm{agg}}$ is generally not equal to the naive average $\mathbb{E}_{\theta\sim q}[\pi^{\star,\theta}(t,x,y)]$ of $\theta$-conditional full-information controls, reflecting the noncommutativity between averaging over $\theta$ and solving a first-order condition.
@@
 Our algorithms and theory target the fixed-$q$, $q$-aggregated projection (23) under the $\theta$-blind Markov feedback restriction.
-2.3 Gaussian references at a fixed decision time
+2.3 Gaussian decision-time references
 This subsection collects Gaussian benchmarks that yield closed-form decision-time reference allocations.
-Fix a decision time $t\in[0,T)$ and suppose an external procedure provides an $\mathcal{F}_{t}$-measurable input law $q_{t}$ for the uncertain premium parameter/state.
-We treat $q_{t}$ as a decision-time input (not a belief process updated within a run): it is held fixed over the remaining horizon $[t,T]$ and the resulting references are used only as validation targets and controlled diagnostics (not as characterizations of the unrestricted optimum).
-If the decision is revisited at a later time $t^{\prime}$, the reference can be recomputed by resolving the decision-time problem with the new input $q_{t^{\prime}}$ (plug-in replanning).
-We present two decision-time references.
-Section 2.3.1 considers static drift uncertainty and provides the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4.
-Section 2.3.2 considers a mean-reverting (OU) premium factor with Gaussian decision-time uncertainty for the current factor level, which induces a horizon-dependent effective Gaussian law for the time-averaged premium and yields a tractable reference used in the hedging-demand recovery study of Section 5.
-Appendix A gives a linear–Gaussian illustration of how an external filtering routine may produce a time-indexed input law $q_{t}$; it is included only as an input mechanism (plug-in replanning), not as a belief-state control formulation.
+Fix a decision time $t\in[0,T)$ and let an external routine provide an $\mathcal{F}_{t}^{\mathrm{obs}}$-measurable input law $q_{t}$ for the uncertain premium parameter/state.
+We treat $q_{t}$ as a decision-time input (not a belief process controlled/updated within a run): it is held fixed over the remaining horizon $[t,T]$.
+The resulting portfolios are used only as validation targets and controlled diagnostics (not as characterizations of the unrestricted optimum).
+If the decision is revisited at a later time $t^{\prime}$, the reference is recomputed by resolving the decision-time problem with the new input $q_{t^{\prime}}$ (plug-in replanning).
+We present two references.
+Section 2.3.1 considers static drift uncertainty and provides the benchmark used in the high-dimensional scaling/geometry experiments of Section 4.
+Section 2.3.2 considers a mean-reverting (OU) premium factor with Gaussian decision-time uncertainty for the current factor level; OU propagation and time-averaging induce a horizon-dependent effective Gaussian premium law, yielding a tractable reference used in the hedging-demand recovery study of Section 5.
+All derivations are deferred to Appendix A.
 2.3.1 Static Gaussian drift uncertainty
-We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter.
-This “unknown mean” specification is a standard device for isolating estimation risk/parameter uncertainty in portfolio choice (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Pástor, 2000).
-Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$.
-The agent chooses a $\theta$-blind policy on $[t,T]$, and all decision-time uncertainty about the (time-constant) premium is summarized by an $\mathcal{F}_{t}$-measurable law $q_{t}(d\theta)$, which we do not update over $[t,T]$.
-Market model (static latent drift).
+Benchmark.
 Let $d$ risky assets satisfy, for $s\in[t,T]$,
@@
 where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\theta\in\mathbb{R}^{d}$ is a latent constant excess-return vector.
-At decision time $t$, uncertainty about $\theta$ is represented by the input law
+At decision time $t$, uncertainty about $\theta$ is summarized by the input law
 $\theta\sim q_{t}(d\theta).$ (25)
@@
 $\frac{dX_{s}^{\pi}}{X_{s}^{\pi}}=\big(r+\pi_{s}^{\top}\theta\big)\,ds+\pi_{s}^{\top}\Sigma^{1/2}\,dW_{s},\hskip 18.49988ptX_{t}^{\pi}=x>0.$ (26)
-We evaluate the (conditional) decision-time objective by mixing over $q_{t}$:
-$J_{t}(\pi):=\mathbb{E}_{\theta\sim q_{t}}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\,\mathcal{F}_{t}\big]\Big].$ (24)
-Under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (Merton, 1969, 1971), which is infeasible here because $\theta$ is latent.
-Analytic $q_{t}$-references via portfolios constant over $[t,T]$.
-To obtain transparent closed-form targets, we temporarily restrict attention to portfolio fractions that are constant in trading time $s$ on $[t,T]$:
-$\pi_{s}\equiv\pi\in\mathbb{R}^{d},\hskip 18.49988pts\in[t,T].$ (25)
-The superscript $\mathrm{const}$” used below refers to this constancy in $s$.
-The resulting reference may still depend on the decision time $t$ through the input $q_{t}$ and on the remaining horizon $\tau=T-t$.
-(Decision-time constant-portfolio targets of this kind are closely connected to myopic/plug-in benchmarks used in the parameter-uncertainty literature; see, e.g., Barberis (2000); Pástor (2000).)
-CRRA ($\gamma>1$): tilted optimality and Gaussian shrinkage.
-For $\gamma>1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal and
-$J_{t}(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)r\tau-\tfrac{1}{2}\gamma(1-\gamma)\tau\,\pi^{\top}\Sigma\pi\Big\}\,M_{q_{t}}\big((1-\gamma)\tau\pi\big),$ (26)
-where $M_{q_{t}}(u):=\mathbb{E}_{\theta\sim q_{t}}[\exp(u^{\top}\theta)]$ is the moment generating function of $q_{t}$.
-Any interior optimizer $\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)$ satisfies the tilted first-order condition
-$\gamma\,\Sigma\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\nabla_{u}\log M_{q_{t}}(u)\Big|_{u=(1-\gamma)\tau\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)}.$ (27)
+We evaluate the decision-time objective by mixing over $q_{t}$:
+$J_{t}(\pi):=\mathbb{E}_{\theta\sim q_{t}}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\,\mathcal{F}_{t}^{\mathrm{obs}}\big]\Big].$ (27)
+Decision-time constant-portfolio reference.
+To obtain a transparent closed-form target, we restrict to portfolio fractions constant over $[t,T]$:
+$\pi_{s}\equiv\pi\in\mathbb{R}^{d},\hskip 18.49988pts\in[t,T],$ (28)
+and take CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ with $\gamma>1$.
 If the decision-time input is Gaussian,
 $q_{t}=\mathcal{N}(m_{t},P_{t}),\hskip 18.49988ptP_{t}\succeq 0,$ (29)
-then $\nabla_{u}\log M_{q_{t}}(u)=m_{t}+P_{t}u$ and the reference reduces to the linear system
-$\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=m_{t},$ (29)
-hence
-$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)^{-1}m_{t}.$ (30)
-For $\gamma>1$, this takes the familiar shrinkage form
-$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt(\gamma>1),$ (31)
-which makes explicit how decision-time uncertainty $P_{t}$ dampens exposures as the remaining horizon grows (a qualitative effect widely emphasized in return-predictability/parameter-uncertainty studies; see, e.g., Barberis (2000); Pástor (2000); Kandel and Stambaugh (1996)).
+then the constant-portfolio reference is
+$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt\tau:=T-t,\qquad(\gamma>1).$ (30)
+This exhibits the familiar Gaussian shrinkage: decision-time uncertainty $P_{t}$ dampens exposures, increasingly so as the remaining horizon $\tau$ grows.
 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
-We next replace the static premium by a mean-reverting Gaussian premium factor, a standard reduced-form device for return predictability and intertemporal hedging (Campbell and Viceira, 2002; Xia, 2001).
-Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$.
-At time $t$ we take as input a Gaussian law for the current factor level $Y_{t}$ (rather than a point estimate), capturing decision-time statistical uncertainty (e.g., interval-type uncertainty quantification, resampling variability, or asymptotic approximations).
-Holding this input fixed over $[t,T]$, OU propagation and time-averaging induce a horizon-dependent effective premium law; when $\rho\neq 0$, return–factor shock correlation adds a further quadratic term relevant for hedging demand (Campbell and Viceira, 2002; Xia, 2001).
-OU premium dynamics.
-Let $Y_{s}\in\mathbb{R}^{m}$ be a premium factor following
+OU premium factor.
+Let $Y_{s}\in\mathbb{R}^{m}$ follow
 $dY_{s}=K(\bar{y}-Y_{s})\,ds+\Xi\,dW_{s}^{Y},\hskip 18.49988ptK\in\mathbb{R}^{m\times m}\ \text{Hurwitz and invertible},$ (31)
-where $\bar{y}\in\mathbb{R}^{m}$ and $\Xi\in\mathbb{R}^{m\times m}$. Risky excess returns satisfy
+and let risky excess returns satisfy
 $dR_{s}:=\frac{dS_{s}}{S_{s}}-r\mathbf{1}\,ds=BY_{s}\,ds+\Sigma^{1/2}\,dW_{s},$ (32)
@@
 $d\langle W,W^{Y}\rangle_{s}=\rho\,ds,\hskip 18.49988pt\rho\in\mathbb{R}^{d\times m}.$ (33)
-Decision-time input law for the factor.
-At decision time $t$, we take as input an $\mathcal{F}_{t}$-measurable Gaussian law for the current factor level,
+Decision-time input law and induced effective premium law.
+At decision time $t$, we take as input a Gaussian law for the current factor level,
 $Y_{t}\sim q_{t}(dy)=\mathcal{N}(m_{t},P_{t}),\hskip 18.49988ptP_{t}\succeq 0.$ (34)
-A common special case is purely model-implied predictive propagation from an initial uncertainty $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$ without conditioning on observations:
-$m_{t}=\bar{y}+e^{-Kt}(m_{0}-\bar{y}),$
-$P_{t}=e^{-Kt}P_{0}e^{-K^{\top}t}+\int_{0}^{t}e^{-K(t-s)}\,\Xi\Xi^{\top}\,e^{-K^{\top}(t-s)}\,ds,$ (36)
-which is an OU transition (prior-predictive) law and should not be interpreted as an information update.
-(References for return predictability and OU-style investment-opportunity dynamics include Campbell and Viceira (2002); Xia (2001).)
-Induced Gaussian law for the time-averaged premium over $[t,T]$.
-Define the remaining-horizon integrated factor and the associated horizon-averaged effective premium:
-$I_{t,T}:=\int_{t}^{T}Y_{s}\,ds\in\mathbb{R}^{m},\hskip 18.49988pt\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\,B\,I_{t,T}\in\mathbb{R}^{d}.$ (37)
-Let
-$A(\tau):=K^{-1}\big(I-e^{-K\tau}\big)\in\mathbb{R}^{m\times m}.$
-Under the linear–Gaussian OU model, $I_{t,T}$ is Gaussian under the input law (35):
-$I_{t,T}\sim\mathcal{N}\big(m_{I}(t,\tau),C_{I}(t,\tau)\big),$
-with
-$m_{I}(t,\tau)=\mathbb{E}[I_{t,T}]=\tau\bar{y}+A(\tau)\,(m_{t}-\bar{y}),$ (38)
-$C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})=A(\tau)\,P_{t}\,A(\tau)^{\top}+\int_{0}^{\tau}A(u)\,\Xi\Xi^{\top}\,A(u)^{\top}\,du.$ (39)
-Consequently, the horizon-averaged effective premium is Gaussian,
-$\bar{\theta}_{t,\tau}\sim\mathcal{N}\!\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big),\hskip 18.49988ptm_{\bar{\theta}}(t,\tau)=\frac{1}{\tau}B\,m_{I}(t,\tau),\hskip 18.49988ptP_{\bar{\theta}}(t,\tau)=\frac{1}{\tau^{2}}B\,C_{I}(t,\tau)\,B^{\top}.$ (40)
-The pair $\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big)$ is the horizon-dependent effective Gaussian law that we use as a controlled analytic input.
-Closed-form decision-time references under portfolios constant over $[t,T]$.
-As in Section 2.3.1, we define an analytic reference by restricting to portfolio fractions that are constant in trading time $s$ over $[t,T]$, i.e.
-$\pi_{s}\equiv\pi\in\mathbb{R}^{d}$.
-The superscript $\mathrm{const}$” below refers to this constancy in $s$; the resulting reference may still depend on the decision time $t$ through $(m_{t},P_{t})$ and on the remaining horizon $\tau$ through OU averaging.
-CRRA ($\gamma>1$).
-When $\rho\neq 0$, return shocks and factor shocks are contemporaneously correlated, producing an additional quadratic term in the constant-portfolio objective (hedging-demand channel; see, e.g., Campbell and Viceira (2002); Xia (2001)).
-Define the cross-covariance between $I_{t,T}$ and the return Brownian increment $W_{T}-W_{t}$:
-$C_{IW}(\tau):=\mathrm{Cov}\!\big(I_{t,T},\,W_{T}-W_{t}\big)=\int_{0}^{\tau}A(u)\,\Xi\,\rho^{\top}\,du\;\in\;\mathbb{R}^{m\times d},$ (41)
-and the induced symmetric cross term
-$M_{\mathrm{cross}}(\tau):=B\,C_{IW}(\tau)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(\tau)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (42)
-Then the Gaussian decision-time reference under constant portfolios is characterized by
-$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1).$ (43)
-Equivalently,
-$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)^{-1}B\,m_{I}(t,\tau).$
-When $\rho=0$, we have $C_{IW}(\tau)=0$ and $M_{\mathrm{cross}}(\tau)=0$, and the reference reduces to the independence-case shrinkage form obtained by plugging the effective Gaussian law (40) into the static Gaussian formula of Section 2.3.1:
-$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{\bar{\theta}}(t,\tau)\big)^{-1}m_{\bar{\theta}}(t,\tau),\hskip 18.49988pt(\rho=0,\ \gamma>1).$
+Define the remaining-horizon integrated factor and horizon-averaged effective premium:
+$I_{t,T}:=\int_{t}^{T}Y_{s}\,ds\in\mathbb{R}^{m},\hskip 18.49988pt\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\,B\,I_{t,T}\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (35)
+Under the linear–Gaussian OU model, holding the input law (34) fixed over $[t,T]$ induces a horizon-dependent Gaussian law
+$\bar{\theta}_{t,\tau}\sim\mathcal{N}\!\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big),$ (36)
+where explicit formulas for $m_{\bar{\theta}}(t,\tau)$ and $P_{\bar{\theta}}(t,\tau)$ are given in Appendix A.
+Decision-time constant-portfolio reference (CRRA, $\gamma>1$).
+Restricting to constant portfolio fractions $\pi_{s}\equiv\pi$ on $[t,T]$ as above, the Gaussian decision-time reference under CRRA ($\gamma>1$) is characterized by the linear system
+$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1),$ (37)
+where $m_{I}(t,\tau):=\mathbb{E}[I_{t,T}]$, $C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})$, and the cross term $M_{\mathrm{cross}}(\tau)$ captures the effect of instantaneous return–factor correlation $\rho\neq 0$ (hedging-demand channel).
+Closed-form expressions for $m_{I}(t,\tau)$, $C_{I}(t,\tau)$, and $M_{\mathrm{cross}}(\tau)$ are collected in Appendix A.
+When $\rho=0$, we have $M_{\mathrm{cross}}(\tau)=0$ and the reference reduces to the independence-case shrinkage form obtained by plugging the effective Gaussian law (36) into the static Gaussian formula (30).
 3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
@@
 Uncertainty enters only through sampling $\theta$ inside the simulator.
-Our solver has two components.
-Stage 1 (PG–DPO) performs stochastic gradient ascent on $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$ by differentiating through an Euler simulator via BPTT.
-Stage 2 (P–PGDPO) aggregates the BPTT costate blocks across $\theta\sim q$ and enforces the $q$-aggregated Pontryagin first-order condition, yielding a deployable projected control.
-Section 3.1 establishes the conditional BPTT–PMP correspondence and specifies the adjoint blocks needed for projection.
+We use a deterministic policy network $\pi_{\varphi}:\mathcal{D}\to\mathbb{R}^{d}$, parameterized by $\varphi$, where $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ denotes a working domain inside the state space.
+Algorithmic overview and terminology.
+Our full solver is Pontryagin–guided direct policy optimization (PG–DPO) and consists of two stages.
+Stage 1 (DPO) performs stochastic gradient ascent on the ex–ante objective $J(\varphi)=\mathbb{E}\big[U\big(X_{T}^{\pi_{\varphi},\theta}\big)\big]$ by differentiating through an Euler simulator via BPTT.
+Stage 2 (Pontryagin guidance / projection) aggregates the BPTT adjoint (costate) blocks across $\theta\sim q$ and enforces the $q$-aggregated Pontryagin first-order condition, yielding a deployable projected control.
+We refer to Stage 1 alone as DPO and to the two-stage procedure (Stage 1+2) as PG–DPO.
+Section 3.1 establishes the conditional BPTT–PMP correspondence and specifies the adjoint blocks needed for the projection stage.
 Section 3.2 develops the $q$-aggregated projection and a residual-based policy-gap bound.
-Practical coupling mechanisms (residual-form projection and interactive distillation) are deferred to Appendix F.
-3.1 PG–DPO as stochastic gradient ascent and a conditional BPTT–PMP correspondence
+Practical coupling mechanisms (residual-form projection and interactive distillation) are deferred to Appendix G.
+3.1 DPO via stochastic gradient ascent and a conditional BPTT–PMP correspondence
 Setup and objectives (frozen $\theta$, deployable $\theta$-blind feedback).
@@
 A deployable portfolio policy is a $\theta$-blind Markov feedback rule represented by a neural network
-$\pi_{\varphi}:\ [0,T]\times(0,\infty)\times\mathbb{R}^{m}\to\mathbb{R}^{d},\hskip 18.49988pt(t,x,y)\mapsto\pi_{\varphi}(t,x,y),\hskip 18.49988pt\varphi\in\mathbb{R}^{p},$ (44)
-which does not take $\theta$ as an input. Conditional on a fixed frozen $\theta$, the controlled state $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ evolves as
-$\frac{dX_{t}^{\pi,\theta}}{X_{t}^{\pi,\theta}}$ $=\Big(r+\pi_{t}^{\top}b\big(Y_{t}^{\theta},\theta\big)\Big)\,dt+\pi_{t}^{\top}\sigma\big(Y_{t}^{\theta},\theta\big)\,dW_{t},\hskip 18.49988ptX_{0}=x>0,$ (45)
+$\pi_{\varphi}:\ \mathcal{D}\to\mathbb{R}^{d},\hskip 18.49988pt\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m},\hskip 18.49988pt(t,x,y)\mapsto\pi_{\varphi}(t,x,y),\hskip 18.49988pt\varphi\in\mathbb{R}^{p},$ (38)
+which does not take $\theta$ as an input. Conditional on a fixed frozen $\theta$, the controlled state $(X_{t}^{\pi_{\varphi},\theta},Y_{t}^{\theta})_{t\in[0,T]}$ evolves as
+$\frac{dX_{t}^{\pi_{\varphi},\theta}}{X_{t}^{\pi_{\varphi},\theta}}$ $=\Big(r+\pi_{t}^{\top}b\big(Y_{t}^{\theta},\theta\big)\Big)\,dt+\pi_{t}^{\top}\sigma\big(Y_{t}^{\theta},\theta\big)\,dW_{t},\hskip 18.49988ptX_{0}=x>0,$ (39)
 $dY_{t}^{\theta}$ $=a\big(Y_{t}^{\theta},\theta\big)\,dt+\beta\big(Y_{t}^{\theta},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m},$ (40)
@@
 Thus $\sup_{\varphi}J(\varphi)$ is a stochastic optimization problem in which $\theta$ is sampled inside the simulator while the policy remains $\theta$-blind.
-Discretization, sampling over $\theta$, and baseline PG–DPO update.
+Discretization, sampling over $\theta$, and the baseline DPO update.
 We discretize the horizon using an Euler scheme.
@@
 Given $\pi_{\varphi}$ and Brownian increments, the mapping
-$\big(x^{(i)},y^{(i)},\theta^{(i)},\{\Delta W_{k}^{(i)},\Delta W_{k}^{Y,(i)}\}_{k=0}^{N-1},\varphi\big)\longmapsto U\big(X_{N}^{(i)}\big)$
+$\big(x_{0}^{(i)},y_{0}^{(i)},\theta^{(i)},\{\Delta W_{k}^{(i)},\Delta W_{k}^{Y,(i)}\}_{k=0}^{N-1},\varphi\big)\longmapsto U\big(X_{N}^{(i)}\big)$
 is a finite computational graph; automatic differentiation thus computes exact discrete-time gradients $\nabla_{\varphi}U(X_{N}^{(i)})$ for the chosen discretization.
-A typical PG–DPO update samples a mini-batch $\{z_{0}^{(i)}\}_{i=1}^{M}\sim\nu$, samples latent parameters either independently per episode ($\theta^{(i)}\sim q$) or shared within the batch (one $\theta\sim q$ reused across $i$), simulates Euler rollouts, and performs stochastic gradient ascent.
+A typical DPO update samples a mini-batch $\{z_{0}^{(i)}\}_{i=1}^{M}\sim\nu$, samples latent parameters either independently per episode ($\theta^{(i)}\sim q$) or shared within the batch (one $\theta\sim q$ reused across $i$), simulates Euler rollouts, and performs stochastic gradient ascent.
 Suppressing episode indices and writing $\theta$ for the frozen draw used in the rollout, the Euler recursion is
@@
 $X_{k+1}$ $=X_{k}+X_{k}\Big(r+\pi_{\varphi}\big(t_{k},X_{k},Y_{k}\big)^{\top}b\big(Y_{k},\theta\big)\Big)\Delta t+X_{k}\,\pi_{\varphi}\big(t_{k},X_{k},Y_{k}\big)^{\top}\sigma\big(Y_{k},\theta\big)\Delta W_{k},$
-starting from $(t_{0},X_{0},Y_{0})=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$. The episode reward is
+starting from $(t_{0},X_{0},Y_{0})=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$.
+Remark 4 (Positivity of wealth in the discrete simulator).
+In continuous time, the wealth process in (39) stays strictly positive for any admissible portfolio.
+A naive additive Euler step may nevertheless generate negative $X_{k}$ for coarse steps or aggressive controls, making CRRA utilities (and the higher-order derivatives used in BPTT) ill-defined.
+In implementations we therefore use a positivity-preserving discretization (e.g.
+exponential Euler on log-wealth) or apply a mild truncation/clipping so that $X_{k}\geq\underline{x}>0$ on the computation graph.
+The correspondence results below should be read for such a positivity-preserving simulator; equivalently, they apply on the event $\{\min_{0\leq k\leq N}X_{k}>0\}$.
+The episode reward is
 $J^{(i)}(\varphi):=U\big(X_{N}^{(i)}\big),$ (43)
@@
 Concretely, for a fixed frozen $\theta$ and the simulator filtration
-$G_{t_{k}}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t_{k}\}\big)\qquad\text{(with the usual augmentation)},$
+$\mathcal{G}_{t_{k}}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t_{k}\}\big)\qquad\text{(with the usual augmentation)},$
 we define the adapted (projected) discrete adjoints
-$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid G_{t_{k}}^{\theta}].$
+$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid\mathcal{G}_{t_{k}}^{\theta}].$
 Theorem 2 concerns the convergence of these adapted discrete objects (together with the associated one-step BSDE martingale coefficients) to the continuous-time Pontryagin adjoints.
-Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, including second-adjoint blocks; uniform on compacts)).
-Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (45)–(46) under the $\theta$-blind policy $\pi_{\varphi}$, and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle, including the wealth-component first adjoint and the relevant blocks of the second adjoint needed for the portfolio Hamiltonian gradient.
-Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote these continuous-time Pontryagin objects under $\pi_{\varphi}$.
-In smooth Markov regimes, they coincide (along the controlled state process) with $(V_{x},V_{xx},V_{xy})$ for the $\theta$-conditional value function.
-Let $(p_{k},p_{x,k},p_{y,k})$ be the discrete pathwise quantities computed by BPTT for the Euler discretization with step $\Delta t$, as defined in (51)–(52).
-Let $(p_{k}^{\Delta t,\theta},p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$ be their adapted projections onto $G_{t_{k}}^{\theta}$ as defined above, and let the associated discrete-time BSDE martingale coefficients be defined via the one-step conditional $L^{2}$ projection on each step.
-Then, as $\Delta t\to 0$, the projected (adapted) BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories).
-Moreover, for any compact set $K\subset\Theta$, the constants in the convergence bounds can be chosen uniformly for all $\theta\in K$.
-Proof.
-See Appendix C. ∎
+Theorem 2 (BPTT–BSDE/PMP correspondence (conditional on $\theta$, including second-adjoint blocks; uniform on compacts)).
+Fix a compact $\Theta_{0}\subset\Theta$ and a $\theta$-blind policy network $\pi_{\varphi}$.
+Assume Assumption 1 holds for $\pi=\pi_{\varphi}$ and every $\theta\in\Theta_{0}$, and that the discrete simulator used for BPTT preserves positivity as in Remark 4.
+For each $\theta\in\Theta_{0}$, let $(p_{t}^{\theta},z_{t}^{\theta},\tilde{z}_{t}^{\theta})$ denote the wealth-component adjoint BSDE solution (15) under $\pi_{\varphi}$, and let $(p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the blocks (16) (equivalently, the relevant blocks of the second adjoint along the controlled trajectory).
+Let $(X_{k},Y_{k})_{k=0}^{N}$ be the Euler simulator on a grid $(t_{k})_{k=0}^{N}$ with step size $\Delta t$, and let $(p_{k},p_{x,k},p_{y,k})$ be the (pathwise, generally non-adapted) BPTT adjoint blocks (45)–(46).
+Define their adapted projections by conditional expectations under the simulator filtration $\mathcal{G}_{t_{k}}^{\theta}$:
+$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid\mathcal{G}_{t_{k}}^{\theta}].$
+Let $(Z_{k}^{\theta},\tilde{Z}_{k}^{\theta})$ be the one-step conditional $L^{2}$ projection coefficients of $p_{k+1}$ onto $\mathrm{span}\{\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$ (so that the orthogonal residual is $L^{2}$-uncorrelated with the increments).
+Form the piecewise-constant interpolations on $t_{k},t_{k+1})$:
+$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\qquad z_{t}^{\Delta t,\theta}:=Z_{k}^{\theta},\qquad\tilde{z}_{t}^{\Delta t,\theta}:=\tilde{Z}_{k}^{\theta}.$
+Then, as $\Delta t\to 0$,
+$\sup_{\theta\in\Theta_{0}}\sup_{t\in[0,T]}\mathbb{E}\Big[\,|p_{t}^{\Delta t,\theta}-p_{t}^{\theta}|^{2}+|p_{x,t}^{\Delta t,\theta}-p_{x,t}^{\theta}|^{2}+\|p_{y,t}^{\Delta t,\theta}-p_{y,t}^{\theta}\|^{2}\,\Big]$ $\to 0,$
+$\sup_{\theta\in\Theta_{0}}\mathbb{E}\!\left[\int_{0}^{T}\Big(\|z_{t}^{\Delta t,\theta}-z_{t}^{\theta}\|^{2}+\|\tilde{z}_{t}^{\Delta t,\theta}-\tilde{z}_{t}^{\theta}\|^{2}\Big)\,dt\right]$ $\to 0.$
+In particular, the convergence bounds can be taken uniform over $\theta$ in compact subsets of $\Theta$.
+Proof sketch.
+Fix $\theta\in\Theta$ and work conditionally on $\theta$ under the augmented simulator filtration $\mathcal{G}_{t}^{\theta}:=\sigma(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\})$ (usual augmentation).
+The argument follows the deterministic-parameter BPTT–BSDE/PMP correspondence of Huh et al.
+(2025), applied conditionally on $\theta$; uniformity on compact $\Theta_{0}\subset\Theta$ is obtained by assuming the required Lipschitz/growth/nondegeneracy/smoothness constants are uniform over $\theta\in\Theta_{0}$.
+  1. 1.
+    Pathwise BPTT adjoints on the Euler computation graph.
+For fixed $\theta$, simulate the Euler discretization of (39)–(40).
+BPTT computes the pathwise derivatives of the terminal utility along the full discrete computation graph, producing the pathwise adjoint blocks $(p_{k},p_{x,k},p_{y,k})$ as in (45)–(46).
+These quantities are generally $\mathcal{G}_{T}^{\theta}$-measurable (they depend on future increments) and hence are not adapted.
+  2. 2.
+    One-step conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables.
+On each step, project $p_{k+1}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ in conditional $L^{2}$ given $\mathcal{G}_{t_{k}}^{\theta}$:
+    $p_{k+1}=\mathbb{E}[p_{k+1}\mid\mathcal{G}_{t_{k}}^{\theta}]+Z_{k}^{\theta}\,\Delta W_{k}+\tilde{Z}_{k}^{\theta}\,\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
+    where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to the span. Define the adapted discrete adjoints by conditional expectations,
+    $p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid\mathcal{G}_{t_{k}}^{\theta}].$
+    By construction, $(p_{k}^{\Delta t,\theta},Z_{k}^{\theta},\tilde{Z}_{k}^{\theta})$ satisfy a discrete-time BSDE representation.
+  3. 3.
+    Identification of the discrete driver; second-adjoint blocks and projection residual.
+Substituting the projection in Step 2 into the one-step BPTT recursion and taking conditional expectations yields that the discrete BSDE driver coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE under the fixed policy $\pi_{\varphi}$, up to one-step remainder terms.
+In particular, in a smooth Markov regime (spelled out in Appendix D), the continuous-time adjoint admits a decoupling field $p_{t}^{\theta}=u^{\theta}(t,X_{t}^{\theta},Y_{t}^{\theta})$ and the sensitivity blocks needed in the portfolio Hamiltonian gradient coincide along trajectories with the spatial derivatives $\partial_{x}u^{\theta}$ and $\partial_{y}u^{\theta}$, yielding the interpretation of $(p_{x,t}^{\theta},p_{y,t}^{\theta})$.
+The projection residual $\varepsilon_{k+1}^{\theta}$ corresponds to higher-order chaos components and is controlled by the same one-step Taylor/chaos estimates as in Huh et al.
+(2025), contributing only higher-order discretization error (uniformly on compacts).
+  4. 4.
+    Convergence as $\Delta t\to 0$ and uniformity on compacts.
+Standard Euler stability for the forward SDE together with convergence of the associated discrete BSDE scheme yields, as $\Delta t\to 0$, mean-square convergence of the adapted discrete adjoint blocks $(p^{\Delta t,\theta},p_{x}^{\Delta t,\theta},p_{y}^{\Delta t,\theta})$ to $(p^{\theta},p_{x}^{\theta},p_{y}^{\theta})$ along the controlled trajectory, and convergence of the martingale coefficients $(Z^{\Delta t,\theta},\tilde{Z}^{\Delta t,\theta})$ to $(Z^{\theta},\tilde{Z}^{\theta})$ in the natural $L^{2}([0,T]\times\Omega)$ sense.
+If the coefficient/policy regularity and nondegeneracy constants are uniform for $\theta\in\Theta_{0}$, then the convergence bounds hold with constants independent of $\theta\in\Theta_{0}$.
+A complete proof (including explicit regularity assumptions, the one-step remainder control, and the uniform-on-compacts bookkeeping) is provided in Appendix D. ∎
 Across $\theta\sim q$, these Pontryagin objects form a $\theta$-indexed family.
-Baseline PG–DPO trains against the ex–ante objective (48) by repeatedly sampling $\theta\sim q$ inside the simulator, while the deployable policy remains $\theta$-blind.
-3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection, diagnostics, and a residual-based policy-gap bound
-Stage 2 is a projection step.
-Given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin adjoint objects under frozen $\theta\sim q$ by BPTT/Monte Carlo, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated stationarity condition derived in Section 2.2.
-Equivalently, Stage 2 can be viewed as a plug-in approximation of one application of the population projection map $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ (defined below) evaluated at $\pi=\pi^{\mathrm{warm}}$ on the working domain.
+Baseline DPO trains against the ex–ante objective (42) by repeatedly sampling $\theta\sim q$ inside the simulator, while the deployable policy remains $\theta$-blind.
+3.2 Stage 2 of PG–DPO under latent $\theta$: $q$-aggregated projection, diagnostics, and a residual-based policy-gap bound
+Stage 2 (Pontryagin guidance) is a projection step.
+Given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ obtained from Stage 1 (DPO), we estimate $\theta$-conditional Pontryagin adjoint objects under frozen $\theta\sim q$ by BPTT/Monte Carlo, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated stationarity condition derived in Section 2.2.
+Together with Stage 1, this projection completes PG–DPO.
+Equivalently, Stage 2 can be viewed as a plug-in approximation of one application of the population projection map $\mathcal{T}(\pi):=-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}$ (defined below) evaluated at $\pi=\pi^{\mathrm{warm}}$ on the working domain.
 The key point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from (estimated) adjoint objects to a portfolio rule.
@@
 Working domain and norms.
-Fix a measurable working state domain $D\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ (e.g.
-a training/evaluation band) and a reference measure $\mu$ on $D$ (e.g.
+Fix a measurable working state domain $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ (e.g.
+a training/evaluation band) and a reference measure $\mu$ on $\mathcal{D}$ (e.g.
 an empirical state distribution induced by rollouts).
-For $h:D\to\mathbb{R}^{n}$ we write
-$\|h\|_{L^{2}(\mu)}:=\Big(\int_{D}\|h(z)\|^{2}\,\mu(dz)\Big)^{1/2},\hskip 18.49988ptz=(t,x,y),$
+For $h:\mathcal{D}\to\mathbb{R}^{n}$ we write
+$\|h\|_{L^{2}(\mu)}:=\Big(\int_{\mathcal{D}}\|h(z)\|^{2}\,\mu(dz)\Big)^{1/2},\hskip 18.49988ptz=(t,x,y),$
 and for $\theta$-indexed families (used when tracking frozen-$\theta$ quantities in analysis/inspection),
-$\|f\|_{L^{2}(q\otimes\mu)}:=\bigg(\int_{\Theta}\int_{D}\|f^{\theta}(z)\|^{2}\,\mu(dz)\,q(d\theta)\bigg)^{1/2}.$ (53)
+$\|f\|_{L^{2}(q\otimes\mu)}:=\bigg(\int_{\Theta}\int_{\mathcal{D}}\|f^{\theta}(z)\|^{2}\,\mu(dz)\,q(d\theta)\bigg)^{1/2}.$ (47)
 Mixed-moment $q$-aggregation under a warm-up policy.
@@
 In the portfolio Hamiltonian (11), this stationarity is equivalent to a statewise linear system and hence to the projected form (23) on the working domain (under invertibility of the aggregated curvature term).
-P–PGDPO constructs a practical approximation of this projection by estimating the relevant aggregated Pontryagin objects under a fixed warm-up policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$.
-Fix a query state $z=(t,x,y)\in D$ and a frozen parameter $\theta$.
+Stage 2 implements a practical approximation of this projection by estimating the relevant aggregated Pontryagin objects under the fixed warm-up policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$.
+Fix a query state $z=(t,x,y)\in\mathcal{D}$ and a frozen parameter $\theta$.
 We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute discrete-time adjoint blocks by autodiff/BPTT on the Euler simulator; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
@@
 $\widehat{g}_{t}^{\mathrm{mix}}(t,x,y)$ $:=\mathbb{E}_{\theta\sim q}\Big[\widehat{g}_{t}^{\theta}(t,x,y)\Big].$ (52)
-Whenever $\widehat{A}_{t}(t,x,y)$ is invertible on $D$, we obtain the mixed-moment projected policy
+Whenever $\widehat{A}_{t}(t,x,y)$ is invertible on $\mathcal{D}$, we obtain the mixed-moment projected policy
 $\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y):=-\,\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{g}_{t}^{\mathrm{mix}}(t,x,y).$ (53)
 Stabilized inversion and projection diagnostics (implementation).
-In practice, finite-sample estimates can make $\widehat{A}_{t}$ ill-conditioned or even singular on parts of $D$.
-Our default implementation therefore uses standard stabilizers (Appendix E.5), including: (i) ridge regularization $\widehat{A}_{t,\lambda}:=\widehat{A}_{t}+\lambda I_{d}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\lambda}:=-\widehat{A}_{t,\lambda}^{-1}\widehat{G}_{t}^{\mathrm{mix}}$, (ii) condition-number or eigenvalue diagnostics to skip/clip projection on unreliable states (falling back to $\pi^{\mathrm{warm}}$), and (iii) residual-form projection (Section F.1) to reduce variance.
-The analysis below is stated for the population projection $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ and for small estimator perturbations, which is precisely the regime enforced by these stabilizers.
-Remark (decoupled aggregation heuristic).
-The mixed-moment aggregation (57)–(58) is the object appearing in the exact $q$-aggregated first-order condition.
-In some high-dimensional experiments, we additionally consider a variance-reduction heuristic that replaces mixed moments by products of empirical means, e.g.
-$\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}\,b(y,\theta)]\approx\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}]\cdot\mathbb{E}_{\theta}[b(y,\theta)]$, and similarly for the other terms.
-All theoretical statements in this subsection are for the mixed-moment projection (59).
+In practice, finite-sample estimates can make $\widehat{A}_{t}$ ill-conditioned or even singular on parts of $\mathcal{D}$.
+Our default implementation therefore uses standard stabilizers (Appendix F.5), including: (i) ridge regularization $\widehat{A}_{t,\lambda}:=\widehat{A}_{t}+\lambda I_{d}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\lambda}:=-\widehat{A}_{t,\lambda}^{-1}\widehat{g}_{t}^{\mathrm{mix}}$, (ii) condition-number or eigenvalue diagnostics to skip/clip projection on unreliable states (falling back to $\pi^{\mathrm{warm}}$), and (iii) residual-form projection (Section G.1) to reduce variance.
+The analysis below is stated for the population projection $\mathcal{T}(\pi):=-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}$ and for small estimator perturbations, which is precisely the regime enforced by these stabilizers.
 Residual diagnostic and a slab-wise small-gain policy-gap bound.
@@
 the objects in (21) evaluated using the $\theta$-conditional Pontryagin adjoint blocks generated by $\pi$).
-Define the warm-up aggregated stationarity residual on $D$ by
-$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}(t,x,y):=A_{\pi^{\mathrm{warm}}}(t,x,y)\,\pi^{\mathrm{warm}}(t,x,y)+G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}(t,x,y),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\big\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\big\|_{L^{2}(\mu)}.$ (60)
-Note that $r^{\mathrm{warm}}_{\mathrm{FOC,mix}}=0$ is equivalent to $\pi^{\mathrm{warm}}$ being a fixed point of $T$ on $D$. More generally, the population projection admits the residual form
-$T(\pi^{\mathrm{warm}})\;=\;\pi^{\mathrm{warm}}\;-\;A_{\pi^{\mathrm{warm}}}^{-1}\,r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$ (61)
-In practice we monitor the empirical analogue $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$, and use it as a diagnostic to flag states where projection is unreliable (cf.
-Appendix E.5).
-(Equivalently, $\widehat{\pi}^{\mathrm{agg,mix}}=\pi^{\mathrm{warm}}-\widehat{A}_{t}^{-1}\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}$, or its stabilized variant, which motivates the residual-form implementation.)
+Define the warm-up aggregated stationarity residual on $\mathcal{D}$ by
+$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}(t,x,y):=A_{\pi^{\mathrm{warm}}}(t,x,y)\,\pi^{\mathrm{warm}}(t,x,y)+g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}(t,x,y),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\big\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\big\|_{L^{2}(\mu)}.$ (54)
+Note that $r_{\mathrm{FOC,mix}}^{\mathrm{warm}}=0$ is equivalent to $\pi^{\mathrm{warm}}$ being a fixed point of $\mathcal{T}$ on $\mathcal{D}$.
+More generally, the population projection admits the residual form
+$\mathcal{T}(\pi^{\mathrm{warm}})\;=\;\pi^{\mathrm{warm}}\;-\;A_{\pi^{\mathrm{warm}}}^{-1}\,r_{\mathrm{FOC,mix}}^{\mathrm{warm}}.$ (55)
+In practice we monitor the empirical analogue $\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{g}_{t}^{\mathrm{mix}}$, and use it as a diagnostic to flag states where projection is unreliable (cf.
+Appendix F.5).
+(Equivalently, $\widehat{\pi}^{\mathrm{agg,mix}}=\pi^{\mathrm{warm}}-\widehat{A}_{t}^{-1}\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}$, or its stabilized variant, which motivates the residual-form implementation.)
 BPTT/Monte Carlo error term (one explicit choice). To make the discretization/Monte Carlo dependence in (59) explicit, one may take for example
@@
 Huh et al.
-(2025b, Appendix B)), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab.
-Concretely, assume $D$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$.
-Let
-$D_{k}:=D\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
-We write $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map.
+(2025, Appendix B)), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab.
+Concretely, assume $\mathcal{D}$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{L}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$.
+For notational convenience, write $\mathcal{D}\subset[0,T]\times\mathcal{S}$ for some spatial band $\mathcal{S}\subset(0,\infty)\times\mathbb{R}^{m}$, and define
+$\mathcal{D}_{k}:=\mathcal{D}\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{\mathcal{D}_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
+We write $\mathcal{T}(\pi):=-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map.
 Theorem 3 below shows that, under a mild slab-wise local stability regime (i.e.
-a short-time contraction of $T$ on each $D_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error.
-The proof combines a projection-map stability bound (Appendix D.1) with a slab-wise closure (Appendix D.2), in the same spirit as the slab analyses used in Huh et al.
-(2025b).
-Stage 1: PG–DPO
+a short-time contraction of $\mathcal{T}$ on each $\mathcal{D}_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error.
+The proof combines a projection-map stability bound (Appendix E.1) with a slab-wise closure (Appendix E.2), in the same spirit as the slab analyses used in Huh et al.
+(2025).
+Stage 1: DPO
 Sample $\theta\sim q$ in the simulator; Euler rollout; BPTT; update $\varphi$.
 Output: $\varphi^{\mathrm{warm}}$.
-Stage 2: $q$-aggregated projection
+Stage 2: Pontryagin guidance
 Under $\pi_{\varphi^{\mathrm{warm}}}$, estimate adjoint blocks; aggregate over $\theta$; solve $A\pi=-g$.
@@
 Residual-form projection; interactive distillation (teacher refresh).
-Figure 1: Two-stage pipeline of Section 3. A detailed schematic (including the coupling mechanisms and implementation notes) is provided in Appendix F, Fig. LABEL:fig:sec3-pipeline-sideways.
-Theorem 3 (Residual-based ex–ante $\theta$-blind policy-gap bound for P–PGDPO (mixed-moment, deployable, slab-wise local)).
-Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix D.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$ (in particular, that $\widehat{A}_{t}$ remains a small perturbation of $A_{\pi^{\mathrm{warm}}}$ on $D$ under the adopted stabilizers).
+Figure 1: Two-stage PG–DPO pipeline of Section 3: Stage 1 (DPO) trains a $\theta$-blind policy by BPTT, and Stage 2 applies a $q$-aggregated Pontryagin projection.
+A detailed schematic (including coupling mechanisms and implementation notes) is provided in Appendix G, Fig.
+4.
+Theorem 3 (Residual-based ex–ante $\theta$-blind policy-gap bound for PG–DPO (mixed-moment, deployable, slab-wise local)).
+Let $\pi^{\mathrm{warm}}$ be the Stage 1 (DPO) warm-up policy and let $(\widehat{A}_{t},\widehat{g}_{t}^{\mathrm{mix}})$ be the Stage 2 mixed-moment estimators computed under $\pi^{\mathrm{warm}}$.
+Assume Proposition 1 (Appendix E.1) applies on the working domain $\mathcal{D}$ with $A=A_{\pi^{\mathrm{warm}}}$ and $g=g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}$; equivalently, there exist constants $\kappa,M_{g}>0$ such that
+$\|A_{\pi^{\mathrm{warm}}}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa,\hskip 18.49988pt\|g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}\|_{L^{\infty}(\mathcal{D})}\leq M_{g},\hskip 18.49988pt\|\widehat{A}_{t}-A_{\pi^{\mathrm{warm}}}\|_{L^{\infty}(\mathcal{D})}\leq(2\kappa)^{-1},$
+where the last condition is the stable inversion regime enforced in practice by ridge/diagnostics/skip.
 Let $\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem.
-Assume there exists a neighborhood $\mathcal{U}$ of $\pi^{\star,\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in\mathcal{U}$,
-$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G},$
-and assume the slab-wise Lipschitz gain of Appendix D.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
+Assume there exists a neighborhood $\mathcal{U}$ of $\pi^{\star,\mathrm{blind}}$ in $L^{2}(\mu)$ such that:
+  1. (B1)
+    $\pi^{\mathrm{warm}}\in\mathcal{U}$ and, for all $\pi\in\mathcal{U}$, $\|A_{\pi}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$ and $\|g_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(\mathcal{D})}\leq M_{g};$
+  2. (B2)
+    the slab-wise Lipschitz gain of Appendix E.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{g}>0$ such that for every slab $\mathcal{D}_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
     $\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|g_{\pi_{1}}^{\mathrm{mix}}-g_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{g}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k}.$ (57)
 Define
-$\rho(\tau):=\big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\big)\tau^{1/2},\hskip 18.49988pt\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k}).$ (64)
-Assume the slab partition is chosen so that $\rho_{*}<1$.
+$\varrho(\tau):=\big(\kappa\bar{L}_{g}+\kappa^{2}M_{g}\bar{L}_{A}\big)\tau^{1/2},\hskip 18.49988pt\varrho_{*}:=\max_{1\leq k\leq L}\varrho(\tau_{k}),$ (58)
+and assume the slab partition is chosen so that $\varrho_{*}<1$.
 Let $\widehat{\pi}^{\mathrm{agg,mix}}$ be the mixed-moment projected policy (53) computed from BPTT/Monte Carlo estimates under $\pi^{\mathrm{warm}}$ (with stabilized inversion as needed), and let $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ be the population residual (54).
 Then there exists $C_{2}>0$ such that
-$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$ (65)
-where one admissible explicit choice of $\delta_{\mathrm{BPTT}}$ is given in (62). Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
-Proof.
-See Appendix D.3. ∎
+$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$ (59)
+where one admissible explicit choice of $\delta_{\mathrm{BPTT}}$ is given in (56). Moreover, under the stable inversion regime above one may take, for example, $C_{2}:=2\kappa+2\kappa^{2}M_{g}.$
+Remark 5 (Population vs. empirical residual; scope of the bound).
+The residual $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ in (59) is defined using population mixed-moment objects $(A_{\pi^{\mathrm{warm}}},g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}})$.
+In practice we monitor the empirical analogue $\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{g}_{t}^{\mathrm{mix}}$ and enforce the stable inversion regime by ridge/diagnostics/skip (Appendix F.5).
+Relating $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ to $\|\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)}$ amounts to controlling the estimator error in $(\widehat{A}_{t},\widehat{g}_{t}^{\mathrm{mix}})$ and the size of $\pi^{\mathrm{warm}}$ on $\mathcal{D}$; we do not pursue nonasymptotic concentration bounds here.
+Proof sketch.
+Define the population mixed-moment projection map on $\mathcal{D}$ by
+$\mathcal{T}(\pi)\;:=\;-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}.$
+Let the population projected policy be
+$\pi^{\mathrm{proj}}\;:=\;\mathcal{T}(\pi^{\mathrm{warm}})\;=\;-A_{\pi^{\mathrm{warm}}}^{-1}g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}.$
+Notation (for the proof only). To avoid clutter in this proof sketch, we write
+$\widehat{A}:=\widehat{A}_{t},\hskip 18.49988pt\widehat{g}:=\widehat{g}_{t}^{\mathrm{mix}},$
+as functions on the working domain $\mathcal{D}$ (with the dependence on $(t,x,y)$ suppressed), so that $\widehat{\pi}^{\mathrm{agg,mix}}=-\widehat{A}^{-1}\widehat{g}$ is exactly (53).
+Write the stage-2 (stabilized) estimator-based projected policy as
+$\widehat{\pi}^{\mathrm{agg,mix}}\;:=\;-\widehat{A}^{-1}\widehat{g}$,
+[…]
+  1. 1.
+    Estimator-to-population stability (Appendix E.1).
+By the stated uniform stability/invertibility assumptions (in particular, $\widehat{A}$ is an $L^{\infty}(\mathcal{D})$-small perturbation of $A_{\pi^{\mathrm{warm}}}$ under the adopted stabilizers), Proposition 1 applies with $A=A_{\pi^{\mathrm{warm}}}$ and $g=g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}$, yielding
+    $\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\big\|_{L^{2}(\mu)}\;\leq\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
+    for some constant $C_{2}$ depending only on the uniform inverse bound $\kappa$ and uniform $L^{\infty}$ bounds on the population $g$ terms (one explicit admissible $\delta_{\mathrm{BPTT}}$ is given in (56)).
+  2. 2.
+    Slab-wise contraction of the population projection map (Appendix E.2).
+On each slab $\mathcal{D}_{k}$, the slab-wise Lipschitz gain assumption (57) together with the uniform bounds on $\mathcal{U}$ implies (via Proposition 2) that for all $\pi_{1},\pi_{2}\in\mathcal{U}$,
+    $\|\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})\|_{k}\;\leq\;\varrho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}\;\leq\;\varrho_{*}\,\|\pi_{1}-\pi_{2}\|_{k}.$
+  3. 3.
+    Fixed-point identity for the local interior optimum.
+Since $\pi^{\star,\mathrm{blind}}$ is a locally optimal interior deployable $\theta$-blind policy, it satisfies the aggregated stationarity condition on $\mathcal{D}$, equivalently the fixed-point relation
+    $\pi^{\star,\mathrm{blind}}=\mathcal{T}(\pi^{\star,\mathrm{blind}}).$
+    Therefore, for each slab $\mathcal{D}_{k}$,
+    $\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\;=\;\|\mathcal{T}(\pi^{\mathrm{warm}})-\mathcal{T}(\pi^{\star,\mathrm{blind}})\|_{k}\;\leq\;\varrho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$
+  4. 4.
+    Residual identity and small-gain closure.
+Let $r_{\mathrm{FOC,mix}}^{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}}\pi^{\mathrm{warm}}+g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}$ and $\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k}$.
+By construction of $\pi^{\mathrm{proj}}$,
+    $\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\;=\;A_{\pi^{\mathrm{warm}}}^{-1}\,r_{\mathrm{FOC,mix}}^{\mathrm{warm}},\hskip 18.49988pt\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\;\leq\;\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$
+    Using the triangle inequality and the contraction estimate above,
+    $\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}+\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\varrho_{*}\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$
+    Since $\varrho_{*}<1$, rearranging gives
+    $\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\frac{\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}},\hskip 18.49988pt\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$
+    Summing over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k}\|f\|_{k}^{2}$ yields
+    $\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\;\leq\;\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$
+  5. 5.
+    Conclusion. Finally, apply the triangle inequality:
+    $\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\big\|_{L^{2}(\mu)}+\big\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)},$
+    and combine the bounds from Steps 1 and 4 to obtain (59).
+A complete proof (including the precise stable-inversion conditions implied by the stage-2 stabilizers and an explicit admissible choice of $\delta_{\mathrm{BPTT}}$) is provided in Appendix E.3. ∎
 Coupling and amortization.
-Figure 1 summarizes the overall two-stage design.
-Stage 2 is used purely as a warm-started post-processing map: it reuses the current stage 1 policy, estimates the adjoint blocks under frozen $\theta\sim q$, and outputs a single deployable $\theta$-blind projected rule.
+Figure 1 summarizes the overall two-stage PG–DPO design.
+Stage 2 is used purely as a warm-started post-processing map: it reuses the current Stage 1 (DPO) policy, estimates the adjoint blocks under frozen $\theta\sim q$, and outputs a single deployable $\theta$-blind projected rule.
 For numerical stability and amortized deployment, we additionally consider (i) a residual/control-variate form of the projection and (ii) interactive distillation of the projected rule into the policy network.
-These practical coupling mechanisms are described in Appendix F.
+These practical coupling mechanisms are described in Appendix G.
 4 Breaking the Dimensional Barrier under Drift Uncertainty
@@
 Because the benchmark admits a transparent decision-time reference, we can measure accuracy directly via a decision-time Euclidean RMSE to the analytic reference, rather than relying only on realized utility.
-Our goal is to test whether Pontryagin-guided learning and projection remain stable as the number of assets grows.
+Our goal is to test whether the two-stage PG–DPO pipeline remains stable as the number of assets grows.
 We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions.
-We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection), and (when applicable) amortized variants via interactive distillation, under Monte Carlo budgets that scale linearly with $d$.
+We report results for (i) Stage 1 (DPO) warm-up policies, (ii) the Stage 2 Pontryagin projection (i.e., the post-processed output completing PG–DPO), and (when applicable) amortized variants via interactive distillation, under Monte Carlo budgets that scale linearly with $d$.
 4.1 Benchmark market and evaluation protocol
@@
 $\theta$-blind deployability (and what uses $\theta$).
-Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $(t,X_{0})$ with fixed horizon $T$), and never takes the realized latent premium $\theta$ as an input.
-The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (notably in Stage 2 projection).
+Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, the observable decision-time state is $(t,x)$ with $t=0$ and a fixed horizon $T$), and never takes the realized latent premium $\theta$ as an input.
+The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (in Stage 1 for stochastic gradients, and in Stage 2 for aggregated adjoint/mixed-moment inputs).
 Any $\theta$-indexed objects (when referenced elsewhere) are used only for offline diagnostics and are not part of the deployable decision rule.
@@
 We write excess returns as
-$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=\theta\,dt+B\,\Sigma_{f}^{1/2}\,dW_{t}^{f}+\mathrm{diag}(\sqrt{D})\,dW_{t}^{\varepsilon},$ (66)
-with $B\in\mathbb{R}^{d\times k_{\Sigma}}$, $\Sigma_{f}\succ 0$, and $D\in(0,\infty)^{d}$. This implies
-$\Sigma=B\,\Sigma_{f}\,B^{\top}+\mathrm{diag}(D)\;=\;FF^{\top}+\mathrm{diag}(D),$ (67)
+$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=\theta\,dt+B\,\Sigma_{f}^{1/2}\,dW_{t}^{f}+\mathrm{diag}(\sigma_{\varepsilon})\,dW_{t}^{\varepsilon},$ (60)
+with $B\in\mathbb{R}^{d\times k_{\Sigma}}$, $\Sigma_{f}\succ 0$, and $\sigma_{\varepsilon}\in(0,\infty)^{d}$. This implies
+$\Sigma=B\,\Sigma_{f}\,B^{\top}+\mathrm{diag}(\sigma_{\varepsilon}^{2})\;=\;FF^{\top}+\mathrm{diag}(\sigma_{\varepsilon}^{2}),$ (61)
 where $F:=B\,\mathrm{chol}(\Sigma_{f})$. We generate the mean premium in an APT-like form by drawing a factor price vector $\lambda_{m}\in\mathbb{R}^{k_{\Sigma}}$ and setting
@@
 One-shot generation and fairness across methods.
-For each dimension $d$, we generate a single market instance $(B,\Sigma_{f},D,\lambda_{m})$ once (using a fixed random seed) and hold it fixed across all algorithmic comparisons and MC-budget variants.
+For each dimension $d$, we generate a single market instance $(B,\Sigma_{f},\sigma_{\varepsilon},\lambda_{m})$ once (using a fixed random seed) and hold it fixed across all algorithmic comparisons and MC-budget variants.
 Within a fixed $d$, we change only the uncertainty covariance $P$ (aligned vs.
@@
 Misaligned: uncertainty factors are rotated away from the market factor space,
-$P=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}+s\,\mathrm{diag}(D),$ (70)
+$P=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}+s\,\mathrm{diag}(\sigma_{\varepsilon}^{2}),$ (64)
 where $\widetilde{B}$ is generated independently of $B$ (and can be orthogonalized against the span of $B$ to enforce large principal angles).
-The factor term is rescaled to match the aligned factor magnitude under the same $s$; concretely, writing $\Sigma_{\mathrm{fac}}:=B\Sigma_{f}B^{\top}$ and $\widetilde{\Sigma}_{\mathrm{fac}}:=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}$, we rescale $\widetilde{\Sigma}_{\mathrm{fac}}$ so that $\mathrm{tr}(\widetilde{\Sigma}_{\mathrm{fac}})=\mathrm{tr}(s\,\Sigma_{\mathrm{fac}})$ while keeping the diagonal term fixed as $s\,\mathrm{diag}(D)$.
+The factor term is rescaled to match the aligned factor magnitude under the same $s$; concretely, writing $\Sigma_{\mathrm{fac}}:=B\Sigma_{f}B^{\top}$ and $\widetilde{\Sigma}_{\mathrm{fac}}:=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}$, we rescale $\widetilde{\Sigma}_{\mathrm{fac}}$ so that $\mathrm{tr}(\widetilde{\Sigma}_{\mathrm{fac}})=\mathrm{tr}(s\,\Sigma_{\mathrm{fac}})$ while keeping the diagonal term fixed as $s\,\mathrm{diag}(\sigma_{\varepsilon}^{2})$.
 This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
@@
 We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$ for both aligned and misaligned geometries.
-To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ trajectory budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ rollouts and a high regime with $N_{\mathrm{MC}}=400\cdot d$ rollouts.
-Here $N_{\mathrm{MC}}$ denotes the number of independent simulated trajectories used to approximate the relevant $q$-expectations (in Stage 1 for stochastic gradient estimates, and in Stage 2 for costate/mixed-moment estimators at queried states).
+To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ trajectory budgets: a base regime with $M_{\mathrm{MC}}=100\cdot d$ rollouts and a high regime with $M_{\mathrm{MC}}=400\cdot d$ rollouts.
+Here $M_{\mathrm{MC}}$ denotes the number of independent simulated trajectories used to approximate the relevant $q$-expectations (in Stage 1/DPO for stochastic gradient estimates, and in Stage 2 for adjoint/mixed-moment estimators at queried states).
 Each trajectory draws its own latent $\theta\sim q$ at time 0 and its own Brownian path thereafter; we do not use a separate inner loop of $\theta$-samples per trajectory.
@@
 We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint.
-For $\gamma>1$ we use the CRRA reference (31).
+For $\gamma>1$ we use the CRRA reference (30) at $(t,\tau)=(0,T)$, i.e., $\pi_{q,\gamma}^{\mathrm{const}}(0,T)$.
 We evaluate each method at $t=0$ with a fixed horizon $T=1.5$ on a fixed grid $\{X_{0}^{(i)}\}_{i=1}^{N_{\mathrm{eval}}}$ and report Euclidean RMSE to the analytic reference:
-$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T)-\pi_{q,\gamma}^{\mathrm{const}}(T)\big\|^{2}\bigg)^{1/2},$ (71)
-where $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$ and $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output).
-With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (68)–(70), the remaining subsections compare baseline Stage 1 PG–DPO, post-hoc Stage 2 P–PGDPO projection, and interactive distillation under matched simulation budgets.
+$\mathrm{RMSE}\!\big(\pi_{0}^{\mathrm{meth}},\pi_{q,\gamma}^{\mathrm{const}}\big):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|\pi_{0}^{\mathrm{meth}}\!\big(X_{0}^{(i)};T\big)-\pi_{q,\gamma}^{\mathrm{const}}(0,T)\big\|^{2}\bigg)^{1/2},$ (65)
+where $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$ and $\pi_{0}^{\mathrm{meth}}(\cdot;T)$ denotes the method’s deployable decision-time action at time 0 for the fixed horizon $T$ (a $\theta$-blind output rule).
+With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (62)–(64), the remaining subsections compare (i) the Stage 1 (DPO) warm-up policy, (ii) the Stage 2 Pontryagin projection output (completing PG–DPO), and (iii) interactive distillation variants, under matched simulation budgets.
 4.2 High-dimensional CRRA benchmark: projection and amortization
 Mixed-moment estimation and a decoupling approximation.
-A practical issue throughout our experiments (both aligned and misaligned) is the estimation of mixed moments across the latent parameter, such as $\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]$ (and analogous products entering $\widehat{G}_{t}^{\mathrm{mix}}$), because the costate $p_{t}^{\theta}(z)$ is $\theta$-dependent and high-dimensional, and finite-sample covariance between $p_{t}^{\theta}$ and $\theta$ can lead to large Monte-Carlo variance once the subsequent linear solve is applied.
+A practical issue throughout our experiments (both aligned and misaligned) is the estimation of mixed moments across the latent parameter, such as $\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]$ (and analogous products entering $\widehat{g}_{t}^{\mathrm{mix}}$), because the costate $p_{t}^{\theta}(z)$ is $\theta$-dependent and high-dimensional.
+Finite-sample covariance between $p_{t}^{\theta}$ and $\theta$ can create large Monte Carlo variance, which is then amplified by the subsequent linear solve in the projection step.
 For numerical stability and a uniform protocol across geometries, we therefore use a simple decoupling (independence) approximation for these mixed moments,
 $\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]\;\approx\;\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)]\,\mathbb{E}_{\theta\sim q}[\theta],$
-(and similarly for other mixed products), which is exact when the relevant Pontryagin objects are effectively $\theta$-invariant and is accurate whenever $\mathrm{Cov}_{q}(p_{t}^{\theta}(z),\theta)$ is small relative to marginal scales.
-While this approximation is most valuable under misalignment—where direction mixing can amplify mixed-moment noise—it also performs well in aligned regimes (where mixed moments are typically easier to estimate), and in the CRRA benchmark below it does not alter the qualitative scaling conclusions: projection remains stable, and the observed misaligned degradation is consistent with residual growth and curvature mismatch rather than catastrophic mixed-moment blow-ups.
-[^1] We note, however, that in extreme uncertainty/misalignment—where $\theta$–costate dependence becomes pronounced—the decoupling can break down, in which case one should revert to full mixed-moment estimation (possibly with larger budgets and/or regularized/certified projection).
+(and similarly for other mixed products).
+This approximation is exact when the relevant Pontryagin objects are effectively $\theta$-invariant and is accurate whenever $\mathrm{Cov}_{q}(p_{t}^{\theta}(z),\theta)$ is small relative to marginal scales.
+In this CRRA benchmark it does not change the qualitative scaling conclusions: projection remains stable in aligned regimes, and the misaligned degradation is consistent with residual growth and curvature mismatch rather than catastrophic mixed-moment blow-ups.
+[^1] In extreme uncertainty/misalignment—where $\theta$–costate dependence becomes pronounced—the decoupling can break down.
+In that case one should revert to full mixed-moment estimation (with larger budgets and/or regularized/certified projection).
 Figure 2: Decision-time Euclidean RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window.
@@
 misaligned geometry.
-Curves compare Stage 1 (deployable) and Stage 2 (post-hoc projection), with and without interactive distillation.
+Curves compare the Stage 1 warm-up policy (DPO) to the Stage 2 output (Pontryagin projection; completing PG–DPO), with and without interactive distillation (amortization).
 Solid vs.
@@
 We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (30).
-We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (71) with fixed horizon $T=1.5$.
+We track (i) the Monte Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (65) with fixed horizon $T=1.5$.
 Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window.
-Unless stated otherwise, the projection/teacher direction uses the $p_{\theta}$ aggregation (Section 3.2).
+Unless stated otherwise, the Stage 2 projection / on-the-fly teacher uses the mixed-moment ($p$-weighted) aggregation of Section 3.2 (implemented with the decoupling approximation above).
 What is compared in Figure 2.
-Stage 1 (PG–DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients.
-Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section F.1.
-Interactive distillation (Section F.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
-Thus Figure 2 separates projection quality (Stage 2: post-hoc projected, still $\theta$-blind) from amortized deployable quality (Stage 1 distilled: single forward pass).
+Stage 1 (DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients through the Euler simulator.
+Stage 2 (Pontryagin projection; Section 3.2) post-processes a Stage 1 checkpoint by estimating (aggregated) adjoint inputs and applying the statewise linear solve; we use the residual form of Section G.1.
+The resulting projected control is the output of the full PG–DPO pipeline.
+Interactive distillation (Section G.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable policy network.
+Thus Figure 2 separates projection quality (Stage 2 / PG–DPO: post-hoc projected, still $\theta$-blind) from amortized deployable quality (distilled Stage 1: single forward pass).
 Figure 3: Pathwise sanity check at $d=100$ under common random numbers.
 Black: analytic Gaussian constant-fraction benchmark evaluated with the remaining horizon $\tau=T-t$ (fixed $q$).
-Blue: Stage 1 PG–DPO.
-Orange: on-the-fly Stage 2 P–PGDPO teacher (computed via Monte Carlo projection at visited states).
+Blue: Stage 1 (DPO) warm-up policy.
+Orange: on-the-fly Stage 2 teacher obtained by Pontryagin projection at visited states (PG–DPO).
 Top/bottom: aligned/misaligned.
 Stage 2 projection versus amortization: scaling with dimension.
-For the aligned geometry, small and moderate uncertainty ($s=10^{-3},10^{-2}$) yield a sharp Stage 2 reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$.
-Interactive distillation consistently improves the deployable policy (Stage 1 (distill.) below Stage 1) while leaving Stage 2 essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation reduces the policy-class approximation/optimization gap by injecting that signal into $\pi_{\varphi}$.
+For the aligned geometry, small and moderate uncertainty ($s=10^{-3},10^{-2}$) yield a sharp Stage 2 reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 (DPO) policies remain around $10^{-3}$.
+Interactive distillation consistently improves the deployable policy (distilled Stage 1 below basic Stage 1) while leaving the Stage 2 projection essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation injects that signal into $\pi_{\varphi}$ to reduce the remaining approximation/optimization gap.
 In contrast, under the misaligned geometry the picture becomes more heterogeneous.
 For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$.
-For the largest uncertainty scale ($s=10^{-1}$), Stage 1 becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed.
+For the largest uncertainty scale ($s=10^{-1}$), Stage 1 (DPO) becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed.
 Across settings, the base and high MC budgets tend to yield similar tail-median RMSE curves, suggesting that linear-in-$d$ scaling of simulation budgets is sufficient for stable comparisons in this benchmark.
 Mechanism: why misalignment can reduce projection gains.
-To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix G; see Figures 47.
-The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting.
-These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix uncertainty directions that are poorly aligned with risk directions, which can amplify estimation noise in the projected correction as $d$ grows.
+To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix H; see Figures 58.
+The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa_{\mathrm{curv}}$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting.
+These patterns are consistent with a geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in the projection mix directions of uncertainty that are misaligned with risk directions, thereby amplifying estimation noise in the projected correction as $d$ grows.
 Pathwise sanity check.
 Figure 3 complements the decision-time RMSE with a trajectory-level view under common random numbers.
-In the aligned case, the on-the-fly Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy.
+In the aligned case, the on-the-fly Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 (DPO) policy.
 In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 2 and motivating amortization/reliability mechanisms in interactive distillation.
+Aligned
+$s$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $8.76\text{\times}{10}^{-03}$ $3.74\text{\times}{10}^{-03}$ $1.81\text{\times}{10}^{-03}$ $2.58\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $4.95\text{\times}{10}^{-05}$ $1.10\text{\times}{10}^{-05}$ $8.50\text{\times}{10}^{-06}$ $5.00\text{\times}{10}^{-06}$ $10^{-3}$ DPO (distill.) $4.10\text{\times}{10}^{-03}$ $2.48\text{\times}{10}^{-03}$ $1.31\text{\times}{10}^{-03}$ $1.70\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $5.55\text{\times}{10}^{-05}$ $1.10\text{\times}{10}^{-05}$ $8.00\text{\times}{10}^{-06}$ $5.50\text{\times}{10}^{-06}$ $10^{-3}$ PPO (baseline) $2.76\text{\times}{10}^{-01}$ $1.14\text{\times}{10}^{-01}$ $1.39\text{\times}{10}^{-01}$ $1.67\text{\times}{10}^{-01}$ $10^{-2}$ DPO (basic) $9.19\text{\times}{10}^{-03}$ $3.91\text{\times}{10}^{-03}$ $1.79\text{\times}{10}^{-03}$ $2.57\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $5.57\text{\times}{10}^{-04}$ $1.09\text{\times}{10}^{-04}$ $9.20\text{\times}{10}^{-05}$ $6.00\text{\times}{10}^{-05}$ $10^{-2}$ DPO (distill.) $4.64\text{\times}{10}^{-03}$ $2.55\text{\times}{10}^{-03}$ $1.31\text{\times}{10}^{-03}$ $1.80\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $5.60\text{\times}{10}^{-04}$ $1.09\text{\times}{10}^{-04}$ $9.15\text{\times}{10}^{-05}$ $6.00\text{\times}{10}^{-05}$ $10^{-2}$ PPO (baseline) $3.00\text{\times}{10}^{-01}$ $7.88\text{\times}{10}^{-02}$ $1.50\text{\times}{10}^{-01}$ $1.58\text{\times}{10}^{-01}$ $10^{-1}$ DPO (basic) $1.04\text{\times}{10}^{-02}$ $4.21\text{\times}{10}^{-03}$ $1.81\text{\times}{10}^{-03}$ $2.50\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (basic) $5.29\text{\times}{10}^{-03}$ $1.03\text{\times}{10}^{-03}$ $8.68\text{\times}{10}^{-04}$ $5.76\text{\times}{10}^{-04}$ $10^{-1}$ DPO (distill.) $6.23\text{\times}{10}^{-03}$ $2.76\text{\times}{10}^{-03}$ $1.33\text{\times}{10}^{-03}$ $1.53\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (distill.) $5.33\text{\times}{10}^{-03}$ $1.03\text{\times}{10}^{-03}$ $8.73\text{\times}{10}^{-04}$ $5.76\text{\times}{10}^{-04}$ $10^{-1}$ PPO (baseline) $2.70\text{\times}{10}^{-01}$ $9.65\text{\times}{10}^{-02}$ $1.55\text{\times}{10}^{-01}$ $1.78\text{\times}{10}^{-01}$
+Misaligned
+$s$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $7.49\text{\times}{10}^{-03}$ $3.65\text{\times}{10}^{-03}$ $1.85\text{\times}{10}^{-03}$ $2.56\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $1.94\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $4.30\text{\times}{10}^{-04}$ $8.89\text{\times}{10}^{-04}$ $10^{-3}$ DPO (distill.) $4.49\text{\times}{10}^{-03}$ $2.70\text{\times}{10}^{-03}$ $1.38\text{\times}{10}^{-03}$ $1.02\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $1.95\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $4.30\text{\times}{10}^{-04}$ $8.89\text{\times}{10}^{-04}$ $10^{-3}$ PPO (baseline) $2.87\text{\times}{10}^{-01}$ $8.25\text{\times}{10}^{-02}$ $1.51\text{\times}{10}^{-01}$ $1.69\text{\times}{10}^{-01}$ $10^{-2}$ DPO (basic) $7.87\text{\times}{10}^{-03}$ $3.50\text{\times}{10}^{-03}$ $2.11\text{\times}{10}^{-03}$ $2.68\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $85\text{\times}{10}^{-03} 1.1$$19\text{\times}{10}^{-03}$ $2.07\text{\times}{10}^{-03}$ $2.62\text{\times}{10}^{-03}$ $10^{-2}$ DPO (distill.) $4.86\text{\times}{10}^{-03}$ $2.69\text{\times}{10}^{-03}$ $2.04\text{\times}{10}^{-03}$ $1.46\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $85\text{\times}{10}^{-03} 1.1$$19\text{\times}{10}^{-03}$ $2.07\text{\times}{10}^{-03}$ $2.62\text{\times}{10}^{-03}$ $10^{-2}$ PPO (baseline) $2.58\text{\times}{10}^{-01}$ $8.99\text{\times}{10}^{-02}$ $1.59\text{\times}{10}^{-01}$ $1.38\text{\times}{10}^{-01}$ $10^{-1}$ DPO (basic) $1.44\text{\times}{10}^{-01}$ $4.81\text{\times}{10}^{-01}$ $5.94\text{\times}{10}^{-01}$ $3.74\text{\times}{10}^{-01}$ $10^{-1}$ PG–DPO (basic) $2.40\text{\times}{10}^{-02}$ $1.33\text{\times}{10}^{-02}$ $2.29\text{\times}{10}^{-02}$ $9.31\text{\times}{10}^{-03}$ $10^{-1}$ DPO (distill.) $9.67\text{\times}{10}^{-03}$ $2.26\text{\times}{10}^{-01}$ $2.34\text{\times}{10}^{-01}$ $1.24\text{\times}{10}^{-01}$ $10^{-1}$ PG–DPO (distill.) $1.23\text{\times}{10}^{-02}$ $6.63\text{\times}{10}^{-03}$ $1.34\text{\times}{10}^{-02}$ $6.93\text{\times}{10}^{-03}$ $10^{-1}$ PPO (baseline) $2.78\text{\times}{10}^{-01}$ $8.84\text{\times}{10}^{-02}$ $1.47\text{\times}{10}^{-01}$ $1.77\text{\times}{10}^{-01}$
+Table 1: Decision-time Euclidean RMSE at $t=0$ (tail median over the late-training window; last six evaluation snapshots), computed using the common Euclidean-norm RMSE definition (65).
+DPO rows report the deployable policy output (Stage 1).
+PG–DPO rows report the post-hoc Pontryagin projection output (Stage 2; residual form).
+“Distill.” rows correspond to the amortized deployable policy trained via interactive distillation.
+PPO is a model-free baseline trained under the same benchmark setting (same simulator, horizon $T=1.5$, and $\theta$-blind restriction).
+Bold entries indicate the best (lowest) RMSE among the five methods for each $(s,d)$ and geometry.
 4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
@@
 This comparison is especially informative in our static Gaussian benchmark because the target decision-time rule is structurally simple (constant and available in closed form), so performance gaps primarily reflect optimization difficulty and credit assignment rather than policy-class expressiveness.
-Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section B), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure.
-For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our PG–DPO pipeline, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective.
+Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section C), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure.
+For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our DPO/PG–DPO setup, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective.
 We match simulation budgets at the level of total environment interactions (simulated trajectories $\times$ time steps); implementation and PPO tuning details are deferred to the appendix and code release.
 Empirical outcome.
-Table LABEL:tab:crra_stage12_distill_ppo_landscape shows that PPO remains far from the analytic decision-time $q$-reference across essentially all conditions, with RMSE typically on the order of $10^{-1}$.
-In contrast, the Pontryagin-based pipeline attains substantially smaller errors: in aligned regimes Stage 2 projection reaches the $10^{-5}$$10^{-4}$ range for small and moderate uncertainty, while in misaligned regimes the projection advantage narrows but remains systematic.
-Distillation improves the deployable Stage 1 policy relative to basic PG–DPO, but does not eliminate the remaining gap to the post-hoc projection, consistent with the amortization interpretation in Section 4.2.
+Table 1 shows that PPO remains far from the analytic decision-time $q$-reference across essentially all conditions, with RMSE typically on the order of $10^{-1}$.
+In contrast, the Pontryagin-based pipeline attains substantially smaller errors: in aligned regimes the Stage 2 projection (PG–DPO) reaches the $10^{-5}$$10^{-4}$ range for small and moderate uncertainty, while in misaligned regimes the projection advantage narrows but remains systematic.
+Distillation improves the deployable Stage 1 policy relative to basic Stage 1 (DPO), but does not eliminate the remaining gap to the post-hoc projection, consistent with the amortization interpretation in Section 4.2.
 Why PPO underperforms in this benchmark.
 The gap is not evidence that PPO is intrinsically weak; rather, it reflects that our benchmark stresses regimes where a generic likelihood-ratio policy gradient is statistically disadvantaged compared to pathwise/adjoint-based updates.
-Terminal-reward credit assignment.
-With terminal utility as the only reward signal, PPO faces a long-horizon credit-assignment problem: most intermediate transitions carry zero reward, so advantage estimates inherit high variance and weak learning signals, especially in higher dimensions.
 One contributor is terminal-reward credit assignment: with terminal utility as the only reward signal, most intermediate transitions carry no direct learning signal, so advantage estimates become noisy and the gradient signal weakens as the horizon and dimension grow.
 A second contributor is high-dimensional action noise and variance scaling: in continuous control, exploration is typically implemented through stochastic policies (e.g., diagonal Gaussians), and as $d$ increases the effective exploration noise and likelihood-ratio gradient variance can inflate rapidly, making it difficult to consistently recover the precise constant-fraction structure of the $q$-optimal $\theta$-blind rule under fixed simulation budgets.
-In contrast, Stage 1 exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction that is tailored to the $\theta$-blind ex–ante objective.
+In contrast, Stage 1 (DPO) exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction tailored to the $\theta$-blind ex–ante objective.
 5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
-Sections 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic.
+Section 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic.
 Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (Campbell and Viceira, 2002; Xia, 2001).
 We adopt the mean-reverting Gaussian premium benchmark of Section 2.3.2.
-At decision time ($t=0$), the investor is assumed to have access only to an interval/posterior uncertainty description $q(dy)$ for the initial premium state $Y_{0}$ (e.g., $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$), and we restrict attention to deployable $Y$-blind policies that depend only on observable wealth and time-to-go.
+At decision time ($t=0$), the investor is assumed to have access only to an interval/posterior uncertainty description $q_{0}(dy)$ for the initial premium state $Y_{0}$ (e.g., $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$), and we restrict attention to deployable $Y$-blind policies that depend only on observable wealth and time-to-go.
 When return and factor shocks are correlated ($\rho\neq 0$), the analytic OU decision-time reference (37) contains the cross term $M_{\mathrm{cross}}(\tau)$ (with $\tau:=T-t$) that encodes intertemporal hedging demand.
-We evaluate the two-stage PG–DPO / $q$-aggregated Pontryagin projection pipeline (and interactive distillation) against a model-free PPO baseline, all under the same $Y$-blind deployability restriction.
+We evaluate: (i) Stage 1 DPO (a $Y$-blind warm-up trained by pathwise/BPTT gradients), (ii) Stage 2 Pontryagin projection completing PG–DPO (a post-hoc $q_{0}$-aggregated stationarity correction), and (iii) interactive distillation (amortization), against a model-free PPO baseline—all under the same $Y$-blind deployability restriction.
 Performance is measured by decision-time RMSE at $t=0$ relative to the analytic constant-portfolio OU reference (37).
@@
 Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes as input (i) a point estimate of $Y_{0}$, (ii) the realized initial premium $Y_{0}$, or (iii) the factor path $(Y_{t})$ (including the PPO baseline).
-At deployment, the investor is assumed to have access only to an uncertainty description $q(dy)$ for $Y_{0}$ (interval/posterior output of an external estimation pipeline), and commits to a single policy optimized ex ante under $Y_{0}\sim q$ (without filtering/updating over $[0,T]$).
+At deployment, the investor is assumed to have access only to an uncertainty description $q_{0}(dy)$ for $Y_{0}$ (interval/posterior output of an external estimation pipeline), and commits to a single policy optimized ex ante under $Y_{0}\sim q_{0}$ (without filtering/updating over $[0,T]$).
 The latent premium factor is sampled and propagated only inside the simulator to generate trajectories and to form Monte Carlo averages used by the Stage 2 projection (and by the teacher in distillation).
@@
 We use the two-stage pipeline of Section 3.
-Stage 1 trains a deployable policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1).
-Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section F.1).
-Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section F.2).
-As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table LABEL:tab:ou_rmse_s0_sweep_landscape.
+Stage 1 (DPO) trains a deployable $Y$-blind policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1).
+Stage 2 (Pontryagin projection) applies the $q_{0}$-aggregated stationarity correction computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section G.1); the resulting projected rule is the output of PG–DPO.
+Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section G.2).
+As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table 2.
 We sweep $d\in\{5,10,50,100\}$ (one fixed market instance per $d$, following the same market-generation protocol used in Section 4), train for $5000$ epochs, and evaluate every $100$ epochs.
@@
 (A simple choice is $Q=[I_{m};0]\in\mathbb{R}^{d\times m}$ when $m\leq d$; any fixed orthonormal-column $Q$ yields an equivalent correlation magnitude.) We fix the horizon to $T=1.5$ in this section.
-In addition to the full allocation error, we use the myopic+hedging decomposition induced by the OU structure, and report component-wise diagnostics for the projected (Stage 2) rules.
+In addition to the full allocation error, we use the myopic+hedging decomposition induced by the OU structure, and report component-wise diagnostics for the projected (Stage 2 / PG–DPO) rules.
 To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
+Aligned $P_{0}$
+$s_{0}$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $6.31\text{\times}{10}^{-03}$ $5.19\text{\times}{10}^{-03}$ $4.01\text{\times}{10}^{-03}$ $3.54\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $4.71\text{\times}{10}^{-05}$ $5.10\text{\times}{10}^{-05}$ $1.39\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$ $10^{-3}$ DPO (distill.) $2.46\text{\times}{10}^{-03}$ $3.57\text{\times}{10}^{-03}$ $3.58\text{\times}{10}^{-03}$ $3.22\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $4.43\text{\times}{10}^{-05}$ $5.39\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$ $10^{-3}$ PPO (baseline) $7.78\text{\times}{10}^{-02}$ $1.03\text{\times}{10}^{-01}$ $4.20\text{\times}{10}^{00}$ $2.41\text{\times}{10}^{00}$ $10^{-2}$ DPO (basic) $5.61\text{\times}{10}^{-03}$ $4.83\text{\times}{10}^{-03}$ $3.83\text{\times}{10}^{-03}$ $3.50\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $5.03\text{\times}{10}^{-05}$ $4.54\text{\times}{10}^{-05}$ $1.45\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$ $10^{-2}$ DPO (distill.) $3.08\text{\times}{10}^{-03}$ $3.16\text{\times}{10}^{-03}$ $3.53\text{\times}{10}^{-03}$ $3.21\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $4.48\text{\times}{10}^{-05}$ $5.11\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$ $10^{-2}$ PPO (baseline) $7.30\text{\times}{10}^{-02}$ $9.15\text{\times}{10}^{00}$ $4.27\text{\times}{10}^{00}$ $2.43\text{\times}{10}^{00}$ $10^{-1}$ DPO (basic) $6.93\text{\times}{10}^{-03}$ $4.53\text{\times}{10}^{-03}$ $3.96\text{\times}{10}^{-03}$ $3.47\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (basic) $5.50\text{\times}{10}^{-05}$ $4.16\text{\times}{10}^{-05}$ $2.63\text{\times}{10}^{-04}$ $2.97\text{\times}{10}^{-04}$ $10^{-1}$ DPO (distill.) $2.85\text{\times}{10}^{-03}$ $3.70\text{\times}{10}^{-03}$ $3.65\text{\times}{10}^{-03}$ $3.28\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (distill.) $4.77\text{\times}{10}^{-05}$ $5.50\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.88\text{\times}{10}^{-04}$ $10^{-1}$ PPO (baseline) $7.34\text{\times}{10}^{-02}$ $1.00\text{\times}{10}^{-01}$ $4.22\text{\times}{10}^{00}$ $2.64\text{\times}{10}^{00}$
+Misaligned $P_{0}$
+$s_{0}$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $6.11\text{\times}{10}^{-03}$ $5.40\text{\times}{10}^{-03}$ $3.87\text{\times}{10}^{-03}$ $3.62\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $5.12\text{\times}{10}^{-05}$ $5.11\text{\times}{10}^{-05}$ $1.35\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$ $10^{-3}$ DPO (distill.) $3.51\text{\times}{10}^{-03}$ $2.93\text{\times}{10}^{-03}$ $3.56\text{\times}{10}^{-03}$ $3.21\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $4.36\text{\times}{10}^{-05}$ $5.48\text{\times}{10}^{-05}$ $1\text{\times}{10}^{-04} 1.44×1$$4\text{\times}{10}^{-04}$ $10^{-3}$ PPO (baseline) $8.96\text{\times}{10}^{-02}$ $9.28\text{\times}{10}^{-02}$ $4.37\text{\times}{10}^{00}$ $2.46\text{\times}{10}^{00}$ $10^{-2}$ DPO (basic) $6.17\text{\times}{10}^{-03}$ $4.89\text{\times}{10}^{-03}$ $3.89\text{\times}{10}^{-03}$ $3.57\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $4.99\text{\times}{10}^{-05}$ $5.96\text{\times}{10}^{-05}$ $1.54\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$ $10^{-2}$ DPO (distill.) $2.84\text{\times}{10}^{-03}$ $3.97\text{\times}{10}^{-03}$ $3.68\text{\times}{10}^{-03}$ $3.21\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $4.63\text{\times}{10}^{-05}$ $6.31\text{\times}{10}^{-05}$ $1.54\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$ $10^{-2}$ PPO (baseline) $8.02\text{\times}{10}^{-02}$ $7.75\text{\times}{10}^{00}$ $4.47\text{\times}{10}^{00}$ $2.53\text{\times}{10}^{00}$ $10^{-1}$ DPO (basic) $6.09\text{\times}{10}^{-03}$ $4.17\text{\times}{10}^{-03}$ $3.97\text{\times}{10}^{-03}$ $3.56\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (basic) $5.63\text{\times}{10}^{-05}$ $2.44\text{\times}{10}^{-04}$ $3.24\text{\times}{10}^{-04}$ $3.29\text{\times}{10}^{-04}$ $10^{-1}$ DPO (distill.) $3.22\text{\times}{10}^{-03}$ $3.08\text{\times}{10}^{-03}$ $3.65\text{\times}{10}^{-03}$ $3.18\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (distill.) $5.31\text{\times}{10}^{-05}$ $2.46\text{\times}{10}^{-04}$ $3.22\text{\times}{10}^{-04}$ $3.21\text{\times}{10}^{-04}$ $10^{-1}$ PPO (baseline) $6.30\text{\times}{10}^{-02}$ $1.08\text{\times}{10}^{-01}$ $4.26\text{\times}{10}^{00}$ $2.46\text{\times}{10}^{00}$
+Table 2: Decision-time RMSE at $t=0$ in the OU premium benchmark (tail median over the last six evaluation checkpoints), sweeping the decision-time uncertainty scale $s_{0}$ in $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$ under both aligned (66) and misaligned (67) geometries.
+DPO rows report the Stage 1 warm-up (deployable) policy output.
+PG–DPO rows report the Stage 2 post-hoc $q_{0}$-aggregated Pontryagin projection output (residual/control-variate implementation; Section G.1) computed under the corresponding warm policy.
+“Distill.” rows use interactive distillation (Section G.2) to amortize the projected teacher into a deployable network.
+RMSE is computed identically for all methods (no method-specific or post-hoc scale correction).
+The PPO baseline is frequently unstable in this terminal-only, $Y$-blind setting; very large entries (e.g., $s_{0}=10^{-2},\,d=10$) reflect genuine training divergence rather than typographical error.
 Metrics and diagnostics.
-Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE at $t=0$ for all deployable policies, including the PPO baseline.
+Table 2 reports the full decision-time RMSE at $t=0$ for all deployable policies, including the PPO baseline.
 To isolate the economic hedging channel, we use the OU-induced decomposition of the analytic reference
@@
 where $\pi^{\mathrm{myo}}$ is the independence-case reference obtained by setting $\rho=0$ (equivalently, dropping $M_{\mathrm{cross}}(\tau)$ in (37)), and $\pi^{\mathrm{hedge}}:=\pi^{\mathrm{ref}}-\pi^{\mathrm{myo}}$ is the residual $\rho$-dependent term.
-We report the hedging-component RMSE for the projected (Stage 2) rules in Table 1; the myopic-component RMSE and hedging-direction cosine similarity are deferred to Appendix H.
-Since Stage 1 and PPO do not expose a compatible myopic/hedging split under our diagnostic protocol, component-wise diagnostics are reported only for Stage 2.
-$s_{0}$ Method $d=5$ $10$ $50$ $100$
-Aligned $P_{0}$
-$10^{-3}$ Stage 1+Stage 2 (Basic) $4.59\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
-Stage 1+Stage 2 (Distill.) $4.39\text{\times}{10}^{-05}$ $5.25\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$
-$10^{-2}$ Stage 1+Stage 2 (Basic) $98\text{\times}{10}^{-05} 4.4$$47\text{\times}{10}^{-05}$ $1.44\text{\times}{10}^{-04}$ $1.74\text{\times}{10}^{-04}$
-Stage 1+Stage 2 (Distill.) $4.43\text{\times}{10}^{-05}$ $4.95\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
-$10^{-1}$ Stage 1+Stage 2 (Basic) $5.27\text{\times}{10}^{-05}$ $3.99\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.95\text{\times}{10}^{-04}$
-Stage 1+Stage 2 (Distill.) $4.72\text{\times}{10}^{-05}$ $5.36\text{\times}{10}^{-05}$ $2.58\text{\times}{10}^{-04}$ $2.86\text{\times}{10}^{-04}$
-Misaligned $P_{0}$
-$10^{-3}$ Stage 1+Stage 2 (Basic) $5.00\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.34\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$
-Stage 1+Stage 2 (Distill.) $4.30\text{\times}{10}^{-05}$ $5.33\text{\times}{10}^{-05}$ $1.40\text{\times}{10}^{-04}$ $1.43\text{\times}{10}^{-04}$
-$10^{-2}$ Stage 1+Stage 2 (Basic) $4.93\text{\times}{10}^{-05}$ $5.65\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$
-Stage 1+Stage 2 (Distill.) $4.56\text{\times}{10}^{-05}$ $6.09\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$
-$10^{-1}$ Stage 1+Stage 2 (Basic) $5.45\text{\times}{10}^{-05}$ $1.55\text{\times}{10}^{-04}$ $3.19\text{\times}{10}^{-04}$ $3.28\text{\times}{10}^{-04}$
-Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-05}$ $1.57\text{\times}{10}^{-04}$ $3.13\text{\times}{10}^{-04}$ $3.20\text{\times}{10}^{-04}$
-Table 1: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the post-hoc projected (Stage 2) policies (tail median over the last six evaluation checkpoints).
-We use the decomposition $\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau)$ where $\pi^{\mathrm{myo}}$ is the $\rho=0$ reference and $\pi^{\mathrm{hedge}}$ is the residual $\rho$-dependent term (cf.
-(43)).
-Component-wise diagnostics are reported for Stage 2 since Stage 1 and PPO do not explicitly output a compatible myopic/hedging split under our protocol.
+We report the hedging-component RMSE for the projected (Stage 2 / PG–DPO) rules in Table 3; the myopic-component RMSE and hedging-direction cosine similarity are deferred to Appendix I.
+Since Stage 1 (DPO) and PPO do not expose a compatible myopic/hedging split under our diagnostic protocol, component-wise diagnostics are reported only for the projected rules.
 Projection and economic hedging-demand recovery.
-Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (Stage 1+Stage 2) substantially reduces decision-time RMSE relative to the deployable Stage 1 policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
-For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, Stage 1 attains $3.54\text{\times}{10}^{-03}$ whereas Stage 1+Stage 2 achieves $1.56\text{\times}{10}^{-04}$.
+Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (PG–DPO) substantially reduces decision-time RMSE relative to the deployable Stage 1 warm-up (DPO) policy (Table 2).
+For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, DPO attains $3.54\text{\times}{10}^{-03}$ whereas PG–DPO achieves $1.56\text{\times}{10}^{-04}$.
 The component-wise diagnostics indicate that the remaining discrepancy is largely driven by the hedging channel: in the same setting, the hedging RMSE is $1.55\text{\times}{10}^{-04}$ (basic) and $1.42\text{\times}{10}^{-04}$ (distill.) (Table 3), while the myopic RMSE is an order of magnitude smaller (Appendix Table 4).
@@
 Amortization, robustness, and the PPO baseline.
-Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
-This matches the intended division of labor in Section F: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
-As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables LABEL:tab:ou_rmse_s0_sweep_landscape1).
+Interactive distillation improves the deployable Stage 1 policy relative to basic Stage 1 (DPO), while the most accurate object remains the post-hoc projected policy (PG–DPO) (Table 2).
+This matches the intended division of labor in Section G: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
+As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables 23).
 Misalignment has a limited effect for small and moderate uncertainty scales, but can induce noticeable deterioration in the hardest settings, where the direction-of-hedge diagnostic can also weaken (Appendix Table 5).
-Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, and can be unstable in this terminal-only setting (Table LABEL:tab:ou_rmse_s0_sweep_landscape).
+$s_{0}$ Method $d=5$ $10$ $50$ $100$ Aligned $P_{0}$ $10^{-3}$ PG–DPO (basic) $4.59\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $10^{-3}$ PG–DPO (distill.) $4.39\text{\times}{10}^{-05}$ $5.25\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (basic) $98\text{\times}{10}^{-05} 4.4$$47\text{\times}{10}^{-05}$ $1.44\text{\times}{10}^{-04}$ $1.74\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (distill.) $4.43\text{\times}{10}^{-05}$ $4.95\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (basic) $5.27\text{\times}{10}^{-05}$ $3.99\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.95\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (distill.) $4.72\text{\times}{10}^{-05}$ $5.36\text{\times}{10}^{-05}$ $2.58\text{\times}{10}^{-04}$ $2.86\text{\times}{10}^{-04}$ Misaligned $P_{0}$ $10^{-3}$ PG–DPO (basic) $5.00\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.34\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$ $10^{-3}$ PG–DPO (distill.) $4.30\text{\times}{10}^{-05}$ $5.33\text{\times}{10}^{-05}$ $1.40\text{\times}{10}^{-04}$ $1.43\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (basic) $4.93\text{\times}{10}^{-05}$ $5.65\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (distill.) $4.56\text{\times}{10}^{-05}$ $6.09\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (basic) $5.45\text{\times}{10}^{-05}$ $1.55\text{\times}{10}^{-04}$ $3.19\text{\times}{10}^{-04}$ $3.28\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (distill.) $5.20\text{\times}{10}^{-05}$ $1.57\text{\times}{10}^{-04}$ $3.13\text{\times}{10}^{-04}$ $3.20\text{\times}{10}^{-04}$
+Table 3: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the projected (Stage 2 / PG–DPO) policies (tail median over the last six evaluation checkpoints).
+We use the decomposition $\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau)$ where $\pi^{\mathrm{myo}}$ is the $\rho=0$ reference and $\pi^{\mathrm{hedge}}$ is the residual $\rho$-dependent term (cf.
+(37)).
+Component-wise diagnostics are reported for projected rules since Stage 1 (DPO) and PPO do not explicitly output a compatible myopic/hedging split under our protocol.
+Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, and can be unstable in this terminal-only setting (Table 2).
 This is consistent with PPO facing a difficult credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline.
@@
 6 Conclusion
-We studied continuous-time portfolio choice in diffusion markets whose coefficients are estimated and therefore subject to statistical uncertainty (Section 2.1).
-We model this uncertainty by an exogenous law $q(d\theta)$ over a latent parameter $\theta$ that is drawn once at time 0 and remains fixed over the investment horizon, while the investor must deploy a single $\theta$-blind Markov feedback policy evaluated under an ex–ante CRRA objective (Remark 1, Section 2.1).
-This information structure shifts the relevant optimality notion from $\theta$-conditional (full-information) criticality to a $q$-aggregated Pontryagin first-order condition that is enforceable within the deployable $\theta$-blind policy class (Section 2.2, Theorem 1).
-Methodologically, we extended Pontryagin–Guided Direct Policy Optimization (PG–DPO) to the latent-parameter setting by sampling $\theta$ only inside the simulator and computing exact discrete-time gradients via BPTT (Section 3.1), and we leveraged the BPTT–PMP correspondence to extract the costate objects needed for structured control updates (Theorem 2).
-Building on the $q$-aggregated stationarity, we proposed uncertainty-aware projected PG–DPO (P–PGDPO), which aggregates Monte Carlo Pontryagin quantities across $\theta\sim q$ and projects them onto the deployable first-order condition to obtain a single $\theta$-blind rule (Section 3.2).
-We established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3).
-In experiments, projection provides the dominant gains—improving stability and decision-time accuracy in high-dimensional drift-uncertainty scaling tests and recovering intertemporal hedging demand in factor-driven markets—while interactive distillation amortizes projection into a fast deployable network (Sections 4 and 5; Section F).
+We studied continuous-time CRRA portfolio choice in diffusion markets whose coefficients are estimated and therefore statistically uncertain (Section 2.1).
+Our modeling choice is to treat estimation risk as an exogenous decision-time input, represented by a law $q(d\theta)$ over a latent parameter $\theta$: Nature draws $\theta$ once at time 0 and keeps it fixed over the investment horizon, while the investor does not observe $\theta$ and must deploy a single $\theta$-blind Markov feedback policy evaluated under an ex–ante objective (Remark 1).
+Under this information structure, the appropriate optimality notion is not the infeasible $\theta$-conditional (full-information) criticality, but rather a $q$-aggregated Pontryagin first-order condition that is enforceable within the deployable $\theta$-blind policy class (Section 2.2, Theorem 1).
+Methodologically, we proposed a simulation-only two-stage solver, PG–DPO, that matches this deployability constraint.
+In Stage 1, DPO optimizes the ex–ante objective by sampling $\theta$ only inside the simulator and performing stochastic gradient ascent via exact discrete-time pathwise gradients computed by BPTT through an Euler discretization (Section 3.1).
+We then leverage a conditional BPTT–PMP correspondence to extract the costate objects needed for structured control updates, including the second-adjoint blocks entering the portfolio Hamiltonian gradient (Theorem 2).
+In Stage 2, we apply a $q$-aggregated Pontryagin projection: we aggregate Monte Carlo Pontryagin quantities across $\theta\sim q$ and project onto the deployable $q$-aggregated stationarity condition to obtain a single $\theta$-blind rule (Section 3.2).
+On the theory side, we established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3).
+Empirically, across benchmarks the projection step provides the dominant gains, improving stability and decision-time accuracy in high-dimensional drift-uncertainty scaling tests and recovering intertemporal hedging demand in factor-driven markets, while interactive distillation amortizes the projected rule into a fast deployable network (Sections 4 and 5; Section G).
 Several extensions are natural.
-A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to belief-aware decision rules (Remark 2, Appendix A).
-A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section B; Section 4.2; Appendix E.5).
+A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to decision-time plug-in replanning and belief-aware decision rules (Remark 3, Appendix B).
+A second direction is to incorporate realistic frictions and constraints—such as transaction costs, leverage, and short-sale limits—and develop certified or regularized projection steps for regimes where mixed-moment estimation becomes fragile (Section C; Section 4.2; Appendix F.5).
 Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
@@
 This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
-Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
+Appendix A Derivations for Gaussian decision-time references
+This appendix derives the closed-form decision-time constant-portfolio references stated in Section 2.3.
+Throughout, we fix a decision time $t$ and hold the input law $q_{t}$ fixed over the remaining horizon $[t,T]$ (plug-in replanning).
+A.1 Static latent drift: tilted FOC and Gaussian shrinkage
+Consider the static latent-drift model (24)–(26) and restrict to constant portfolio fractions $\pi_{s}\equiv\pi$ on $[t,T]$. Let $\tau:=T-t$ and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$.
+Step 1: conditional evaluation under fixed $\theta$.
+For constant $\pi$, conditional on $\theta$ the wealth satisfies
+$\log\frac{X_{T}^{\pi}}{x}=\Big(r+\pi^{\top}\theta-\tfrac{1}{2}\pi^{\top}\Sigma\pi\Big)\tau+\pi^{\top}\Sigma^{1/2}(W_{T}-W_{t}),$
+hence $X_{T}^{\pi}$ is lognormal and
+$\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\mathcal{F}_{t}\big]=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)r\tau+(1-\gamma)\tau\,\pi^{\top}\theta-\tfrac{1}{2}\gamma(1-\gamma)\tau\,\pi^{\top}\Sigma\pi\Big\}.$ (68)
+Step 2: ex-ante mixing over $q_{t}$.
+Mixing over $\theta\sim q_{t}$ yields
+$J_{t}(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)r\tau-\tfrac{1}{2}\gamma(1-\gamma)\tau\,\pi^{\top}\Sigma\pi\Big\}\,M_{q_{t}}\big((1-\gamma)\tau\pi\big),$ (69)
+where $M_{q_{t}}(u):=\mathbb{E}_{\theta\sim q_{t}}[\exp(u^{\top}\theta)]$ is the mgf of $q_{t}$.
+Step 3: tilted first-order condition.
+Maximizing $J_{t}(\pi)$ is equivalent to maximizing $\log J_{t}(\pi)$. Differentiating (69) and using interior optimality gives the tilted FOC
+$\gamma\,\Sigma\,\pi=\nabla_{u}\log M_{q_{t}}(u)\Big|_{u=(1-\gamma)\tau\,\pi}.$ (70)
+Step 4: Gaussian input law.
+If $q_{t}=\mathcal{N}(m_{t},P_{t})$, then $M_{q_{t}}(u)=\exp\{u^{\top}m_{t}+\tfrac{1}{2}u^{\top}P_{t}u\}$ and $\nabla_{u}\log M_{q_{t}}(u)=m_{t}+P_{t}u$.
+Substituting into (70) yields the linear system
+$\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)\,\pi=m_{t},$ (71)
+and hence
+$\pi=\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)^{-1}m_{t}=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt(\gamma>1),$ (72)
+which matches (30) in the main text.
+A.2 OU premium factor: induced effective law and constant-portfolio reference
+We derive the OU-based reference stated in (37).
+Fix $t$ and let $\tau:=T-t$.
+Under (31)–(33), we take as decision-time input $Y_{t}\sim\mathcal{N}(m_{t},P_{t})$ and hold this input fixed over $[t,T]$.
+A.2.1 Integrated factor: mean and covariance
+Define $I_{t,T}:=\int_{t}^{T}Y_{s}\,ds$. Assume $K$ is positive stable (equivalently, K-K is Hurwitz), so that $e^{-Ku}$ decays as $u\to\infty$. For $u\geq 0$ define
+$\mathcal{A}(u):=K^{-1}\big(I_{m}-e^{-Ku}\big),$
+where $I_{m}$ is the $m\times m$ identity matrix. The OU solution is
+$Y_{t+u}=\bar{y}+e^{-Ku}(Y_{t}-\bar{y})+\int_{0}^{u}e^{-K(u-s)}\Xi\,dW_{t+s}^{Y}.$
+Integrating over $u\in[0,\tau]$ gives
+$I_{t,T}=\tau\bar{y}+\mathcal{A}(\tau)(Y_{t}-\bar{y})+\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\,dW_{t+u}^{Y}.$
+Taking expectation under the input law yields
+$m_{I}(t,\tau):=\mathbb{E}[I_{t,T}]=\tau\bar{y}+\mathcal{A}(\tau)\,(m_{t}-\bar{y}),$ (73)
+and the covariance decomposition gives
+$C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})=\mathcal{A}(\tau)\,P_{t}\,\mathcal{A}(\tau)^{\top}+\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\Xi^{\top}\,\mathcal{A}(u)^{\top}\,du.$ (74)
+A.2.2 Effective premium law
+Define the horizon-averaged effective premium
+$\bar{\theta}_{t,\tau}:=\frac{1}{\tau}B\,I_{t,T}.$
+Then $\bar{\theta}_{t,\tau}$ is Gaussian with
+$m_{\bar{\theta}}(t,\tau)=\frac{1}{\tau}B\,m_{I}(t,\tau),\hskip 18.49988ptP_{\bar{\theta}}(t,\tau)=\frac{1}{\tau^{2}}B\,C_{I}(t,\tau)\,B^{\top},$ (75)
+matching (36).
+A.2.3 Cross term induced by return–factor correlation
+Let $\Delta W:=W_{T}-W_{t}$. Using $d\langle W,W^{Y}\rangle_{s}=\rho\,ds$, we obtain
+$C_{IW}(\tau):=\mathrm{Cov}\!\big(I_{t,T},\,\Delta W\big)=\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\,\rho^{\top}\,du\;\in\;\mathbb{R}^{m\times d},$ (76)
+and the induced symmetric cross term
+$M_{\mathrm{cross}}(\tau):=B\,C_{IW}(\tau)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(\tau)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (77)
+A.2.4 Constant-portfolio CRRA objective and FOC
+Restrict to constant $\pi$ over $[t,T]$. Write the excess-return integral as
+$\int_{t}^{T}\pi^{\top}dR_{s}=\pi^{\top}BI_{t,T}+\pi^{\top}\Sigma^{1/2}\Delta W.$
+As in the static case, CRRA evaluation reduces to the log-mgf of a jointly Gaussian vector.
+Collecting the terms involving $\pi$ and applying the interior first-order condition yields the linear system
+$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1),$ (78)
+which matches (37) in the main text.
+Independence case.
+If $\rho=0$, then $C_{IW}(\tau)=0$ and $M_{\mathrm{cross}}(\tau)=0$. In this case, (78) reduces to the static Gaussian shrinkage formula applied to the effective law (75).
+Appendix B Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
 Purpose and scope.
@@
 In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (Barberis, 2000; Pástor, 2000; Xia, 2001).
-This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., Bensoussan and van Schuppen (1985); Pham and Wei (2017)), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$.
+This appendix records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter, and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$.
 We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization.
-In particular, our simulation-based Pontryagin-guided solvers developed later (Section 3) can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
-A linear–Gaussian hidden-premium model (OU state, observed returns).
-We use a stylized linear–Gaussian counterpart of the mean-reverting premium setting, but now assume the premium factor is not directly observed.
-Let $Y_{t}\in\mathbb{R}^{m}$ be a latent premium factor following an OU dynamics
-$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(\hat{y}_{0},P_{0}),$ (74)
-where $K\in\mathbb{R}^{m\times m}$ is stable, $\bar{y}\in\mathbb{R}^{m}$, and $\Xi\in\mathbb{R}^{m\times m}$. Risky assets satisfy
+In particular, the simulation-based Pontryagin-guided solvers developed in Section 3 can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
+A linear–Gaussian hidden-premium model (OU state, observed returns). Let $Y_{t}\in\mathbb{R}^{m}$ be a latent premium factor following an OU dynamics
+$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$ (79)
+where $K$ is positive stable (equivalently K-K is Hurwitz), $\bar{y}\in\mathbb{R}^{m}$, and $\Xi\in\mathbb{R}^{m\times m}$. Risky assets satisfy
 $\frac{dS_{t}}{S_{t}}=r\mathbf{1}\,dt+BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},\hskip 18.49988pt\Sigma\in\mathbb{R}^{d\times d}\ \text{s.p.d.},$ (80)
@@
 $d\hat{Y}_{t}$ $=K(\bar{y}-\hat{Y}_{t})\,dt+P_{t}B^{\top}\Sigma^{-1}\Big(dZ_{t}-B\hat{Y}_{t}\,dt\Big),$ (83)
-$\dot{P}_{t}$ $=KP_{t}+P_{t}K^{\top}+\Xi\Xi^{\top}-P_{t}B^{\top}\Sigma^{-1}BP_{t},\hskip 18.49988ptP_{0}\ \text{given}.$ (79)
-Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (79); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
-From a posterior on $Y_{t}$ to a Gaussian uncertainty description for decision-time optimization.
-To mirror the decision-time perspective of the OU benchmark, we consider the remaining-horizon time-averaged premium
-$\bar{\theta}_{t,T}:=\frac{1}{T-t}\int_{t}^{T}BY_{s}\,ds\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (80)
-For the OU dynamics (74), one has the decomposition
-$\int_{t}^{T}Y_{s}\,ds=\tau\bar{y}+K^{-1}\big(I-e^{-K\tau}\big)(Y_{t}-\bar{y})+\int_{t}^{T}K^{-1}\big(I-e^{-K(T-u)}\big)\Xi\,dW_{u}^{Y}.$ (81)
-Conditioning on $\mathcal{F}_{t}^{Z}$, the random variable $Y_{t}$ is distributed as $\mathcal{N}(\hat{Y}_{t},P_{t})$ by (77), while the future increments $(W_{u}^{Y}-W_{t}^{Y})_{u\geq t}$ are independent of $\mathcal{F}_{t}^{Z}$ in the independent-noise case.
-Hence $\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}$ is Gaussian:
-$\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}\!\big(m_{t,T},\,P_{t,T}\big),\hskip 18.49988ptm_{t,T}:=\frac{B\,m_{I}(t,T)}{\tau},\hskip 18.49988ptP_{t,T}:=\frac{1}{\tau^{2}}\,B\,C_{I}(t,T)\,B^{\top},$ (82)
+$\dot{P}_{t}$ $=-KP_{t}-P_{t}K^{\top}+\Xi\Xi^{\top}-P_{t}B^{\top}\Sigma^{-1}BP_{t},\hskip 18.49988ptP_{0}\ \text{given}.$ (84)
+Thus, even though $q_{t}$ is distribution-valued, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (84).
+From a posterior on $Y_{t}$ to a Gaussian uncertainty description for decision-time optimization. Define the remaining-horizon time-averaged premium
+$\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\int_{t}^{T}BY_{s}\,ds\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (85)
+Let $\mathcal{A}(u):=K^{-1}(I_{m}-e^{-Ku})$, where $I_{m}$ is the $m\times m$ identity matrix. The integrated factor admits the representation
+$\int_{t}^{T}Y_{s}\,ds=\tau\bar{y}+\mathcal{A}(\tau)\,(Y_{t}-\bar{y})+\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\,dW_{t+u}^{Y}.$ (86)
+Conditioning on $\mathcal{F}_{t}^{Z}$, we have $Y_{t}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}(\hat{Y}_{t},P_{t})$ by (82), and the future increments $(W_{t+u}^{Y}-W_{t}^{Y})_{u\geq 0}$ are independent of $\mathcal{F}_{t}^{Z}$ in the independent-noise case.
+Hence $\bar{\theta}_{t,\tau}\mid\mathcal{F}_{t}^{Z}$ is Gaussian:
+$\bar{\theta}_{t,\tau}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}\!\big(m_{t,\tau},\,P_{t,\tau}\big),\hskip 18.49988ptm_{t,\tau}:=\frac{B\,m_{I}(t,\tau)}{\tau},\hskip 18.49988ptP_{t,\tau}:=\frac{1}{\tau^{2}}\,B\,C_{I}(t,\tau)\,B^{\top},$ (87)
 where
-$m_{I}(t,T)$ $:=\tau\bar{y}+K^{-1}\big(I-e^{-K\tau}\big)(\hat{Y}_{t}-\bar{y}),$ (83)
-$C_{I}(t,T)$ $:=K^{-1}\big(I-e^{-K\tau}\big)P_{t}\big(I-e^{-K\tau}\big)^{\top}K^{-\top}+\int_{0}^{\tau}K^{-1}\big(I-e^{-Ks}\big)\Xi\Xi^{\top}\big(I-e^{-Ks}\big)^{\top}K^{-\top}\,ds.$ (84)
-Equation (82) provides a concrete example of an online-updated Gaussian uncertainty description $q_{t,T}:=\mathcal{L}(\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(m_{t,T},P_{t,T}).$
-A plug-in decision-time benchmark (receding-horizon fixed-$q_{t,T}$).
-Given $(m_{t,T},P_{t,T})$ from (82), a simple decision-time rule is obtained by treating $q_{t,T}$ as fixed over the remaining horizon and applying the Gaussian constant-allocation benchmark of Section 2.3.1 with horizon $\tau$:
-$\pi_{t}^{\mathrm{plug}}:=\Big(\gamma\,\Sigma+(\gamma-1)\tau\,P_{t,T}\Big)^{-1}m_{t,T}.$ (85)
-One may interpret (85) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (Barberis, 2000; Pástor, 2000; Xia, 2001).
+$m_{I}(t,\tau)$ $:=\tau\bar{y}+\mathcal{A}(\tau)\,(\hat{Y}_{t}-\bar{y}),$ (88)
+$C_{I}(t,\tau)$ $:=\mathcal{A}(\tau)\,P_{t}\,\mathcal{A}(\tau)^{\top}+\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\Xi^{\top}\,\mathcal{A}(u)^{\top}\,du.$ (89)
+Equation (87) thus provides an explicit example of an online-updated Gaussian uncertainty description $q_{t,\tau}:=\mathcal{L}(\bar{\theta}_{t,\tau}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(m_{t,\tau},P_{t,\tau}).$
+A plug-in decision-time benchmark (receding-horizon fixed-$q_{t,\tau}$).
+Given $(m_{t,\tau},P_{t,\tau})$ from (87), a simple decision-time rule is obtained by treating $q_{t,\tau}$ as fixed over the remaining horizon and applying the Gaussian constant-allocation benchmark of Section 2.3.1 with horizon $\tau$:
+$\pi_{t}^{\mathrm{plug}}:=\Big(\gamma\,\Sigma+(\gamma-1)\tau\,P_{t,\tau}\Big)^{-1}m_{t,\tau}.$ (90)
+One may interpret (90) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (Barberis, 2000; Pástor, 2000; Xia, 2001).
 Remarks (relation to belief-aware control).
-The plug-in rule (85) is intentionally decision-time: it conditions on the current uncertainty description and does not attempt to optimize over how the posterior will evolve.
-In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (Barberis, 2000; Pástor, 2000; Xia, 2001).
+The plug-in rule (90) is intentionally decision-time: it conditions on the current uncertainty description and does not optimize over how the posterior will evolve.
 Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
 Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (Haussmann, 1987; Li and Tang, 1995; Baghery and Øksendal, 2007).
-Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in $(x,y,\hat{Y},P)$-space (or its sufficient-statistic analogues) is an important direction that we defer to future work.
-Appendix B Why dynamic programming and deep PDE surrogates are not primary baselines under the $\theta$-blind fixed-$q$ interface
-This appendix explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted in this paper.
-DP is conceptually sound in low-dimensional Markovian settings (Fleming and Soner, 2006; Pham, 2009), but two obstructions dominate in our regime: (i) value-function approaches require accurate derivatives (often including mixed sensitivities) to recover optimal portfolios, and this becomes prohibitive as dimension and nonlinearity grow; (ii) principled parameter uncertainty magnifies both the state dimension and the aggregation difficulties.
-These points are standard for grid-based HJB solvers (Bellman, 1961; Kushner and Dupuis, 2001) and remain relevant for modern deep PDE surrogates (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), especially under partial information/belief-state formulations (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
-B.1 The value-function target for latent $\theta$ is a belief-state functional
+Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in belief-state space is an important direction that we defer to future work.
+Appendix C Why dynamic programming and deep PDE surrogates are not primary baselines under the $\theta$-blind fixed-$q$ interface
+This appendix clarifies why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as primary baselines for the high-dimensional uncertain markets targeted in this paper.
+The main obstruction is not that PDE solvers are “invalid”; rather, under our $\theta$-blind fixed-$q$ deployability interface, the natural DP/PDE/BSDE target is typically not a low-dimensional Markov value function.
+Derivative instability and high-dimensional conditioning issues are secondary to this target mismatch.
+C.1 The value-function target for latent $\theta$ is a belief-state functional
 Fixed-$q$ ex–ante objective and $\theta$-blind deployability.
 Our decision-time interface is
-$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}J(\pi)\;:=\;\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right]\right],$ (86)
-where $\theta\sim q$ is latent, fixed along each episode, and $\mathcal{A}^{\mathrm{blind}}$ denotes admissible controls measurable with respect to the observable filtration (so the deployed policy cannot take $\theta$ as an input).
-Why $(t,X_{t},Y_{t})$ is not a closed DP state.
-In general, the observable history carries information about $\theta$, so the continuation value at time $t$ depends on the conditional law
+$\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}J(\pi)\;:=\;\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}\mathbb{E}_{\theta\sim q}\!\left[\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right]\right],$ (91)
+where $\theta\sim q$ is latent, fixed along each episode, and $\mathcal{A}^{\mathrm{obs}}$ denotes the class of deployable controls, i.e.
+controls measurable w.r.t.
+the observable filtration (so deployed policies cannot condition on $\theta$).
+(One may further restrict to the feedback subclass $\mathcal{A}^{\mathrm{fb}}$ as defined in Section 2.1.)
+Why $(t,X_{t},Y_{t})$ is not a closed DP state under latent $\theta$.
+Even if the prior $q$ is fixed at decision time (and even if the controller is constrained to be $\theta$-blind), the conditional law of $\theta$ given observations typically evolves as data arrive.
+Formally, the continuation value at time $t$ depends on the conditional law
 $q_{t}(\cdot):=\mathcal{L}\!\big(\theta\mid\mathcal{F}_{t}^{\mathrm{obs}}\big),$ (92)
 not only on $(t,X_{t},Y_{t})$.
-This is the standard partially observed control phenomenon: DP becomes Markov only after augmenting the state by a belief (or a finite-dimensional sufficient statistic when available) (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+This is the standard partially observed control phenomenon: the unconditional (mixed-over-$\theta$) law of the observed state is generally not Markov in $(X_{t},Y_{t})$, and DP becomes Markov only after augmenting the state by a belief (or a finite-dimensional sufficient statistic when available) (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
 Belief-state DP is infinite-dimensional under our interface.
-The principled DP value function is therefore a functional of a probability measure:
-$V(t,x,y,\mu):=\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}^{t,x,y,\mu}\!\left[U(X_{T}^{\pi})\right],\hskip 18.49988pt\mu\in\mathcal{P}(\Theta),$ (88)
-so the associated HJB equation is posed on a space of measures in general (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
-Under our inference-agnostic interface, $q$ need not belong to a conjugate family and we do not specify a filtering model that would keep $q_{t}$ in a tractable finite-dimensional class.
-Consequently, a DP/PDE baseline would be forced to approximate $V$ over an infinite-dimensional belief state, which is not the computational target of this paper.
-B.2 Even with deterministic parameters, HJB solvers face a derivative bottleneck in high dimension
-Classical HJB: curse of dimensionality and full nonlinearity.
-With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (Fleming and Soner, 2006; Pham, 2009).
-Grid-based solvers scale exponentially in the state dimension (Bellman, 1961; Kushner and Dupuis, 2001).
-In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points.
-Moreover, realistic features (constraints, transaction costs, non-affine dynamics) typically yield fully nonlinear HJBs.
-Recovering the optimal portfolio from a value function also requires derivatives (schematically, in Merton-type problems)
+The principled value object is therefore a functional of a probability measure:
+$V(t,x,y,\eta):=\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}\mathbb{E}^{t,x,y,\eta}\!\left[U(X_{T}^{\pi})\right],\hskip 18.49988pt\eta\in\mathcal{P}(\Theta),$ (93)
+where $\mathbb{E}^{t,x,y,\eta}$ denotes the expectation for a system started from $(t,x,y)$ with latent parameter $\theta\sim\eta$.
+Hence the associated DP/HJB is posed on a (typically infinite-dimensional) space of measures in general (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
+Under our inference-agnostic interface, $q$ need not be conjugate and we do not specify a filtering model that would keep $q_{t}$ in a tractable finite-dimensional family.
+Consequently, any value-function baseline that assumes a low-dimensional Markov target is, by construction, solving a different object from the $\theta$-blind fixed-$q$ problem studied in this paper.
+C.2 Even with deterministic parameters, value-function baselines are indirect under our interface
+When parameters are deterministic, DP leads to an HJB for a Markov value function $V(t,x,y)$ (Fleming and Soner, 2006; Pham, 2009).
+In portfolio problems, however, optimal feedback rules are typically recovered through nonlinear operations on derivatives of $V$, e.g.
 $\pi^{\star}(t,x,y)\;\propto\;-\frac{V_{x}(t,x,y)}{V_{xx}(t,x,y)}\times(\text{risk model})^{-1}\times(\text{drift model}),$ (94)
-and in factor-driven settings intertemporal hedging demands are driven by mixed sensitivities (e.g.
-$V_{xy}$).
-Thus the numerical burden is not only approximating $V$ but also obtaining stable first/second/mixed derivatives in high dimension.
-B.3 Deep PDE surrogates remove grids but not the value-function derivative bottleneck
-PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities that drive intertemporal hedging.
-In practice this can induce nonconvex and ill-conditioned training objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable high-dimensional portfolio policies.
-This is not a claim that deep PDE methods are “invalid”; rather, under our high-dimensional and $\theta$-blind interface, they are not practical baselines relative to direct, simulation-only policy optimization methods.
-B.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
-A natural alternative is to condition on $\theta$ and solve a family of full-information problems.
-$\theta$-conditional HJB.
+with intertemporal hedging driven by mixed sensitivities such as $V_{xy}$.
+This “derivative bottleneck” is a practical issue for grid-based solvers (Bellman, 1961; Kushner and Dupuis, 2001) and also affects neural surrogates, but it is not the core reason we exclude DP/PDE baselines here.
+The core reason is the belief-state mismatch in Section C.1.
+C.3 Deep PDE surrogates (PINNs / deep BSDE) do not resolve the target mismatch
+PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019; Huré et al., 2020).
+However, both remain value-function-based PDE/BSDE solvers: they implicitly assume a (finite-dimensional) Markov target for the value/decoupling field.
+Under latent $\theta$, a solver must therefore choose between:
+  • learning $\theta$-conditional objects (making the policy/value depend on $\theta$, hence not deployable); or
+  • enforcing a $\theta$-agnostic Markov representation (e.g.
+a single $V(t,x,y)$) by sampling $\theta$ only inside the fitting loss, which generally does not correspond to the $\theta$-blind ex–ante optimum (91).
+The next two subsections make this mismatch explicit.
+C.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
+$\theta$-conditional HJB / BSDE.
 If $\theta$ were observed, the full-information value function $V^{\theta}$ solves
 $0=\partial_{t}V^{\theta}(t,x,y)+\sup_{\pi\in\mathbb{R}^{d}}\Big\{\mathcal{L}^{\pi,\theta}V^{\theta}(t,x,y)\Big\},\hskip 18.49988ptV^{\theta}(T,x,y)=U(x),$ (95)
-for the controlled generator $\mathcal{L}^{\pi,\theta}$ (Fleming and Soner, 2006; Pham, 2009).
-This yields a $\theta$-dependent optimizer $\pi^{\star}(t,x,y,\theta)$, which is infeasible in our deployable $\theta$-blind setting.
-Supremum–expectation non-commutativity.
-More fundamentally, even perfect knowledge of the $\theta$-conditional optimizer does not solve the $\theta$-blind objective, because
-$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[J(\pi,\theta)\right]\;\leq\;\mathbb{E}_{\theta\sim q}\!\left[\sup_{\pi}J(\pi,\theta)\right],\hskip 18.49988ptJ(\pi,\theta):=\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right],$ (91)
-and the inequality is typically strict. The right-hand side corresponds to the (infeasible) $\theta$-adaptive controller.
-B.5 Why “sample $\theta$ only in the PDE loss” is mismatched to the ex–ante objective
-One might try to train a single value network $V_{\varphi}(t,x,y)$ that does not take $\theta$ as input, and sample $\theta\sim q$ only inside a collocation/pathwise residual loss:
+(Fleming and Soner, 2006; Pham, 2009), yielding a $\theta$-dependent optimizer $\pi^{\star}(t,x,y,\theta)$. This is infeasible under our $\theta$-blind deployability restriction.
+Supremum–expectation non-commutativity is the structural obstruction.
+Even perfect access to $\theta$-conditional optimizers does not solve the $\theta$-blind objective because
+$\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}\mathbb{E}_{\theta\sim q}\!\left[J^{\theta}(\pi)\right]\;\leq\;\mathbb{E}_{\theta\sim q}\!\left[\sup_{\pi}J^{\theta}(\pi)\right],\hskip 18.49988ptJ^{\theta}(\pi):=\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right],$ (96)
+and the inequality is typically strict.
+The right-hand side corresponds to an (infeasible) $\theta$-adaptive controller.
+Equivalently, our Pontryagin stationarity is $q$-aggregated over $\theta$-blind variations, not enforced parameter-by-parameter.
+C.5 Why “sampling $\theta$ only in the PDE/BSDE loss” is mismatched to the ex–ante objective
+A common workaround is to learn a single value/decoupling network that does not take $\theta$ as input, and sample $\theta\sim q$ only inside a training loss.
+For PINNs this appears as an averaged collocation residual; for deep BSDE methods it appears as an averaged pathwise BSDE/HJB fitting objective.
+Schematically,
 $\min_{\varphi}\;\mathbb{E}_{\theta\sim q}\!\Big[\big\|\mathrm{HJB}_{\theta}\!\big(V_{\varphi}\big)(t,x,y)\big\|^{2}\Big],\hskip 18.49988pt\text{(schematic)}$ (97)
-where $\mathrm{HJB}_{\theta}(\cdot)$ is the full-information HJB operator for parameter $\theta$. This is appealing but generally mismatched to (86):
-  • Different $\theta$ impose different (often conflicting) optimality conditions at the same $(t,x,y)$, so there is generally no single $V$ satisfying $\mathrm{HJB}_{\theta}(V)=0$ for many $\theta$ simultaneously.
-  • Minimizing an expected squared PDE residual (92) is not equivalent to maximizing expected terminal utility under $\theta$-blind controls.
-  • Portfolio feedback rules are typically extracted through nonlinear operations on derivatives (ratios, inverses), so small value-function approximation errors can induce large control errors.
-B.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
-The aggregation issue is visible even in a one-asset Merton model with static latent drift. Let
+where $\mathrm{HJB}_{\theta}(\cdot)$ denotes the full-information HJB operator for parameter $\theta$.
+The issue is conceptual (target mismatch), not merely numerical:
+  • Conflicting optimality conditions.
+Different $\theta$ impose different local optimality conditions at the same $(t,x,y)$, so there is generally no single low-dimensional Markov $V$ (or decoupling field) that simultaneously satisfies $\mathrm{HJB}_{\theta}(V)=0$ for many $\theta$.
+  • Loss mismatch. Minimizing an averaged PDE/BSDE fitting loss such as (97) is not equivalent to maximizing expected terminal utility under $\theta$-blind controls in (91).
+  • Wrong stationarity notion.
+Our deployable optimum is characterized by a $q$-aggregated first-order condition with $\theta$-blind admissible variations; averaging full-information HJB/BSDE conditions across $\theta$ targets a different stationarity concept.
+Therefore, while $\theta$-sampling can be a sensible approximation heuristic, it does not produce a principled DP/PDE/BSDE baseline aligned with the fixed-$q$, $\theta$-blind optimality notion studied in this paper.
+C.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
+The aggregation issue is visible even in a one-asset Merton model with static latent drift:
 $\frac{dS_{t}}{S_{t}}=r\,dt+\theta\,dt+\sigma\,dW_{t},\hskip 18.49988pt\theta\sim\mathcal{N}(m,P),$
-and restrict to constant fractions $\pi$ over a remaining horizon $\tau$.
-For CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ with $\gamma>1$, the $\theta$-conditional full-information constant Merton rule is
+and constant fractions $\pi$ over remaining horizon $\tau$. For CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$, $\gamma>1$, the $\theta$-conditional full-information rule is
 $\pi^{\star}(\theta)=\frac{\theta}{\gamma\sigma^{2}}.$ (98)
-Averaging this infeasible rule gives $\mathbb{E}[\pi^{\star}(\theta)]=m/(\gamma\sigma^{2})$. By contrast, the deployable $\theta$-blind ex–ante optimizer within the constant-fraction class is
+Averaging gives $\mathbb{E}[\pi^{\star}(\theta)]=m/(\gamma\sigma^{2})$. By contrast, the deployable $\theta$-blind ex–ante optimizer within the constant-fraction class is
 $\pi_{q}^{\mathrm{const}}(\tau)=\frac{m}{\gamma\sigma^{2}+(\gamma-1)\tau\,P},\hskip 18.49988pt(\gamma>1),$ (99)
-which exhibits shrinkage due to estimation uncertainty ($p>0$) and depends on the horizon $\tau$. Thus $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\mathrm{const}}(\tau)$ whenever $p>0$.
-B.7 Implication for baselines in this paper
-The discussion above motivates our baseline choices in the main text:
-  • We validate on controlled drift/premium benchmarks where closed-form decision-time Gaussian targets are available and match the fixed-$q$ deployability restriction.
-  • For learning-based comparisons, we emphasize simulation-only direct policy search baselines (e.g., PPO) that operate under the same observation restriction (no access to $\theta$; $\theta$ sampled only inside the simulator), rather than value-function PDE/BSDE solvers whose natural DP target is belief-augmented.
-Appendix C Proof of Theorem 2
-Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in Huh et al.
-(2025b)).
+which exhibits shrinkage and horizon dependence. Thus $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\mathrm{const}}(\tau)$ whenever $P>0$.
+C.7 Implication for baselines in this paper
+The discussion above motivates our baseline choices:
+  • We validate on controlled decision-time Gaussian benchmarks that match the fixed-$q$ deployability restriction.
+  • For learning-based comparisons, we emphasize simulation-only direct policy optimization methods that operate under the same observation restriction (no access to $\theta$; $\theta$ sampled only inside the simulator), rather than value-function PDE/BSDE solvers whose natural DP target is belief-augmented.
+Appendix D Proof of Theorem 2
+Proof of Theorem 2.
+Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see Huh et al.
+(2025)).
 Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.
-(This appendix replaces the current Appendix C in the draft.)
 Important remark (what this proof does not use).
 This proof concerns the $\theta$-conditional Pontryagin adjoint/costate for the fixed-$\theta$ control problem induced by (39)–(40).
-It does not use the $\theta$-blind $q$-aggregated stationarity condition (Theorem 1 in Section 3.2).
-Those constructions affect only stage 2 projection targets and are irrelevant to the BPTT–PMP convergence itself.
+It does not use the $q$-aggregated $\theta$-blind stationarity condition (e.g.
+Theorem 1), which enters only the stage 2 projection target.
 Notation, filtration, and the pathwise-vs-adapted issue. Fix $\theta\in\Theta$. We work conditionally on this $\theta$ and consider the augmented (simulator) filtration
-$\mathbb{G}^{\theta}:=(G_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988ptG_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
-All conditional expectations and $L^{2}$ projections below are taken with respect to $G_{t_{k}}^{\theta}$.
+$\mathbb{G}^{\theta}:=(\mathcal{G}_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988pt\mathcal{G}_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
+All conditional expectations and conditional $L^{2}$ projections below are taken w.r.t. $\mathcal{G}_{t_{k}}^{\theta}$.
 Key point.
 The raw BPTT “adjoints” are pathwise derivatives of the terminal utility along the full discrete computation graph.
-Hence, for $k<N$, they are typically $G_{T}^{\theta}$-measurable and need not be $G_{t_{k}}^{\theta}$-measurable (they can depend on future increments).
+Hence, for $k<N$, they are typically $\mathcal{G}_{T}^{\theta}$-measurable and need not be $\mathcal{G}_{t_{k}}^{\theta}$-measurable (they can depend on future increments).
 In contrast, the stochastic maximum principle (SMP/PMP) costate is $\mathbb{G}^{\theta}$-adapted.
 Therefore, the correspondence is formulated after a standard one-step conditional $L^{2}$ projection (equivalently, a discrete-time BSDE representation), which turns pathwise objects into adapted discrete adjoints.
-Let $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, $N\Delta t=T$.
-For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind but evaluated along the $\theta$-conditional trajectory.
-Step 1: Conditioning on $\theta$ and uniformity of bounds.
-Fix a compact set $K\subset\Theta$.
-Assume the coefficients in (45)–(46) satisfy the usual Lipschitz and linear-growth conditions uniformly over $\theta\in K$, and that the block covariance structure of $(W,W^{Y})$ (including instantaneous correlation) is uniformly nondegenerate on $K$.
-Assume also the regularity/integrability needed for the $\theta$-conditional SMP (including the blocks relevant for the portfolio Hamiltonian gradient) to be well posed.
-Then, for each fixed $\theta\in K$, the controlled SDE system is well posed and admits uniform-in-time $L^{2}$ moment bounds.
+Let $W$ be $d_{W}$-dimensional and $W^{Y}$ be $d_{Y}$-dimensional (allowing instantaneous correlation).
+Fix $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, with $N\Delta t=T$.
+For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\Delta t,\theta},Y_{k}^{\Delta t,\theta})$, where $\pi_{\varphi}$ is $\theta$-blind.
+Standing assumptions (smooth Markov regime; uniform on compacts). Fix a compact set $\Theta_{0}\subset\Theta$. Assume the following hold uniformly for $\theta\in\Theta_{0}$:
+  • Uniform well-posedness and moments.
+The coefficients in (39)–(40) satisfy the usual Lipschitz and linear-growth conditions (in the state variables), yielding unique strong solutions and uniform $L^{2}$ moment bounds.
+  • Uniform smoothness.
+The coefficients are sufficiently smooth in state variables so that the $\theta$-conditional adjoint BSDE is well posed and the state-derivative blocks used in the portfolio Hamiltonian gradient exist and are square-integrable.
+(It is sufficient to assume $a,b,\beta,\sigma$ are $C^{2}$ in state variables with derivatives obeying uniform-in-$\theta$ growth/Lipschitz bounds on $\Theta_{0}$.)
+  • Policy regularity. The deployed feedback $(t,x,y)\mapsto\pi_{\varphi}(t,x,y)$ is $C^{2}$ in $(x,y)$ on the working domain, with the required derivatives uniformly bounded there.
+  • Smooth decoupling field (identifying the blocks used here).
+In the smooth Markov regime, the $\theta$-conditional adjoint admits a decoupling field $p_{t}^{\theta}=u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, and we may identify $p_{x,t}^{\theta}:=\partial_{x}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, $p_{y,t}^{\theta}:=\partial_{y}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$.
+These are precisely the sensitivity blocks needed for the portfolio Hamiltonian gradient.
+  • One-step projection nondegeneracy.
+The conditional covariance matrix of $(\Delta W_{k},\Delta W_{k}^{Y})$ given $\mathcal{G}_{t_{k}}^{\theta}$ is uniformly nondegenerate on $\Theta_{0}$, so that the one-step conditional $L^{2}$-projection onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ is well defined with unique coefficients.
+These are the deterministic assumptions used in Huh et al. (2025), now imposed conditionally on $\theta$ and uniformly over $\theta\in\Theta_{0}$.
+Step 1: Conditioning on $\theta$ and uniformity of forward Euler error.
+Under the standing assumptions, for each fixed $\theta\in\Theta_{0}$ the controlled SDE system is well posed and admits uniform-in-time $L^{2}$ moment bounds.
 Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
-$\sup_{t\in[0,T]}\mathbb{E}\!\left[\left\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\right\|^{2}\right]^{1/2}\leq C_{K}\,\Delta t^{1/2},$
-with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are exactly the deterministic assumptions used in Huh et al. (2025b), now stated uniformly on $K$.
-Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme on the grid $(t_{k})$:
-$Y_{k+1}^{\theta}=Y_{k}^{\theta}+a(Y_{k}^{\theta},\theta)\Delta t+\beta(Y_{k}^{\theta},\theta)\Delta W_{k}^{Y},$
-$X_{k+1}^{\theta}=X_{k}^{\theta}+X_{k}^{\theta}\Big(r+\pi_{k}^{\top}b(Y_{k}^{\theta},\theta)\Big)\Delta t+X_{k}^{\theta}\,\pi_{k}^{\top}\sigma(Y_{k}^{\theta},\theta)\Delta W_{k},$
-with terminal reward $U(X_{N}^{\theta})$. Define the discrete pathwise wealth costate
-$p_{k}^{\mathrm{pw},\theta}:=\frac{\partial}{\partial X_{k}^{\theta}}\,U(X_{N}^{\theta}),\hskip 18.49988ptk=0,\dots,N,$
+$\sup_{t\in[0,T]}\mathbb{E}\!\left[\left\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\right\|^{2}\right]^{1/2}\leq C_{\Theta_{0}}\,\Delta t^{1/2},$
+with a constant $C_{\Theta_{0}}$ independent of $\theta\in\Theta_{0}$.
+Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Fix $\theta\in\Theta_{0}$. Consider the Euler scheme on the grid $(t_{k})$:
+$Y_{k+1}^{\Delta t,\theta}=Y_{k}^{\Delta t,\theta}+a(Y_{k}^{\Delta t,\theta},\theta)\Delta t+\beta(Y_{k}^{\Delta t,\theta},\theta)\Delta W_{k}^{Y},$
+$X_{k+1}^{\Delta t,\theta}=X_{k}^{\Delta t,\theta}+X_{k}^{\Delta t,\theta}\Big(r+\pi_{k}^{\top}b(Y_{k}^{\Delta t,\theta},\theta)\Big)\Delta t+X_{k}^{\Delta t,\theta}\,\pi_{k}^{\top}\sigma(Y_{k}^{\Delta t,\theta},\theta)\Delta W_{k},$
+with terminal reward $U(X_{N}^{\Delta t,\theta})$.
+Define the discrete pathwise wealth costate
+$p_{k}^{\mathrm{pw},\theta}:=\frac{\partial}{\partial X_{k}^{\Delta t,\theta}}\,U(X_{N}^{\Delta t,\theta}),\hskip 18.49988ptk=0,\dots,N,$
 and the additional pathwise blocks
-$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}}.$
+$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\Delta t,\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\Delta t,\theta}}.$
 Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph.
-The one-step algebraic form of the backward recursion coincides with the deterministic-parameter analysis in Huh et al.
-(2025b), with the replacements
-$\mu\mapsto b(\cdot,\theta),\hskip 18.49988pt\sigma\mapsto\sigma(\cdot,\theta),$
-and with the factor block handled exactly as in the wealth–factor extension therein.
-All one-step remainder terms are controlled by standard Taylor/Euler estimates with constants uniform in $\theta\in K$.
-Step 3: Conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables (fixed $\theta$).
-Fix $\theta\in K$.
-Let $\Delta W_{k}:=W_{t_{k+1}}-W_{t_{k}}$ and $\Delta W_{k}^{Y}:=W^{Y}_{t_{k+1}}-W^{Y}_{t_{k}}$.
-Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ implies that the conditional $L^{2}$-projection onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ is well defined (and yields unique coefficients).
-(a) One-step projection (defining the discrete martingale coefficients).
-Define $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ by the conditional orthogonality relations (equivalently, the $L^{2}$-projection) such that
-$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
-where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $G_{t_{k}}^{\theta}$.
-(b) Adapted discrete adjoints (the objects compared to PMP). Define the adapted discrete-time costates by conditional expectation:
-$p_{k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{x,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{y,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right].$
-These are $G_{t_{k}}^{\theta}$-measurable and are the discrete-time counterparts of the $\theta$-conditional Pontryagin objects.
-(c) Discrete BSDE form and identification of the driver.
-Substituting the projection in (a) into the one-step BPTT recursion from Step 2 and then taking conditional expectations with respect to $G_{t_{k}}^{\theta}$ yields a canonical discrete-time BSDE representation for the adapted triple $(p_{k}^{\Delta t,\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$.
-The resulting drift coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE associated with (45)–(46) under the fixed policy $\pi_{\varphi}$, up to standard one-step remainder terms of higher order in $\Delta t$.
-The same argument applies to the blocks $(p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$: they satisfy linearized discrete backward recursions (obtained by differentiating the discrete adjoint equations) and hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$.
-All estimates and remainder controls are identical to Huh et al.
-(2025b) once we condition on $\theta$, and the constants can be chosen uniformly on $K$ by the uniform-in-$\theta$ assumptions from Step 1.
-Step 4: Passage to continuous time and convergence to the $\theta$-conditional PMP costate.
-For each fixed $\theta\in K$, the forward SDE and the $\theta$-conditional adjoint BSDE form a standard FBSDE with coefficients parametrized by $\theta$.
-Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under $\pi_{\varphi}$, with terminal condition $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
-Define the piecewise-constant interpolations of the adapted discrete variables:
-$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
-Then, by the same stability and convergence arguments as in Huh et al.
-(2025b) (Euler convergence for the forward equation plus convergence of the discrete-BSDE scheme for the backward equation), we obtain, for each fixed $\theta\in K$,
+Conditionally on $\theta$, this is exactly the deterministic-parameter BPTT construction of Huh et al.
+(2025).
+Step 3: One-step conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables.
+Fix $\theta\in\Theta_{0}$.
+On each step, perform the conditional $L^{2}$-projection of $p_{k+1}^{\mathrm{pw},\theta}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$.
+Equivalently, define regression coefficients $z_{k}^{\theta}\in\mathbb{R}^{d_{W}}$ and $\tilde{z}_{k}^{\theta}\in\mathbb{R}^{d_{Y}}$ such that
+$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right]+(z_{k}^{\theta})^{\top}\Delta W_{k}+(\tilde{z}_{k}^{\theta})^{\top}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
+where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$.
+Define the adapted discrete-time adjoint blocks by conditional expectation:
+$p_{k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{k}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right],\qquad p_{x,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{x,k}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right],\qquad p_{y,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{y,k}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right].$
+By construction, $(p_{k}^{\Delta t,\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ satisfy a discrete-time BSDE representation.
+Moreover, substituting the projection into the one-step BPTT recursion and taking conditional expectations yields that the discrete driver coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE under the fixed policy $\pi_{\varphi}$, up to standard one-step remainder terms controlled by Taylor/Euler estimates.
+Under the standing smoothness assumptions, the projection residual $\varepsilon_{k+1}^{\theta}$ contributes only higher-order discretization error; all bounds can be chosen uniformly for $\theta\in\Theta_{0}$ (as in Huh et al.
+(2025)).
+Step 4: Convergence as $\Delta t\to 0$ and uniformity on compacts.
+Let $(p_{t}^{\theta},z_{t}^{\theta},\tilde{z}_{t}^{\theta})$ denote the continuous-time $\theta$-conditional adjoint triple under $\pi_{\varphi}$, and let $(p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the corresponding sensitivity blocks (in smooth Markov regimes, value-derivative blocks).
+Define the piecewise-constant interpolations of the adapted discrete variables on $[0,T]$:
+$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\qquad z_{t}^{\Delta t,\theta}:=z_{k}^{\theta},\qquad\tilde{z}_{t}^{\Delta t,\theta}:=\tilde{z}_{k}^{\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
+Standard stability of Euler for the forward SDE together with convergence of the corresponding discrete BSDE scheme implies, for each fixed $\theta\in\Theta_{0}$,
 $\sup_{t\in[0,T]}\mathbb{E}\big[\,|p_{t}^{\Delta t,\theta}-p_{t}^{\theta}|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{x,t}^{\Delta t,\theta}-p_{x,t}^{\theta}\|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{y,t}^{\Delta t,\theta}-p_{y,t}^{\theta}\|^{2}\,\big]\to 0,$
+and, for the martingale coefficients,
+$\mathbb{E}\!\left[\int_{0}^{T}\|z_{t}^{\Delta t,\theta}-z_{t}^{\theta}\|^{2}\,dt\right]\to 0,\hskip 18.49988pt\mathbb{E}\!\left[\int_{0}^{T}\|\tilde{z}_{t}^{\Delta t,\theta}-\tilde{z}_{t}^{\theta}\|^{2}\,dt\right]\to 0,$
 as $\Delta t\to 0$.
-Because all Lipschitz, growth, ellipticity, and block-covariance constants were assumed uniform on $K$, the constants in these convergence bounds can be chosen independently of $\theta\in K$.
-This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$, completing the proof.
+Because all Lipschitz, growth, nondegeneracy, and smoothness constants were assumed uniform on $\Theta_{0}$, the constants in these convergence bounds can be chosen independently of $\theta\in\Theta_{0}$.
+This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.
 ∎
-Appendix D Auxiliary results for Theorem 3
-D.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
+Appendix E Auxiliary results for Theorem 3
+E.1 Stability of the projection map $(A,g)\mapsto-A^{-1}g$
 Proposition 1 (Stability of the projection map $(A,g)\mapsto-A^{-1}g$).
-Let $D$ be a measurable domain and let $\mu$ be a reference measure on $D$. Let $A,\widetilde{A}:D\to\mathbb{R}^{d\times d}$ and $G,\widetilde{G}:D\to\mathbb{R}^{d}$ be measurable. Assume:
+Let $\mathcal{D}$ be a measurable domain and let $\mu$ be a reference measure on $\mathcal{D}$.
+Let $A,\widetilde{A}:\mathcal{D}\to\mathbb{R}^{d\times d}$ and $g,\widetilde{g}:\mathcal{D}\to\mathbb{R}^{d}$ be measurable.
+Assume:
   1. (i)
-    $A(z)$ is invertible for $\mu$-a.e. $z\in D$ and $\|A^{-1}\|_{L^{\infty}(D)}\leq\kappa$ for some $\kappa>0$;
+    $A(z)$ is invertible for $\mu$-a.e. $z\in\mathcal{D}$ and $\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$ for some $\kappa>0$;
   2. (ii)
-    $\|G\|_{L^{\infty}(D)}\leq M$ for some $M>0$;
+    $\|g\|_{L^{\infty}(\mathcal{D})}\leq M_{g}$ for some $M_{g}>0$;
   3. (iii)
-    $\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq(2\kappa)^{-1}$.
+    $\|\widetilde{A}-A\|_{L^{\infty}(\mathcal{D})}\leq(2\kappa)^{-1}$.
 Define $\pi:=-A^{-1}g$ and $\widetilde{\pi}:=-\widetilde{A}^{-1}\widetilde{g}$.
 Then $\widetilde{A}(z)$ is invertible for $\mu$-a.e.
-$z\in D$ with $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$, and
-$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (95)
+$z\in\mathcal{D}$ with $\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq 2\kappa$, and
+$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{g}-g\|_{L^{2}(\mu)}\;+\;2\kappa^{2}M_{g}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (100)
 Proof.
-Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm.
+Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm, and $\|\cdot\|_{L^{\infty}(\mathcal{D})}$ denotes the $\mu$-essential supremum of the corresponding pointwise norm on $\mathcal{D}$.
 Step 1: invertibility and uniform bound for $\widetilde{A}^{-1}$. For $\mu$-a.e. $z$, write
-$\widetilde{A}(z)\;=\;A(z)\Bigl(I+A(z)^{-1}\bigl(\widetilde{A}(z)-A(z)\bigr)\Bigr).$
-By (i) and (iii),
-$\bigl\|A^{-1}(z)\bigl(\widetilde{A}(z)-A(z)\bigr)\bigr\|\;\leq\;\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\;\leq\;\kappa\cdot(2\kappa)^{-1}\;=\;\tfrac{1}{2}.$
-Hence $I+A^{-1}(\widetilde{A}-A)$ is invertible and $\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}\bigr\|\leq(1-\tfrac{1}{2})^{-1}=2$. Therefore $\widetilde{A}(z)$ is invertible and
-$\|\widetilde{A}^{-1}(z)\|\;=\;\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}A^{-1}(z)\bigr\|\;\leq\;2\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa,$
-so $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$.
+$\widetilde{A}(z)\;=\;A(z)\Bigl(I_{d}+A(z)^{-1}\bigl(\widetilde{A}(z)-A(z)\bigr)\Bigr),$
+where $I_{d}$ is the $d\times d$ identity matrix. By (i) and (iii),
+$\bigl\|A^{-1}(z)\bigl(\widetilde{A}(z)-A(z)\bigr)\bigr\|\;\leq\;\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|\widetilde{A}-A\|_{L^{\infty}(\mathcal{D})}\;\leq\;\kappa\cdot(2\kappa)^{-1}\;=\;\tfrac{1}{2}.$
+Hence $I_{d}+A^{-1}(\widetilde{A}-A)$ is invertible and $\bigl\|\bigl(I_{d}+A^{-1}(\widetilde{A}-A)\bigr)^{-1}\bigr\|\leq(1-\tfrac{1}{2})^{-1}=2$. Therefore $\widetilde{A}(z)$ is invertible and
+$\|\widetilde{A}^{-1}(z)\|\;=\;\bigl\|\bigl(I_{d}+A^{-1}(\widetilde{A}-A)\bigr)^{-1}A^{-1}(z)\bigr\|\;\leq\;2\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\;\leq\;2\kappa,$
+so $\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq 2\kappa$.
 Step 2: difference of inverses. Using the identity $\widetilde{A}^{-1}-A^{-1}=\widetilde{A}^{-1}(A-\widetilde{A})A^{-1}$ and Hölder ($L^{\infty}\times L^{2}\times L^{\infty}\to L^{2}$),
-$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\;\leq\;\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\,\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
+$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\;\leq\;\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\,\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\;\leq\;2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
 Step 3: control error bound. Since $\pi=-A^{-1}g$ and $\widetilde{\pi}=-\widetilde{A}^{-1}\widetilde{g}$,
@@
 Taking $L^{2}(\mu)$ norms and applying Hölder ($L^{\infty}\times L^{2}\to L^{2}$) yields
-$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\,\|G\|_{L^{\infty}(D)}.$
+$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|\widetilde{g}-g\|_{L^{2}(\mu)}+\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\,\|g\|_{L^{\infty}(\mathcal{D})}.$
 Using Step 1, Step 2, and (ii),
-$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\|\widetilde{A}-A\|_{L^{2}(\mu)},$
+$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{g}-g\|_{L^{2}(\mu)}+2\kappa^{2}\,M_{g}\,\|\widetilde{A}-A\|_{L^{2}(\mu)},$
 which is (100). ∎
-D.2 Slab-wise small-gain for the $q$-aggregated projection inputs
+E.2 Slab-wise small-gain for the $q$-aggregated projection inputs
 Time-slab decomposition.
-Assume the working domain carries a time coordinate and, for concreteness, $D\subset[0,T]\times S$ and $\mu(dt,d\xi)=dt\otimes\nu(d\xi)$ for some reference measure $\nu$ on $S$.
-Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
-$D_{k}:=D\cap([t_{k-1},t_{k}]\times S),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
-Then $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{K}\|f\|_{k}^{2}$.
+Assume the working domain carries a time coordinate and, for concreteness, $\mathcal{D}\subset[0,T]\times\mathcal{S}$ and $\mu(dt,d\xi)=dt\otimes\nu(d\xi)$ for some reference measure $\nu$ on $\mathcal{S}$.
+Fix a partition $0=t_{0}<t_{1}<\cdots<t_{L}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
+$\mathcal{D}_{k}:=\mathcal{D}\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{\mathcal{D}_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
+Then $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{L}\|f\|_{k}^{2}$.
 Proposition 2 (Short-time (slab) Lipschitz gain).
-Let $U$ be a neighborhood of a locally optimal deployable $\theta$-blind policy $\pi^{\star}_{\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in U$,
-$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi}\|_{L^{\infty}(D)}\leq M_{G}$
-for some constants $\kappa,M_{G}>0$. Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
-$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\overline{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi_{1}}-G^{\mathrm{mix}}_{\pi_{2}}\|_{k}\leq\overline{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (96)
-for all $\pi_{1},\pi_{2}\in U$, with constants $\overline{L}_{A},\overline{L}_{G}>0$ depending only on band data.
-Define the (population) projection map $T(\pi):=-A_{\pi}^{-1}G^{\mathrm{mix}}_{\pi}$ and
-$\rho(\tau):=\Bigl(\kappa\,\overline{L}_{G}+\kappa^{2}M_{G}\,\overline{L}_{A}\Bigr)\tau^{1/2}.$
-Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in U$,
-$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}.$
+Let $\mathcal{U}$ be a subset of the admissible policy class that is an $L^{2}(\mu)$-neighborhood of a locally optimal deployable $\theta$-blind policy $\pi^{\star,\mathrm{blind}}$ (in particular $\pi^{\star,\mathrm{blind}}\in\mathcal{U}$), and on which the following uniform bounds hold for all $\pi\in\mathcal{U}$:
+$\|A_{\pi}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa,\hskip 18.49988pt\|g^{\mathrm{mix}}_{\pi}\|_{L^{\infty}(\mathcal{D})}\leq M_{g}$
+for some constants $\kappa,M_{g}>0$. Assume that on each slab $\mathcal{D}_{k}$ the $q$-aggregated projection inputs satisfy
+$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|g^{\mathrm{mix}}_{\pi_{1}}-g^{\mathrm{mix}}_{\pi_{2}}\|_{k}\leq\bar{L}_{g}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (101)
+for all $\pi_{1},\pi_{2}\in\mathcal{U}$, with constants $\bar{L}_{A},\bar{L}_{g}>0$ depending only on band data.
+Define the (population) projection map $\mathcal{T}(\pi):=-A_{\pi}^{-1}g^{\mathrm{mix}}_{\pi}$ and
+$\varrho(\tau):=\Bigl(\kappa\,\bar{L}_{g}+\kappa^{2}M_{g}\,\bar{L}_{A}\Bigr)\tau^{1/2}.$
+Then for each slab $\mathcal{D}_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
+$\|\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})\|_{k}\leq\varrho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}.$
 In particular, if the partition is chosen so that
-$\rho^{\ast}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1,$
-then $T$ is a strict contraction on every slab with constant at most $\rho^{\ast}$.
+$\varrho_{*}:=\max_{1\leq k\leq L}\varrho(\tau_{k})<1,$
+then $\mathcal{T}$ is a strict contraction on every slab with constant at most $\varrho_{*}$.
 Proof.
-Fix $k$ and $\pi_{1},\pi_{2}\in U$. Set $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G^{\mathrm{mix}}_{\pi_{i}}$. Then
-$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},\hskip 18.49988ptA_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
-Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$ yields
-$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\kappa\|G_{1}-G_{2}\|_{k}+\kappa^{2}M_{G}\|A_{1}-A_{2}\|_{k}.$
-Applying (96) gives the claim with $\rho(\tau_{k})$. ∎
-Remark 3 (Where the $\tau^{1/2}$ gain comes from; relation to huh2025a).
+Fix $k$ and $\pi_{1},\pi_{2}\in\mathcal{U}$. Set $A_{i}:=A_{\pi_{i}}$ and $g_{i}:=g^{\mathrm{mix}}_{\pi_{i}}$. Then
+$\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})=-A_{1}^{-1}(g_{1}-g_{2})-(A_{1}^{-1}-A_{2}^{-1})g_{2},\hskip 18.49988ptA_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
+Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $\mathcal{D}_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$ and $\|g_{2}\|_{L^{\infty}(\mathcal{D})}\leq M_{g}$ yields
+$\|\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})\|_{k}\leq\kappa\|g_{1}-g_{2}\|_{k}+\kappa^{2}M_{g}\|A_{1}-A_{2}\|_{k}.$
+Applying (101) gives the claim with $\varrho(\tau_{k})$. ∎
+Remark 6 (Where the $\tau^{1/2}$ gain comes from; relation to Huh et al. (2025)).
 The short-time factor $\tau^{1/2}$ in (101) is the same parabolic smoothing mechanism used in our earlier slab-wise PGDPO analysis:
@@
 A full PDE derivation of this short-time contraction philosophy is given in Huh et al.
-(2025b, Appendix B); see also the companion supplementary note to Huh et al.
-(2025b) (“On Proof of Theorem 4 (Policy Gap Bound)”) for a step-by-step slab contraction argument in a closely related setting.
+(2025, Appendix B); see also the companion supplementary note to Huh et al.
+(2025) (“On Proof of Theorem 4 (Policy Gap Bound)”) for a step-by-step slab contraction argument in a closely related setting.
 In the present paper the same smoothing principle is applied not to the costate update map directly, but to the induced $q$-aggregated projection inputs $(A_{\pi},g^{\mathrm{mix}}_{\pi})$.
-D.3 Proof of Theorem 3
+E.3 Proof of Theorem 3
 Proof of Theorem 3.
@@
 Define the population projected policy
-$\pi^{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
-Let $\widehat{A}$ and $\widehat{G}$ be the stage-2 (BPTT/Monte Carlo) estimators of $A_{\mathrm{warm}}$ and $G_{\mathrm{warm}}$ used to form the deployable stage-2 control
-$\widehat{\pi}_{\mathrm{agg,mix}}:=-\widehat{A}^{-1}\widehat{G}$.
-Let $\pi^{\star}_{\mathrm{blind}}$ denote a locally optimal interior deployable $\theta$-blind policy, so in particular it satisfies the fixed-point (stationarity) identity
-$\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}}).$
+$\pi^{\mathrm{proj}}:=\mathcal{T}(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}g_{\mathrm{warm}}.$
+Let $\widehat{A}$ and $\widehat{g}$ be the stage-2 (BPTT/Monte Carlo) estimators of $A_{\mathrm{warm}}$ and $g_{\mathrm{warm}}$ used to form the deployable projected control
+$\widehat{\pi}^{\mathrm{agg,mix}}:=-\widehat{A}^{-1}\widehat{g}$,
+where $\widehat{A}$ is understood as the effective matrix after any inversion stabilizers (ridge/diagnostics/skip) so that $\widehat{A}$ is well-defined and invertible on the working domain.
+Let $\pi^{\star,\mathrm{blind}}$ denote a locally optimal interior deployable $\theta$-blind policy, so in particular it satisfies the fixed-point (stationarity) identity
+$\pi^{\star,\mathrm{blind}}=\mathcal{T}(\pi^{\star,\mathrm{blind}}).$
 Step 1 (Estimator-to-population projection stability).
-Assume we are in the stable inversion regime (as enforced in practice by ridge/diagnostics) so that the hypotheses of Proposition 1 apply on $D$ with $A=A_{\mathrm{warm}}$ and $G=G_{\mathrm{warm}}$.
-Then there exists a constant $C_{2}>0$ (depending only on the uniform inverse bound $\kappa$ and uniform $L^{\infty}$ bounds on the relevant $G$ terms) such that
-$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}},$ (97)
-where one convenient choice is
+To apply Proposition 1 with $A=A_{\mathrm{warm}}$ and $g=g_{\mathrm{warm}}$, we assume we are in the stable inversion regime enforced by the stage-2 stabilizers, i.e.
+$\|\widehat{A}-A_{\mathrm{warm}}\|_{L^{\infty}(\mathcal{D})}\leq(2\kappa)^{-1},$ (102)
+where $\kappa$ is such that $\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$. Define the $L^{2}$ estimation error level
 $\delta_{\mathrm{BPTT}}:=\|\widehat{A}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}+\|\widehat{g}-g_{\mathrm{warm}}\|_{L^{2}(\mu)}.$
+Then Proposition 1 yields $\|\widehat{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq 2\kappa$ and
+$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widehat{g}-g_{\mathrm{warm}}\|_{L^{2}(\mu)}+2\kappa^{2}\,\|g_{\mathrm{warm}}\|_{L^{\infty}(\mathcal{D})}\,\|\widehat{A}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}},$ (103)
+where one may take, for instance,
+$C_{2}:=2\kappa+2\kappa^{2}\,\|g_{\mathrm{warm}}\|_{L^{\infty}(\mathcal{D})}.$
 […]
 Step 2 (Slab-wise contraction for the population projection map).
-Assume $\pi^{\mathrm{warm}},\pi^{\star}_{\mathrm{blind}}\in U$ and let $\rho^{\ast}<1$ be the slab contraction constant of Proposition 2.
-Since $\pi^{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}})$, for each slab $D_{k}$ we have
-$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star}_{\mathrm{blind}})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$ (98)
+Assume $\pi^{\mathrm{warm}},\pi^{\star,\mathrm{blind}}\in\mathcal{U}$ and let $\varrho_{*}<1$ be the slab contraction constant of Proposition 2.
+Since $\pi^{\mathrm{proj}}=\mathcal{T}(\pi^{\mathrm{warm}})$ and $\pi^{\star,\mathrm{blind}}=\mathcal{T}(\pi^{\star,\mathrm{blind}})$, for each slab $\mathcal{D}_{k}$ we have
+$\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}=\|\mathcal{T}(\pi^{\mathrm{warm}})-\mathcal{T}(\pi^{\star,\mathrm{blind}})\|_{k}\leq\varrho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\varrho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$ (104)
 Step 3 (Residual identity and slab-wise warm-up deviation). Define the population mixed-moment aggregated stationarity residual of the warm-up policy
-$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}.$
+$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}:=A_{\mathrm{warm}}\pi^{\mathrm{warm}}+g_{\mathrm{warm}},\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)},\hskip 18.49988pt\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k}.$
 By definition of $\pi^{\mathrm{proj}}$,
-$\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}=\pi^{\mathrm{warm}}-(-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}})=A_{\mathrm{warm}}^{-1}\bigl(A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}}\bigr)=A_{\mathrm{warm}}^{-1}r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$
+$\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}=A_{\mathrm{warm}}^{-1}r_{\mathrm{FOC,mix}}^{\mathrm{warm}}.$
 Hence, on each slab,
-$\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\leq\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\,\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (99)
+$\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\leq\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (105)
 Step 4 (Slab-wise closure). Combine the triangle inequality with (104) and (105):
-$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}+\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$
-Since $\rho^{\ast}<1$, we can close slab-wise:
-$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (100)
+$\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\varrho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$
+Since $\varrho_{*}<1$, we can close slab-wise:
+$\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\frac{\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (106)
 Plugging (106) into (104) yields
-$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (101)
-Step 5 (Global bound and completion). Sum over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k}\|f\|_{k}^{2}$:
-$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}.$
+$\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (107)
+Step 5 (Global bound and completion). Sum over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{L}\|f\|_{k}^{2}$:
+$\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\leq\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$
 Finally, combine with (103) and the triangle inequality:
-$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}+\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}},$
-which is the claimed slab-wise residual-based $\theta$-blind policy-gap bound. ∎
-Appendix E Implementation details for Section 3
+$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}+\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}},$
+which is the claimed slab-wise residual-based $\theta$-blind policy-gap bound (under the stable inversion condition (102)). ∎
+Appendix F Implementation details for Section 3
 This appendix provides reproducible step-by-step templates for the methods in Section 3.
-The high-level pipeline is summarized in Figure LABEL:fig:sec3-pipeline-sideways.
+The high-level pipeline is summarized in Figure 4.
 For compactness we present one template per subsection of Section 3.
-E.1 Stage 1 (PG–DPO) template for Section 3.1
+Conventions.
+Throughout, the deployed policy is $\theta$-blind.
+The latent parameter $\theta\sim q$ is sampled only inside the simulator.
+Stage 2 computes projection ingredients on a working domain $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ with reference measure $\mu$.
+F.1 Stage 1 (DPO) template for Section 3.1
 Stage 1 performs stochastic gradient ascent on the fixed-$q$ ex–ante objective (42), with latent $\theta\sim q$ sampled inside the simulator while the policy remains $\theta$-blind.
@@
   2. 2.
-    Sample latent environment parameter.
-Sample $\theta\sim q$ inside the simulator (unseen by $\pi_{\varphi}$).
-(Variant: sample $\theta^{(i)}\sim q$ independently per episode; both are unbiased for $\nabla_{\varphi}J(\varphi)$.)
+    Sample latent parameter(s) inside the simulator.
+Sample $\theta\sim q$ and keep it frozen along the rollout(s).
+(Variant: sample $\theta^{(i)}\sim q$ independently per episode; both are unbiased for the discretized gradient of $J(\varphi)$.)
   3. 3.
-    Simulate Euler rollouts. For each episode $i$, simulate the Euler scheme in (45)–(46) under the $\theta$-blind policy $\pi_{\varphi}$ and collect terminal utilities $\{U(X_{T}^{(i)})\}_{i=1}^{M}$.
+    Simulate Euler rollouts.
+For each episode $i$, simulate the Euler scheme for (39)–(40) under the $\theta$-blind policy $\pi_{\varphi}$, and compute terminal utilities $\{U(X_{T}^{(i)})\}_{i=1}^{M}$.
   4. 4.
-    Backpropagation through time (BPTT). Compute the Monte Carlo gradient estimator
+    BPTT (pathwise gradient). Compute the Monte Carlo gradient estimator
     $\widehat{g}\;\leftarrow\;\frac{1}{M}\sum_{i=1}^{M}\nabla_{\varphi}U(X_{T}^{(i)}).$
@@
     Checkpoint. Periodically save a warm-up checkpoint $\varphi^{\mathrm{warm}}$ for stage 2 projection.
-E.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
-Stage 2 is a post-processing map: given a warm-up $\theta$-blind policy $\pi_{\varphi^{\mathrm{warm}}}$, it estimates Pontryagin sensitivity objects by Monte Carlo and constructs a deployable projected control on a working-domain sample $z\sim\mu$.
+F.2 Stage 2 (Pontryagin projection; mixed-moment $q$-aggregation) template for Section 3.2
+Stage 2 is a post-processing map: given a warm-up $\theta$-blind policy $\pi_{\varphi^{\mathrm{warm}}}$, it estimates Pontryagin sensitivity objects by Monte Carlo and constructs a deployable projected control on queried working-domain states $z\sim\mu$.
 The aggregation used here matches the mixed-moment $q$-aggregation in (51)–(52), yielding the projected control (53).
 Inputs.
-Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; working-domain sampler $\mu$ on $D$; budgets $(M_{z},M_{\theta},M_{\mathrm{MC}})$.
+Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; working-domain sampler $\mu$ on $\mathcal{D}$; budgets $(M_{z},M_{\theta},M_{\mathrm{MC}})$.
 Template (constructing projection targets on a batch of query states).
@@
   2. 2.
-    For each query state $z_{j}$, sample latent parameters. Sample $\{\theta_{\ell}\}_{\ell=1}^{M_{\theta}}\sim q$.
+    For each query state $z_{j}$, sample frozen parameters. Sample $\{\theta_{\ell}\}_{\ell=1}^{M_{\theta}}\sim q$.
   3. 3.
-    For each frozen $\theta_{\ell}$, estimate costates at $z_{j}$. For each $\ell=1,\dots,M_{\theta}$:
+    For each frozen $\theta_{\ell}$, estimate costate blocks at $z_{j}$. For each $\ell=1,\dots,M_{\theta}$:
     1. (a)
@@
   5. 5.
-    Solve the projection (mixed-moment aggregation). Whenever $\widehat{A}_{t}(z_{j})$ is invertible and the solve is numerically stable, compute the deployable projected control
+    Solve the projection (deployable $\theta$-blind control). Whenever $\widehat{A}_{t}(z_{j})$ is numerically invertible and passes stability diagnostics, compute
     $\widehat{\pi}^{\mathrm{agg,mix}}(z_{j})\;\leftarrow\;-\big(\widehat{A}_{t}(z_{j})\big)^{-1}\widehat{g}_{t}^{\mathrm{mix}}(z_{j}),$
-    which matches (59).
-E.3 Coupling I: residual/control-variate projection (Section F.1)
+    which matches (53). Otherwise, fall back to $\pi_{\varphi^{\mathrm{warm}}}(z_{j})$ (or a ridge-stabilized solve; see Appendix F.5).
+F.3 Coupling I: residual/control-variate projection (Section G.1)
 This subsection records a variance-reduced implementation of the stage 2 map using the residual identity (110).
@@
 Inputs.
-Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section E.2.
+Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 ingredients $(\widehat{A}_{t}(z),\widehat{g}_{t}^{\mathrm{mix}}(z))$ from Appendix F.2.
 Template (statewise residual projection; mixed-moment aggregation).
@@
     Apply the residual correction. Compute
-    $\widehat{\pi}^{\mathrm{agg,mix}}(z)\;\leftarrow\;\pi_{\varphi^{\mathrm{warm}}}(z)-\big(\widehat{A}_{t}(z)\big)^{-1}\widehat{r}_{\mathrm{FOC}}(z).$
-E.4 Coupling II: interactive distillation (Section F.2)
-This subsection records an implementation template for interactive distillation: the projected output from stage 2 is used as a teacher signal during stage 1 training via the mixed objective (106).
-The teacher is built from the mixed-moment projected rule (possibly evaluated in residual form for variance reduction).
+    $\widehat{\pi}^{\mathrm{agg,mix}}(z)\;\leftarrow\;\pi_{\varphi^{\mathrm{warm}}}(z)-\big(\widehat{A}_{t}(z)\big)^{-1}\widehat{r}_{\mathrm{FOC}}(z),$
+    with ridge/diagnostics/skip as needed (Appendix F.5).
+F.4 Coupling II: interactive distillation (Section G.2)
+Interactive distillation amortizes the stage 2 projected control into the policy network.
+A lagged policy copy $\pi_{\varphi^{-}}$ is used to construct a teacher on a slower timescale, and the student $\pi_{\varphi}$ is trained against a mixed objective (112).
 Inputs.
-Student parameters $\varphi$; teacher refresh interval $K$; distillation schedule $\lambda(n)$; working-domain sampler $\mu$.
+Student parameters $\varphi$; teacher refresh interval $n_{\mathrm{refresh}}$; distillation schedule $\lambda(n)$; working-domain sampler $\mu$.
 Template (training loop with intermittent teacher refresh).
@@
     1. (a)
-      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section E.1.
+      Stage 1 update (DPO step). Perform one DPO ascent step on $J(\varphi)$ as in Appendix F.1.
     2. (b)
-      Teacher refresh (every $K$ steps). If $n\bmod K=0$:
+      Teacher refresh (every $n_{\mathrm{refresh}}$ steps). If $n\bmod n_{\mathrm{refresh}}=0$:
       1. i.
-        Set $\varphi^{-}\leftarrow\varphi$ (lagged copy).
+        Set $\varphi^{-}\leftarrow\varphi$.
       2. ii.
@@
       3. iii.
-        For each $z_{j}$, run stage 2 under $\pi_{\varphi^{-}}$ (mixed-moment aggregation) to compute a projected teacher $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j})$.
-(In practice we compute it in residual form around $\pi_{\varphi^{-}}$ as in (105).)
+        For each $z_{j}$, run stage 2 under $\pi_{\varphi^{-}}$ (mixed-moment aggregation) to compute the projected teacher $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j})$ (typically in residual form (111)).
       4. iv.
-        Optionally filter states using diagnostics (Section E.5) and update the buffer:
+        Optionally filter unreliable states using diagnostics (Appendix F.5) and update the buffer:
         $\mathcal{B}\ \leftarrow\ \{(z_{j},\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j}))\}_{j=1}^{M_{z}}\ \ \text{(after filtering)}.$
@@
       2. ii.
-        Apply a gradient step to minimize the proximity term $\|\pi_{\varphi}(z)-\mathrm{stopgrad}(\pi^{\mathrm{teach}})\|^{2}$ with coefficient $\lambda(n)$, consistent with (106).
-E.5 Engineering notes and stabilizers
-This subsection collects practical stabilizers that we found helpful for reliable training and projection in high dimensions.
+        Apply one optimizer step for the teacher proximity term in (112) (treating $\pi^{\mathrm{teach}}$ as $\mathrm{stopgrad}$-fixed).
+F.5 Engineering notes and stabilizers
+This subsection collects practical stabilizers that are useful for reliable training and projection.
   • Antithetic sampling for $\theta$.
 When $q$ is symmetric (e.g.
-Gaussian in a latent normal parameterization), sample $\theta$ in antithetic pairs by drawing $z\sim\mathcal{N}(0,I)$ and using (z,z)(z,-z) to construct $(\theta^{+},\theta^{-})$.
-This reduces the variance of $q$-averaged quantities and typically improves the stability of stage 2 diagnostics on the working domain.
+Gaussian in a latent-normal parameterization), sample $\theta$ in antithetic pairs.
+This reduces the variance of $q$-averaged quantities and improves stage 2 concentration.
   • Blockwise Monte Carlo and robust aggregation.
-To control rare-tail domination, split Monte Carlo replications into $B$ blocks and compute blockwise averages of costate-driven ingredients (e.g.
-$\widehat{A}_{t}^{\theta}(z)$ and $\widehat{G}_{t}^{\theta}(z)$).
-Aggregate across blocks using a robust statistic such as the median or median-of-means, which makes the projection less sensitive to outlier trajectories.
-  • Curvature/denominator stability checks.
-Because the projection map $(A,G)\mapsto-A^{-1}G$ can be sensitive to near-singularity of $A$, monitor the conditioning of $\widehat{A}_{t}$ (or failure rates of the linear solve).
-When diagnostics indicate ill-conditioning, skip projection-guided updates at that state or increase Monte Carlo budgets locally.
+Split Monte Carlo replications into blocks and aggregate blockwise estimates using robust statistics (median / median-of-means) to reduce tail domination.
+  • Curvature/denominator checks for $\widehat{A}_{t}$. Monitor conditioning of $\widehat{A}_{t}(z)$ (e.g. eigenvalue floors / condition numbers). If unstable, skip projection at $z$ or use ridge:
+    $\widehat{A}_{t,\lambda}(z):=\widehat{A}_{t}(z)+\lambda I_{d},\hskip 18.49988pt\widehat{\pi}_{\lambda}(z):=-\widehat{A}_{t,\lambda}(z)^{-1}\widehat{g}_{t}^{\mathrm{mix}}(z).$
   • Residual magnitude as a reliability diagnostic.
-For the residual form, compute $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{G}_{t}^{\mathrm{mix}}(z)$.
-Small $\|\widehat{r}_{\mathrm{FOC}}(z)\|$ indicates approximate satisfaction of the mixed-moment aggregated first-order condition at $z$ and empirically correlates with more reliable teacher targets.
-  • Diagnostics-based teacher selection on the working domain.
-Rather than applying distillation on all sampled $\{z_{j}\}\sim\mu$, keep only states that pass a reliability predicate.
-In practice, filter using residual-magnitude thresholds together with stable linear-solve diagnostics to prevent a small subset of pathological states from contaminating the teacher buffer.
-  • $\lambda$ schedule and safeguards.
-Use a warm-up period with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after stage 2 diagnostics on the working domain are stable.
-To prevent the teacher term from dominating the ex–ante objective, cap the effective coefficient via
+Track $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi^{\mathrm{warm}}(z)+\widehat{g}_{t}^{\mathrm{mix}}(z)$.
+Small $\|\widehat{r}_{\mathrm{FOC}}(z)\|$ empirically correlates with reliable teacher targets.
+  • Diagnostics-based teacher selection.
+For distillation, keep only $z$ that pass stability predicates (invertibility/conditioning + moderate residual), to prevent noisy stage 2 targets from contaminating the teacher buffer.
+  • Scheduling $\lambda$ (teacher loss).
+Use an initial warm-up with $\lambda=0$ (pure DPO), then increase $\lambda$ only after stage 2 diagnostics stabilize.
+Optionally cap the effective coefficient via
     $\lambda_{\mathrm{eff}}:=\min\Big\{\lambda,\;c\,\frac{|L_{\mathrm{main}}|}{L_{\mathrm{distill}}+\varepsilon}\Big\},$
-    with $c\in(0,1)$ and $\varepsilon>0$.
-  • Initialization and scale control in high dimensions.
-To avoid early-time numerical blow-ups (often through quadratic variation terms of the form $\pi^{\top}\Sigma\pi$), initialize the policy output near zero and/or scale the output by $d^{-1/2}$.
-As a last-resort safety net, a mild log-wealth clamp can prevent overflow, but it should be used conservatively and monitored, since frequent clamping may distort higher-order sensitivities.
-Appendix F Coupling stage 1 and stage 2: residual projection and interactive distillation
+    with $c\in(0,1)$, $\varepsilon>0$.
+  • Initialization/scale control in high dimensions. Initialize policy outputs near zero and/or scale outputs by $d^{-1/2}$ to avoid early-time blow-ups (e.g. via $\pi^{\top}\Sigma\pi$).
+Appendix G Coupling stage 1 and stage 2: residual projection and interactive distillation
 We keep the ex–ante objective (42) and the $\theta$-blind deployability constraint throughout.
-Stage 2 is not a separate optimization problem: it is a (warm-started) post-processing map that reuses the current stage 1 policy as a warm-up control, estimates the projection ingredients under this warm-up policy, and then applies the $q$-aggregated Pontryagin projection as a deterministic transformation of the policy.
+Stage 2 is not a separate optimization problem: it is a (warm-started) post-processing map that reuses the current stage 1 policy as a warm-up control, estimates projection ingredients under this warm-up policy, and then applies a $q$-aggregated Pontryagin projection as a deterministic transformation.
 Concretely, stage 2 takes $\pi^{\mathrm{warm}}$ as input, runs BPTT/Monte Carlo under frozen $\theta\sim q$ to estimate the adjoint blocks required for the Hamiltonian gradient, aggregates them across $\theta$, and outputs a single deployable $\theta$-blind rule $\widehat{\pi}^{\mathrm{agg,mix}}$ (Section 3.2).
-This appendix records two couplings between the two stages, each with a distinct role:
+This appendix records two couplings between the two stages:
   • Residual (control-variate) projection.
-We compute the same projected rule in a residual form around the warm-up policy.
-This is algebraically equivalent to the direct projection, but typically reduces Monte Carlo variance and improves numerical stability in high dimensions.
+The projected rule is evaluated in a residual form around the warm-up policy.
+This is algebraically equivalent to the direct solve, but can reduce Monte Carlo variance and improve numerical stability.
   • Interactive distillation (amortized projection).
-We use the projected rule as a teacher signal at intermittent refresh times.
-Beyond accelerating/stabilizing training, distillation amortizes the cost of projection: stage 2 can be accurate but Monte-Carlo intensive, whereas a distilled student policy approximates the projected rule with a single forward pass at stage 1 inference cost.
-F.1 Control-variate (residual) form of the projected rule
+The projected rule is used as a teacher signal at intermittent refresh times.
+Distillation amortizes the cost of projection by compressing projected controls into the policy network.
+Stage 1: DPO (Section 3.1)
+Goal. Maximize $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$ with a deployable $\theta$-blind policy.
+Loop.
+Sample $(t_{0},x_{0},y_{0})\sim\nu$ and latent $\theta\sim q$ inside the simulator; simulate Euler rollouts; compute $U(X_{T})$; BPTT for $\nabla_{\varphi}U(X_{T})$; update $\varphi$ (e.g.
+Adam).
+Output.
+Warm-up checkpoint $\varphi^{\mathrm{warm}}$.
+Stage 2: $q$-aggregated Pontryagin projection under latent $\theta$ (Section 3.2)
+Input. $\pi_{\varphi^{\mathrm{warm}}}$, working-domain sampler $\mu$ on $\mathcal{D}$, budgets $(M_{\mathrm{MC}},M_{\theta})$.
+Adjoint blocks.
+Simulate under $\pi_{\varphi^{\mathrm{warm}}}$ and compute $\widehat{p}_{t}^{\theta}(z),\widehat{p}_{x,t}^{\theta}(z),\widehat{p}_{y,t}^{\theta}(z)$ by autodiff/BPTT.
+Projection.
+Aggregate $\widehat{A}_{t}$ and $\widehat{g}_{t}^{\mathrm{mix}}$ via (51)–(52); output $\widehat{\pi}^{\mathrm{agg,mix}}(z)=-(\widehat{A}_{t}(z))^{-1}\widehat{g}_{t}^{\mathrm{mix}}(z)$ (53).
+Stabilizers.
+Use ridge/diagnostics/residual form as needed (Appendix F.5).
+Output.
+Accurate but Monte-Carlo intensive projected rule.
+Coupling (Appendix G)
+(G.1) Residual projection.
+Compute $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{g}_{t}^{\mathrm{mix}}(z)$ and evaluate $\widehat{\pi}^{\mathrm{agg,mix}}(z)=\pi_{\varphi^{\mathrm{warm}}}(z)-\widehat{A}_{t}(z)^{-1}\widehat{r}_{\mathrm{FOC}}(z)$ (110).
+(G.2) Interactive distillation (amortized projection).
+Freeze a lagged copy $\varphi^{-}$, compute a teacher via stage 2, and train the student with the mixed objective (112).
+Refresh teacher on a slower timescale and optionally gate teacher loss by projection diagnostics.
+[…]
+Figure 4: Pipeline of Section 3: stage 1 learning, stage 2 $q$-aggregated projection, and the coupling mechanisms in Appendix G.
+Distillation stabilizes training and amortizes projection by compressing projected controls into the policy network.
+G.1 Control-variate (residual) form of the projected rule
 Recall the mixed-moment projected rule (53).
@@
 $\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y)=\pi_{\varphi^{\mathrm{warm}}}(t,x,y)-\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{r}_{\mathrm{FOC}}(t,x,y),$ (110)
-which is an algebraic rewriting of (59) (hence it does not change the target).
+which is algebraically equivalent to (53) (hence it does not change the target).
 Practical value (variance reduction and stability).
@@
 Moreover, the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{g}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and often partially cancel, improving concentration of $\widehat{r}_{\mathrm{FOC}}$.
-In implementation, we combine (104) with the inversion stabilizers of Appendix E.5 (e.g.
+In implementation, we combine (110) with the inversion stabilizers of Appendix F.5 (e.g.
 ridge regularization and diagnostic-based fallback to $\pi_{\varphi^{\mathrm{warm}}}$ on unreliable states).
-F.2 Interactive distillation: projection-guided training and amortized deployment
+G.2 Interactive distillation: projection-guided training and amortized deployment
 Let $\pi_{\varphi}$ be the trainable stage 1 policy network.
-At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher.
+At intermittent refresh times, freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher.
 This coupling serves two purposes:
-  • Training aid. Projection-guided targets can stabilize and accelerate stage 1 optimization, especially once the warm-up policy enters a locally stable regime.
-  • Amortization. Stage 2 projection can be accurate but Monte-Carlo intensive; distillation compresses it into a single forward pass for deployment.
-Teacher definition (computed by stage 2, treated as fixed).
-In residual form (104), the teacher is the $\theta$-blind map
+  • Training aid. Projection-guided targets can stabilize and accelerate stage 1 optimization once the warm-up policy enters a locally stable regime.
+  • Amortization. Stage 2 projection is accurate but Monte-Carlo intensive; distillation compresses it into a single forward pass for deployment.
+Teacher definition (computed by stage 2; treated as fixed).
+In residual form (110), define the teacher as the $\theta$-blind map
 $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y):=\pi_{\varphi^{-}}(t,x,y)-\big(\widehat{A}_{t}^{\varphi^{-}}(t,x,y)\big)^{-1}\,\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}(t,x,y),$ (111)
-where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy.
-In practice, the inversion in (105) is implemented with the same stabilizers as in Section 3.2 (ridge/diagnostics/residual form); for notational simplicity we keep the formal inverse.
-Student objective (hybrid PG–DPO + distillation loss).
-We train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to the teacher on the working domain:
-$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(t,x,y)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y)\big)\big\|^{2}\Big],$ (106)
-where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance.
-The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2: once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
-Refresh schedule, annealing, and diagnostic gating (implementation).
-In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps. A typical and robust configuration is:
-  • Two-time-scale refresh. Hold $\varphi^{-}$ fixed for $K_{\mathrm{refresh}}$ stage 1 updates of $\varphi$ under (106), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher via stage 2.
-  • Anneal $\lambda$.
-Start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain indicate that the stage 2 map is reliable.
-Concretely, we anneal $\lambda$ toward a target value $\lambda_{\max}$ once the warm-up residual $\|\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)}$ and curvature diagnostics remain stable over several refresh cycles.
+where $\widehat{A}_{t}^{\varphi^{-}}$ and $\widehat{g}_{t}^{\mathrm{mix},\varphi^{-}}$ denote the stage 2 estimators computed under the lagged policy $\pi_{\varphi^{-}}$, and $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{g}_{t}^{\mathrm{mix},\varphi^{-}}$.
+(We omit stabilizer notation; in code, $(\widehat{A}_{t}^{\varphi^{-}})^{-1}$ is implemented with ridge/diagnostics/skip as needed.)
+Student objective (hybrid DPO + distillation penalty).
+Train $\pi_{\varphi}$ by combining the ex–ante objective with a proximity penalty to the teacher on the working domain:
+$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{z=(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(z)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z)\big)\big\|^{2}\Big],$ (112)
+where $\lambda\geq 0$ controls guidance strength and $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2.
+Refresh schedule, annealing, and diagnostic gating.
+In practice, $\varphi^{-}$ and the teacher are refreshed on a slower timescale than stage 1 updates:
+  • Two-time-scale refresh. Hold $\varphi^{-}$ fixed for $n_{\mathrm{refresh}}$ stage 1 steps, then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher.
+  • Anneal $\lambda$. Start with $\lambda=0$ (pure DPO) and increase $\lambda$ only after projection diagnostics on the working domain stabilize.
   • Adaptive teacher selection (gating).
-To avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we optionally apply the teacher loss only on states where projection diagnostics certify reliability (e.g.
-bounded condition number of $\widehat{A}_{t}$, eigenvalue/denominator lower bounds, and moderate residual magnitude), falling back to pure PG–DPO elsewhere.
-Implementation details, including the precise diagnostics and stabilizers, are deferred to Appendix E.
-Deployment.
-After training, the deployed controller is simply the student network $\pi_{\varphi}$, which approximates the stage 2 projected rule with a single forward pass, without repeatedly running BPTT/Monte Carlo projection at inference time.
-Appendix G Stage 2 projection diagnostics
-We report Stage 2 diagnostic statistics as a visual supplement to Section 4.2. Each figure summarizes the same tail-median protocol and layout; see captions for definitions and interpretation.
+Apply teacher loss only on states where projection diagnostics certify reliability (stable $\widehat{A}_{t}$ solve and moderate residual), falling back to pure DPO elsewhere.
+Deployment. After training, the deployed controller is the student network $\pi_{\varphi}$, which approximates the stage 2 projected rule with a single forward pass.
+Appendix H Stage 2 projection diagnostics
+We report Stage 2 diagnostic statistics as a visual supplement to Section 4.2. Each figure uses the same tail-median protocol and layout; see captions for definitions.
 Figure 5: Stage 2 stationarity residual (q50).
@@
 This helps rule out “catastrophic inversion” as the primary driver of degradation; instead, residual growth and curvature mismatch (Fig.
-6) provide a more consistent explanation in misaligned/high-$d$ regimes.
+7) provide a more consistent explanation in misaligned/high-$d$ regimes.
 Figure 7: Stage 2 curvature-consistency statistic $\kappa$.
@@
 Near-zero bad-sign fractions across most regimes suggest that the projection typically operates in a locally well-behaved region and that failures are not dominated by sign flips, supporting the main-text conclusion that misalignment primarily increases residual/costate mismatch rather than inducing widespread concavity violations.
-Appendix H Supplementary decomposition diagnostics for Section 5
-Tables 23 report Stage 2 decomposition diagnostics at $t=0$.
+Appendix I Supplementary decomposition diagnostics for Section 5
+Tables 45 report Stage 2 decomposition diagnostics at $t=0$ (tail medians over the last six evaluation checkpoints).
+Diagnostics are reported for the projected (Stage 2 / PG–DPO) rules, since Stage 1 (DPO) and PPO do not expose a compatible myopic/hedging split under our protocol.
 $s_{0}$ Method $d=5$ $10$ $50$ $100$
 Aligned $P_{0}$
-$10^{-3}$ Stage 1+Stage 2 (Basic) $7.17\text{\times}{10}^{-06}$ $1.51\text{\times}{10}^{-05}$ $1.60\text{\times}{10}^{-05}$ $1.23\text{\times}{10}^{-05}$
-Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-06}$ $9.81\text{\times}{10}^{-06}$ $1.64\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
-$10^{-2}$ Stage 1+Stage 2 (Basic) $6.21\text{\times}{10}^{-06}$ $1.38\text{\times}{10}^{-05}$ $1.63\text{\times}{10}^{-05}$ $1.42\text{\times}{10}^{-05}$
-Stage 1+Stage 2 (Distill.) $7.13\text{\times}{10}^{-06}$ $7.13\text{\times}{10}^{-06}$ $1.62\text{\times}{10}^{-05}$ $1.76\text{\times}{10}^{-05}$
-$10^{-1}$ Stage 1+Stage 2 (Basic) $8.18\text{\times}{10}^{-06}$ $1.41\text{\times}{10}^{-05}$ $3.82\text{\times}{10}^{-05}$ $2.42\text{\times}{10}^{-05}$
-Stage 1+Stage 2 (Distill.) $6.72\text{\times}{10}^{-06}$ $9.21\text{\times}{10}^{-06}$ $3.67\text{\times}{10}^{-05}$ $3.16\text{\times}{10}^{-05}$
+$10^{-3}$ PG–DPO (basic) $7.17\text{\times}{10}^{-06}$ $1.51\text{\times}{10}^{-05}$ $1.60\text{\times}{10}^{-05}$ $1.23\text{\times}{10}^{-05}$
+$10^{-3}$ PG–DPO (distill.) $5.20\text{\times}{10}^{-06}$ $9.81\text{\times}{10}^{-06}$ $1.64\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
+$10^{-2}$ PG–DPO (basic) $6.21\text{\times}{10}^{-06}$ $1.38\text{\times}{10}^{-05}$ $1.63\text{\times}{10}^{-05}$ $1.42\text{\times}{10}^{-05}$
+$10^{-2}$ PG–DPO (distill.) $7.13\text{\times}{10}^{-06}$ $7.13\text{\times}{10}^{-06}$ $1.62\text{\times}{10}^{-05}$ $1.76\text{\times}{10}^{-05}$
+$10^{-1}$ PG–DPO (basic) $8.18\text{\times}{10}^{-06}$ $1.41\text{\times}{10}^{-05}$ $3.82\text{\times}{10}^{-05}$ $2.42\text{\times}{10}^{-05}$
+$10^{-1}$ PG–DPO (distill.) $6.72\text{\times}{10}^{-06}$ $9.21\text{\times}{10}^{-06}$ $3.67\text{\times}{10}^{-05}$ $3.16\text{\times}{10}^{-05}$
 Misaligned $P_{0}$
-$10^{-3}$ Stage 1+Stage 2 (Basic) $1.10\text{\times}{10}^{-05}$ $1.41\text{\times}{10}^{-05}$ $70\text{\times}{10}^{-05} 1.1$$18\text{\times}{10}^{-05}$
-Stage 1+Stage 2 (Distill.) $7.71\text{\times}{10}^{-06}$ $5.94\text{\times}{10}^{-06}$ $1.84\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
-$10^{-2}$ Stage 1+Stage 2 (Basic) $7.90\text{\times}{10}^{-06}$ $2.02\text{\times}{10}^{-05}$ $1.46\text{\times}{10}^{-05}$ $1.20\text{\times}{10}^{-05}$
-Stage 1+Stage 2 (Distill.) $6.00\text{\times}{10}^{-06}$ $1.71\text{\times}{10}^{-05}$ $2.21\text{\times}{10}^{-05}$ $1.43\text{\times}{10}^{-05}$
-$10^{-1}$ Stage 1+Stage 2 (Basic) $1.24\text{\times}{10}^{-05}$ $1.93\text{\times}{10}^{-04}$ $5.77\text{\times}{10}^{-05}$ $3.14\text{\times}{10}^{-05}$
-Stage 1+Stage 2 (Distill.) $1.20\text{\times}{10}^{-05}$ $1.90\text{\times}{10}^{-04}$ $7.40\text{\times}{10}^{-05}$ $2.47\text{\times}{10}^{-05}$
-Table 2: Myopic-component RMSE at $t=0$ (tail medians).
+$10^{-3}$ PG–DPO (basic) $1.10\text{\times}{10}^{-05}$ $1.41\text{\times}{10}^{-05}$ $70\text{\times}{10}^{-05} 1.1$$18\text{\times}{10}^{-05}$
+$10^{-3}$ PG–DPO (distill.) $7.71\text{\times}{10}^{-06}$ $5.94\text{\times}{10}^{-06}$ $1.84\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
+$10^{-2}$ PG–DPO (basic) $7.90\text{\times}{10}^{-06}$ $2.02\text{\times}{10}^{-05}$ $1.46\text{\times}{10}^{-05}$ $1.20\text{\times}{10}^{-05}$
+$10^{-2}$ PG–DPO (distill.) $6.00\text{\times}{10}^{-06}$ $1.71\text{\times}{10}^{-05}$ $2.21\text{\times}{10}^{-05}$ $1.43\text{\times}{10}^{-05}$
+$10^{-1}$ PG–DPO (basic) $1.24\text{\times}{10}^{-05}$ $1.93\text{\times}{10}^{-04}$ $5.77\text{\times}{10}^{-05}$ $3.14\text{\times}{10}^{-05}$
+$10^{-1}$ PG–DPO (distill.) $1.20\text{\times}{10}^{-05}$ $1.90\text{\times}{10}^{-04}$ $7.40\text{\times}{10}^{-05}$ $2.47\text{\times}{10}^{-05}$
+Table 4: Myopic-component RMSE at $t=0$ for the projected (Stage 2 / PG–DPO) rules (tail medians).
 $s_{0}$ Method $d=5$ $10$ $50$ $100$
 Aligned $P_{0}$
-$10^{-3}$ Stage 1+Stage 2 (Basic) $0.994$ $0.988$ $0.991$ $0.990$
-Stage 1+Stage 2 (Distill.) $0.995$ $0.986$ $0.990$ $0.987$
-$10^{-2}$ Stage 1+Stage 2 (Basic) $0.993$ $0.989$ $0.992$ $0.988$
-Stage 1+Stage 2 (Distill.) $0.992$ $0.994$ $0.990$ $0.987$
-$10^{-1}$ Stage 1+Stage 2 (Basic) $0.988$ $0.990$ $0.936$ $0.932$
-Stage 1+Stage 2 (Distill.) $0.996$ $0.990$ $0.949$ $0.922$
+$10^{-3}$ PG–DPO (basic) $0.994$ $0.988$ $0.991$ $0.990$
+$10^{-3}$ PG–DPO (distill.) $0.995$ $0.986$ $0.990$ $0.987$
+$10^{-2}$ PG–DPO (basic) $0.993$ $0.989$ $0.992$ $0.988$
+$10^{-2}$ PG–DPO (distill.) $0.992$ $0.994$ $0.990$ $0.987$
+$10^{-1}$ PG–DPO (basic) $0.988$ $0.990$ $0.936$ $0.932$
+$10^{-1}$ PG–DPO (distill.) $0.996$ $0.990$ $0.949$ $0.922$
 Misaligned $P_{0}$
-$10^{-3}$ Stage 1+Stage 2 (Basic) $0.988$ $0.988$ $0.993$ $0.990$
-Stage 1+Stage 2 (Distill.) $0.994$ $0.995$ $0.990$ $0.987$
-$10^{-2}$ Stage 1+Stage 2 (Basic) $0.994$ $0.976$ $0.992$ $0.988$
-Stage 1+Stage 2 (Distill.) $0.994$ $0.980$ $0.988$ $0.989$
-$10^{-1}$ Stage 1+Stage 2 (Basic) $0.990$ $0.005$ $0.668$ $0.851$
-Stage 1+Stage 2 (Distill.) $0.992$ 0.009-0.009 $0.642$ $0.871$
-Table 3: Hedging-direction cosine similarity at $t=0$ (tail medians). Higher is better; negative indicates direction reversal.
+$10^{-3}$ PG–DPO (basic) $0.988$ $0.988$ $0.993$ $0.990$
+$10^{-3}$ PG–DPO (distill.) $0.994$ $0.995$ $0.990$ $0.987$
+$10^{-2}$ PG–DPO (basic) $0.994$ $0.976$ $0.992$ $0.988$
+$10^{-2}$ PG–DPO (distill.) $0.994$ $0.980$ $0.988$ $0.989$
+$10^{-1}$ PG–DPO (basic) $0.990$ $0.005$ $0.668$ $0.851$
+$10^{-1}$ PG–DPO (distill.) $0.992$ 0.009-0.009 $0.642$ $0.871$
+Table 5: Hedging-direction cosine similarity at $t=0$ for the projected (Stage 2 / PG–DPO) rules (tail medians). Higher is better; negative indicates direction reversal.